A Complete Bibliography of Publications in Applied Mathematics and Computation: 2000–2004

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

20 December 2023 Version 1.22

Title word cross-reference

(1 < p < ∞) [MRŽ04]. (2 + 1) [CZ04, ESK04c, LZ04b, LqZ04]. (3 + 1) [ESK04c]. (3, k) [BŞY03]. (β, γ) [ÇD04]. (h, ϕ) [LP03]. (n + 1) [Fen04a]. (p, q, l) [Cin04h]. (S) [WW02a]. (T) [WW02a]. (t, m) [WH00]. (t, n) [YCH04]. (V_{σ}, λ, q) [GEA04]. 0 [JLST04]. 1 [JLST04, SEG03]. 180^{∘} [CMOS01]. 2 [Çel04b, De 02b, KA03b, Mel00, SEG03]. 2ℓ [ÖKOD04]. 2kk^{−1}k^{−n} [Bat04]. 2m [Yan03v]. 2N [SY03, SYK03, Yan03h, Yan03g]. 3 [Juk04, Küç04, Lee04b, Mef02, ÖAD03a, Özk03, ŞBC03]. 3 + 1 [LCXZ04, LJ04]. 3 × 3 [Ana01]. 4 [KA03b]. 6 [HKA04a]. = [dCD03]. ^{+} [MMER03]. ^{2} [XYD02]. _{2} [LCH00, WL02]. _{3} [LCH00]. _{T,S}^{(2)} [Li04a]. A [Bil04, Bil07, LW04e, WZ04e]. A(v^{2}) + Cv = f [MZ04b]. A − ϕ [KML04]. A_{T,S}^{(2)} [WZ04a, ZB04b]. A^{(}d) [Che01c]. A_{MN}^{+} [WWW00]. A_{T} [WW03b, WD03, WW03c, WZ04e, ZW03a]. A_{T} , S^{(2)} [Zha02c]. A_{T,S}^{(2)}, B_{T_{1},S_{1}}^{(2)} [LW04e]. α [CC03a, HZ04c]. AuBu + Cu = f [lLgC02]. Au: = −(k((u′)^{2})u′)′ + g(u) [Has04c]. Ax = b [SZ03a]. B [LW04e]. β [Abb04a, CC03a]. C^{1} [HDZ04, KS02a]. c_{0}^{p}(Λ) [MRŽ04]. c_{2}^{P} (p) [GÇ04]. c_{2}^{P B}(p) [GÇ04]. H_{∞} [PKW04]. D [Tad03]. E^{n} [HIİ04]. E_{2}^{4} [ÇT04b]. exp(z) [GhW04a]. F − γ [MAK04a]. F_{μ} [MAK03]. G [Aru03, SEG03, Ram00b]. GF (2^{m}) [KY04]. H [pFjH03, KML04, Khu02, LHM04, XL04b, XL04c]. H2/H [WBW01]. H_{∞} [PKLW04]. ∞ [ASAI03, Tar02]. k [ASAB03, Aru03, CC01, Kah05, Mil03, TK04, YK03]. k(2, q) [KU03]. K(n, n)

[Waz02d, Waz02e]. K(p, q) [UKŞ03]. L [All04a, MS01]. L^{2} [CY02]. l_{1} [JLS^{+}04c]. L_{2} [YW02, YW03]. L_{p} [Yoo04, dK02]. L_{p}(p ≥ 1) [EsEKEA03]. λ [CD03b]. M [AS03c, AB02b, LW04d, zBxF03, CW02, CL04a, JCZ02, LY02a, LQG04, Ma04a, ÖAD03a, YWY03]. M/G/1 [Mor04]. M/M/2 [AS04g]. \binom{M^{[x]}/G_{1}}{G_{2}/1}

[MANAM04]. M_{T,S}^{(2)} [LW04d]. μ [LRMSVO01]. N [CP04, EA03, SS00, SYK02, Ten00, XC02b, ZC04b, AT04, Aru03, EOM02, EOM03d, EOM08, Guo00, Guo03a, Lac03, MK02, MK03a, ML04a, Mat00, Mil03, MV04, RAH01, Tri04, XCC02, Yan03j]. O(h^{4}) [MSJ04]. O(k^{2} + h^{4}) [Moh03]. O(n^{3}) [LSJR00]. ⊕^{k} [KSL02]. P [Aln04, KS00, LW03a, LW04a, ALO03a, AKÖ04, Dik04, HC02a, HC02b, Kad04, KA03b, KS03b, LOZ02, Noo04a, OK04, UvBP00, WHM04, XX03, Yan03u]. P^{M}∗ [ZY00]. P_{0} [ZG03]. ϕ [CH02]. ϕ_{0} [Sol02c]. q [AGM02]. R^{2} [Zay03e, Zay04a, XD03, Zay02a, ZAH02, Zay03d]. R^{3} [Zay04b, Zay02b]. R_{1}^{3} [Küç04]. R^{n} [DX04, eT04]. S [Abb04b, GLR02, SV02]. S^{(2)} [WW03b, WD03, WW03c, WZ04e, ZW03a]. tanh [FH03, LqZ04]. Θ [KK04, GLR02, XL04b]. t → +∞ [YTM00]. V [RM03, Xu02]. v_{0}^{p}(Λ) [MRŽ04]. |N_{p}|(M, r, q, s) [AET04]. W [RW03, RM03, Wei02a, Wei03a, WWL04]. W^{1,p}(Ω, H^{n}) [JYC04]. w_{0}^{p}(Λ) [MRŽ04]. X + A^{∗}X^{−2^{m}}A = I [ESEA02]. X + A^{∗}X^{−n}A = I [ESAD04]. x = 0 [NM03]. x_{n+1} = (a_{n} + b_{n}x_{n})/c_{n}x_{n−1} [Çin04a]. x_{n+1} = (α − βx_{n})/(γ − x_{n−1}) [YL03b]. x_{n+1} = (α − βx_{n−1})/(γ + x_{n}) [EOAE04]. x_{n+1} = (x_{n−1})/(1 + x_{n}x_{n−1}) [Çin04b]. x_{n+1} = −αx_{n−1}β ± x_{n} [EOAM03]. x_{n+1} = 1/y_{n}, y_{n+1} = y_{n}/(x_{n−1}y_{n−1}) [Çin04d]. x_{n+1} =

a + αx_{n} + αx_{n−1} + ⋯ + αx_{n−k+2}x_{n−k+1}

[EAA03]. X_{n+1} = A/x_{n}^{p} + B/x_{n−1}^{q} + C/x_{n+2}^{s} [EOREA00]. x_{n+1} = α + (x_{n−k}/x_{n}) [HLY04]. x_{n+1} = α + βx_{n} + γx_{n−1}Bx_{n} + Cx_{n−1} [EA04c]. x_{n+1} = α + x_{n−k}x_{n} [EOAM04]. x_{n+1} = ax_{n−1}/(1 + bx_{n}x_{n−1}) [Çin04e]. x_{n+1} = x_{n−1}/(−1 + ax_{n}x_{n−1}) [Çin04g]. x_{n+1} = x_{n−1}/(−1 + x_{n}x_{n−1}) [Çin04c]. x_{n+1} = x_{n−1}/(1 + ax_{n}x_{n−1}) [Çin04f]. X_{n+2} = (1 + X_{n+1})/(X_{n}) [EOEA00]. X ± A^{∗}X^{−n}A = Q [HI04a]. y′′ = f(x, y) [MLG00].

-accretive [JCZ02]. -adic [KS03b].

-approximation [Yoo04]. -ary [AB02b]. -based [KML04]. -body [MV04]. -braids [ŞBC03]. -Coherent [AGM02]. -complementarity [ZY00]. -constant [HKA04a]. -continuity [MAK04a]. -convergence [Bil07, Bil04]. -curve [MS01]. -cycle [Xu02]. -d [De 02b, Lee04b]. -dimensional [CZ04, ESK04c, Fen04a, LCXZ04, LZ04b, Tri04]. -distance [ÇD04, CD03b]. -divergences [LP03]. -Euclid [HIİ04]. -fold [MK02, MK03a, ML04a]. -fractional [SS00]. -function [LqZ04]. -fuzzy [All04a]. -general [GHW04b]. -groups [Dik04]. -irresolute [Abb04a, Abb04b]. -jitter [Ram00b]. -Laplacian [LOZ02, Yan03u, CH02, Yan04d, Yan04f]. -layered [AS03c]. -LCP [ZG03]. -like [Khu02]. -Lucas [Kah05, YK03]. -method [XL04b]. -methods [GLR02]. -Moment [WHM04]. -Monotone [pFjH03]. -norm [JLS^{+}04c]. -open [MAK04a]. -order [Yan03h, Yan03g, EA03]. -orthogonal [AGM02]. -oscillation [LHM04]. -out-of- [Aru03, Mil03, SEG03]. -parametric [Juk04]. -point [zBxF03, LY02a, LQG04, Ma04a, YWY03]. -policy [CP04, Tad03]. -polynomials [HZ04c]. -properties [Cin04h]. -sector [RAH01]. -solution [HDZ04]. -space [Küç04]. -species [AT04, XCC02, SYK02, Ten00, XC02b, ZC04b]. -stability [Sol02c, LRMSVO01, Aln04, GLR02, XL04b, XL04c]. -stable [KS00, LW03a, LW04a]. -Step [ÖAD03a, Özk03, CW02, KA03b]. -strong [MAK03]. -strongly [MAK03]. -summable [GEA04]. -system-based [SV02]. -th [EOM08]. -Torus [BŞY03]. -transform [YW02, YW03]. -type [WA01]. -valent [Noo04a, OK04]. -valently [AKÖ04, Kad04]. -weighted [RW03, Wei02a, Wei03a, WWL04].

145 [dCD06]. 147 [KAK05, KAK06, Kas06]. 149 [Sab18].

2D [BRVI00].

3D [AS01, GZ00].

752 [HES05].

806 [Sab18].

= [ZTD04, dCD06].

Abel [Mar04c]. abnormal [AS03b]. above [AS03a, BS04, Sel03, SM04a, ZS02]. abrupt [KH01]. Absence [Kok03]. absolute [Li02, RS04a, Sav04a]. absolutely [Cho02b]. absorption [ZLS04]. absorptions [CY01c]. Abstention [Sta00]. abstract [Fu03, Fu04, LM04a]. Accelerate [Tia03]. accelerated [HMM03]. accelerating [Thu04a]. Acceleration [SAV04b, ES03a, GH01, RdL01]. according [JLS^{+}04b]. accretive [JCZ02]. accumulation [LWZ00]. accuracy [ASM03, GZ00, HT00, Kho04, WX01]. Accurate [ML03, Kor03, Moh03, MSJ04]. acid [BTBI03]. acidified [Gho03]. acoustic [AM00, AA04d, BA04, Gzy01]. across [Lun00]. action [HZ04a]. actions [Ver00]. activation [ZZC04]. active [ES03b]. active-trust-region [ES03b]. Adaptive [HZL02, Yas03, dlS03, AQD04, HML^{+}02, KM02, NN04b, SK02]. Addendum [dCD06]. additional [AS04g, CP04]. additive [Che01b, Isk04]. ADEOS [KIY00]. ADEOS/POLDER [KIY00]. ADI [Pov02]. Adiabatic [Bis04a]. adic [KS03b]. adjoint [Che01b, DMT02, Ras02d, Sal03a, Sal03b, ZZ04b, de 00]. Adjusted [KM02]. ADM [ESK04a]. admissible [ZN04]. Adomian [Abb03, Abd03b, BB02a, BB02b, BB02c, BJ03, BB03, BD04a, BJ04, BSJ04, BJS04, BBV04d, BBK^{+}03, BTBI03, BBI03, BBI04, BI04a, BEAB04, CL04c, CS03b, Deh04a, ESAA03, ETBAN04, IRR04, IRS04, KES04a, MMR01, Waz00b, WES01, WG04b]. adsorption [LLCC03]. Advanced [SDR03, SDR03]. advantage [SYK02]. advection [BD03, CMOS01, Deh04f, Deh04h, Kha03b]. advection-diffusion [CMOS01]. advective [CJ04a, HB00, HT00]. advective-dispersive [HB00, HT00]. advertising [MT04]. aerosol [AL00, KIY00, LCH00, MTK^{+}00, MSO00, SM00a]. affected [EG04c]. Affectons [Ada03]. affine [LL03, MS04a, NS03]. affinor [MCS04]. Africa [SP03a]. against [EB03c, EBSAG04, MMP03]. age [Far02, LHL04, LLW04, QMWAZK04]. age-dependent [LLW04, QMWAZK04]. age-structure [LHL04]. ageing [Koł01]. agent [TE03, TE04, Vou03]. agent-based [TE03, TE04]. agents [TE03, TE04]. aggregation [BR01]. agreement [CJ04b, HWWM03, LL04b, LLL04, Tse03]. aided [RHB04]. air [WL02]. airspace [LLCC03]. al [XY04]. Alessandro [Ric03]. Alexander [BŞY03, ŞBY04]. algebra [Bay04, II00, PZJ03, UK03]. Algebraic [GhW04a, HC04a, Aya04a, BJ03, BH03, ÇB03a, ÇKB03, ÇB03b, Çel04a, ÇB04c, Çel04b, CG03, ES02b, EM03h, He03c, Hos04, KES04a, Les01, Liu03c, Sha04a, TS02, Wan00b, WJ00, WS01, yWShX02, Zha02a, gZC04]. algebras [BEV02, BFN01, BFN03, ENR04, Ras02a]. Algorithm [SM00a, ADG03, AGS03, BBK^{+}03, BW04, CL04c, CS03b, CPTZ04a, CC04c, DKX00, DDX02, Din00, Din03b, DR01, EME03, EMR03, ES03b, GHW04b, Gar01a, Gar01b, GMGC01, Has06, He04e, Ji02, KSS02, KS03a, KSJ02, LSJR00, Lia03b, LK04, qLzWcC03, MS04b, Mor01, NK04, OASM04, Ram03a, RS03, Ras04a, Sah04a, SH02, Sta03, Sug00, TLX04, Wan04b, Waz00b, Waz01f, Waz01h, Wei00c, WWW00, WW02a, Wu04a, YB03, Yam05a, Yan00a, Yan02a, YD04, YCC02, YL02, Yu04, ZW04a, Zhu04a]. algorithmic [KM03a, Mou04]. Algorithms [Bar04, Din04d, CHL02, DM04, Din04e, EM04a, Fad04, GLWY04, JJ02, KAAD01, LC03b, LCN04, Lun01, Mus00c, NG02b, NN04b, NS04, PPS04, Ram04a, RG03, SK02, Thu04a, WBW01]. Allee [SPH04]. allocation [JLSS04, MT04]. Almost [GB03, ZZ02b, AYW04, Bar03, CC03d, CC03c, CHC03, Fen03, GM00, Kha03a, XCCC04, Zha04c]. almost-periodic [Bar03]. along [AO04b, EB02b, Elb03, GP03, Has03b, KST00]. alternate [BBK^{+}03]. alternating [CP03b, CPTZ04b, gLgW04, MMER03, YD04, wYjSlZ01]. Alternative [BRVI00, Lun01, BS00, JLSS04, Liu04f, VCD04, Yan04f, ZY00]. alternatives [EB03c, MMP03]. ambiguity [YL02]. among [ENR04]. Amplitude [BD04c, DE00, DF02]. analyse [EA00, VCV01]. analyses [MG03a]. Analysis [CY02, CC03f, HT04, Mog04, Mur03, SZ03a, Sil02, Sil03, Sil04, Yan00b, Yan00a, YA03, YDW02, AA00, AL03, Cao03b, CO01, Cha04b, CBK00, EORI01, Gha01, Gha03, GL01, Has04b, HLO02, He04b, II00, JSD04, JMV04a, JAK04a, JMV04b, JLS^{+}04c, JLM04, Jah03, JY04, KS04a, KS03b, KSC02, LRMSVO01, Lia04, LM04b, Liu00, LC04c, ML04d, NN04a, OSL03, Ram04a, RHB04, RG03, SMF04, SB04a, SC02, SWY00, Sou03, SL04c, gWxZ04, Xu04, YB00a, ZC04b, ZW03b, ZW04a, ZW04b, ZClC03]. Analytic [Abd04, Gin04, Sai02, SZK04, SP03b, Waz01b, vAG03, AZ04, BH02, Lia03a, Lia03b, LL04c, Noo04a, Orh03a, Orh03b, Tad01, Waz03b, Waz04a, WG04b, Wol02, Xu04]. Analytical [DD02, DWC04, EEB03a, EE03, EE04a, ENAAM01, KU01, KW04a, NZP^{+}04, Sha02, Dem03b, Dem03c, GM03b, He03a, Kha03b, MK02, ML04a]. analyticity [AS02a]. ancient [He04f, He04e]. angle [OT04, YD04]. angle-ply [OT04]. angles [KST00]. angular [EG03a]. animal [HA02]. anisotropic [ALO04, KNJ04]. annular [CM04a, Ram00a, Ram00b, Ram00c, Ram01a, Zay02b, Zay02a]. annular-bounded [Zay02a]. annulus [RH04]. Anonymous [WC04a, JL04]. anti [BCI03]. anti-maximum [BCI03]. ants [ML04d]. any [AS03a, Kay04g]. AOP [KY04]. AOR [YS03]. apology [The03]. Appell [Pom01]. Appl [AaAaZ00b, AaAaZ00a, Bil07, CC03c, EOM08, KAK05, KAK06, Kas06, Kum09, Sab18, dCD06]. Applicability [LMG00]. Application [Ald04, Bah04, Bay04, ÇB04a, CC04a, CJ04a, DSI04, Deh04a, RES04a, UK03, WC02, YAYA03, Abd01a, AY04a, BN02, Boy03a, De 03, DÖ03a, ER03, ESK04a, EP00, GM04a, HM04, HOL02, HV01, JAK04b, Kar04, Kay04a, Kay04b, KU03, LCXZ04, LZ04a, LC03a, LW00, PdR01, RO01, SS00, STHN02, SP03a, VK04, Wei00a, WZ01b, Yan03t, ZW03a]. Applications [Aya04a, FH03, HL00, LZ04f, ÖAD03b, SZW03, Zed02, AEL04, AS04d, AY04b, AO01a, AGS03, BT00, ÇB04c, CS03a, CCKS02, CLX02, CT04a, De 04, EEYK03, Esc03c, Fen04b, GhW04a, GNS02, HY04, Has02a, He04c, Jam01, Kır04, KÖ04, Lan04b, Lee04a, LC01a, LCN04, Mağ04, MCS04, ML04c, Olu04, Özd03b, Özd03a, ÖK03d, ÖKOD04, RS04b, Sen01, SAV04b, STB03, SS01, Tia04a, WW02b, WL04d, XL04a, Zay02b, Zay02a, Zay03a, Zay03c, Zha02c, gZC04]. Applied [HV08, HES05, MB04, Mus00a, Mus00b, Pan08, Yam05b, AR04a, BBV04a, KES04a, Lee04b, Liu04g, Waz00a, Waz01e, WZ02, Pop04]. Applying [CL03a, CL04b, JWL03]. approach [AO04c, Bel04, BMMRS04, BK02, CM04a, CHL02, Che04a, Cho02a, DG03, Dog04, EBA03, EKE03, EMY04, EM03d, FQ04, GNS02, GM03b, GM04b, Gzy01, GV04, HS02, He03d, He03e, He03f, He03g, He03h, Isk04, JLSS04, KST00, KB04a, Khu04, KM03a, LWT00, LS04f, Lon00, Mat01, Mou04, Mun03, Mus00e, NN04a, yN04c, ÖAÖ03, PKW04, PKLW04, RC04b, SJM04, SV04, Tag00b, Tag01c, UvB01, VCV01, Vou03, Waz01c, XH04, YR01, Zay03e, Zay04b]. Approaches [WW00a, KML04]. appropriate [LS04a]. approval [Ino03, Wan04d]. approximants [AÇB03, BN02, ÇAB03, ÇAB04, GhW04a, ŞÇ03b, Thu03, Thu04b]. approximate [Abd04, CMM02, CC00a, EsEKEA03, Fen04a, Has04c, Sha02, Wu04a, YY04a, YS00a, Zha02b]. Approximating [Bel01, KML04]. Approximation [HZ04c, QY00a, AAAE03b, AD00, ALO03b, AÖE04, APS04, BC04, Boy03a, Bry02, Bİ04b, Çat03, Che01c, Ehr02, EEK03a, EEYK03, Hau04, HZZ03, IR04, IES04a, IES04b, Kho04, Kor03, LOS02, Liu04i, Mef02, Mun03, Nar02, OSL03, Pap00, QY00b, RW03, Sah04a, ŞÇ03b, SY03, WW01, WQ03, WW03c, Wei03b, WZ04e, XL03, Yoo04, ZLZ03, Zhi04, dK02]. approximations [AÇB03, BK04, Cao01, ÇAB03, ÇB04b, ÇAB04, CAY01, De 03, FL02b, Gzy01, Ism04, JY00, SAV04b]. Arbitrary [ÇB03a, Bog04, Cap01, ES00, GH04, SB04a, SP03b, ZW03b]. arc [LBE00, SK02]. arc-length [SK02]. arch [Yan00b]. Archimedean [JK04a]. architectures [KY04, LB01, WH02]. area [BRVI00]. area-perimeter [BRVI00]. areas [Hİ04b]. argument [Aln04, AÇB03, ÇAB03, ÇAB04]. arguments [DFF04, LHM04, LQL03, LG03, LGZ04, LG04c, SW04b]. arising [BE02, HES05, HAM04, Khu03b, MMR01, MS04b, MCC01, Pet03, SD02b, WBW01]. arithmetic [EN01, Gal03, KY04, SZW03]. arithmetic-geometric [EN01]. arm [HkT03, HT04]. ARMA [JKP03]. ARMA-representations [JKP03]. Arnoldi [NK04, Wu04a]. arrival [ASAI03, CP04, CM04b]. Artificial [SWL00, EE03, EE04a, EE04b, Gop04, HIS04, MCS00, dGKG^{+}00]. ary [AB02b]. Ashraf [Pan08]. aspects [Hil04, Tag01a, Tag01c]. assess [Bon02]. assessing [JSdN04]. assessment [Sai02]. assignable [Jol00b]. assigned [Tag02a, Tag02c]. assignment [CHC04, Ram03a]. assisting [KS01]. associated [Bae04b, CCS03, DC03, DGR04, DK04, EM03e, Kad04, KKSY04, LS04d, RS02a, RS02b, Sri03, TS01, Wün03]. association [ECL02]. associative [CD03a, Cao03a, ECL02, LC04b, TCL02]. assumptions [Par02]. assurance [JK04a]. asymmetric [CBK00, Yan03f, Yan03m]. asymmetrical [ECL02]. Asymptotic [ALP03, BG04, EB04a, EBMA03, JY04, LC04a, Ona02, SR02a, TB04, VR03a, WL04a, YG04, AZTK00, Cao03a, EOAM04, EOM04a, GXL04, HC03, KKSY04, Khu03b, Kur02, Lac03, LZ04c, LL00, LC04b, LM02, MR04, NR02a, Sad02, TT04, Tun04c, VR04b, VR02a, XCC02, YLEM04, ZL01, ZC04a]. asymptotic-numerical [NR02a]. asymptotics [Kow02, Wan01]. Asymptotology [He04a]. atmosphere [Mat00, Mat01, MSO00]. atmosphere-ocean [MSO00]. Atmospheric [TM00, TRC03, LCH00, Sel03, ZS02]. attacks [WL04c]. attitude [YL02]. attraction [BDGG04, Din04b, Din04a, Kur02, Moh00]. attractivity [CC03d, CC03c, EOM03c, EOAE04, HLY04, Sak03b, Sak03c, SC03c, XCCC04, XC02b, YL03b, YL03a, YL04c, YCME04, Zha04c]. attractor [WZ01a]. attractors [MZ04a]. attribute [Liu04e]. augmented [GC01]. authenticated [CJ04b, HWWM03, LL04b, LLL04, XY04]. authentication [LHL03, Pei04, WL04c]. Author [Ano01a, Ano01b, Ano01c, Ano01d, Ano01e, Ano01f, Ano01g, Ano01i, Ano02i, Ano02a, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano02g, Ano02h, Ano03m, Ano03h, Ano03i, Ano03l, Ano03j, Ano03k, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano04a, Ano04g, Ano04h, Ano04i, Ano04n, Ano04b, Ano04c, Ano04j, Ano04k, Ano04d, Ano04l, Ano04e, Ano04f, Ano04m]. authorities [WH02]. authority [JL04]. autocatalytor [AGES04]. automata [Ada03, HM04, KKM04, Ver00, Ada01]. automated [HV01]. Automatic [BŞY03, ŞBC03, Asa04c]. automation [IN04]. automorphic [ÖKOD04]. Autonomous [Jam01, CLW02, FZ04a, He00, HMM03, Ten00, TT04, Tun04c]. auxiliary [ADAAM03, Ahm04b]. AV [ZTD04]. availability [Ana03, AG04]. avalanche [GS02, KW04a]. avalanche-semiconductor [GS02]. average [HHK01, Kat01]. Averages [Yan03b, Mař04b]. axially [EEES03, ESEE02]. axiom [He04h]. axioms [All04a, MAE04b]. axis [YÇA04]. axisymmetric [DWC04].

B [Bah04, DSI04, HKA04b, HL04a, KENM03, KE04a]. B-spline [Bah04, KE04a]. B-splines [DSI04, HL04a, KENM03]. Babuska [TM03]. backward [Aln04, YÖ00, YYS03]. bacteria [HES05, HAM04]. Bahadur [ADAMM03]. Baker [KL02]. Balance [Fen04b, Sen01, Ino02a, MWAF02, QBK02]. balanced [BCT04, LRHRD01, Wan02]. Balancing [GLVW00, LC02b]. balking [AS04f, AS04g]. BAM [CHC03, Li04g]. Banach [AAMEST04, BS01, sC03a, CC03b, Dar03b, DG04, Guo00, Guo02, Guo03a, Guo03b, Guo04, Han01, Hua03, JCZ02, KM03b, Liu03f, ZClC03]. banded [GS00, Gar01b]. bank [JAK04b]. bar [AÖE04]. bare [QBK02]. barrier [Liu04e]. baseball [Kat01]. based [AhL00, AL00, AV04b, CMM02, CWS02, CCH04, DG02, GHL00, Has04b, HM00, HLL03, JBS04, JT04, KML04, Kee03a, KLK02, KG03, Kum03b, Kwo03, Lee00, LHL03, Lie04, qLzWcC03, MAR04a, MMP03, NV01, Par02, Par04a, Pov02, Ram04a, SV02, TE03, TE04, TLX04, WLX00, WH00, WH02, YDW02, Zha01, dGKG^{+}00, dlS03]. bases [BFGG04, YAYA03]. Basic [LZL00, AEC03, ÇB04b, RES04a, Sar02, Ver03a]. basins [Din04b, Din04a]. basis [ÇB04a, HS01, LNW03, WHG02, Yoo04, gZC04]. batch [CP04, CM04b]. Bayes [EG04d, PS00, Sar03a]. Bayesian [ASM03, CBK00, EG04e, EG04a, GL01, Jah03]. be [LRHRD01]. beaches [Bon02]. beam [CC00a, CC00b, EA04a, JY00, MdS04, MdOPF04]. beams [TY04, WLX00]. bearing [AR04a, AO01b]. bed [DKV04]. behavior [ALP03, EOM04a, GXL04, KS04a, TE00, TE04, WL04a, YL04b, YTM00, ZS04a]. behaviour [Abd03a, AS03a, DD00, EOAM04, EA00, Has00, HS00, Ona02, Sad02, Tar02, TT04, Tun04c, YG04]. Bellman [AC01, BC04]. Benard [Dem03a, Mur03]. bend [CMOS01]. bending [KNJ04]. Bernoulli [Mus00a, Mus00b, CM04b, JK03b, JKP04, KS03b, MADT03, Mus00d]. Bernstein [Bİ04b, HZ04c, Sah04a]. Bertrand [KG04]. Bessel [DRP04, Kow02, LCS03, YW02].

Best [GG00, HZZ03]. better [AY04a, HLLL03, LM02]. between [AR04b, AAMK01, AH02, BS00, CCY03, CYC03, EM03b, EM03c, EMED03, EEE03, Lay02, MS04d, MRŽ04, NG02a, Ton00]. Beyond [CY01a]. Bézout [LLB03]. bi [CD03a, Cao03a, DGR04, LC04b, Mae01]. bi-criteria [Mae01]. bi-directional [CD03a, Cao03a, LC04b]. bi-orthogonal [DGR04]. biased [Abu00, Dis01, Olu03a]. biased-nonlinearities [Abu00]. bidiagonal [Gar01a]. bidirectional [ECL02]. bifurcating [LwWW03]. Bifurcation [Li03c, BF03, KSC02, Ram01a, SWY00, ZLZ03]. bifurcations [GR03, Sta00]. biharmonic [Moh03, XL03]. bilingualism [EOI02]. binary [Cha04a, SSPA01]. binomial [ASAB03]. bio [Dun02]. bio-reactor [Dun02]. biochemistry [YAOY03]. biological [BB02c]. biorthogonal [RS02a]. Biorthogonality [Khu03a]. biparameter [yWShX02]. birth [LHL04]. bistable [Mog04]. bitopological [MAK04a]. bivariate [ADAJM03]. BK [LqZ04, MRŽ04]. Blasius [Has06, He03d, Wan04b]. blind [Fan03]. Block [KB04b, Sab18, ZZ02a, CC03e, GC01, qJkSlS03, KSJ02, TW03, Wu04a, Zan00, MK03b, MM04b]. block-bordered [Zan00]. block-rank [TW03]. Blood [Els04, ES03a, Mek04]. Blow [CX04, eT04, Waz01a, CT03a, Che04c, JLM^{+}03]. Blow-up [CX04, Waz01a, CT03a, Che04c, JLM^{+}03]. blowflies [SWY00, SZ01]. Blowing [TE02a]. Blowing-up [TE02a]. board [Ano04r, Ano04s, Ano04t, Ano04u, Ano04v, Ano04w, Ano04x, Ano04y, Ano04z, Ano04-27, Ano02j, Ano02k, Ano03n, Ano03o, Ano03p, Ano03q, Ano03r, Ano03s, Ano03t, Ano03u, Ano03v, Ano03w, Ano03x, Ano03y, Ano03z, Ano03-27, Ano03-28, Ano03-29, Ano04o, Ano04p, Ano04q]. bodies [ES04]. body [AAHAD04a, EGE02, EG02, EGE03, EG03a, EG03b, EG04c, EGT04, EG04f, ES03a, MV04, TCL02]. Bond [Ghe00, DDX02]. bonded [OT04]. bone [AAA02, BGWX03]. bones [ENAAM01]. Boolean [YS02]. Bootstrap [Ami01, MPS04]. bordered [Zan00]. both [EA04a, HL04a, WSX03, ZW03c]. Bott [CLX02, WX03a, XC02d]. bottom [Rao02]. bound [DR01, HLLL03, LW03b, Ras02c, Tag02c, Wei02c]. boundaries [CY01c, ESH02, Kru03, Lun00]. Boundary [EKE04, Liu03f, Mas03, AR03, AS02b, AD02, ANS02, ALO03b, AZ04, ASN03, AO01b, AO04c, BPJ03, Bae04a, Bae04b, zBxF03, BH01, CPL00, CH02, CH03, Çağ04, Cap01, CCS04a, CY02, Cha04b, CL01, Che04c, Cho02a, De 02a, De 01, Deh00, Deh02, Deh03d, Deh03e, Deh03g, Deh04a, Deh04c, DMT02, DH04, Dur04, EGBH03, ERN01, ES02a, EEK03a, EESF04, EK04, FL02a, FNO04, FZ04b, FS04b, GGB03, GB02, Guo03a, Has00, HS00, HS02, Has03a, Has04a, Has04b, Has04c, He03g, HZ04b, Hlo04, HA04, HHZ04, IH01, IES04b, Jan04, Jay03, JM02, JGW02, JJ04, KP02a, KA04, KR03a, KR03c, KR03b, KR04, KA03c, Kha04, KA00, Kum02, Kum03a, Kum03b, Kum03c, Kum03d, Kum03e, Kum09, Kuo04, LM04a, Lan04a, Lan04b, Lee04b, LL04a, Li04f, LC03b, LY02a, Liu02a, LY02b, Liu03g, Liu03a, Liu03b, Liu04b]. boundary [Liu04a, Liu04c, LQG04, ML04b, MdS04, Ma04a, MCC01, Mef02, MSJ04, NK03, NR02b, Ode02, Pan08, PB02, Pen04, Ram03d, RM02, RKS04a, RR02a, RC03, RC04b, RC04c, RG03, SH02, Sha03, SR02b, yS01b, SV02, ŞÇ03b, SD02a, SV04, SY03, SL03b, SL04d, TM04b, TS01, VR03a, VR04b, VR02b, VR03b, Waz01a, Waz01g, Waz01h, Wei04, WJ04, YL04a, Yan02a, Yan02f, Yan03i, Yan03h, Yan03l, Yan03u, Yan03v, YD04, YWY03, Yao02, Yao03a, Yao03b, YZ04, YTM00, Yür04a, Yür04b, Zak04a, Zak04c, Zay02a, Zay02c, ZAH02, Zay03b, Zay03a, Zay03d, Zay03e, ZI04, Zay04a, ZLS04, dJ04]. boundary-layer [Kuo04, TS01]. boundary-value [AZ04, zBxF03, KA03c, NR02b, TM04b]. bounded [AAMEST04, BDD00, Din03a, EA00, HSE03, Özd03a, TE03, Yan04i, YA03, Yu04, Zay02a, Zay03b, Zay03a, Zay03c, ZClC03]. Boundedness [Yan02b, Yan03f, Yan03c, Yan03d, Yan03e, Yan04d, Yan04b, Yan04c, ZS04b, EA03, Tun04b]. bounding [Wol02]. Bounds [DW02, Bai00, Ehr02, Gün04, Kol02, LC03b, Olu03a, TB02, ZW03a]. Boussinesq [DD02, ESK04c, Rao02]. Boxed [Har03]. braids [ŞBC03, ŞBY04]. brain [IYM00, MCS00, MIM00, SWL00, dGKG^{+}00, CS00, dGKG^{+}00]. brainware [MCS00, TKK00]. branch [DR01, DNS03, Waz02d, Waz02e]. branches [Boy03a, JAK04b]. Branin [ZG02]. Bratu [Boy03a, Khu04]. break [Ram04c, Ven03]. break-up [Ram04c]. Brezzi [TM03]. broadcasting [WH02]. broanded [Gha01]. Broer [LqZ04, LZ04b]. Broyden [CC03e]. Brusselator [Waz00a]. BSB [TCL02]. BSP [CPTZ04a]. bubbles [BA04]. Building [dGKG^{+}00]. bundle [CS03a]. bundles [MS04a]. Burger [IRR04, IR04]. Burgers’ [Bah03, ÖAÖ03, AÖ04a, Dem04, Dog04, ESK04d, KES04b, Kay04g, KE04c, LM02]. Burgers-like [KE04c]. Burgers-type [Kay04g]. Burridge [He03e]. bursty [EP00]. busy [Tar03]. BVP [CDH01, HA04]. BVPs [LNS04, LS04b, LS04c, ZKSD02].

calculate [ŞBY04]. calculating [BJ04, CS03b, MK02, Waz00b]. Calculation [Yüc04, BŞY03, KV03, SLC04, ŞBC03]. calculus [ESG03, EEAES01b, LTS02, OK03a, RES04b, SS00]. CAM [CS00, dGKG^{+}00]. CAM-Brain [CS00, dGKG^{+}00]. can [LRHRD01]. cancellous [BGWX03]. cannibalism [Sol03b]. Canonical [CP03a, GM00, BRS04, ESEE02]. cantilever [EA04a]. capacity [Tad01]. capillary [Els04]. carbon [WL02]. cardinal [Ehr02]. Carlo [SAV04b]. carrier [GCSS04]. carrier-dependent [GCSS04]. carrying [Tin01]. Cartesian [STHN02]. cascadic [ZH04]. case [ÁR02, AR04c, BC04, Tag02a]. cases [BBK^{+}03, LC02a]. cash [JL04]. Castelnuovo [Ras02c]. catalysed [DXL02, YAYA03]. Cauchy [CKNU00, Gün04, RKS04b, Sev02, Şim04b, SB03, TE00, TE02a, TE02b, TB02]. causes [Jol00b, Sta00]. causing [LLCC03]. Cavalieri [He04h]. cavity [AAAaZ99b, AAHAD04a, AaAaZ00b, AAA02]. cell [AGES04, Ald04, Sal03b]. cellular [CC03d, CC03c, KKM04, LH04b, MG03b, Zha04c, ZC02, ZZC04]. centers [Gin04]. Central [XFL04, Liu04d]. centroidal [DG02]. centrosymmetric [Liu03h]. Certain [LTS02, LCS03, LS04e, Sri03, AY03, BE02, Bil03, EL04a, KA03a, Li02, LM03, Li03b, LHM04, MRŽ04, Noo04a, OK03a, Ras02c, RS02a, Sad02, Sad03, Sad04b, TT04, Tun04a, Tun04b, Tun04c, Yan03q, ZZ04a]. certified [HW04, HL04g, Sha04b, TJC03]. chain [GP03, MG03a, Mun03, SMQ04, WW02b]. chains [HHK01, LCN04]. Chandrasekhar [NHCJLLP04, Khu02, Sug00]. Chandrasekhar-type [NHCJLLP04]. Chandy [NK01]. Chandy-Misra [NK01]. change [Coo04, GL01]. changeable [RKS04b]. Changes [Hİ04b, KH01, LPS02, VCV01]. changing [ALO03a, NG02a, SD01, YWY03]. channel [Cha04b, CP04, CMOS01, Man03, Rao02, Rao04a, YR01]. channels [AS04f, LH04a, Mek04]. chaos [Hab04a]. Chaotic [AE04, Hab04a, Yas02, Yua03]. characterisation [AAH04]. Characterising [FL02b]. characteristic [SLC04, UvB01]. characteristics [KEB03, MSO00]. Characterization [Ino02a, You04, ZB04a, AY03, Gin04, Ino02b, Wei02a, Zha02c]. Characterizations [CGVC04]. characterize [Tia04b]. charts [Jol00a, Jol00b]. cheater [Tse03]. Chebyshev [AD04a, AS03d, ABEM04, Bar04, Boy03a, Dat04, EKE02, EEK03a, EEYK03, HIS04, KR03b, Keş03a, KEE03b, OIO03, PR04a]. Chebyshev-like [Dat04]. chemical [Bao02, ÇKB03, Çel04a, LMG00, MMR01, Thu04c]. chemicals [Hel01]. chemostat [EOM04a, Kan00b, ZL03]. chemotaxis [Dun02]. chemotherapy [Gum02]. Chen [Yas02]. Chengtian [He04c]. China [JK04b]. Chinese [He04f, He04e]. cholera [PdR01]. Cholesky [gWxZ04]. chopping [KS03a]. Chua [Yas03]. CHW [CHC04]. circle [CL03a, MM02a]. circuit [Yas03]. Circulant [qJkSlS03, NG02b]. Circulant-block [qJkSlS03]. circular [AAAAZ01, enNAaA02, AZ02, He03b, Yan00b, Zhu04c, Zhu04b]. clamped [KNJ04]. class [AY04a, AY04b, ÁR02, AR04c, AD04b, CCP03, Che04a, CLC04, CMF04, ER03, ES02a, GKK02, Guo02, Guo04, HZ04a, HS03, HGS04, HL04f, IKS02, Iwa01, JJ02, KR04, KA03b, KS03a, Kok03, Kum02, Kum03c, Kum03d, LC01a, Li03c, lLgC03, LZ04d, gLgW04, Mil03, Noo04c, OOAA04, Orh03a, PJ04, Par04b, RS02a, RCS03, RK03, RKS04a, Ras04f, RKS04b, SW03, WG00, Wan00c, Wan04c, WG04a, Wei04, WW00b, Wu03, lWCH03, lWyCjT03, XD03, YL04c, YGL01, Yan03c, YC03, Yan04h, YH04, Yür04a, Yür04b, ZC03, ZLZ03]. Classes [ZK03, IG03, LTS02, LCS03, Liu04h, Noo04a, RS02b, Sri03, Waz04k]. classic [Deh03g]. classical [AM02b, De 01, Has06, KS02b, VCD04, Wan04b]. Classification [LR04, Pra03, YCC02, Yür04a, Yür04b, ZZ01]. Classifications [Li04b, OLT04]. classifying [BFN01]. Clifford [AEC03]. climbing [JJ02]. close [OK03c]. close-to [OK03c]. closed [AD04b, CO01, HA02]. closed-loop [CO01]. cloud [AL00, BMMRS04, Rad04]. cloud/aerosol [AL00]. cluster [ER02, KM02]. Clusterability [Ino03, Ino02a, Ino04a, Ino04b]. co [AAMEST04, WL02]. co-dimensional [AAMEST04]. coalitions [Ino00b, YIN00a]. coarse [qLzWcC03]. code [Liu03c, Liu03d]. codes [GLM04, GMGC01, LS04f]. coefficient [Ahm04a, Gar01a, Has03a, Has04a, Has04b, gLgW04, NZP^{+}04, Waz01b]. coefficients [AS03d, AV04a, BCI03, ÇB04b, ÇB04d, CSD01, CC02, CC03d, CC03c, CC04b, GV04, HS02, Kad03, KA03a, KÖY03, KM03a, Moh04, OK03a, Orh03a, Orh03b, OK04, Sak03c, Sak03d, SH02, Waz02c, WG04c, WL02, YY04a, YY04b, Yüc04]. Coexistence [KL03, ZL03]. Coherent [AGM02]. coil [KB04a]. coinciding [Tag03a, Tag03c]. collapse [SG01]. collective [Ada01]. collinearity [Wan02]. Collocation [LNW03, EHM03, GHSJ00, HS01, KS02a, Ona02, Wan00b]. color [TRC03]. coloured [NCÁHCLP04]. column [ZTD04]. combination [EEB03b, JK04b]. combinations [ÖÖ04]. combinatorial [HV01]. combined [HEM03]. combining [ADAMM03, ADARAM04]. combustion [BE02, Bao02, Ram00c]. Comment [Fen04b, HV08, HES05, Kah05, Kah06, Pan08]. Comments [DGR04, DRP04, Has06, Pop04, Zha02a]. commercial [JAK04b]. committees [YIN00a]. communicable [PdR01]. communication [Sug00]. communications [WC04a]. commutative [ZZ04b]. Compact [Waz02a, Waz03c, Waz04b, Waz04c, Mak04b, Man03, Waz03f, Waz03g, Waz03j, Waz03k, Waz04f, Waz04h, Waz04k]. compaction [(Mu04, Yan02h]. Compactly [WHG02]. compactness [BRVI00, Sor01a]. Compacton [Waz03e, Waz03d, Waz04d, Waz04a, Waz04i]. compacton-like [Waz04i]. Compactons [Waz02b, Waz03a, Waz04e, Waz02d]. compactum [YÖ00]. Comparability [YIN00a]. comparative [EM04a, YAOY03]. Comparing [ESK04b]. Comparison [HOL02, He04b, AKKN04, CH03, CP03b, DM04, EB02a, ESAA03, MY03, SM02b]. competing [Ten00]. competition [Che04d, Kan00b, ZC04b]. competitions [SYK02]. competitive [Gum02, SC03c, Ten02, Tin01, XCC02]. complementarity [Yu04, YS03, ZG03, ZY00, ZL04a]. complements [Tia04a]. Complete [IN04, Ma01, MS04a, QBK02]. completely [Din04d, LK04]. completeness [Şim04a]. Complex [BFN03, Ram03b, BFN01, ENR04, Jam01, Kad04, LB01, Li02, LW04c, OK03c, Ram03a, YE04]. complexity [CD02]. component [EG04e, EG04d, JAK04b, JAK04a, Rol02]. components [AS02b, Koł01, KS01, SEB03, Sar03b]. componentwise [bSI01]. composite [AAAaZ99a, AaAaZ00a, ALO04, FRRSCS02]. composites [DYH04]. compound [Kay04g, LC02a]. Compression [Bar03, HL04a]. compromise [MJCM03]. Comput [AaAaZ00b, AaAaZ00a, Bil07, CC03c, EOM08, KAK05, KAK06, Kas06, Kum09, Sab18, dCD06]. computable [Liu02b]. Computation [HV08, HES05, Mus00a, Mus00b, Pan08, Yam05b, Aru03, CJB02, CZLZ04, CWZ02, Deh04b, Der03, Fan03, GM04b, HA03, HLL03, KI04, Kor03, sLqZ03b, Liu04i, Mef02, Nys01, ST03b, UvBP00, UvB01, WL04d, XLLL03, Pop04]. Computational [PPS04, SR04a, VR03b, vBU02, BBV04b, BEV02, Fad04, GM03b, Les01, Nys01, VR02b, Waz01c, Yen04]. computationally [De 02a]. computations [KEB03]. compute [ML04a, Ras04a, Ras04c, Ras04b]. Computer [GMGC01, RHB04, Bay04, PZJ03, ŞBY04, UK03]. computerizable [SJM04]. computing [BBK^{+}03, BW04, CM04a, ES02a, GKaM01, IYM00, KST00, LLB03, MCS00, MIM00, NK04, Wei00b, Wei00c, WWW00, WW02a]. concave [Sla03, YD04]. concentrated [AV04a]. concentration [Asa04c]. concept [Ino02b]. concepts [AS04f]. Concerning [HZ04a]. concomitant [AAMK01]. concrete [BB03, WZ01b, WZ02]. Condition [WX03a, WWW03, WW03a, Deh01, Deh02, Deh03b, Deh04c, Deh04d, Deh04g, EK04, KR03a, Khu03a, KE04b, MC00, NM03, Sor01a, TZ04, Zhu04b, dJ04]. conditioned [SEK01, yWShX02]. conditions [CPL00, Çağ04, CCS04a, CY02, CCY03, CYC03, Che04c, CLC04, De 01, Deh03g, DYH04, EA03, EOM03b, Far04, FL02a, FNO04, Gia03, HZ04b, HA04, Jan04, Lee04b, LC01b, LX03, MdS04, MCC01, Ram03d, RG03, Sav04a, SD02a, TM04b, Waz01a, WZ01b, WA01, XC04, YC03, YE04, Zay02a, Zay02c, ZAH02, Zay03b, Zay03a, Zay03d, Zay03e, ZI04, Zay04a, Zay04b, ZC02, ZHZ04, Zhu04c]. conducting [AESS00b, AEG03, EEK03b, Zak04c]. conduction [AR03, CHL02, EEYK03, LC04c, MdL04, XFL04]. conductive [RdL01]. conductive-convective [RdL01]. conductivity [EEK03b, HEM03, KE04a, KW04b]. conductor [KML04]. cone [AR04d, ESSA01, NG02a, Sol04]. cone-relation [NG02a]. cone-valued [ESSA01, Sol04]. conference [WH00]. Confidence [Ana03, Ami01, MPS04]. configuration [Ame01, Sil02, Sil03]. conflict [BHH02]. confluent [ASKT03, CQSP04]. confocal [DNS03]. conforming [SS03]. congestion [JK04b, Kan04]. congruence [AB04a, JKP04]. conic [AR04b]. conical [AR04a]. conjecture [ADAJM03, Mus00e]. conjugate [Kwo03, Lan04a, QS04]. connected [Zay02c, ZAH02, Zay03b, Zay03a, Zay03d, Zay03c, Zay03e, Zay04a, Zay04b]. connection [EM03b, EM03c, EMED03, Lay02, Ton00]. connections [MS04a]. connectivity [CTHK03]. conquer [CPTZ04a]. Consequences [dCD03, dCD06]. conservation [DK04, SM04b, YB00b]. Conservative [HL04a, CC04a, NK01, ZC03]. conservative-distributed [NK01]. considerations [Mus00a, Mus00b, Mus00c, Mus00d]. considering [AS04f, JK04b]. consistent [Ino00b]. consisting [AESS00b]. consists [TE03]. constant [HKA04a, TH04, TM03]. constantness [Özd03c]. constants [CK03]. constrained [AAEA03, AGS03, CP03a, Din02, GHW04b, IN04, ML04d, WQ04, ZWRL03, ZW03a, Zhu04a]. Constraint [MD00, Hos04, KU01, WZ04d]. constraints [FQ04, GV00, HZZ03, yN04c, QY00b, Waz03i]. constricted [Man03]. constructed [LB01]. constructing [CC04c, JGW02, LS04f]. Construction [Kap04, GhW04a, Waz03f]. constructive [CDH01]. consumption [Sla04]. contact [Abd01a, Abd01b, Abd02b, AB02a, Abd02d, AM03, AS04a, d’O03b]. containing [SD02a]. contaminant [SC02]. contingency [MMP03, Olu04]. continuation [GLVW00, KA00, Wu00, ZG02]. continued [Cho00, CK03]. continuities [JYC04]. continuity [CD03b, MAK04a, Özd03b, WW03b]. continuous [ÁND02, Cho02b, Dje00, EB04a, EB04b, EB04e, Fuy04, Gha01, HR01, Hua03, KM03b, Li03a, LwWW03, LQL03, MG03b, Nak02, Nak03, ÖK03d, TET02, TLX04, UvB01, ZZC04]. continuous-time [MG03b, Nak03]. Continuum [AC01]. contour [AD04b]. contraction [ER03, YŞ00b]. Contribution [Rao02]. Control [OT04, AQD04, Ama03, BT00, Bay04, Cao01, Che04b, CLC04, CS00, Deh03a, Deh03c, Deh04b, EA04a, EGE02, EGE03, EG03a, EGB03a, EGAR03, EGB03b, EG04c, EGT04, EKE02, EKE03, Ghe00, GL00, HA03, HkT03, HT04, HML^{+}02, HL04f, Jam01, JM02, Jol00a, Jol00b, KENM03, KAAD01, LS04a, LZ04d, LL03, LC03a, Lie04, LF04, LHL04, LLW04, Mun03, PY02, Par03, Par04b, Shi02, Sub02a, WBW01, WW04, Yas02, Yas03, YS01a]. Controllability [AZ04, BS01, BDS03, Fu03, Fu04, LX03, Aas03a, Cao01, GB02]. Controlled [HHK01, MR03]. controller [Ame01, PJPL04, PJ04, Par04a, Par04b]. controllers [LRHRD01]. controversial [Sou03]. convection [AO04b, AR04d, Dem03a, Dis01, EA04d, ENEA04, Elb01, EB02b, Elb03, EESF04, Ezz04, GZ00, Has03b, HSE03, HEM03, HEM04, HSE04, IH01, IES04a, MdL04, Mur01, Mur03, NM03, Opp00, Pan08, Sam04, SR02a, SY02, Xu04, Zak03a, Zak03b, Zak04b, Zha00, ZZ04a]. convection-diffusion [GZ00, Opp00, Zha00]. convection-dominated [SY02]. Convective [AEG03, RdL01]. conventional [CJ04b]. Convergence [APS04, CP03b, CPTZ04b, Li04f, LK04, Qui03, Shi04, wYjSjS03, ZZN04, ZHD04, AS04h, BB02a, Bil04, Bil07, CX03, CO00, EB04b, ES03b, FS03, GLWY04, Han01, Has04c, KES04c, Kum03d, Li03a, Özb04, PR04b, RdL01, SB04a, SZ03b, Thu04a, WW00b, WXO02, WSX03, Xen03, Yua03, Zan00, ZH04, ZW03c]. convergent [HP02a, JJ02, KC00, PTG03, Zhu04c, Zhu04b]. converters [SSPA01]. Convertible [LH04a]. Convex [ÖYY04, ES04, GV00, KA03a, Kee03a, OK03c, ÖK03d, QS04, SE04, TCL02, TLX04, You04]. Convexity [AA00, Amm04, Tin01]. convolution [MK02, MK03a, ML04a, Sub02b, eT04]. convolutions [Kan00a]. cooperative [AT04]. Coordinate [AESS00a, MR00a]. coordinates [CP03a]. coprocessing [Lew01]. core [Ino02b, YIN00b]. Coriolis [SD01]. corners [RS01a]. correction [GM04a, LLP02, TM00, TRC03]. correctness [BBN04]. corrector [Din04e, GG00, NNAK04]. correlated [NHCLPSR04]. correlation [Taw00]. Corrigendum [Bil07, Mus00a]. cosine [Waz04h]. cosmo [Esc03c]. cosmography [Nys01]. cost [DR01, HB04, Par03, PJPL04, PJ04]. costs [JLSS04]. Cotes [EM03d]. couple [Zak03b]. coupled [AO00a, AF03, Bae04a, zBxF04, FH04, HF03, KI04, Kay04f, Khu02, LM04b, LW00, Mel00, Mel01, (Mu04, PB02, VM03, WBW01, Yan01b, ZLS04, vBU02]. Coupling [WZ01b]. covariance [Nak02, NCÁHCLP03b, Nak03, NCÁHCLP03a, NHCJLLP04, Nak04a, NCÁHCLP04]. cracks [APS04]. Cramer [rWS04, WQ04, Wei02a]. Crandall [ML03]. Crank [Deh01]. created [AS04c]. Credibility [ITN00]. credible [GNS02]. Creeping [SR04b, Khu03a]. criteria [CH04, Che04a, DS03, ESS00a, FS04b, HC04a, HHK01, HHZ04, Jia02, JL03a, JL03b, Juk04, KSM03, Li04d, LQL03, Mae01, Mař04b, McR01, MJCM03, Özb04, PGX00, Pen03, Sak03a, SC04, SW03, SW04a, Sol03a, Sun04c, SM04d, Sun04a, XX03, Yan02e, Yan03a, Yan03r, Yan04g, ZL04b, HT00]. criterion [De 02a, Özd03c, Yan03s, Yan04h]. Critical [ZLS04, BN02, LC02a]. cross [AS01, CM04a, LF04]. cross-derivatives [AS01]. crossing [ŞBC03, Wol04, ZAA01]. Cryptanalysis [CJT03, Pei04, XY04]. cryptographic [CHC04]. cryptography [Lee04a]. cryptosystem [CCH04]. cube [Sil04]. Cubic [ASN03, KA04, AGES04, Bah04, CGG01, DSI04, EBE03, EB03b, Kha04, MMP03, MSJ04, Moo01, RC04c, eT04]. current [AS04c, KML04, KB04b, LRMSVO01, SSPA01]. current-programmed [LRMSVO01]. current-steering [SSPA01]. curvatures [AAAE03a]. curve [CCH04, Kor03, Liu04f, MS01, YÇA04]. curved [BK04, TY04]. curves [CG03, ÇT04b, KST00, QY00a, QY00b]. curvilinear [Abd02c]. cusps [Waz02h]. customers [Mor04]. cycle [Çin04a, EOEA00, Xu02]. cycles [DXL02, RM03]. cyclic [SYK02]. cylinder [AAAAZ01, enNAaA02, AZ02, ALO04, DC03, DWC04, Lee04b, Wün03]. cylindrical [AAA02, OT04, RH04]. cylindrically [AAAaZ99a, AaAaZ00a].

d [De 02b, Lee04b, Mef02, Mel00, SSPA01]. D/A [SSPA01]. DAEM [GG02b]. DAEs [LS03b]. dam [Ven03]. dam-break [Ven03]. damped [Bae04a, CC00a, FZ04a, GG00, GB02, WZ01a, WG04a]. damping [BPJ03, Bae04a, Bae04b, KR01, Li04d, MdOPF04, PB02, SC04, SMQ04, XJM04, Yan03r]. Darboux [YÇA04]. Darcian [Elb01, Elb03]. Darcy [AO04b, EA04d]. data [ADASM04, Bar03, BR01, Bog04, GG02a, GG02b, GV00, HIS04, HP02b, JSD04, JMV04a, JAK04b, JAK04a, JMV04b, JLS^{+}04c, JLM04, KIY00, Pra03, SEB03, Sar03b, Sar04a, Sar04b, SV04, TM00, TRC03, Tro04, WL02]. DEA [HV08, JK03a, JLSS04, JK04a, JLS^{+}04a, JVFM04, JAK04b, JHS^{+}04, JSdN04, JLST04, JK04b, JLM04, JSLS04]. deadlock [Ino00a]. decades [KN00]. decay [BPJ03, Bae04a, Bae04b, CCS04a, MdOPF04, Oqu03]. Decentralized [PJPL04, HML^{+}02]. decision [FI01, Ino00b, Ino00a, Ino03, Oda01, Oht04, OASAE04]. decision-making [FI01, OASAE04]. declared [GM04a]. decoded [SH01]. decomposition [Abb03, BB02b, BB02c, BBV04a, BBV04c, BBV04b, BD04a, BBV04d, BTBI03, BBI04, BI04a, BD03, BEAB04, CAH02, CL04c, Deh04a, ES02b, ESAA03, ESK04c, ESK04d, ETBAN04, GHL00, IRR04, IRS04, Kay03a, Kay03b, Kay04a, KES04a, Kay04b, KES04c, Lee00, Les01, RS01a, RG03, Waz00a, Waz01d, Waz01e, WES01, Waz01g, WG04b, Yan02a]. decoupling [CTZ03]. decreasing [WJ00, Zha02a]. defect [LLP02]. defect-correction [LLP02]. defence [EOM04b]. deficient [Tia03]. defined [AS02b, GEA04, KŞ04b, Noo04a, YXC03]. definite [ESEA02, ESAD04, LWL00]. definition [dCD03]. deflection [AÖE04, He03b]. defocusing [Waz02e]. deformation [CL03a, Wu00]. deformed [ALO04]. degenerate [Aas03b, CCS04a, CY01a, CY01b, CT03a, FZ04a, MdOPF04, SK02]. degree [EBE03, EB03a, ÖKOD04]. Delay [CW04b, Abd03a, Aln04, BDS03, BG04, CLF04, CH04, CWY04, CLC04, CCK04, DFF04, DS03, EHM03, EOI03, EOM03a, Fen03, FL02b, Fu04, FZ04b, FS04b, HC01, HGS04, HL04f, IbS01, Jia02, JL03a, Jia03a, KS04a, KSM03, LZ01a, LZ01b, LZL00, LC01b, LS03a, LZ04d, LwWW03, Lin04b, LL04d, PJPL04, Par04a, PKW04, Par04b, Pen03, PP01, Sad04b, Şah04c, Sak03a, Sak03b, Sak03d, SPH04, THY02, TZ04, WL04a, WL04b, Wan04e, lWyCjT03, XCD04b, XL04b, YY04a, YC03, YY04b, YL04d, ZZ02b, ZLZ03, ZS04b, bSI01]. Delay-dependent [CW04b]. delay-differential [EHM03, Par04a]. delayed [CD03a, Cao03a, EOM04a, Has04e, HC02a, HC02b, HL04e, HL04c, Jan04, Shi02, XC02c, XCD04e, YA03]. delays [CC03d, CC03c, CHC03, Che04a, Che04b, CLD03, GR03, HC04a, HHZ04, HC03, HL04b, HL04d, Li04g, LC04b, LC02b, LG04b, Luo02a, MG03b, PJ04, SC03c, SL04c, Ten02, WL03, XCC02, XC02b, XCD04a, XCD04f, YC04, Zha04c, ZC02, ZZC04]. demographic [d’O03b]. dense [JLMC03]. densities [Rad03, Tag02a]. Density [Yan02h, ASKT03, AL00, Cap01, JD03, Olu03a, TET02, Tag01b, Tag01d, Tag01e, Tag02b]. Density-driven [Yan02h]. denumerability [Mus00c]. departure [ASAI03]. dependence [EB04b, EEKS04, JD03, Olu03b, Olu04]. dependent [AA00, Ami01, Asa04c, CW04b, CWY04, Dun02, EG04e, EG04d, EB04e, EESF04, GCSS04, KE04b, Lee04b, Les01, LZ04d, LLW04, Mat00, ML00, Pan08, QMWAZK04, XC02a, XC02b, XCD04d, XCD04e, d’O03b]. depending [CD03b, LL04d]. depth [LCH00]. Derivation [YB00b, Bar04]. derivative [BS00, HL00, Hua03, LS04d, MM04a, SMF04, SR04a, YS02, ZW03c]. derivatives [AS01, ABP04, CCY04, Özd03a, Ras04a, Ras04b, WW00b, Yoo04]. derive [YAYA03]. derives [TET02]. descent [Ric00]. describe [LLCC03]. described [Yan01a]. describing [TET02]. description [SM00a]. descriptor [CFS02, CFS04]. design [Jol00a, Jol00b, Lie04, NCÁHCLP03b, Nak03, Nak04a, Nak04b, PJPL04, PJ04, Par04a, PKLW04, RT03, RHB04]. designing [DR01]. desirability [YIN00a]. detecting [GL01, JHS^{+}04]. detection [AB02b]. deteriorating [BBA03]. determinant [Liu04i, dlS03]. determinantal [Zha02c]. Determination [BF04, HS02, SH02, CCC01, GG02b, Has04a, LMG00, LBE00, SV04, YL02]. determine [Thu04c]. Determining [JK04a, Fat04, JK03a, JK04b, Waz03i]. deterministic [AFRH02, MAR04a, Tie03]. developable [Küç04]. developments [Noo04d]. deviating [DFF04, LHM04, LQL03, LG03, LGZ04, LG04c, SW04b]. DFOM [ZW04a]. DFT [KV03]. DGDMMI [Tie03]. DGMRES [ZW04b]. Diagonal [CS03a, Yua00]. diagonalizing [Ana01]. diagonally [Yin03]. diameter [WSX03]. diamond [Kan00a, Sil02, Sil03, SK04]. diatomic [GGRS03]. dielectric [AS03c, Sel04b].

Difference [CAY01, Dur04, AO00a, ALP03, AA04c, Asa04b, AB04b, BV02, Bah03, BDD00, BM02, BS00, CC00a, Çin04c, Çin04b, Çin04e, Çin04d, Çin04f, Çin04g, Deh00, Deh01, Deh04h, DMS01, EAA03, EA03, EMED03, ENEA04, EOM02, EOM03c, EOM03b, EOM03d, EOAM04, EOM08, EEK03a, GH03, GhW04a, Gon04, GXL04, HIN04, HLY04, HZ04b, Jia02, JL03a, JL03b, Jia03a, JY04, KR04, KSS02, KKM04, Kum02, Kum03a, Kum03b, Kum03c, Kum03d, Kum03e, Kum09, LS03a, Li04b, LR04, LNS04, LS04b, LZ04c, Li04c, LS04c, McR01, Moh04, NS03, PGX00, Pen03, Pet03, PTG03, RK03, RKS04a, RC04a, Sak03a, Sak03d, Sak03e, SC04, Sal03a, SYK03, SY02, SL03a, SL04d, SM04c, WL04b, WA01, XFL04, YL04c, YGL01, wYjSlZ01, Zha00, ZZ02b, ZKSD02, ZS04a]. differences [HR01, RCS03, Ras04f]. Different [Has02a, CCY03, CYC03, EGB03a, KML04, Rad03, TB04, Zay04a]. differentiability [SE04]. differentiable [Kır04, KÖ04]. Differential [Has04d, SDR03, Abd03a, AO00b, AS03d, AS04h, AV04a, AV04b, Aya03, Aya04a, Aya04b, AÇB03, BH02, BH03, BSJ04, zBxF04, BDD00, BH01, BG04, BFGG04, BBI04, BJD^{+}03, Boy03b, CCP03, Cao01, CLF04, CJB02, CAH02, ÇB03a, ÇKB03, ÇB03b, ÇAB03, ÇAB04, Çel04a, ÇB04c, Çel04b, CC04a, CJ04a, CZLZ04, Che04d, CCK04, CW03, CAY01, CLD03, De 01, DKX00, Deh04b, DF04, DH04, DG04, DS03, EB04d, EHM03, EKE03, EOM03a, ER03, ESAA03, ESES04, ESS00a, ESS00b, ESSA00, ESSA01, ETBAN04, EM03h, Fad04, FH04, FL02b, FNO04, Fu03, Fu04, Gia03, GHL00, Guo00, Guo02, Guo03a, Guo03b, GR04, Guo04, GKaM01, GM03b, HZ04a, HDZ04, HES05, Has02a, Has02b, HAM04, He00, HOL02, HLO02, HL03, Hos04, HL04f, IG03, IbS01, JCL00, JCL01, JWL03, Jan04, JL03c, qJkSlS03, KP02b]. differential [KP03, KS04a, KAVM00, Kap04, Keş03a, Keş03b, Keş04d, Khu03b, KLK02, KR01, KM03a, KH01, KSM03, Kuo04, LM04a, Lac03, LZ01a, LZ01b, LCH02, LL04a, LC01a, LC01b, Li03c, LM03, Li03b, Li04d, Li04e, LHM04, LZ04e, LZ04d, LTS02, Lin04a, Lin04b, Lin04c, LY02a, Liu03f, LL04d, LG03, LG04a, LGZ04, LG04c, LS02, Luo02b, LF04, MM03, MMA04, MK04, Mař04b, Meh00b, MA04b, OLT04, Par03, Par04a, PKW04, PP01, RA03, Ras03b, Ras04c, Ras04d, Ras04f, Ron03, Sad02, Sad03, Sad04a, Sad04b, Şah04c, Sal00, SR02a, SR02b, SR04a, Sha02, SW03, SW04a, SW04b, SZK04, Sol02b, Sol02c, Sol03a, Sun04a, TE02b, THY02, TT04, Tun04a, Tun04b, Tun04c, VR03a, VR04b, VR02a, VR02b, VR03b, Wan00b, Wan01, WS01, WL04a, WG04a, Wan04e, Waz00a, Waz01b, Waz01f, Waz01h, Waz02c, Waz02f, Waz03b, Waz03h]. differential [Waz03k, Waz03i, lWyCjT03, WX04, WHM04, XH04, YY04a, YS00a, YL04a, Yan02d, Yan02c, Yan02e, Yan02g, Yan03e, Yan03g, Yan03j, Yan03a, Yan03q, Yan03r, Yan03v, Yan04e, Yan04h, Yan04g, YY04b, YW02, YW03, ZL01, ZLZ03, ZS04b, ZZN04, ZY04, Zhi04, ZL04b, bSI01, HS03]. differential-algebraic [Aya04a, BH03, ÇB03a, ÇKB03, ÇB03b, EM03h, Hos04, Wan00b, WS01]. differential-difference [BDD00]. differentiation [Aln04, Asa04c]. differintegral [LSY03]. diffuse [QS04]. diffusion [AESS00a, Abd04, AAAE03b, AL03, Asa04c, BBA03, BD03, Çat03, CMOS01, Deh01, Deh03d, Deh04f, Deh04h, DLYC04, EESF04, GZ00, GV04, Hel01, IES04a, Jay03, KNU00, KL03, LM04b, MP03, Opp00, Pan08, RS01a, SM02a, SR02a, SR04a, SK02, SY02, SM04c, TZ04, VR02a, VM03, Waz00a, Waz01d, Waz03i, WFT03, Zha00, ZZ04a, ZL03]. diffusion-thermo [EESF04, Pan08]. diffusive [Ram04b, Ram04c, SWY00, SZ01, SPH04]. Digital [TJC03, Der03, Sha04b, Sug00]. Dimension [WZ01a, BFN03, yN04c, WJ00, Zha02a]. dimensional [AAH04, ALO03a, AST04, AAMEST04, AEK04, Aya03, BE02, Bah03, BF04, BR01, Boy03a, CLW02, CZ04, CWZ02, Dar03b, Deh00, Deh01, Deh03a, Deh03c, Deh03d, Deh03f, Deh04a, Deh04d, Deh04f, Deh04g, Deh04h, Din02, ESK04c, ESK04d, Fat04, Fen04a, GHSJ00, GM04b, Hab04b, IOAB01, JCL01, Jay03, Kay04b, KHL02, LWZ00, LqZ04, LCXZ04, LZ04d, LZ04b, LJ04, LOZ02, MdL04, Moh03, Pom01, QS04, RS01a, Ram03d, Ram04b, Ram04c, RM03, Rao04a, RCS03, SKM04, SH01, Sil04, SM04b, Sor01b, Tri04, Waz02c, Waz02d, Waz02e, Yür04a, Zay02c, Zay03b, Zay03a, Zed02, ZK03]. dimensions [IES04b, LWT00, LZ04a, Man03, Moh04, Sal03a, Waz02a, Waz02b, Waz02h]. dipole [AS03c, BS04, ZS02]. Dirac [All04b]. Direct [Aru03, Gla04, BE02, EEES03, ESEE02, GG02b, RS03]. directional [CD03a, Cao03a, LC04b]. directions [GK02, KC00, ZZ04c]. directly [Yan01a]. Dirichlet [AGES04, ERN03, GM03b, Yan02a]. Dirichlet-type [GM03b]. discipline [AS04g]. discontinues [RKS04b]. discontinuities [CAH02, He04d]. Discontinuous [MM04a, AV04a, Bog04, CPL00, CH02, CH03, CCP03, LW00, RCS03, RK03, RKS04a, Ras04f, YX04]. Discrete [AFRH02, Mel01, Tag00a, Tag01a, ALO03b, AQD04, ÁND02, BRS04, ÇG04, Che04a, CWZ02, CFS02, CFS04, EP00, GM04a, GM00, HC01, HL04b, JKP03, KM03b, KG03, LOS02, LHL03, LZL00, LC04a, LS04b, LS04c, LC03a, Mel00, MG03b, NCÁHCLP03b, Nak04a, Nak04b, PJPL04, Par04b, San00, SL03a, SL04d, SL04c, TET02, TE03, Tag00b, Tag02c, TB04, TM03, UvBP00, YB03, YB04, Yam05a, Yam05b, wYjSjS03, YH04]. discrete-delay [HC01, PJPL04, Par04b]. Discrete-time [AFRH02, BRS04, CFS02, CFS04, EP00, MG03b, NCÁHCLP03b, Nak04a, Nak04b, YB03, YB04, Yam05a, Yam05b, YH04]. discretisations [DD00]. discretization [AESS00a, BBN04, KKM04, Moh03, Zha04b]. discretizations [Dis01, LNW03, Sal03b]. discretized [WZ01a, Yua03]. discriminant [MR04]. Discussion [HC02a]. disease [XC02a, d’O03b]. disease-dependent [d’O03b]. diseases [GCSS04, PdR01, d’O03b]. disk [KST00]. dispersal [XCD04e]. dispersed [RA03]. dispersion [EA04d, Kha03b, SC03c, Waz03d, XCD04a]. dispersive [CJ04a, HB00, HT00, LZ04a, Waz02d, Waz02e, Waz03b, Waz03f, Waz03e, Waz03g, Waz03k, Waz04b, Waz04d, Waz04e, Waz04g]. Displacement [CW04a, Li04a, AB02a, OT04]. displacements [KU01]. disposal [Sai02]. dissipation [eT04]. Dissipative [All04b, Tsi01, Dem03b, Gül04a, Gül04b, Oqu03]. Dissipativity [ZL01]. dissociation [GGRS03]. distance [AS03a, CCKS02, CD03b, ÇD04, TH04]. Distinct [Waz04f]. distinguished [WH02]. distributed [CC03d, CC03c, Che04a, DFF04, KNJ04, Li04g, LC04b, LwWW03, NV01, NK01, SPH04, XC02b, YL04d, Zha04c, d’O04]. distribution [AAMAE03, ADARAM04, ASAB03, AL00, ÇG04, MK02, MK03a, ML04a, ML04d, MSO00, RAH01, SM00a, Tag00b, WH00, WX04, WHT04]. distributional [Tri04]. Distributions [ÁND02, ADAMM03, AY03, AY04a, AY04b, Far02, KASK01, MMT00, Tag00a, Tag01a, Tag02c, Tag03a, Tag03c, TV04, vAG03]. disturbance [LC04c]. disturbances [CBK00]. divergence [Par02]. divergences [LP03, MMP03]. diversity [WW00a]. divide [CPTZ04a]. divided [EMED03, HR01]. DMU [JLS^{+}04b]. DMUs [JLST04, JMV04b, JSLS04]. do [ABP04, KU01]. Domain [BD03, RS01a, De 02a, GR03, GHL00, Guo02, HA04, LMS02, Lee00, Rao04a, RG03, TE03, Yan02a, YD04, Zay02a, Zay03b, Zay03a, Zay03e, ZZ04b]. Domains [Kur02, BK04, Liu03f, MM04a, Özd03c]. dominant [Yin03, Yua00]. dominated [SY02]. dominating [HL04c]. Double [CHL02, EA04d, CL03a, CL04b, GÇ04, Hai00, Zhu04c]. double-step [Zhu04c]. doubly [HF03, ZAH02]. Douglas [ML03]. doxa [Ada01]. Drazin [BW04, CC03b, Che01c, Dia04, Ji02, KV03, Kol02, LW03b, LW04b, LW04c, RW03, Wei00a, Wei00c, Wei02a, Wei02b, Wei02c, WWW03, Wei03a, WQ03, WWL04]. driftless [LL00]. driven [BDGG04, KB04b, Yan02h]. drum [Zay03c]. drying [NM03]. dual [Iwa01]. duct [AS03a, HNA04b, Sel03, Sel04a, ZS02]. due [VCV01]. Duffin [CLX02, WX03a, XC02d]. Duffing [MR03, NO03, Yan02b]. Duke [WL02]. duopoly [AE04, Has04e, YA03]. Durand [Zhu04c]. during [EG03a, EGB03a, EGAR03, EGB03c]. Dusty [AQBT04]. Dynamic [FK03, JBMR02, YL04b, AS04d, DWC04, ER02, Iwa01, JBS04, Jia03b, Lew01, LSK00, Mel01, Oda01, PJPL04, PJ04, PZZF02, Ram04a, RK03, Sak04, SM00b, SL04b, TLX04, TLLKB03, Tie03, Yüc04, ZG02]. Dynamical [MG03a, BMR00, BMR03, Din04c, Has04e, Noo03a, SP03a, SN01, UvBP00, UvB01, Yas02]. dynamically [NG02a]. Dynamics [ABP04, EBE03, AE04, AST04, BA04, Bis04a, Dua00, HM04, Hel01, HL04a, KU01, Li04c, LLW04, Ram03b, Ram03c, SYK02, SM04b, Sta00, WLX00, WC02, vBU02]. Dziok [LS04e]. e-cash [JL04]. E-convex [You04]. e-voting [CJT03]. earned [Kat01]. earned-run [Kat01]. earth [EE04a, EE04b, AS04c, EE03]. ecological [XCCC04, ZC04a]. economic [BF03, LSK00]. economizition [KR03b]. eddy [KML04, Lay02]. edge [CTHK03, TH04]. editor [NCK^{+}00]. Editorial [Ano04r, Ano04s, Ano04t, Ano04u, Ano04v, Ano04w, Ano04x, Ano04y, Ano04z, Ano04-27, The03, Ano02j, Ano02k, Ano03n, Ano03o, Ano03p, Ano03q, Ano03r, Ano03s, Ano03t, Ano03u, Ano03v, Ano03w, Ano03x, Ano03y, Ano03z, Ano03-27, Ano03-28, Ano03-29, Ano04o, Ano04p, Ano04q]. Effect [ED02, EEE03, Gho03, Mur01, RPT04, Waz04g, CO01, EEES03, EOM04b, EB04e, GCSS04, HSE03, Kat01, Mek04, NM03, SPH04, SD01, Waz03g]. effective [LCH00]. Effects [AEE04, DLYC04, EEE02, XC01, AEK04, EA04d, EE03, EE04a, EE04b, ENEA04, ESEE02, EESF04, Ezz04, Has03b, HEM03, HNA04b, JD03, JDV04, Pan08, Zak03a]. Efficiency [XLLL03, JMV04a, JSdN04, JAK04a, JLS^{+}04b, MR04, Sah04b]. Efficient [Dis01, HW04, Kwo03, QY00b, Gar01a, JF04, JLST04, JMV04b, KSS02, Lie04, ML03, UvB01, You04, ZN04]. Eigenfrequencies [Cap01]. eigenproblem [Kay04h]. eigenprojections [CGVC04]. eigensystems [AA00]. eigenvalue [AA00, All04b, De 01, DNS03, Has02b, gLgW04, Ram03a, Wu04a, ZI04, Zay04a, Zay04b]. Eigenvalues [AA04d, Boy03b, NK04, TM04b]. eigenvectors [GC01, Wu04a]. eight [Sil04]. eight-point [Sil04]. El-Saka [Pan08]. elastic [AAAaZ99b, AAA03, AaAaZ00b, BD04b, BD04c, Dem02b, ENAAF04, HZL02, HkT03, HT04, RWC03, SK04, WZ01b, YÇAM04]. elasticity [ÇB04b, ÇB04d, EEKS04, SL04a]. Eldabe [Pan08]. electric [AS03c, BS04, DD00, SM04a]. electrical [EEK03b, KE04a, KW04b]. electrically [AESS00b, AEG03]. electro [ENAAM01]. electro-mechanical [ENAAM01]. electrochemical [Ald04]. electromagnetic [AS02a, AS02b, AS03b, ASEM03, AS04c, KEB03, Sei03, Sel03, Sel04a, Sel04b, Zak04b, ZS02]. electromagnetics [JLMC03]. Electromechanical [AAA02]. electron [SK03]. element [BK04, Bah04, BC04, CY02, Che01b, CC02, CX03, CC04b, CC00b, De 01, Dog04, JK04a, JY00, KH02, KS02a, KNJ04, KE04a, Kwo03, LM04b, Liu00, Mef02, OSL03, Ona02, ÖAÖ03, SS03, yS01b, SC02, TM03, Xen03, XL03, Yan00b, Yan02a, ZaYD04]. elemental [WX03b]. Elements [BMR00, BMR01, EEES03, HT00, KL00b, KLK02, KHL02, Xu02]. ElGamal [HL04g]. ElGamal-like [HL04g]. elliptic [AD02, AS01, Afr04, AZTK00, AF03, AV04b, CC02, CCH04, CC04b, GKK02, GL00, GHL00, Has00, Kap04, KEE03b, KL00a, KH02, KS02a, LM04a, LS04a, LYT04, PS04, Sal03a, Sal03b, SS04b, SV04, SW01, XJM04, Yan00a, Yan02a, Yan04i, YD04, YZ04, ZI04, ZZ02a, Zha04a, Zha04b, ZH04]. elliptic-type [AZTK00]. Ellipticity [Bou02]. embedded [EME03, EMR03, Elb01, EB02b, Elb03, Has03b, SP03b, TS02, Xu04]. Emden [He03f, Lia03b, Waz01f]. Emergence [SG01]. emotional [Ada03]. emotions [ITN00]. Empirical [Sar03a, HP02b]. empty [Tar02, Tar03]. Encke [LBE00]. enclosed [Ram00c]. enclosing [WSX03]. enclosure [QS04]. encountered [Dob00]. encryption [XY04]. endolymph [EEAES01a, EEAES01b]. endoscope [AEE04, EEE02]. energy [Deh04e, Mel00, QBK02, QY00a, QY00b, RAH01]. engineering [HS02]. enhance [BK02]. enhancement [Kul03, LS04a]. enriched [Bon02]. enrichment [WL02]. enthalpy [EK04]. entire [DX04, HEM03, Koz03, LYT04, PS04, XYD02, XD03, Yan04i]. entirely [CC02]. entrant [RS01a]. Entropic [CD02, DG03]. Entropy [Tag01b, Tag02a, Gzy02, HR02, Tag00b, Tag01c, Tag02c]. envelopment [JSD04, JAK04a, JMV04b, JLS^{+}04c]. Environment [Oda01, CWY04, DLYC04, FI01, QBK02, WFT03]. environmental [GCSS04]. environments [KNJ04, Liu01]. enzymatic [YB00a]. enzyme [ÇB04a, DXL02, SV02, YAYA03]. enzyme-catalysed [DXL02]. epidemic [EGB03c, Mog04, d’O04]. epidemics [ÖM02]. equal [CGVC04, DW04]. equalities [CT04a, LW04e, LW04d, Tia04b]. equality [CP03a, Din02, FQ04]. equality-constrained [CP03a]. equation [Aas03a, Aas03b, AR03, Abd00, Abd01a, Abd02b, Abd02a, AB02a, Abd02d, Abd03c, AM03, AMI03, AN03, Abd03d, AS04a, AS04b, Abd04, ALO03a, AÖ04a, Asa04a, Asa04b, AÇB03, Bad01, BPJ03, BF04, BG04, BFGG04, BI04a, BJD^{+}03, Boy03a, Boy03b, BGT04, ÇB03a, ÇAB03, ÇAB04, ÇB04c, Cha04b, Che01a, CJ04a, CX04, CZ04, Çin04c, Çin04b, Çin04e, Çin04d, Çin04f, Çin04g, CCK04, DSI04, Dar03b, DD02, DD00, DKX00, Deh00, Deh02, Deh03e, Deh04a, Deh04b, Deh04e, Deh04f, Deh04g, Deh04h, Dem02a, Dem03b, Dem03c, Dem04, Dis01, Dog04, EAA03, EA03, EKE04, EOAM04, ER03, ER02, ESEA02, ESEBD03, ESK04b, ESK04c, ESSEF04, ESAD04, ETBAN04, EEAES01a, EEAES01b, Fad04, FL03, Fat04, Fen04a, GGB03, GM00, Gül04a, Gül04b, GZ00, GKaM01, GM04b, Gzy01, HHJ01, HZL02, HIN04, HES05, Has06]. equation [HAM04, He03d, He03e, He03f, He03h, HLY04, IES04a, IRS04, JM02, JLM^{+}03, KS00, KSL02, KSS02, Kay03a, Kay03b, Kay04a, Kay04b, KES04b, KES04c, Kay04d, Kay04f, Kha03b, Khu03b, KR01, KBÖ00, LMS02, LM04a, LZ01a, LZ01b, lLgC02, LCZ03, sLqZ03a, sLqZ03b, Li04f, LCXZ04, Lia03a, LwWW03, LM02, LJ04, LL04c, LNW03, LG03, LG04a, LGZ04, LG04c, Luo02a, LF04, MZ04a, MdS04, MP03, MM04b, MK04, MS04c, Mař04b, Meh03, MdOPF04, MZ04b, Moh04, ML04d, NZP^{+}04, NO03, Ona02, Oqu03, ÖAÖ03, PB02, PZJ03, RA03, Ram01b, Ram02, RM02, RCS03, RKS04b, Sab18, Sad04a, SS00, SK03, SKM04, SB04a, yS01b, SLC04, SZK04, SWY00, SZ01, Sub02a, Sub02b, SL04c, TE00, TE02a, TE02b, TZ04, eT04, TW03, Tro04, VM03, WZ01a, WL04a, rWS04, Wan04b, Waz01c, Waz02a, Waz03e, Waz03g, Waz03j, Waz03i, Waz04a]. equation [WG04b, Waz04b, Waz04f, lWyCjT03, XZ03, XL03, XL04c, Yan01a, YGL01, Yan02a, YÖ00, YS01a, YYS03, YW02, YW03, Zak04c, ZAH02, Zay03b, Zay03e, Zay04a, Zay04b, ZC03, ZTD04, ZS04a, Zha04b, ZB04a, ZB04b, ZaYD04, de 00]. Equations [SDR03, Abb03, AA04b, Abd03a, AAAE03b, AS01, Afr04, AO00a, AO00b, Ahm04a, AD04a, AS03d, AKKN04, All04c, Aln04, ABEM04, AD04b, AF03, AS04h, AV04a, AL03, AV04b, Aya04a, Aya04b, BH02, BB02b, BJ03, BH03, BBV04a, BBV04c, BSJ04, BJS04, BBV04d, BV02, BK04, Bae04a, Bae04b, BE02, Bah03, zBxF04, BDD00, BH01, BÇ02, BBI03, BBI04, BC04, Bou02, BAB01, CCP03, Çağ04, Cao01, CLF04, CJB02, CAH02, CKNU00, ÇKB03, ÇB03b, Çel04a, Çel04b, CY01a, CY01b, CY02, CL01, CC03e, Che04c, CZLZ04, CC00a, CC00b, CAY01, Dar03a, Dar04, DK04, DX04, Deh03a, Deh03b, Deh04d, DF04, DW02, DH04, Dob00, DG04, DS03, EBMA03, EB04c, EB04d, EE03, EE04b, EHM03, EKE03, EOM02, EOM03c, EOM03a, EOM03b, EOM03d, EOM08, ESEE02, ESI01]. equations [ES02b, ESAA03, ESK04a, ESES04, ESK04d, ESS04, ESS00a, ESS00b, ESSA00, EsEKEA03, El 03, FH04, FH03, FZ04a, Far04, Fen03, FL02a, FL02b, FNO04, FS04a, Gal03, GH03, GHW04b, Gia03, GF00, GS02, GHL00, Guo00, Guo02, Guo03a, Guo03b, GR04, Guo04, GHSJ00, GM03b, HZ04a, Han01, HDZ04, HM04, HI04a, Has02a, Hau04, HOL02, HLO02, He03b, He03c, HL03, HZ04b, He04e, HS03, Hos04, HA04, IOAB02, IG03, IbS01, IRR04, IR04, JCL01, Jan04, Jia02, JL03a, JL03b, JL03c, Jia03a, Jia03b, qJkSlS03, JCZ02, KP02b, KP03, KS04a, KAVM00, Kap04, KES04a, KI04, Kay04g, Keş03a, Keş03b, Keş04d, KEE03b, Khu02, KLK02, KS02a, KM03a, KEB03, KS02b, KH01, KKM04, KSM03, KM03b, KE04c, Lac03, Lan04b, LLP02, LCH02, LSY03, LL04a, LC01b, Li03c, lLgC03, LM03, Li03b, LS03a, Li04b]. equations [LR04, LqZ04, LZ04c, Li04c, Li04d, Li04e, LZ04a, LHM04, LZ04e, LZ04d, Lia03b, LTS02, Lin04a, Lin04b, Lin04c, LM04b, LW00, Liu03f, LQL03, LDKU04, LL04d, LYT04, LS02, Luo02b, MM04a, MMR01, MK03b, MM03, MS04b, MMA04, Man03, Mar04c, Meh00b, MY03, Min03, Min04, (Mu04, NS03, yN04b, OLT04, PY02, PGX00, Pen03, Pet03, PP01, RS01a, Rao04b, Ras03a, Ras03c, Ras03b, Ras04c, Ras04d, Ras04e, RK03, RKS04a, Ras04f, Sad02, Sad03, Sad04b, Şah04c, Sak03a, Sak03d, Sak03e, SC04, Sak04, Sal03a, Sal03b, Sal00, Sha04a, SR02a, SR02b, SR04a, Sha02, SL04a, SK02, She03, SW03, STHN02, SM04b, Sol02c, SYK03, SY02, SM04c, SL04b, SMQ04, THY02, TT04, Tun04a, Tun04b, Tun04c, VR03a, VR04b, VR02a, VR02b, VR03b, Wan00b, WG00, Wan00c, WJ00, Wan01, WBW01, WS01, WG04a, WL04b]. equations [Wan04e, WFC04, WQ04, Waz00a, Waz01b, Waz01a, Waz01d, Waz01f, Waz01h, Waz02b, Waz02c, Waz02d, Waz02e, Waz02f, Waz02g, Waz02h, Waz03b, Waz03f, Waz03c, Waz03a, Waz03h, Waz03k, Waz03i, Waz04d, Waz04e, Waz04c, Waz04g, WG04c, Waz04i, Waz04k, Waz04j, WW02a, WA01, WHG02, yWShX02, WSX03, WX04, WHM04, XH04, XYD02, XD03, XJM04, XL04b, YY04a, YS00a, Yal02, YG04, YL04c, YL04a, Yan02b, Yan02d, Yan02c, Yan02e, Yan02g, Yan03f, Yan03e, Yan03g, Yan03j, Yan03a, Yan03q, Yan03s, Yan03r, Yan03v, Yan04i, Yan04e, Yan04h, Yan04g, YY04b, YAYA03, wYjSjS03, Yür04a, Yür04b, ZAA01, Zed02, Zed03, ZL01, Zha02a, ZKSD02, ZLZ03, ZZN04, ZY04, ZZ04a, Zhi04, ZW03c, ZL04b, bSI01, dK02, dJ04]. equatorial [BN02]. equilibria [Mog04]. equilibrium [AAAE03b, EGE02, GLWY04, LC04a, yN04c, Noo04c]. equitable [JLSS04]. equity [CCY04]. equivalence [BMR01, BMR03, Sar02, SARAEG04]. equivalent [ASEM03, NG02a, WWQ04, XC04]. equivalent-relation [NG02a]. eradication [d’O03b]. Erlang [HM04]. Erratum [AaAaZ00b, AaAaZ00a, CC03c, EOM08, KAK05, KAK06, Kas06, Kum09, Mus00b, Sab18, Yam05a, Yam05b]. Error [ALO03b, Has04b, Kol02, LC03b, OSL03, SB04a, AAAE03b, AO00c, Ama03, Bai00, Cao03b, CCKS02, Ehr02, HS03, LSJR00, Liu00, Mor01, OIO03, XFL04, Yan04a, ZW04a]. error-free [LSJR00, Mor01]. errors [BK04, EB04b, FL02a, JCZ02, Lun00]. escape [SD01]. essential [EA04b]. Establishment [AESS00b]. estimate [Bel01, Che04c, Juk04, LC03b, PdR01, Sev04, Şim04a, Sub02b, Tag01b, Tag02a, TM03, XFL04, YYS03]. estimated [MR04]. estimates [ALO03b, HI04a, Mel00, Sar03a, TB04, dlS03]. Estimating [JSD04, Dob00, Mut03, RT03]. Estimation [AS04c, ASAB03, HA02, SEB03, Sar03b, WL02, WHT04, ADASM04, AAMK01, CQL01, EG04e, EG04d, GM04a, HV08, HP02b, HS03, JLS^{+}04a, JVFM04, JLS^{+}04b, KIY00, KB04a, MC00, MSO00, Nak02, NCÁHCLP04, Olu03a, Sah04b, ST03a, SM02a, Sug00, Taw00, WBW01, Yan04a, YAOY03]. Estimations [EG04b, EG04a, SEG03, Sar04a, Sar04b]. estimator [AAAE03b, AO00c, AYW04, OIO03, Sah04b]. estimators [ADAAM03, Ahm04b, KC04, LP03, NCÁHCLP03b, NCÁHCLP03a, SM02b, Wan02]. ETS [BMR03]. Euclid [Hİ04b, HIİ04]. Euclidean [ÇT04b, IOAB01, YÇAM04]. Euler [Ana01, CLF04, DP01, FL02a, Thu04c, ZZN04]. evaluating [JJ04]. Evaluation [Li02, RM03, SM04a, Sel04a, GMGC01, Ino04a, Ino04b]. evaluations [Wu03, Wu04b]. evaporation [Sel04a]. even [LC01a, Lin04c, WL04b, Waz03f]. even-order [WL04b]. eventually [YH04, ZY04]. everywhere [Kha03a]. evidence [KL02]. EVJ [ZTD04]. Evolution [EB04c, Ald04, BDS03, DE00, DF02, EBMA03, Hau04, LX03, PY02, TET02, Yan02h]. evolutionary [XLLL03]. Evolving [CS00]. Exact [Aas03a, ADAMM03, ER02, El 03, GB02, HKA04b, Kay04c, KI04, LCZ03, Liu00, LCN04, Waz01d, Waz02c, WG04c, YY04a, BAB01, Cao01, CZLZ04, EEAES01a, KERG04, KM03a, lLgC03, sLqZ03b, Waz03i]. Exactification [Kow02]. examples [BB03]. exchange [Ino03, NG02a]. excitable [Ram03d]. excitation [EA04a, EL04b]. excited [EESESS03, EB02a, EB03b, EEB03b]. exhaustive [MAR04a]. exhibiting [EOM04b]. Existence [Afr04, ALO03a, zBxF03, CCS04a, CC03d, CC03c, CHC03, DXL02, Din00, Din04c, FL03, FNO04, GS02, HL04b, Kha03a, LNS04, LS04b, Li04g, LW00, LY02a, Liu03g, LYT04, LG04b, LG04a, MZ04a, ML04b, MdOPF04, PS04, QMWAZK04, RCC04, San00, SL04d, Tag00b, WL03, WA01, XCCC04, Yan02d, Yan02c, Yan03i, Yan03h, Yan04i, Yan04e, ZKSD02, ZY04, Zha04a, Zha04c, zBxF04, CLC04, DX04, Din02, EOM03b, ESEA02, Fen03, GF00, Hai00, Has04c, Juk04, Li04b, LR04, OLT04, PGX00, TZ04, Waz03h, WJ04, Yan01b, Yan01c, Yan02g, Yan03g, YC03, YTM00]. exotic [dCD03, dCD06]. Expansion [RR02b, ALO04, EKE02, KEE03b, Khu03b, LCXZ04, Ras03a, ZI04]. expansions [Boy03a, Ma01]. expected [ASM03, ÇG04, Tag03b]. experience [PPS04]. experiences [Mun03]. Experimentation [Sha04a]. experiments [CMOS01]. Explicit [EM03f, LSY03, Deh03d, HT00, KS03a, Kay03a, KES04b, LW03a, LW04a, Lia03a, Moo01, Ram04a, TM04a, Tsi03, Xu04]. exploration [MD00]. explorations [vBU02]. explosive [ERN03, LYT04]. exponent [KA03b]. Exponential [CD03a, EGE02, EGE03, LH04b, MG03b, Qui04, Ana03, GL01, HP02b, Juk04, LC03a, MK02, MK03a, ML04a, MWAF02, PS00, QMWAZK04, Sar03a, ZC02, ZZC04]. exponential-sum [MK02, MK03a, ML04a]. exponentiality [EB03c, EBSAG04]. exponentially [BN02, KS00, RC04a, ZL01]. exponents [CJB02, UvB01, Waz03c, ZLS04]. exposed [KEB03]. expression [LW04b, XC02d]. Extended [CQSP04, CEM04, FT00, DF02, FH03, Fen04b, LqZ04, LZ04a, LZ04f, Sen01]. extensible [CC00a]. Extension [YE04, CC04c, EESS04, Sha04a, Tra00, Yil04, gZC04]. exterior [Özd03c, PPS04]. external [EA04a]. extinction [LCH00, LC02b]. extra [JLS^{+}04a]. extracellular [Ezz04]. extrapolated [Cao04b]. Extrapolation [AA04c, GHSJ00]. Extremal [AM02b, IG03, JGW02]. Extremality [CH03, CH02]. extreme [JF04]. extremity [Ver03b].

face [NM03]. facilities [SD02b]. factor [APS04, ERN03, (Mu04, RS04b]. factorial [Liu04d, Mus00a, Mus00b, Mus00d].

Factorization [Ron03, gWxZ04, Zha01]. factorized [CMM02]. factors [AV04a, HV08, JVFM04, Qua03, d’O03b]. failure [EG04e, EG04d, Gra03, Sar04b]. Falkner [Asa04a, Asa04b, MLG00]. falsi [WSX03]. Families [Tsi03, CCS03, LS04d]. family [ABP04, CGG01, DC03, GL01]. Far [AM00, AA04d]. Far-field [AM00]. farmers [VCV01]. Fast [Cao04a, CWZ02, ABEM04, EM04a, KSJ02, SEK01, Waz03i, XLLL03]. fast-GAs [XLLL03]. fault [CTHT04]. fault-tolerant [CTHT04]. FDH [JMV04a]. feasible [GK02, KC00, ZZ04c]. features [CT03a]. feedback [Ame01, CLC04, HA03, HkT03, HL04f, HL04c, LZ04d, PJPL04, YL04d]. Fekete [OK03b]. fermentation [BTBI03]. Fermi [He03h, Lia03a]. ferromagnet [SMQ04]. fibers [DYH04]. Fibonacci [Kah05, Kah06, DÖ03a, KA03b, KY03, Kar04, ÖAD03b, ÖAD03a, Özk03, SYY04, YK03]. field [AAHAD04b, AEL04, AR04a, AS02a, AS02b, AS03b, AS03a, AS04c, AM00, AA04d, BA04, DD00, EEE03, Ezz04, Mek04, Mel00, Sel03, SM04a, Sel04a, Sel04b, Zak03a, ZS02, gZC04]. fields [ASEM03, AS03c, CGG01, EB04a, NK03, RH04]. fifth [Kay03a, Sad02, Sad03, Tun04c, Waz02b, Waz02h, Waz04i]. fifth-order [Kay03a, Sad02, Waz02b, Waz02h]. filiform [BEV02, BFN01, BFN03, ENR04]. filled [BD04b, BD04c, Dem02b, HH01, Liu02b, Liu02c, Liu04e, Liu04h]. films [AR04b]. filter [FT00, HB04, NHCJLLP04, PKLW04, WN04]. filtering [PKW04, Qua03]. filters [Kee03a]. filtration [(Mu04, ZZ01]. finance [SAV04b]. financial [MCC01, Sev04]. find [MMER03]. Finding [Deh03a, JF04, GNS02, HP02a, JLST04, SV02, Zhu04b]. Finite [AAMEST04, BV02, BK04, CC00a, CC00b, Coo04, ENEA04, RK03, RKS04a, SC02, Abd02c, ADAAM03, AO00c, Ahm04b, Asa04b, Bah03, Bah04, BC04, BS00, CY02, CC04b, De 01, Deh00, Deh01, Deh04h, Dik04, Dog04, EG03a, EGAR03, EGB03c, EEK03a, EESS04, Gon04, GR03, HIN04, HT00, Ji02, JY00, JY04, Kan04, KR04, KSS02, KL00b, KH02, KS02a, KNJ04, Kum02, Kum03c, Kum03d, Kum03e, Kum09, KE04a, Kwo03, LCS03, LM04b, Liu00, Mat00, Mef02, OSL03, Ona02, ÖAÖ03, ÖAD03b, PTG03, Ram01b, RCS03, Ras04f, RC04a, Sal03a, SS03, Sor01b, Sta03, SY02, SM04c, TET02, Tag01c, TM03, Xen03, XL03, Yan00b, ZAH02, Zay03b, Zha00, ZaYD04]. finite-difference [Asa04b, Bah03, PTG03]. finite-dimensional [Sor01b]. Finite-sample [Coo04]. finite-series [LCS03]. finitely [Ras02a]. fins [CM04a]. firms [TET02]. first [Abd02b, AS04b, BP01, BBI03, CPL00, FNO04, GhW04a, GG00, GL00, Guo03b, HZ04b, JL03a, KSM03, Özd03b, Pen04, PP01, Ras03c, Ras04f, Tro04, Wol04, dlS03]. first-kind [Tro04]. first-order [BP01, GL00, dlS03]. fish [Gho03]. Fisher [IRR04, IR04, KES04b, WG04b]. fishery [JD03]. fishing [Sol03b]. fit [EBA03, Par02]. fitness [RS01b]. fitted [KS00, RC04a, VR02a]. fitting [HIS04, HP02b]. Five [KN00, Sil03, Wu04b]. five-point [Sil03]. Fixed [NHCLPSR04, AO01a, AS04f, BMR03, CGVC04, HM04, LMG00, LW00, Nak03, Nak04a, Nak04b, NM03, ÖA04, Sla04, VCD04]. Fixed-interval [NHCLPSR04, Nak03]. fixed-lag [Nak04a, Nak04b]. fixed-point [HM04]. flat [Elb01, IH01, Kuo04, TS01, Xu04]. flexible [CL03b, GB02, Ino00b, WLX00]. flexural [GB02]. Flow [AR04a, AESS00b, AAAE03a, AR04b, AG03, Ald04, AQBT04, AR04d, BEAB04, CCY03, CYC03, ENEA04, ES03a, ED02, EESF04, Els04, EEE03, EG04g, HSE03, HEM04, HSE04, HKA04b, HNA04b, HKA04a, IH01, Khu03a, Lee04c, LWZ00, Mas03, MM04c, Mek04, OAB04, OASM04, Pan08, Pet02, Pop04, RPT04, Sam04, SM00a, SKM04, SR04b, Vaj01, VR04a, Ven03, Waz01e, Zak03a, Zak03b, Zak04a, Zak04b, Zak04c, de 00]. flows [Din04b, Gon04, HNA04a, Hlo04, RP00, Rao02, RM02, RM03, Rao04a, TM04a, YR01]. fluctuations [PdR01]. Fluid [AH02, AEE04, AR04b, AEG03, AG03, AR04d, BD04b, BD04c, CCY03, CYC03, Dem03a, Dem02b, ESH02, EEE02, EESF04, EEE03, Ezz04, HEM04, HKA04b, HNA04b, HKA04a, HNA04a, HAS04f, He04g, IH01, LWZ00, Mas03, MM04c, Pan08, SR04b, VR04a, Zak03a, Zak03b, Zak04b, Zak04c]. fluid-filled [BD04b, BD04c, Dem02b]. fluids [AESS00b, Mur01, Mur03, Qui04, Rad03, YTM00, Yür04a, Yür04b]. flux [EB02b, Esc02, Esc03b, Esc03c, Has03b, HEM04, Mas03, NM03, ZLS04]. fluxes [QBK02]. focusing [Waz02d]. fold [MK02, MK03a, ML04a]. folding [Dua00]. FOM [CY04]. food [EORI01, MG03a]. food-chain [MG03a]. food-web [EORI01]. force [SD01, Yan01a]. Forced [FZ04b, LM03, Li03b, Min04, Yan03j, DF02, EB03a, EESF04, Pan08, Ram01a, Yan03a]. forces [KU01]. forcing [DE00, EL04b, LOS02]. Forest [WL02]. form [KASK01, MLG00, Qua03]. formal [BK03, BL01, HV01]. formation [Ram04c]. formed [Waz03c]. forms [BRS04, Isı03, LLL04, Mar04c, ÖKOD04, WX03b]. formula [APS04, Kır04, KÖ04, Liu04d, RES04b, SJM04, Tar02, Tar03, Wei00a, Zha02c]. formulae [DP01, EM03d, FS04a, SB04a, WW00b, Wu03]. formulas [AZTK00, HY04, He04f, KKSY04, Liu04i, Tia04b]. Formulation [ZWRL03, EKE04, KE04b, Rao04a, Sam04, dCD06, de 00, Ram00a]. formulations [MCC01, Mel01, TM03]. Fouad [Pan08]. Four [Liu04f, EME03, Liu04c, Sal03b, Sil02, Wu03, ZWRL03]. four-objective [ZWRL03]. four-parameter [EME03]. four-point [Liu04c, Sil02]. Fourier [CSD01, He01, RS04a, SEK01]. fourth [AF03, EGBH03, JGW02, KS03a, KEE03b, KNJ04, Kum02, Liu04a, Man03, PR04b, Sad03, Sad04a, Sad04b, SR02a, SR02b, SR04a, Tun04b, Waz01b, Waz02c, Wei04, ZKSD02]. fourth-order [EGBH03, JGW02, KS03a, KEE03b, Kum02, Liu04a, Sad04a, SR02b, Waz01b, Waz02c, ZKSD02]. FPGA [dGKG^{+}00]. Fractional [Deh03b, ESG03, OK03a, RES04b, AL03, zBxF04, DRS03, Dat03, DF04, EB04d, ER03, ES00, ESI01, ESES04, EEAES01b, Gia03, HNA04a, IPPT03, LSY03, LTS02, LS04d, MP03, Nar01, Rid04, SS00, SK03, Sha02, SS01, TV03, Tag03c, Ton00, Ven03]. fractional-orders [ESI01]. fractional-step [Ven03]. Fractionally [VM03]. fractions [Cho00, CK03]. Fracture [DYH04]. fragile [Par04b]. Fredholm [Sab18, Abd00, Abd02b, Abd02a, Abd03c, AM03, AMI03, AS04b, BBV04a, CKNU00, Dar04, HA04, MM03, MM04b, Tro04, Waz02g, YS00a, Yal02, Yan04f]. Free [Ezz04, Zak03a, AEG03, ALO03b, AO01b, AO04c, Bel04, Cha04b, Cho02a, ENEA04, EESF04, LSJR00, Mor01, Pan08, Pet02, Pop04, RP00, RM02, RM03, Sam04, SK03, WL02, Xu04, Zak03b, Zak04b, ZW03c]. free-air [WL02]. free-convection [ENEA04]. free-derivative [ZW03c]. free-forced [EESF04, Pan08]. freedom [EBE03, EB03a]. front [Ram04c]. FSAL [EME03, EMR03]. fuel [Sai02]. Full [ZTD04, Tan03]. Full-column [ZTD04]. Fully [LOS02, ALO03b, Bah03, JMV04b]. function [ASKT03, AL00, Boy03a, CJ04b, CCS04b, Deh03c, Deh04b, ESSA01, HH01, HZ04c, JM02, Juk04, KKSY04, Kin01, KS01, Li03a, LqZ04, LZ04a, LLCC03, Liu02b, Liu04g, Liu04d, Mat00, MR04, Mus00a, Mus00b, Mus00d, Nak02, Ras04a, Ras04b, Ras02a, RT03, Sol04, Sor02, Tag01e, VCV01, Wu03, Wu04b, YY04a, Yan03k, Yoo04, ZZC04]. functional [CH02, CH03, CCP03, Dar03b, DH04, ESES04, ESSEF04, ESS00a, ESS00b, Fu03, Fu04, HOL02, HLO02, HL03, Lac03, Li03b, Li04e, LZ04e, LZ04d, LL04d, LG04a, LS02, MB04, Ras04d, Ras04e, SW03, SZK04, Sol02a, WG04a, XH04, XC02c, ZL01, ZZN04]. functional-differential [ZL01]. functional-integral [Dar03b]. functionals [AS04h, Din01]. Functions [SDR03, Abb04a, AA00, AEC03, AKÖ04, Asa04a, BM02, Bil04, Bil07, BN02, CCS03, CBK00, CQSP04, CEM04, CSA03, CCS04b, Chu03, Cic01, DC03, DGR04, DRP04, GAZK03, GEA04, HP02a, HS01, Kad03, Kad04, KA03a, Kap04, KŞ04b, KÖY03, Koł01, Kow02, Koz03, KK04, Kur02, LOS00, LS04a, Li03a, LCS03, LS04d, Liu02c, Liu04e, Liu04h, LNW03, MK02, MK03a, ML04a, MP03, MK03b, MM04b, MMA04, MK04, MR03, MM02b, McR01, Mil03, Moo01, Noo04a, OK03a, Orh03a, Orh03b, OK03c, OK04, Özd03c, ÖYY04, Pap00, PSS00, RS02b, RCS03, RK03, RKS04a, Ras04f, RKS04b, RO01, Rid04, Sab18, SE04, Sla03, Sri03, Tag01d, Tag02b, Wol02, WHG02, WX03b, Wün03, Zay03d]. Fundamental [Abd02c, TM04b, Li04f]. fundamentals [XLLL03]. fungal [BJD^{+}03]. Further [GAZK03, HC02b, LL04b, WZ04d, LqZ04, LZ04f, Nys01, Zay02b]. fuzziness [OASAE04]. Fuzzy [Abb04a, Abb04b, sC03a, Lee01, Mae01, MAE04b, MJCM03, AA04b, AA04a, AS04d, All04a, All04c, Amm04, BSJ04, ES04, FSLMC03, FI01, Isk04, Iwa01, JSdN04, LDKU04, MAK04a, MAE04a, NN04a, Shi02, ZWRL03].

G [Pan08]. gain [Par04b]. Galerkin [CL01, CC00b, Dog04, ESAA03, GR04, HB00, KLK02, KB04b, KG03, MK03b, MK04, Ona02, SZ03a, TM03, Ven03, YX04]. Galerkin/minimal [SZ03a]. game [AE04, Pet03, YA03]. games [ITN00, Ino02b, YIN00b, YIN00c]. gamma [KASK01, CSA03, CCS04b]. gamma-type [KASK01]. Gammel [KL02]. gap [BHH02]. Gardner [NZP^{+}04]. gas [AQBT04, El 03, GGRS03, OAB04, RA03, Rad04, RK03, SM04b, Waz01e, Yüc04, Zed02, ZK03, Che04b, XLLL03]. gate [Pet02, Pop04]. Gauge [Pra03]. Gauss [Ahn03, GMGC01, Özb04]. gaussian [MMT00, AL00, EG04c, LC03a, NG04]. GCD [STB03]. Gegenbauer [Rid04]. Gelfand [Kor03]. General [Dik04, KA03b, Waz02d, Waz02e, AGS03, BCI03, EMED03, EME03, EM04c, ES03b, GHW04b, IRS04, Khu03b, Kol02, LQL03, MK02, MM04a, Mor04, Noo03c, NNAK04, NN04b, Noo04d, RR02a, Sar04a, TM03, rWS04, XL04b, XL04c, Zay02b, Zay02a, ZAH02, Zay03b, Zay03d, Zay03c, Zay03e, Zha01, Zha02b]. Generalised [Mel00]. generalization [AO01a, CPL00, EN01, OK04, SK03]. Generalizations [Ino02b]. Generalized [ANS02, Ahn03, AZTK00, Cho00, DRS03, Din01, ECL02, IOAB01, KAVM00, LCXZ04, LDKU04, LF04, Noo04b, TY04, TW03, Wün03, ZB04b, Abd02a, AL00, ÁR02, AR04c, Amm04, AEK04, Bis03, BCT04, BGT04, Bİ04b, CC03a, CLX02, CZ04, CK03, Cin04h, CP03b, Din00, Din03b, Din04d, Din04e, DW04, EKE04, EMY04, EMED03, EM03f, ES04, ESS04, EEE02, EAA04, EEKS04, GAZK03, GGB03, GKK02, Gin04, GG02a, GKaM01, HL00, HF03, IR04, IN04, JK03b, JJ02, KV03, KES04b, Kay04d, Kay04f, LCZ03, Li04a, LCS03, LK04, Liu04i, LW04e, LW04d, MK03a, MS04d, MG03a, Rak04, RS02b, RES04b, Rid04, Sta03, SW02, Tag00a, TK04, Wan02, WH00, gWxZ04, WZ04a, WZ04b, Waz04b, Waz04c, Waz04k, WWW00, WW03b, WD03, WW03c, XC02d, Yan01b, Yan03s, YTM00, Zha02c, ZW03a, ZY00, ZB04a]. generate [qLzWcC03]. generated [Din04a, Ras02a]. generating [CCS03, LOS00, LCS03, PSS00]. generation [DG02, EB03d, EB04e, EB04f]. generator [ASK04]. generators [DK04, Tan03]. Genetic [Shi02, CHL02, LC03b, NS04, OASM04, TLX04]. Geng [He04h]. genital [EGB03c]. gentleman [Nat00]. Geodesics [MS04a]. Geometric [GNS02, AAH04, EN01, SZ04]. Geometrical [AAMAE03, Ana01]. geometries [QS04]. geometry [Hİ04b, Käm02b, Ras02b]. Gilbert [BAB01, SB04b]. Ginzburg [BGT04, GGB03]. given [CCY03, CYC03, ŞBY04]. glassy [AESS00a]. gliding [HES05, HAM04]. Global [BDGG04, Cao03a, EG02, EOI03, EOM04a, EOAE04, HLY04, JT04, LZ04c, Li04c, LC04b, SZ04, WZ04c, XCC02, XCD04b, YL03b, YL03a, YL04c, YCME04, YLEM04, Zan00, ZS04a, ZC04a, AGS03, Bel04, EOM03c, EOM04b, ES03b, FT00, Far00, FQ04, GF00, HH01, HA04, HC03, HL04b, Kor03, LH04b, Liu02b, Liu04g, MZ04a, Moh00, QW03, RS03, Sak03b, Sak03c, SM00a, SZW03, SC03c, WZ01a, WL03, WXO02, XCCC04, XC02c, XCD04d, Zha04c, ZZC04]. Globally [ZC02, d’O03b, HP02a]. glucose [BG04]. GMA [STHN02]. GMA-system [STHN02]. GMRES [CY04, GC01, Zha04b]. Goal [dOVS03, Cha04a, Isk04, SS04a]. Gompertz [Jah03, MR04, WHT04]. Goodness [Par02, EBA03]. Goodness-of-fit [Par02]. Gordon [Fen04a, Kay03b, Kay04b, KES04c, Ram01b, WZ01a, Waz03j]. Gordon-type [Waz03j]. Goursat [HHT04]. governed [Sub02a]. governing [De 02b]. Gower [HL04e]. GPL [CCK04]. GPL-stability [CCK04]. GPS [CC03f, Lun01, YL02]. grade [CCY03, Mas03, MM04c, VR04a]. gradient [JSLS04, Kwo03, Mar01, QS04, SD01]. gradient-like [Mar01]. gradients [Ric00]. Gramian [BCT04]. graph [BHH02, Ino02a, Kan04, KU03, Ton00]. graphic [AD01]. Graphical [WC04b, WC02, WC04c]. graphs [CTHT04, Ino04a, Ino04b]. gravitating [Rad04]. gravitation [Esc02, Esc03b, Esc03c]. gravity [AAA03, AAHAD04b, Cha04b, Esc03c]. Gray [LS04f]. Green [Yan03k]. grey [QS04, Tie03]. Grid [DG02, LS04a, Moh03, RM03, VCD04, YR01, Zha01]. grid-based [Zha01]. grid-point [Moh03]. grids [GH04, Sil02, Sil03, Sil04]. Gröbner [ÇB04a, YAYA03, gZC04]. ground [ÇB04b, ÇB04d, ZaYD04]. groundwater [SC02]. Group [Ver00, Yür04a, Yür04b, ZZ01, CZT04, CT03b, EOM04b, Ino00b, Ino03, Isı03, LCH02, ÖAD03b, Yen04, YCC02, Yil04, Zed02, ZX04]. group-preserving [LCH02]. groups [DÖ03a, Dik04, Ino03, KA03b, KY03, Özk03, Pra03]. growth [BJD^{+}03, EOM04a, GB03, TS01, XC01]. guaranteed [Par03, PJPL04, PJ04, PR04b]. Guest [NCK^{+}00].

Haar [DEK04, MMA04, RO01]. habitats [BRVI00]. half [AAHAD04b, Li04e, Liu03g, Mař04b, SP03b, SL04b, Wan01, YL04a]. half-line [Liu03g, YL04a]. half-linear [Li04e, Mař04b, SL04b, Wan01]. half-plane [SP03b]. half-space [AAHAD04b]. Hall [HNA04b]. Halley [GH01]. Hamilton [BC04]. Hamiltonian [CTHT04, ESEE02, Men02, Sun04c, SM04d]. Hammerstein [EsEKEA03, Lan04b, MMR01]. Hankel [EESS04, Gün04, Kow02, SB03, Tri04, TB02]. Hardy [ÇD04]. Harmonic [AL03, AM00, AA04d, EBE03, KÖY03, Özd03c, ÖYY04, RS02b]. harmonically [EESESS03, EB03b]. harvested [KSC02]. harvesting [JK04c, LLW04, WW04]. Hausdorff [IPPT03, Tag01c]. having [Abd02c, NR02b, Tag02a, Tag02c, Waz02h, ZW03c]. hazard [Sar04a]. hearing [Zay03d]. heart [KEB03]. Heat [EB03d, EB04f, Ram00a, Ram00b, Ram00c, AR03, AEG03, BF04, CM04a, CHL02, Che04c, Deh00, DKV04, EA04d, EEYK03, ED02, EB02b, EB04e, Fat04, Has03b, HEM03, HEM04, JM02, KE02, Lee04c, LC04c, Mas03, MWAF02, NM03, QBK02, Ram03b, WG04c, XFL04, YÖ00, YYS03, Zak04a]. heat-conduction [EEYK03]. heat-like [WG04c]. heavy [Tag02b, TV04]. heavy-tailed [Tag02b, TV04]. helical [RH04]. Helmholtz [ESK04b, Li04f, RH04]. hematopoiesis [Sak03b, Sak03c]. Hemisphere [TRC03]. Hermite [BN02, DC03, Qua03, Wün03, YW03, dBGM02]. Hermitian [Bai00]. herpes [EGB03c]. Hessian [Zhu04a]. heterogeneity [LLCC03]. heterogeneous [AE04, TM00]. Heun [Ron03]. Heuristic [LLB03, LSJR00, LS04f]. hierarchy [WC04a]. High [GZ00, TS02, AS03d, CW03, DYH04, HM04, Kho04, LW04a, Lin04a, MM03, TCS01, Tsi01, WLX00, YS00a]. high-order [AS03d, MM03, YS00a]. high-speed [WLX00]. Higher [KR04, Zay03b, Zay03a, AB03b, CLW02, Cic01, EM03h, Has04d, JK03b, JKP04, Li04b, LZ04d, LZ04b, Lin04b, OLT04, Sev02, Vul00, WG04a, Wan04e, Waz01h, Waz02a, Waz02b, Waz02c, Waz02d, Waz02e, Waz02h, Waz04d, YL04c, Yan03k, Yan03q, Yan03n, YWY03]. higher-dimensional [CLW02, LZ04d, Waz02c, Waz02d]. higher-order [Has04d, Li04b, LZ04b, Lin04b, OLT04, Sev02, Vul00, Waz01h, Yan03k, Yan03q, Yan03n]. highest [SR04a]. Highly [KSS02, ÖM02]. Hilbert [DW02, Din02, Liu04f, RW03, STB03, WC01, WQ03, Wei03b, Xia03, XC04]. hill [JJ02, SLC04]. Hirota [HF03, Kay04f]. HIV [Gum02, SP03a]. Hölder [HR01, JYC04, Yan03o, Yan03t]. hole [Abd02c]. hollow [DWC04, Lee04b]. homogeneity [NG04, PS00]. Homogeneous [Fen04b, AAAaZ99a, AAA03, AaAaZ00a, AZ02, Cic01, DWC04, EGAR03, KM03a, LLB03, Mat00, Mil03, ÖKOD04, Sen01]. Homology [Hab04a]. Homotopy [He03a, GLVW00, He04a, He04b, He04d, Lia04, WBW01, Xu04, ZL04a, LMG00]. hopefulness [YIN00a]. Hopf [KSC02, Ram01a, SWY00, ZLZ03]. Hopfield [WFC04]. Hopfield-type [WFC04]. horizon [Der03]. Horizontal [MCS04, Elb01, HSE03, IH01]. horizontally [d’O03b]. HPD [GNS02]. HPD-credible [GNS02]. human [BT00]. Hurwitz [KKSY04, LS04d]. Huxley [IRR04]. Hybrid [Lee04b, MK03b, MM04b, MK04, MR03, Sab18, She03]. hydride [DKV04]. hydrocarbon [Bon02]. hydrodynamic [Rao04b]. hydrodynamics [BK02]. hydrogen [Ald04]. Hydromagnetic [VR04a]. hyperbolic [Aas03b, AAAE03a, CLD03, DFF04, FZ04b, HL03, Kha03a, LM03, LHM04, LQL03, LC04c, Luo02b, MZ04a, Min04, Moh04, RCS03, RKS04a, WG00]. hyperconvex [ÖA04]. hypercube [KSS02]. hypergeometric [ASKT03, CCS03, CQSP04, CEM04, GAZK03, RES04a]. hyperplane [Ras02d]. hyperruled [AAE04]. hypersurfaces [YÇAM04]. hypothesis [AY04a, AY04b, Els04, JBS04]. Hypothetical [TKK00]. hysteresis [Dar04]. hysteretic [Tad01].

ID [WH00, WH02]. ID-based [WH00, WH02]. ideal [gZC04]. ideals [Ras02d]. idempotency [ÖÖ04]. idempotent [ÖÖ04]. Identification [CCC01, Moh00, Cho02a, CQL01, Has03a, JLS^{+}04a, JWL03, NS04, SN01, Tse03, WFC04]. Identifying [Deh03c]. identity [Wan04d]. IFRA [EB03c]. II [Liu03a, AR04c, BMR00, EEES03, EE04a, EM03c, HS02, LCH00, Ram00b, Thu04b]. III [LY02b, BMR01, Ram00c]. ill [SEK01, WN04, yWShX02]. ill-conditioned [SEK01, yWShX02]. ill-posed [WN04]. image [SH01]. images [SH01]. imaginary [BN02]. imaging [Kee03a]. imbedding [CKNU00, KNU00, NU00a]. immiscible [Qui04]. immobile [AR04b]. immobilized [SV02]. impedance [Kru03]. impelling [Liu04g]. implementable [De 02a]. Implementation [AD02, MLG00, BD04a, CS03b, GH04, KNJ04, LSJR00, RAH01, WWQ04, Xu02]. implementations [Kwo03]. Implicit [CH02, MR00a, Noo03a, AESS00a, Bah03, CH03, CCP03, Din00, Din04d, MR00b, Ram02]. imposing [JSdN04]. imposition [Çağ04]. impressing [Abd01b]. improve [GLVW00, NK01]. Improved [CJ04b, Fan03, Özb04, ÁR02, AR04c, RM02, Sah04a, Thu04a, WSX03]. Improvement [HWWM03, HWW03, Jay03, Sha04b, JLS^{+}04b, Kah06, LL04b, SYY04]. Improving [Abb03, JK04a, JK04b]. impulse [Nak02]. impulses [Gop04, LY02a, Luo02b, THY02, lWCH03, lWyCjT03]. impulsive [AO00b, BDD00, BH01, CAY01, CLD03, DH04, FZ04b, FS04b, Guo02, Guo03b, HZ04b, LL04a, LS02, Pen03, Sol02b, Sol03a, Sol04, Sun04a, XH04, ZS04b]. incidence [GM03a, PdR01]. incidences [SP03a]. inclined [ÇT04b, EEE03, HSE04]. including [CO01, YY04a]. inclusion [Sav04a, SP03b]. inclusions [AO01a, sC03a, Din01, Din03b, Din04d, pFjH03, GXL04, LX03, LDKU04]. incomes [GM04a]. incomplete [WZ03, WW02b, Zha01]. incompressible [BEAB04, Gon04, LLP02, SKM04, She03, SM04c]. inconsistent [WW02a]. incorrectness [CHC04]. increasing [Sla03, SZ03b]. indefinite [Afr04, Che01b, CX03, LF04, WZ03, WZ04d, Zha04b]. independent [ADAMM03, ADARAM04, Ami01, EG04e, EG04d]. Index [Ano00a, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano00j, Ano00k, Ano01c, Ano01h, Ano01i, Ano02i, Ano03h, Ano03i, Ano03l, Ano03j, Ano03k, Ano04a, Ano04g, Ano04h, Ano04i, Ano04n, Ano04b, Ano04c, Ano04j, Ano04k, Ano04d, Ano04l, Ano04e, Ano04f, Ano04m, CC03b, Ano01a, Ano01b, Ano01d, Ano01e, Ano01f, Ano01g, Ano02a, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano02g, Ano02h, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, BH03, Çel04b, DRP04, EM03h, Hos04, LS03b, Mog04, ZW03b, Ano03m]. index- [Çel04b]. index-3 [LS03b]. indicator [HIİ04]. indices [CFS04]. individual [Hür04]. induced [Ram03b, d’O03b]. induction [Mef02]. industry [JK04b]. Inequalities [Kır04, Olu03a, AM02b, CT04a, Din00, Din04e, KÖ04, LC01a, LK04, ML04c, Mou04, Noo03a, Noo03b, Noo03c, NWX03, Noo04b, NNAK04, NN04b, Noo04d, Olu03b, Olu04, Yan03p, ZZ02a, ZY00]. inequality [Aas03b, AY04a, ÇD04, He04c, HZZ03, Xia03, Yan03q, Yan03o, Yan03t]. inertia [RA03]. inertial [FT00, Far00, HB04]. Inexact [CC03e]. inf [TM03]. infant [BT00]. infeasibility [Käm02b]. infection [Yan01a]. infections [Yan01a]. infectious [GCSS04, d’O03b, d’O04]. inference [EB04a]. infinite [AAAAZ01, Abd02c, AZ02, AO01b, CY01b, Dar03b, De 02a, Din02, EG03a, EGAR03, EGB03c, ENAAF04, HSE03, KR03c, Kuo04, LL04d, Mat01, NM03, RS02b, TET02, TZ04, Waz02h, XCC02, YD04]. infinity [GG00, TE04]. influence [AAA03, AAHAD04b, GV00, Zed02]. influential [JHS^{+}04]. Information [HR02, ADAAM03, Ahm04b, BHH02, CD02, ITN00, Ino03, LCH00, Nak02, NCÁHCLP03b, Nak03, NCÁHCLP03a, NHCJLLP04, Nak04a, NCÁHCLP04]. information-gap [BHH02]. inhomogeneous [AR03, AM00, AA04d, Waz03h]. initial [AAHAD04b, CH03, De 03, Deh04d, Deh04g, ERN01, FH04, Gia03, Has04d, IES04b, JCL00, JBMR02, KS03a, KW04a, LCH02, LW03a, McR01, MLA01, RKS04a, RC04b, STHN02, VR03a, VR04b, Waz02f]. initial-boundary [CH03, ERN01]. initial-value [ERN01, JCL00, JBMR02, RC04b, STHN02]. initialization [Bou02]. injection [AO04b, Elb01, Elb03, EB04f, TS01]. inner [KR03a, wYjSjS03, ZLS04]. Innovation [WFT03]. Input [JLS^{+}04a, Che04b, JK03a, JSdN04, JK04b, Ram03a]. Inputs [HV08, JVFM04, JLS^{+}04a, JAK04b, JAK04a, JK04b, Tie03]. Inputs/outputs [HV08, JVFM04]. inquiry [GM04a]. Insights [HL00]. Instability [RA03, Sad03, BD04b, BT00, HSE03, RH04, Rad04, Tun04a]. instantaneous [HL04c, Kan00b]. instantaneously [WL04a]. insulin [BG04]. integer [Pap00, SS04a, YL02]. integrability [Gin04]. integrable [LqZ04]. Integral [Abd01a, AS04a, Bat04, Dat04, ES02a, Mař04b, Wei03a, Abd00, Abd01b, Abd02b, Abd02a, AB02a, Abd02d, Abd03c, AM03, AMI03, AN03, Abd03d, AS04b, AO01a, AD04a, Aln04, AD04b, BBV04a, BBV04c, BJS04, Bad01, BBI03, Bry02, CKNU00, Cha04b, CL01, Dar03a, Dar03b, Dar04, DD00, Deh01, Dob00, EKE04, ESI01, ESEBD03, ESES04, EsEKEA03, GHSJ00, HHJ01, HA04, KKSY04, KEB03, Lan04b, LS04d, LW00, MMR01, MK03b, MM04b, MS04b, MS04c, Meh03, ML04c, MWAF02, Noo04a, Ras03a, Ras03c, Ras04c, Ras04d, Ras04e, Sab18, SL04a, Thu03, Thu04b, Waz02g, Waz03h, WD03, Yal02, Yan01a, ZZ04b, dK02]. integrals [AZTK00, GKK02, JJ04, Li02]. integrated [FT00, Far00]. integration [Bar04, GLM04, HHJ01, KS00, LWT00, Lun00, MM02b, MR00a, OOAA04, Ram04a, RR02a, Sal00, SS01, Tsi03]. integrations [yS01b]. integro [BFGG04, Che04d, DKX00, ESAA03, ESES04, Guo00, Guo02, Guo03a, Guo03b, GR04, Guo04, MM03, MMA04, MK04, Ras03b, Ras04c, Ras04d, TE02b, Waz01h, YS00a, HS03]. integro-differential [BFGG04, Che04d, DKX00, ESAA03, ESES04, Guo00, Guo02, Guo03a, Guo03b, GR04, Guo04, MM03, MMA04, MK04, Ras03b, Ras04c, Ras04d, Waz01h, YS00a, HS03]. integro-partial [TE02b]. integrodifferential [BS01, BDS03, ZAA01]. intelligent [MB04]. intensities [NU00a]. intensity [APS04, TLLKB03]. interacting [BA04]. Interaction [AG03, BG04, EB03a, EB03b, NG02a]. interactions [AAHAD04a, Ada03, Pla03, Ver03a, Ver03c]. interactive [AD01]. interconnection [CTHK03]. interdependent [MB04]. interface [AH02]. interfaces [RH04]. interfacial [WZ01b]. interior [DG03, Kru03]. intermediate [ESG03]. Intermodulation [Abu00]. internal [EB03d, EB04e, EB04f, NU00a]. internally [GB02]. interperception [ITN00, Ino00a]. Interpolation [Far03, EMED03, EM03e, He04f, Moo01, Ras04c, Ras04b, Yoo04]. interpolatory [QY00b]. interpretation [YIN00b]. interpreter [YS02]. interstellar [BMMRS04]. interstitial [BG04]. Intertwined [Din04a, Din04b]. Interval [Li04d, Li04e, Yan03a, ZL04b, JK04a, JMV04a, JMV04b, JLM04, Moh00, Nak03, NHCLPSR04, SZW03, YB03, YB04, Yam05a, Yam05b, ZW03c]. intervals [Ana03, EG03a, EGB03a, EGAR03, EGB03c, KR03c, KM03b]. intervalwise [IbS01]. intratrophic [JB02]. Introduction [De 02b, MCS00, Thu04a, NCK^{+}00]. Invariant [CG03, NU00a, Cho02b, Hİ04b, KNU00, KŞ04b]. invariants [CFS02, ENR04, KG04]. inventory [AB02b, AB03a, BBA03]. Inverse [Gzy01, Kwo01, MSO00, TV04, Wan00a, Zay03c, AL00, AA04d, BMMRS04, CC03a, CC03b, Che01c, CLX02, CWS02, Deh03c, Deh03f, Dia04, EM03f, EM04c, Fat04, GKaM01, GM04b, Hai00, Has03a, Has04b, IN04, JLS^{+}04a, Ji02, KV03, Kol02, LW03b, Li04a, LW04b, LW04c, Liu04i, LW04d, (Mu04, OOAA04, Rak04, RW03, RWC03, Sev04, SD02b, TLLKB03, WLX00, WZ04a, Wan04a, Wei00a, Wei00b, Wei00c, WWW00, WW01, Wei01, Wei02a, Wei02b, Wei02c, WW03b, WX03a, WWW03, WW03a, WD03, Wei03a, WQ03, WW03c, Wei03b, WWL04, WZ04e, XFL04, XC02d, YG04, Zay02b, Zay02a, Zay02c, ZAH02, Zay03b, Zay03a, Zay03e, Zay04a, Zay04b, Zha02c, Zha02b, ZB04b, ZHZ04]. inversely [dlS03]. inverses [BW04, CZT04, CMM02, DW04, FK03, LW04e, Sta03, SW02, TW03, Tia04b, WZ04b, ZW03a, ZB04a]. inversion [CSD01, GG02a, Tag01d, Tag02b, Tag03b, TV03, WZ02]. invertibility [BMR03]. invertible [HZ04a]. investigate [HC02b]. Investigating [Mut03]. investigation [DKV04, Les01]. investment [QWY02]. investments [ySW01]. inviscid [Ram03c]. involutory [CT03b]. involving [ASKT03, Bat04, BM02, CCS04b, CMF04, Li02, LCS03, Liu04d, MLA01]. ionosphere [AS03b]. irregular [Gha03]. irresolute [Abb04a, Abb04b]. Ishikawa [Din03b]. Isochronicity [CGG01]. isotherm [KE04b]. isotone [Wan00c]. isotropic [AAAAZ01, enNAaA02, AS02a, AZ02]. issue [MCS00]. items [BBA03]. Iterated [KG03, GHSJ00]. iterates [Nar01]. iteration [He00, He03c, SB04a, Ton00, WW00b, ZW03c, Zhu04c, Zhu04b]. iterations [Har03, Özb04, wYjSjS03]. Iterative [CL03b, JCZ02, LW04c, MdS04, SW01, yWShX02, AO01b, ABP04, BE02, BM02, CC03a, Cao03b, Cao04a, Cao04b, CDH01, CP03b, CPTZ04b, Din03b, Din04e, HLO02, HZ04b, JLMC03, Kam02a, KLK02, LWL00, LK04, Mor01, Sah04a, SZK04, Wu04a, Yao02, Yao03a, YY02, ZZ02a, Zha00]. IV [BMR03, Liu03b]. IVPs [LW04a].

Jacobi [BC04, Isı03, Özb04]. jet [RA03, Ram00c]. jets [Ram00a, Ram00b, Ram00c, Ram01a]. Jimbo [LJ04]. jitter [Ram00b]. Johnson [HAS04f]. joint [EA04b, EEES03, EE04b, HT04]. Jones [UK03]. journal [AO01b]. Jumbo [LCXZ04]. jumps [WHM04].

Kadomstev [sLqZ03a, sLqZ03b]. Kadomtsev [Waz01c]. Kalman [FT00, HB04]. Kamenev [Sun04c, XX03]. Kamenev-type [XX03]. Karman [He03b]. Kauffman [GMGC01]. Kaup [LqZ04, LZ04b]. Kawahara [Waz03e]. Kawahara-type [Waz03e]. KDV [IRS04, Dem04, Kay03a, Kay04f, Kay04g, PZJ03, Waz02a, Waz02b, Waz02h, Waz03a, Waz04b, Waz04d, Waz04c, Waz04f, Waz04i, Waz04k]. KdV-like [Waz04d, Waz04i]. KdV-type [Kay04g]. KdVB [Kay04a]. Kelvin [RH04]. Kepler [Mus00e, SB04a]. Kernel [FZ04a, Abd00, Abd01a, Abd02a, Abd02d, Abd03c, HHJ01, Kan00a, lLgC03, Olu03a]. Kerner [Zhu04c]. Kerr [Bis03, Bis04a, Bis04c]. key [CHC04, CJ04b, HLLL03, HWWM03, LHL03, LL04b, LLL04, Pei04, Tse03, WH00, WL04c]. keys [HW04, HL04g, Sha04b, TJC03]. Khintchine [CK03]. kills [KL02]. kind [Abd02b, AN03, AS04b, BBV04a, BBI03, EM04b, LG03, LG04a, MK03b, Moh03, Ras03c, SW04b, TLX04, Tro04]. kinematically [EEB03b]. Kinematics [WC04c]. kinetic [BTBI03, GG02b]. kinetics [ÇB04a, Kay04e, LMG00, YB00a]. Kirchhoff [BPJ03, Bae04b, FZ04a, PB02, dJ04]. Kirchhoff-type [FZ04a]. Klein [KES04c, Waz03j]. Knopoff [He03e]. knot [KU03, ŞBY04]. knots [BŞY03, Bel04, UKŞ03]. Korteweg [Dem02a, Dem03c, KBÖ00]. Kotara [ESK04a]. KP [CZ04, ESK04c, Waz03a, Waz04c]. Krishnamoorthi [AS04g]. Kronecker [CT04a, DMT02, MLA01]. Krylov [Bai00, ZZ04a]. Kummer [JKP04]. Kuramoto [Waz04a]. Kutta [EMR03, ÁR02, AR04c, EME03, HT00, IbS01, KS03a, MS04c, OOAA04, RP00, Wu03, Wu04b, XL04c, bSI01]. Kuznetsov [LCZ03].

L [Has06]. labyrinth [Yüc04]. lactic [BTBI03]. Lacunary [Bil04, Bil07, KŞ04b]. lag [LW04a, Nak04a, Nak04b, Tsi01, TS02]. Lagrange [KS02a, Ras04c, Ras04b]. Lagrangian [He03b, RA03]. Laguerre [DSC01, ES00, Koz03, LOS00, PSS00, RS02a]. lake [Gho03]. Laminar [Elb01]. laminated [OT04]. Lanczos [HL00, MS01, Ode02, OIO03]. Landau [SB04b, BAB01, BGT04, GGB03]. Landscapes [BRST02, RS01b]. Lane [He03f, Lia03b, Waz01f]. Laplace [BBV04b, Meh03, yS01b, Tag01d, Tag03b, TV03, TV04]. Laplacian [ALO03a, CH02, KSL02, LOZ02, XX03, Yan03u, Yan04d, Yan04f]. Large [Boy03b, MIM00, AAEA03, Cao03b, CL03a, EOM04b, HLO02, He03b, JLMC03, Koł01, KS01, PJPL04, Par04b, PS04, TE03, Wu04a]. Large-scale [MIM00, AAEA03, PJPL04, Par04b]. largest [AA00, UvBP00]. laser [SK03]. lasers [SS02]. last [Rol02]. latent [d’O04]. lateral [Şim04b]. latitude [BN02, KST00]. lattice [Din04c]. Lavrentiev [SS04b]. law [AR04a, Bis03, Bis04a, Bis04c, CZT04, EB04b, SYK02, SW02, WZ04a]. laws [DK04, SM04b, Ver03a]. layer [BS04, De 02b, EESF04, IH01, KS04a, Kuo04, Mas03, NR02b, OT04, Pan08, SM04a, Sel04b, TS01, Yür04a, Yür04b, Zak04a, Zak04c]. layered [AS03c, EG04g, RWC03]. layers [AH02, BD03, Rad03]. LCEs [UvBP00]. LCM [STB03]. LCP [ZG03]. leading [HS02, Has03a, Has04a, SH02]. leads [BF03]. leakage [Yüc04]. Learning [SN01, KAAD01, SM00b]. Least [Bel04, GR04, JY00, KL00a, AAEA03, CY02, Din02, JW03, Juk04, Liu00, MLA01, Wan04a, WC01, WW02a, WW03a, WHT04, Yan00b]. Least-squares [GR04, JY00, KL00a, AAEA03, CY02, Din02, GL00, WW02a]. Lee [Koz03]. Legendre [DSC01, MK03b, Ode02]. Leggett [AO01a]. Leighton [Tra00]. Leja [De 04]. length [Olu03a, SK02]. length-biased [Olu03a]. Lerch [LS04d]. Leslie [HL04e]. less [NR02b]. level [Deh04g, JD03, Lee01, OASAE04, XL03, ZAA01]. levels [JSdN04]. Levin [Thu04a]. Levin-type [Thu04a]. LHL [Pei04]. LHL-key [Pei04]. Liapunov [ESSA01, Sol02a, Yan03q]. Liapunov-type [Yan03q]. liberation [CM04a]. Lidstone [Yao02, Yao03a, Yao03b]. Lie [BEV02, BFN01, BFN03, ENR04]. Liénard [Yan03s]. life [AY04a, AY04b, SEB03]. Lifshitz [BAB01, SB04b]. lift [CS03a, MCS04]. lifting [GLVW00]. lifts [MS04a]. like [AS04h, CZ04, Dat04, Din01, Din04e, HL04g, KAK06, Khu02, Küç04, KE04c, sLqZ03a, LqZ04, Mar01, Noo04b, Waz02b, Waz02h, Waz04d, WG04c, Waz04i, Wu00, Zhu04b]. likelihood [Bel01, SM02b]. liming [Gho03]. Limit [Cic01, DXL02, Koł01, KS01, Mil03]. limitation [Sol03b]. limits [Ami01]. Lindstedt [EL04a]. line [AB04a, AS04c, GLWY04, JSLS04, Liu03g, MM02b, Ram01b, Shi04, Tag01d, Tag02b, Tag03b, YL04a, YÇAM04]. Linear [BFGG04, GM04a, GV00, HY04, AEL04, Abd03d, AD04a, AS03d, All04c, AD01, Asa04a, AV04b, BBV04c, Bai00, BP01, BRS04, CC03a, CW02, Cao03b, CY04, CH04, Cao04b, CC03b, CBK00, CG03, CC04a, Che04c, CTZ03, CPTZ04a, CW03, Deh04c, DG03, DW02, Din03a, EEB03a, EESESS03, EB02a, EL04a, EM03g, EOM03a, Far04, Far00, GHW04b, Gia03, GR03, Gül04a, GM03a, HES05, HAM04, HC04a, HHZ04, JLST04, JW03, JLMC03, JL03b, JJ04, Käm02b, Keş03a, Keş03b, Keş04d, KERG04, KM03a, LSJR00, LSY03, LLL04, LRHRD01, LZL00, Li04d, Li04e, LL03, LC03a, Lie04, LS03b, LYT04, LLW04, MM04a, MZ04a, Mae01, MM04b, MS04b, MMA04, MK04, Mař04b, Mar04c, Men02, Moh03, Moh04, MS01, MR04, Nak02, NCÁHCLP03b, Nak03, Nak04a, Nak04b, NN04a, NS03, OSL03, OASAE04, ÖÖ04, Pap00]. linear [Par03, Par04a, PP01, RA03, RW03, SS04a, Sab18, Sar04b, SZ04, ŞÇ03b, SK04, Sub02a, Sun04c, SM04d, SL04b, Tia03, TM03, WJ00, Wan01, Waz01a, WES01, Waz03i, WW02a, WWW03, WW03a, WQ03, WLW03, yWShX02, XL04b, YS00a, YZ04, YS01a, YS03, Zha02a, ZWRL03, ZW03a, ZW03b, ZW04a, ZClC03, de 00]. linear-exponential-quadratic-Gaussian [LC03a]. linearities [EBE03]. linearization [De 02b, JBMR02, MM04c, Ram04a]. Linearized [EOM03a, GLR02, Hlo04, KE04c, Luo02a]. Linearly [AGS03, Ram02]. link [HT04]. Liouville [AhL00, HS02, Has03a, Has04a, Has04b, MA04b, SH02, WJ04, Yan03i]. Lipschitz [Hua03, ZZC04]. Lipschitz-continuous [ZZC04]. liquid [RA03, Ram00a, Ram00b, Ram00c, Ram01a, Ram03b, Ram03c]. lithium [ZaYD04]. Littlewood [ÇD04]. living [BB02c]. LL [SMQ04]. LMI [Che04a, Che04b, PKW04, PKLW04]. loading [AAHAD04a, FRRSCS02]. Local [ADARAM04, IH01, RS04a, Yan04a, CT03a, Deh00, Deh03d, Deh03e, HA04, JBMR02, Olu04, Özd03b, SM04b]. localized [CX04].

Locally [Deh03d, AAH04]. location [Liu03c, Liu03d]. logarithms [LHL03]. Logistic [ADASM04, ADAMM03, FL03, GM00, Luo02a, SL04c, WL04a, Fen03]. lognormal [Ana03, AG04]. Long [FSLMC03, LBE00, EB04a, EB04b, ENAAM01, Kay04d, LZ04a]. long-memory [EB04a]. long-range [EB04b]. Long-term [FSLMC03]. long-wave [Kay04d]. longer [AS04g]. longitudinal [enNAaA02, HSE04]. look [CQL01]. loop [CO01]. Lorentz [Küç04, YÇA04]. Lorenz [EGB03a]. loss [ASM03, CBK00, ySW01, Wan02]. Lotka [AT04, EOI03, HL04d, LC02b, Ten00, Tin01, XCC02, XCD04b, XCD04a, YC04, ZJ04]. Low [HB04, Fan03, HLL03, KST00, WX01]. low-computation [Fan03]. Lower [Gün04, CPL00, CHL01, CDH01, NO03, Yan03l, Yan03w]. LQ [LF04]. LRT [ADAJM03]. LSE [EB04a, EB04b]. LU [WZ03, Zha01]. lubrication [AO04c, He04g]. Lucas [Kah05, TK04, YK03]. Lyapunov [AS04h, CJB02, McR01, Sol04, UvB01, WBW01, Yan03p]. Lyapunov-like [AS04h].

M [ASAI03, Tar02, ASAI03, Tar02]. M/M/ [ASAI03]. M/M/1/ [Tar02]. Macdonald [Abd01a, Abd02d]. Machine [dGKG^{+}00]. macro [Esc03c]. magnetic [AR04a, ASEM03, AS03a, EEE03, Ezz04, KB04a, Mef02, Mek04, RH04, Zak03a, ZS02]. Magneto [AAHAD04a, Rad03, enNAaAd04, Bis04b, ESH02, EEE03, Mur01]. magneto-convection [Mur01]. magneto-fluid [ESH02]. magneto-optic [Bis04b]. Magneto-selfgravitational [Rad03]. Magneto-thermo-viscoelastic [AAHAD04a, enNAaAd04]. Magnetoelastic [enNAaA02, AAHAD04b]. Magnetohydrodynamic [HKA04a, RH04, Zak03b, Zak04a, Sam04, Zak04c]. magnetostatics [Mef02]. Magnetothermoelasticity [EEK03b]. majorant [Han01]. majority [WW00a]. majorization [Kad04]. makers [Ino00b, Ino00a]. making [FI01, Ino00b, Ino03, OASAE04]. management [FSLMC03, Ghe00, Oda01]. Manifestations [ML00]. manifold [Din04a]. manifolds [Hab04a]. manufacturing [CL03b, NV01]. many [AAH04, FQ04, Pla03]. many-particle [Pla03]. map [Lee04a, HMM03]. MAPLE [JKP03, YB00a, YB00b, FH04, GKaM01, KU03]. mapping [ER03, MB04, Wan00c, YŞ00b]. Mappings [Sor01b, Abb04b, Kır04, KÖ04, LDKU04, Özd03a, ÖK03d, ZL04a]. maps [Amm04, JYC04, ÖA04, WL04d]. Marcel [Kan00a]. marginal [BD04b, Bel01]. mark [HA02]. mark-resighting [HA02]. Markov [AB02b, CD02, EG04a, EG04b, EP00, GP03, Gra03, HHK01, LCN04, Mun03, NU00b, Oht04, WW02b]. Markov-modulated [AB02b]. Maruyama [CLF04]. masked [SEB03, Sar03b, Sar04a, Sar04b]. mass [Cap01, DKV04, EA04d, EESF04, HEM03, Pan08, Ram00a, Ram00b, Ram00c, Ram01a, Ram03b]. masses [EGT04, EG04f]. Master [Ano01i, TE00, TE02a]. material [AAHAD04b, HL04a]. Math [AaAaZ00b, AaAaZ00a, Bil07, CC03c, EOM08, KAK05, KAK06, Kas06, Kum09, Sab18, dCD06]. Mathematica [KL02]. Mathematical [EORI01, Hil04, KE02, Pet02, Pop04, QWY02, QW03, Abd03b, Dje00, EOI02, Far00, Gho03, MT04, NN04a, yN04c, NS04, OAB04, Sev02, SAV04b, VK04]. mathematical-programming [NN04a]. Mathematics [HV08, HES05, Mus00a, Mus00b, Pan08, SDR03, Yam05b, Esc02, Esc03b, He04f, MCC01, MB04, Pop04]. matrices [Ana01, BCT04, BW04, CWL00, CZT04, CGVC04, CWS02, CEM04, CT04a, CT03b, EM03b, EM03c, EM03e, Gün04, HZZ03, Kar04, LWL00, Liu03h, gLgW04, LW04e, ML00, ÖÖ04, STB03, SB03, SEK01, Sta03, ST04, TB02, WZ03, Wei01, WZ04d, WN04, WC04b, YB03, YB04, Yam05a, Yam05b, Yin03, Yua00, Zha01, Zha02b, ZZ04b, ZHZ04]. Matrix [Bil03, MS04d, MRŽ04, Ahm04a, Bay04, CF03, CL04a, CC04c, EM03f, EM03a, EM04c, EM04b, ESEA02, ESAD04, Gar01a, HI04a, Ji02, KV03, KB04b, KU03, LW04b, LW04c, Liu04i, MMER03, NK04, ÖKOD04, Rak04, SW04a, Sun04c, SM04d, Tad01, Taw00, Tia04b, rWS04, WQ04, WZ04b, Wei00a, Wei00c, WWW00, Wei02b, WZ04e, Yan04g, ZTD04]. max [Ghe00]. max-min [Ghe00]. maxentropic [GG02a]. maximal [EA04b, Tia04a]. Maximum [BCI03, ÇG04, Gzy02, SM02b, Tag00b, Tag01c]. Maxwell [BV02, HNA04a, SB04b]. McKendrick [Far04]. mean [ADAAM03, Ahm04b, EN01, EBA03, Gal03, Gzy02, Hür04, Mut03, RS04a, Sah04b, SD01]. meaning [Gia03]. means [Ame01, Kır04, KÖ04, LTS02, MS04d, Özd03b, Özd03a, ÖK03d, PS00, Xu04]. measles [AST04]. measurable [Li03a]. Measure [JSdN04, ASM03, IKS02, Kin01]. measurement [BRVI00, Has04b, JAK04b]. measurements [HS02, Has03a, Has04a, LCH00, MTK^{+}00, SH02]. measures [CD02, Cho02b]. Measuring [Hür04, JAK04a]. mechanical [ENAAM01, IN04, RHB04]. mechanics [AH02]. mechanism [KE02]. mechanisms [DYH04]. media [AQBT04, HEM03, LPS02, (Mu04, Qui04, Ram02, Ram03d, Ram04b, Ram04c, Yan02h]. medical [Kee03a]. medium [AAAaZ99b, AAA03, AaAaZ00b, AS02a, AS03c, AG03, AO04b, AM00, AA04d, BS04, CX04, EA04d, ES03a, Elb01, EB02b, Elb03, EB04f, EEE03, EEK03b, Has03b, HSE03, HEM04, HSE04, HNA04b, LWZ00, SM04a, Waz01e, Xu04, Zak03b, Zak04b]. meeting [BL01]. Meetings [Ino00a]. MEI [JLMC03]. Melfi [SDR03]. Mellin [Tag01b, Tag01e, Tag02b]. membrane [Zay02b, Zay02c, Zay03d]. membranes [Cap01]. memetic [DM04]. Memory [Ric03, BPJ03, Bae04a, CD03a, Cao03a, CCS04a, EB04a, ECL02, JT04, Kay04h, KNJ04, LC04b, MZ04a, PB02, TCL02, Taw00, dJ04]. Menten [XC02c]. Merton [Mun03]. mesh [KS03a, Kum03b, VR02a, Xen03]. mesh-chopping [KS03a]. meshes [Bog04, qLzWcC03]. meshing [SS03]. Meshless [CM04a]. message [Sha04b, TJC03, WC04a]. metabolic [Bay04, YB00b]. metal [DKV04]. metallic [Yen04]. Method [Fen04b, Kru03, RC03, Abb03, AA04b, AEL04, AB02a, AD02, AAMK01, ANS02, AZ04, ASN03, Aln04, AO01b, AB03b, ABEM04, Ana01, Aru03, Asa04b, Aya03, Aya04a, Aya04b, AÇB03, BH02, BB02a, BB02b, BB02c, BJ03, BB03, BBV04a, BBV04c, BBV04b, BD04a, BJ04, BSJ04, BJS04, BBV04d, BV02, BH01, Bao02, BTBI03, BBI03, BBI04, BI04a, BFN01, BEAB04, CPL00, CHL01, CC03a, CLF04, Cao04b, CJB02, CAH02, ÇB03a, ÇAB03, ÇB04b, ÇAB04, CY02, CL03a, CP03a, CL04b, Cha04b, Che01b, CHL02, CC02, CX03, CL04c, CC04a, CC04b, CDH01, CP03b, CTZ03, De 01, Deh00, Deh01, Deh04a, Dob00, EGBH03, EKE02, ER03, ES02b, ESAA03, ESK04c, ESK04d, ESSA01, ETBAN04, Esc03a, EK04, FH03, FS03, Gal03, Gla04, Gon04, GK02, GC01, Gum02, GG02b, GL00]. method [GHL00, GHSJ00, GH01, GKaM01, Han01, HHJ01, HZL02, Has03a, He00, He01, HOL02, HP02a, HLO02, He03a, He03c, He04a, He04b, He04d, HR01, HT00, Hos04, HMM03, Hua03, IOAB02, IbS01, IRR04, IRS04, IN04, JHS^{+}04, JLST04, JCL00, JWL03, Jan04, JGW02, JLMC03, JBMR02, JJ04, Kah05, Kah06, KS00, KR03a, KR03c, KR03b, KR04, KS03a, Kay03a, Kay03b, Kay04a, KES04a, Kay04b, KES04c, Kay04e, KEE03b, KL00a, KH02, KLK02, KS02a, KEB03, KU03, KC00, Kul03, Kum02, Kum03a, Kum03c, Kum03d, Kum03e, Kum09, Kuo04, KE04a, Kwo03, LLP02, Lee00, Lee04b, Les01, LW03a, Li04f, LqZ04, LCXZ04, LZ04a, LW04a, Lia04, Lie04, LC03b, Liu00, Liu04g, qLzWcC03, LZ04f, LBE00, MMR01, MMA04, MS04c, Man03, MLG00, MM04c, MR00b, MY03, MIM00, MSJ04, Moh04, MD00, MS01, MWAF02, (Mu04, MLA01]. method [NK04, NR02a, NR02b, NU00b, yN04b, yN04a, Ode02, OIO03, PR04a, PR04b, PdR01, QS04, Ras03a, RCS03, RKS04a, RR02a, RC04a, RC04c, Sal03a, SS00, Sal00, SKM04, SS03, SR02b, yS01b, She03, SV02, SH01, ŞÇ03b, Sol04, SEK01, ST04, SYK03, SYY04, SZ03b, SM04c, Sun04b, SMQ04, TM04a, Thu04c, VR04b, VR02a, VR02b, VCD04, Vul00, Wan00b, Wan04c, Waz00a, Waz01d, Waz01e, WES01, Waz01g, Waz02f, WG04b, Waz04h, Waz04j, Wu00, WC02, WSX03, WHT04, Xen03, XL03, Xu04, XL04b, Yan00b, Yan03l, YX04, Yen04, YK03, Yua03, wYjSjS03, YW02, YW03, ZG03, ZaYD04, Zhi04, ZH04, ZW03c, Zhu04c, Zhu04b, ZZ04c, Zhu04a, ZG02, bSI01, HS03, LMG00, Sen01]. methodology [JSdN04]. methods [AO00c, AKKN04, ASM03, All04c, ÁR02, AR04c, ABP04, AA04c, AV04b, BH03, Bad01, BE02, Bai00, Bog04, BK02, Cao03b, CLL00, CL01, CC01, CC03e, CL03b, CQL01, CPTZ04b, CCK04, CO00, Dat03, Deh03b, Dis01, DNS03, EB02a, EHM03, ES02a, ESK04b, FS04a, GLVW00, GLR02, GG00, GH04, HH01, JY04, KA00, KG03, Kwo03, Lay02, LLB03, LW04c, LWL00, LS03b, MK03b, MK04, Mar01, MC00, MAD04, NWX03, NNAK04, Ona02, PGyL03, PZZF02, Ram02, Rui03, SR02a, SR04a, SZ03a, SL04a, Shi04, SY03, SY02, Tia03, Tsi01, TS02, Tsi03, VR03a, VR03b, VK04, Wan00c, WS01, WW02b, WX01, WXO02, WC04c, Wu04b, XL04c, YAOY03, YY02, YS03, Zan00, ZZ02a, Zha00, ZZ04a, Zha04b, ZW04a, dK02]. MFSK [WW00a]. MHD [ENEA04, RPT04]. Michaelis [XC02c]. microbiology [VK04]. micropolar [AEG03, HEM04, IH01, Mur01, Mur03, SR04b, Zak04b]. middle [Waz04g]. midpoint [Kır04, KÖ04]. Mie [ML00]. migration [KE04b]. min [Ghe00]. Mindlin [YX04]. mineralogy [LPS02]. Minimal [Guo00, LLB03, SZ03a, Tia04a]. Minimax [PY02]. minimization [Liu02b]. Minimum [LP03, Din03a, DR01, QY00a, QY00b, ŞBC03, WW02a, ZClC03]. minimum-norm [WW02a]. Misra [NK01]. missing [SV04]. Mitigation [Lun00]. mitigators [Liu02c]. Miwa [LCXZ04, LJ04]. Mixed [Mef02, Noo03b, d’O04, AAAE03b, ANS02, AO04b, AR04d, BJD^{+}03, Cha04a, Che01b, De 01, Din00, Din04d, Din04e, ESES04, Elb01, EB02b, Elb03, EESF04, Has03b, HSE03, HEM03, HSE04, IH01, JCZ02, KL00a, KH02, KNJ04, Lay02, LW00, LDKU04, MM04a, Noo03c, Noo04b, NNAK04, OSL03, Pan08, Pap00, Waz01a, Waz02g, Yan02a, Zay02a, Zay03b, Zay03e]. mixed-integer [Pap00]. mixed-type [BJD^{+}03]. mixing [De 02b, LLCC03]. mixing/adsorption [LLCC03]. mixture [AESS00b, MR04]. MLC [TLLKB03]. mobile [CS00, ÖM02, SWL00]. Möbius [CSD01]. Modal [EB03a]. mode [Bou02, Boy03b, EB03b]. Model [Kan00b, AGES04, Abd03b, AD00, AB03a, Ald04, AQBT04, Bao02, BHH02, BBA03, BG04, BJD^{+}03, BCT04, CBK00, Dje00, Dun02, EB04b, EG04b, EORI01, EOI02, EOM04a, ESK04b, Far00, GB03, Gar01a, Gho03, GR03, Gum02, GM03a, HIN04, HNA04a, He04g, HB00, Hür04, IYM00, JK03a, JK04a, Jah03, JB02, JD03, JBS04, JDV04, Jol00a, KE02, KL03, KEB03, KSC02, LWZ00, LC04a, LG04b, MG03a, Mog04, NG02a, OAB04, PR04a, Pet02, PTG03, Pop04, QBK02, QWY02, QW03, Rao04b, Sak03b, Sak03c, Sar03a, Sar04a, Sar04b, SM00b, SV02, SP03a, SPH04, TET02, TE03, TLLKB03, Tie03, Ven03, Waz00a, WFT03, XCCC04, XC01, XC02a, XCD04b, XCD04d, XCD04e, XCD04a, XCD04c, XCD04f, YL04d, ZC04a, ZL03, d’O04]. Modeling [BT00, Yan01a, BTBI03, KW04a, MB04, RM03, Rao04a, WBW01, WW02b]. modelisation [ALDP01]. Modelling [GCSS04, ML04d, Vou03]. models [Ada01, Ada03, AB02b, AFRH02, Bel01, BBN04, CLC04, EB04a, EGAR03, EGB03b, EG04a, HV01, HL04e, JLS^{+}04a, JHS^{+}04, Kay04c, LMG00, Mel00, Mel01, NS04, SM02a, TE04, WZ01b, YC03, YAOY03]. Modification [ER03, WES01, dlS03]. modifications [SM02b]. Modified [AS01, CHC04, Kat01, SK02, YS03, Abb03, AÇB03, BH02, BB02b, ÇAB04, Coo04, Dem03c, EL04a, ES02b, HWWM03, JJ04, Kay03b, Kay04b, KE04a, LCH02, LL04b, Li04f, MAD04, PSS00, SS00, Wan00b, Waz01e, Waz01g, Waz04i, Wei01, Wei02b, Wu04a, Yas03]. modular [Isı03, NV01, ÖKOD04, Yil04]. modulated [AB02b, TLLKB03]. Modulation [BD04b, BD04c]. modulo [ÖAD03a]. modulus [Bil04, Bil07, Din03a, EEKS04, KŞ04b, Özd03b, ZClC03]. MOLP [JF04]. Moment [AY04a, BGVHN02, IPPT03, Tag00a, Tag01a, Tag01c, WHM04]. moments [AYW04, IPPT03, Tag02a, Tag02c, TV03, Tag03a, Tag03c, Tag03d, YDW02]. money [SG01]. monitoring [OAB04]. monochromatic [Zed02]. monomiality [Che03, DGR04, SC03d]. Monotone [BM02, CHL01, HZ04b, Jan04, Kam02a, Yao02, Yao03a, AY04b, AD00, CDH01, GLWY04, Has04a, JGW02, Li03a, LW00, PGX00, ZZ02a, ZL04a, pFjH03]. monotonicity [WL04d]. Monte [SAV04b]. Moore [Rak04, Tia04b, Wan04a, Wei00b, WW01, Wei01, WW03a, Wei03b]. morphology [ML00]. morphology-dependent [ML00]. mortar [CX03]. most [JK03a]. motion [AEE04, EEB03a, EE03, EE04b, EG02, EGE03, EG03a, EG03b, EG04c, EGT04, ESEE02, EEE02, El 03, HES05, HAM04, WC02, Zed02, ZK03]. motions [AAH04, RCC04]. moving [EGT04, EB04e, Rao04a, TE03]. MRI [PZZF02]. MS [CLF04]. MS-stability [CLF04]. Mu [CJT03]. Multi [DF04, JAK04b, LLL04, Tse03, CWZ02, HC04b, HL04g, JAK04a, JL04, Koł01, LCH00, LLB03, LY02b, Liu03a, Liu03b, LL04c, MB04, OASAE04, Pom01, Ram03a, Ten02, YCH04, You04, Zay02c, Zay03b, Zay03a, Zay03d, Zay03c, Zay03e, Zay04a, Zay04b]. multi-authority [JL04]. Multi-component [JAK04b, JAK04a]. multi-connected [Zay02c, Zay03b, Zay03a, Zay03d, Zay03c, Zay03e, Zay04a, Zay04b]. multi-dimensional [CWZ02, Pom01]. multi-functional [MB04]. multi-homogeneous [LLB03]. multi-input [Ram03a]. multi-level [OASAE04]. multi-linear [LLL04]. multi-objective [OASAE04, You04]. Multi-order [DF04]. multi-pantograph [LL04c]. Multi-party [LLL04, Tse03]. multi-point [LY02b, Liu03a, Liu03b]. multi-proxy [HC04b]. multi-secret [YCH04]. multi-signature [HC04b, HL04g]. multi-species [Ten02]. multi-spectral [LCH00]. multi-state [Koł01]. multiattribute [Jol00b]. multicast [WC04a]. Multigrid [Che01b, AS01, AV04b, Dis01, GZ00, HZL02, qLzWcC03, MdL04, RdL01, She03, Xu02, ZH04]. multilayers [Lee04b]. multilevel [Che01b, SKM04, Zha01]. multimachine [HML^{+}02]. multinomial [Ma01, MPS04]. multiobjective [AD01]. Multiple [AS04d, AO00b, CSA03, Guo03b, Guo04, Lan04a, Lan04b, LL04a, LS04c, LOZ02, Ma04a, SL03a, SL03b, zBxF03, CCS03, CZ04, CJ04b, CCS04b, ECL02, HC04a, HHZ04, Jol00a, Lee01, LG03, LG04b, LG04c, Qui04, RM03, Tan03, Tie03, VCV01, YR01, Zhu04b]. multiple-key [CJ04b]. multiplexing [EP00]. Multiplicative [Rui03, Tri04]. multiplicity [Jol00b]. multiplier [SLC04]. multipoint [SY03]. multiprocessor [Kay04h]. multiprojections [CCKS02]. multiresolution [Ama03]. multiscale [CC02]. multiserver [Tar03]. multisignature [CCH04, LC04d, ZX04]. multisignatures [WH02]. multisingular [Has04b]. multisoliton [LZ04b]. multispectral [MTK^{+}00]. multisplitting [CW02, Wan00c]. multistep [GLM04]. multisymplectic [WWQ04]. multivalued [IG03, ÖA04]. multivariable [LRHRD01, RR02b]. Multivariate [YXC03, Ahm04b, EB04a, EB04b, HZ04c, Jol00a]. multiwavelets [CC04c]. musings [Nys01]. MVLUE [HM00].

N [Pan08, WA01, Tra00]. Nagumo [Abd04]. narrow [SS02]. narrow-stripe [SS02]. Natural [HEM04, CL01, EA04d, OAB04, SL04a]. naturally [AQBT04, TY04]. nature [Ver03a, Ver04]. Navier [Çağ04, Cao04a, ESS04, Hlo04, KSS02, LLP02, Man03, She03, SM04c, Zed03]. navigation [FT00, Far00]. NBUFR [EBSAG04]. NDEs [HC02a, HC02b]. near [BD04b, GS00, Gar01b, NG02b]. near-circulant [NG02b]. nearest [AA04a, MMER03]. necessarily [dlS03]. Necessary [EA03, LC01b, Sav04a, EOM03b, Juk04, YE04]. negative [ASAB03, HL04c, Ino04a, Ino04b, Kad03, KA03a, KÖY03, OK03a, Orh03a, Orh03b, OK04, Sak03d, WL04a, Waz03c]. nest [ML04d]. nets [NV01, NK01]. network [ASEM03, qLzWcC03, MIM00, Opp00, RHB04, SM00b, TCS01, TCL02, WFC04, Zha04c, ZHD04]. networks [Bay04, CD03a, Cao03a, CC03d, CC03c, CHC03, CTHK03, Gop04, HM04, HC03, LH04b, Li04g, LC04b, Mak04b, MG03b, TLX04, Vou03, Wei00b, YB00b, ZC02, ZZC04]. Neumann [CHL01, EK04, GGB03, KE04b, SL03b]. Neumann-type [EK04]. neural [Cao03a, CC03d, CC03c, CHC03, Gop04, HC03, LH04b, Li04g, MIM00, MG03b, PKLW04, SM00b, TCS01, TCL02, TLX04, WFC04, Wei00b, Zha04c, ZC02, ZZC04, ZHD04]. neurocomputer [SN01]. Neurofuzzy [OAB04]. neurons [ZHD04]. neutral [Abd03a, CH04, Che04a, Che04b, CLC04, Fu03, Fu04, GXL04, HC04a, HHZ04, Jia02, JL03c, Jia03a, KSM03, Lac03, LC01a, LC01b, LS03a, Li04b, Lin04a, Lin04b, LG04b, LG04a, LGZ04, LG04c, Min04, OLT04, Par03, Par04a, PKW04, PP01, Şah04c, Sak03a, Sak03d, SW03, SW04b, WX04, YC03, YC04, ZAA01, ZL01, ZY04, ZL04b]. Neutrality [RS01b]. neutron [wYjSjS03]. Newcomb [Ino02a]. Newton [Abb03, AA04b, EMED03, EM03d, FS03, Gal03, GH01, Hua03, Sha04a, Ver03a, Wu00, ZG03, Zhu04b]. Newton-arithmetic [Gal03]. Newton-like [Wu00, Zhu04b]. Newtonian [Pan08, AR04a, EEE02, EESF04, HNA04b, He04g, YTM00, Yür04a, Yür04b]. Nicholson [SWY00, SZ01]. Nicolson [Deh01]. nilpotency [KA03b]. nilpotent [Gin04, KA03b, KY03, ÖAD03b, Özk03]. No [CY01b, LCH00]. node [Nar01]. noise [NHCLPSR04, Waz03h]. noises [EG04c, NCÁHCLP04]. noisy [GG02a, YDW02]. nominally [AQD04]. Non [DEK04, Far00, MMT00, Ram04a, Ras02a, SM04b, AAAaZ99a, AAA03, AEL04, AESS00b, AaAaZ00a, Abd03a, AR04a, AZ02, AO04b, AS04g, BH02, Bai00, BP01, BT00, Bis03, Bis04a, Bis04c, CLW02, Che01b, CX03, CC04a, Che04c, De 02b, Deh00, Deh03b, Deh03d, Deh03e, Deh03g, Deh04c, DWC04, EEB03a, EESESS03, EA04d, EB02a, EBE03, EGAR03, EG04e, EG04d, EL04a, EME03, EMR03, EOM03a, Elb01, Elb03, EESF04, EEB03b, FZ04a, Far04, GLWY04, GM03a, HES05, HNA04b, HAM04, He04g, Hel01, Hlo04, HA04, JK04a, JJ04, Kum03b, LL03, Mar04c, Mek04, MR04, Mur03, Mus00c, OASAE04, Pan08, Par04b, QS04, RA03, SS03, SZW03, SE04, ŞÇ03b, SK04, SMQ04, TLX04, Tar02, Tar03, Ten00, TT04, Tun04c, XFL04, YZ04, YTM00, Yür04a, Yür04b, de 00, dlS03]. non-analytic [BH02]. non-Archimedean [JK04a]. non-autonomous [CLW02, FZ04a, Ten00, TT04, Tun04c]. non-classic [Deh03g]. non-conforming [SS03]. non-convex [QS04, TLX04]. non-damping [SMQ04]. non-Darcian [Elb01, Elb03]. non-Darcy [AO04b, EA04d]. non-degenerate [FZ04a]. non-denumerability [Mus00c]. non-differentiability [SE04]. non-electrically [AESS00b]. non-empty [Tar02, Tar03]. Non-finitely [Ras02a]. non-fragile [Par04b]. non-FSAL [EME03, EMR03]. Non-gaussian [MMT00]. non-Hermitian [Bai00]. non-homogeneous [AAAaZ99a, AAA03, AaAaZ00a, AZ02, DWC04, EGAR03]. non-independent [EG04e, EG04d]. non-Kerr [Bis03, Bis04a, Bis04c]. Non-linear [Far00, AEL04, BP01, CC04a, Che04c, Deh04c, EEB03a, EESESS03, EB02a, EL04a, EOM03a, Far04, GM03a, HES05, HAM04, JJ04, LL03, Mar04c, MR04, OASAE04, RA03, ŞÇ03b, SK04, YZ04, de 00]. non-linearities [EBE03]. Non-local [SM04b, Deh00, Deh03d, Deh03e, HA04]. non-monotone [GLWY04]. non-necessarily [dlS03]. non-Newtonian [Pan08, AR04a, EESF04, HNA04b, He04g, YTM00, Yür04a, Yür04b]. non-oscillatory [Abd03a]. non-periodic [EEB03b]. non-polar [Hel01]. non-REM [BT00]. non-self-adjoint [Che01b]. non-selfadjoint [CX03]. non-smooth [SZW03]. Non-standard [Ram04a, Deh03b, XFL04]. non-stationary [Hlo04]. non-truncated [AS04g]. Non-uniform [DEK04, Kum03b, Mek04, Mur03]. non-uniformity [De 02b]. nonautonomous [AT04, AD00, EOM03b, LC02b, SC03c, XCC02, ZJ04]. noncentral [vAG03]. noncoercive [BC04]. noncompact [HHJ01, Waz02a, Waz03f, Waz03c, Waz03g, Waz03j, Waz03k, Waz04b, Waz04f, Waz04h, Waz04k]. nonconforming [CLL00, KL00b, KLK02, KHL02]. noncontinuous [EN01]. nonconvex [Din01, Noo04c]. Nondeterministic [Lew01]. nondifferentiable [Han01]. Nonexistence [YGL01, ZY04]. nonfragile [PJ04]. nonhomogeneous [Li04d, Ram02, Sad04a]. nonisothermal [GG02b]. noniterative [LSJR00]. Nonlinear [CC01, HML^{+}02, Yan03m, Abb03, AA04b, Abd03d, AD00, ALP03, AE04, AZ04, AD04b, AV04b, BB02b, BBV04d, Bad01, BPJ03, Bae04a, Bae04b, zBxF03, zBxF04, BD04b, BD04c, BDS03, BDGG04, BM02, CPL00, CH02, CW04b, CJB02, CY01c, CT03a, CP03a, CL04b, Che01a, CHL02, CC03e, CZLZ04, Dar04, DK04, DX04, DKX00, Deh04g, Dem02b, Dem03b, DMS01, DE00, DF02, EB04d, EGBH03, EKE02, EOM03c, ERN01, ES02b, ESSEF04, ESAD04, ESS00b, ESSA00, ESSA01, ES03b, FH03, FQ04, FNO04, FS04a, Gal03, GHW04b, Guo03b, Guo04, GHSJ00, HA03, HDZ04, Has04c, He03a, He04d, He04e, HGS04, IOAB02, JL03a, JL03c, Jia03a, Jia03b, JLM^{+}03, JCZ02, KES04a, Kay04c, KI04, Kay04g, Kha03a, Khu03b, KR01, KSM03, Lee00, lLgC03, Li04b, LR04, LNS04, LS04b, Li04c, LZ04e, Lia04, LC02a, LC03b, Liu02a, LQL03]. nonlinear [LK04, Liu04c, MdS04, Ma04a, MM03, MdOPF04, MY03, Min03, MZ04b, MSJ04, yN04a, OLT04, PY02, PB02, PGX00, PGyL03, Ram02, Ram04a, RCS03, Ras04f, RKS04b, RT03, Sak03a, Sak03e, Sak04, Sha04a, Sha02, SW04a, SD02a, Sol02c, SC03c, SN01, SW01, TE02b, TY04, eT04, Wan00c, WZ04c, Wan04c, Wan04e, Waz00b, Waz01a, Waz01d, WES01, Waz02d, Waz02e, Waz03f, Waz03c, Waz03e, Waz03d, Waz03g, Waz03j, Waz03k, Waz04b, Waz04d, Waz04e, Waz04g, Waz04j, WSX03, lWCH03, XYD02, XD03, XL03, Yal02, YL04c, YGL01, Yan02d, Yan02c, Yan02e, Yan02f, Yan03b, Yan03d, Yan03e, Yan03i, Yan03j, Yan03a, Yan03r, Yan04d, YAOY03, Zan00, ZZ01, ZL01, ZC03, ZY04, Zha04a, ZS04a, ZY00, ZC04b, ZLS04, ZW03c, ZZ04c, ZL04b]. nonlinearities [Abu00, ALO03a, EB03b, KH01, YWY03]. nonlinearity [Waz04g]. nonlinearly [Vaj01]. nonlocal [CT03a, CMF04, LX03, Liu04b]. Nonmonotone [Sun04b, Zan00]. nonmonotonic [Zhu04a]. nonnegative [AO00b, Guo00, HDZ04]. Nonoscillation [Yan02e]. Nonparametric [AY04b]. nonplanar [CC00b]. nonquadratic [HA03]. Nonresonance [Yan03n]. nonselfadjoint [CC04b]. Nonseparable [Ama03]. nonsimilarity [IH01]. Nonstandard [EM03h, Deh04a, Deh04d, SD02a]. nonstationary [GF00]. nonsymmetric [CY04, NK04, WZ04d, Zha04b]. norm [JLS^{+}04c, WW02a]. normal [ADAMM03, ADAJM03, ADARAM04, Bou02, Bry02, EB04b, FRRSCS02, Sah04b]. normalization [Ras02c]. normals [HIİ04]. norms [Gün04, SB03, TB02]. Note [WXO02, YY02, ASM03, CWL00, CF03, CY04, CTZ03, DCS03, Dem02a, Dem03c, EM03a, EM04b, EOEA00, GG02b, Hau04, JW03, KSJ02, LW03b, Lin04a, Liu04i, Liu04b, Tag03a, Tun04b, UvBP00, WZ04d, WWL04, WZ04e, Xen03, Xia03, Yan03o, Yin03]. Notes [LSK00, Rol02, ZZ04b]. novel [ML04a, Thu04c, Wan04c, Yen04]. NP [dCD06, dCD03]. nuclear [Sai02]. nuisance [ADARAM04]. null [yN04b, yN04c]. null-space [yN04c]. number [AA04a, Boy03b, LLB03, MC00, TE03, TE04, Tan03, WX03a, WWW03, ZAH02, Zay03b]. numbers [EMED03, JK03b, JKP04, Kah05, Kar04, KS03b, Kır04, KÖ04, Liu04d, Mus00a, Mus00b, Mus00d, RS02b, ŞBC03, TK04, WX03a, WWW03, WW03a, YK03]. Numerical [AKKN04, All04c, Asa04a, BD04a, BEAB04, Cao01, CO01, Çat03, Çat04, ÇKB03, ÇB04c, CMOS01, Deh02, Deh03f, Deh03e, Deh04b, Deh04c, Deh04d, Deh04e, Deh04f, DKV04, DNS03, EGBH03, FH04, Fat04, Gül04a, HA03, HS03, KP02a, KS04a, Khu03b, KW04b, LC04c, LPS02, MM04b, MR03, MM04c, PGyL03, QBK02, Ras03c, Ras03b, Ras04d, Ras04e, RKS04b, RR02a, RC04c, Sab18, SWY00, Tag01c, Tag01d, Tag02b, Tag03b, TV03, WWQ04, XZ03, ZaYD04, AMI03, AN03, AZ04, AÖ04a, AO04c, BH02, BMMRS04, Bog04, BJD^{+}03, BBN04, CJB02, ÇB03a, ÇB03b, Çel04a, Çel04b, ÇB04d, Cha04b, Cho02a, DSI04, DD02, DF04, EEB03a, EA04b, ERN01, ESK04b, ESK04d, EEE03, EK04, FL02a, FL02b, GLVW00, GLM04, Gar01b, Gül04b, Gum02, HZL02, HC01, IOAB02, Jay03, JLM^{+}03, KP02b, KS00]. numerical [KR03c, KR03b, Kay03a, Kay03b, Kay04c, KI04, KES04b, KES04c, Kay04d, Kay04e, KA03c, KW04a, KH01, KE04c, LWT00, Lee04b, LS03b, LL04c, Lun00, MS04c, Man03, MR00b, MG03a, Moh03, Mun03, (Mu04, NZP^{+}04, Nar02, NR02a, OOAA04, Pet03, RdL01, Ras04c, RG03, SMF04, SR02a, SR02b, yS01b, SV04, SB04b, Tag01a, VR02a, Waz01g, WL04d, ZC03, ZLZ03]. numerical-analytic [AZ04]. Numerically [BSJ04]. Numerov [TS02]. Numerov-type [TS02]. nutrient [Bon02, GR03, JB02]. nutrient-plankton [GR03]. nutrient-prey-predator [JB02]. NWUFR [EBSAG04]. Nystrom [EMR03].

O [LCH00]. object [Gra03]. objective [AS04d, OASAE04, VCV01, You04, ZWRL03]. objects [BR01]. oblateness [EE03, EE04a]. Oblique [CT03b, MM04a]. observability [CFS04]. observation [JHS^{+}04]. observations [Ami01, Gha03, GM04a, NCÁHCLP03b, NCÁHCLP03a, NHCJLLP04, NHCLPSR04, NCÁHCLP04]. observed [KST00]. observer [Lie04, Par04a]. observer-based [Lie04, Par04a]. obstacle [NAS04]. obtain [CHL02]. obtained [Waz01d]. obtaining [Waz04h]. occupancy [WW00a]. occurring [AQBT04]. ocean [MTK^{+}00, MSO00, TRC03]. ODE [BK03, GhW04a, TM04a, WXO02, lWCH03, Yan03k]. ODEs [MR00b, Sha03]. odor [LLCC03]. offensive [Kat01]. offs [BS00]. offsets [KG04]. oil [Bon02, NK03]. Oldroyd [HKA04b, HKA04a]. Oldroyd-B [HKA04b]. One [AST04, AAH04, ALO03a, Boy03a, CG03, CJ04b, DRP04, Deh03a, Deh03f, Deh04d, Deh04g, Deh04h, Jay03, KSC02, KE04b, LOZ02, Moh03, PS00, RM03, RR02b, SB04a, SM04b, TM04b, VCD04, WBW01, Zak03b, Zed02, ZK03, ZG03]. One-dimensional [AST04, AAH04, ALO03a, Boy03a, Deh03a, Deh03f, Deh04d, Deh04g, Deh04h, LOZ02, SM04b, Zed02, ZK03]. one-phase [KE04b, VCD04]. one-predator [KSC02]. one-space [Moh03]. one-step [ZG03]. one-way [CJ04b]. only [Wan04d]. onset [HSE04]. Open [Sor01c, MAK04a, Rao02, Rao04a, YR01]. operating [Ana03]. operation [Bay04, WC02, WC04b, WC04c]. Operational [Dat03, AYW04, DSC01, Hür04]. operations [Liu03d]. operator [AA04d, BCI03, CC03b, CMF04, DW02, ESSEF04, pFjH03, HS02, Has03a, Has04a, Has04b, Has04c, KSL02, lLgC03, LS04e, Mağ04, MZ04b, MA04b, Moh04, Moh00, Noo04a, Ron03, Sah04a, SS00, SH02, WQ03, ZC03, ZClC03]. Operators [DW04, SS01, Abd01b, DRS03, Dat03, Din03a, HA03, JCZ02, Kok03, LW00, RW03, SK04, Waz00b, WES01, XC04, Yan03b, YXC03]. optic [Bis04b]. optical [KIY00, LCH00, MTK^{+}00]. Optimal [CCY04, EG03a, EGB03a, EGAR03, EGB03b, EG03b, EGB03c, EGT04, EKE03, JM02, LHL04, LLW04, MdL04, Oht04, RAH01, WW04, Cao01, EKE02, Ghe00, GL00, HA03, JBS04, JK04c, Jol00b, ML04d, OASM04, RHB04, Sub02a, Yas02, YS01a]. optimality [Sla03, Sla04, YE04]. optimistic [FI01]. Optimization [RdL01, SH01, AGS03, Bel04, CP03a, DG02, FQ04, GHW04b, GHL00, HH01, HV01, Kah06, LWT00, Lee00, Liu01, Liu04g, Mae01, Moh00, PKLW04, QWY02, QW03, SZW03, SZ04, Shi04, SSPA01, SYY04, Sun04b, TLX04, WZ04c, WXO02, YAOY03, ZWRL03, Zhu04a]. optimization-based [GHL00, Lee00]. optimizations [JT04]. Optimized [HB00, EMR03]. Optimum [Meh00a, Meh01]. option [AA04c]. optional [MANAM04, MAR04a]. options [ML03]. orbit [CCC01, LBE00, Lun00]. Order [Li03a, Abd03a, AEC03, AO00b, AS03d, ASN03, AB03b, ABP04, AF03, Asa04b, BB02a, BDD00, BP01, BCI03, CPL00, CHL01, CH03, CCP03, ÇG04, CZT04, ÇB03a, CDH01, Cic01, CW03, DRS03, DF04, DH04, DS03, EA03, EGBH03, EOM02, EOM03d, EOM08, ESK04a, ESS00a, FNO04, FS03, FS04a, GhW04a, GG00, Gia03, GL00, Guo00, Guo02, Guo03a, Guo03b, GM03b, HM04, Has04d, HC02a, HC02b, He03g, HZ04b, JK03b, JKP04, Jan04, Jia02, JGW02, JL03a, JL03b, JL03c, Jia03a, Kad04, KR04, KS03a, Kay03a, Kay04g, Keş03a, Keş04d, KEE03b, KA03c, Khu03b, KL00a, KH02, KLK02, KR01, KM03a, KSM03, KNJ04, Kum02, Kum03d, Kum03e, Kum09, LM04a, Lac03, LL04a, LRHRD01, LC01a, Li03c, LS03a, LW03a, Li04b, LNS04, LS04b, Li04d, Li04e, LS04c, LZ04b, LW04a, Lin04a, Lin04b, Lin04c, LY02a, Liu03f]. order [Liu04a, LQG04, LG03, LG04a, MM04a, MdS04, MM03, Man03, Mat00, MJCM03, NCÁHCLP03a, NAS04, OK03c, OLT04, PGX00, Pen03, Pen04, PR04b, PP01, Ras04f, RKS04b, RC03, Sad02, Sad03, Sad04a, Sad04b, Şah04c, Sak03a, Sak03e, Sal03b, Sev02, SR02a, SR02b, SR04a, SB04a, SW03, SW04a, SW04b, SYK03, SY03, SW02, SL03b, SZ03b, SL04b, TK04, Thu04c, Tsi01, TS02, Tun04a, Tun04b, Tun04c, VR03a, VR04b, VR02a, VR02b, VR03b, Vul00, Wan01, WBW01, WG04a, WL04b, Wan04e, WZ04a, Waz01b, Waz01g, Waz01h, Waz02b, Waz02c, Waz02f, Waz02h, Waz03b, Waz03f, Waz03g, Waz04d, Waz04i, Wei04, Wu03, lWCH03, lWyCjT03, Wu04b, YY04a, YS00a, YL04c, YL04a, Yan02e, Yan02a, Yan02g, Yan03h, Yan03g, Yan03j, Yan03k, Yan03a, Yan03q, Yan03n, Yan03v, Yan04g, YWY03, YY04b, ZKSD02, ZW03c, ZL04b, dlS03, Tra00]. order- [TK04]. ordered [MMP03]. ordering [Liu03c, Liu03d]. orderings [Olu03b]. orders [EME03, EMR03, ES00, ESI01, ESES04, Liu03e, Rid04]. Ordinary [KH01, BH02, BBI04, CJB02, CAH02, ÇAB03, DGR04, EKE03, ESSA00, FNO04, HZ04a, He00, qJkSlS03, KP02b, LCH02, LSY03, Li03c, LTS02, Sad02, Sal00, SR02a, SR02b, SR04a, VR03a, VR04b, VR02a, VR02b, VR03b, Waz02f, Yan03a]. ordinate [wYjSjS03]. organized [Pla03]. orientation [EG04f, HB04, JK03a]. orifice [BEAB04]. origin [BDGG04]. Orlicz [Far03, GEA04]. ortho [ZHZ04]. ortho-symmetric [ZHZ04]. Orthogonal [Fuy04, Pom01, AM02b, AGM02, CC04c, DGR04, Ehr02, HR02, IKS02, KS02b, MM02a, RR02b, YDW02, dBGM02]. Orthogonality [SC03d, IKS02]. orthomin [CC01]. orthotropic [AAAaZ99b, AAAaZ99a, AAA03, AAHAD04b, AaAaZ00b, AaAaZ00a]. oscillating [Tsi03]. Oscillation [BDD00, DS03, ESS00a, FS04b, JL03a, JL03b, JL03c, Jia03a, LS03a, Lin04b, Luo02b, Luo02a, Men02, MY03, Pen03, Sak03b, Sak03c, Sak03d, Sak03e, SC04, Sak04, SW04a, WG00, Wan01, WG04a, WL04b, Wan04e, lWCH03, lWyCjT03, XX03, XJM04, Yan03r, Yan04h, Yan04g, CLD03, EOM03a, FZ04b, HL03, KR01, KSM03, LC01b, Li03c, LM03, Li03b, Li04c, Li04d, Li04e, LHM04, Lin04a, Lin04c, Mař04b, PGX00, Şah04c, Sak03a, Sun04c, SM04d, Yan03j, Yan03a, Yan03s, ZL04b]. oscillations [EB03a, Min03, Min04, Yan03d, Yan03i, Yan03m]. oscillator [BDGG04, EL04b, Mar04c, MR03]. oscillator-equations [Mar04c]. oscillators [EL04a, He04d, HMM03, vBU02]. Oscillatory [Abd03a, DFF04, Jia02, LQL03, Ram03c, SW03, CC04b, CC00b, GLM04, GXL04, TS02, Yan03a, Zak03a]. Ossicini [Ric03]. Ostrowski [Yan03t]. other [LS04d]. outer [CWS02, WZ04e]. output [JK03a, JSdN04, PJPL04]. outputs [HV08, JK04a, JVFM04, JAK04b, JAK04a, JLS^{+}04b, JK04b]. overcoming [CHC04]. overhead [AS04c]. overlapped [CC03e]. overlapping [qLzWcC03]. overlay [WZ02]. overlays [WZ01b]. overrelaxation [Tia03]. overspecified [Deh02]. own [Ver04]. oxygen [Çat03].

P [dCD06, WA01, Yan04d, Yan04f, dCD03]. packet [Gla04]. Padé [BN02, AÇB03, ÇB03a, ÇB03b, ÇAB03, ÇB04b, ÇAB04, Çel04a, IR04, Ism04, Kho04, Liu04i, ŞÇ03b, Thu03, Thu04b, Ven03]. Page [Ano02k]. pair [SZ03a]. pairing [Lee04a]. pairs [AGM02]. pairwise [MAE04a]. panel [JAK04b, OT04]. pantograph [LL04c, XL04b, XL04c]. paper [Pop04]. Parabolic [CY01c, AKKN04, AS04h, BD03, Cao01, CY01a, CY01b, CT03a, DC03, Deh02, Deh03a, Deh03b, Deh03c, Deh03f, Deh03e, Deh04a, Deh04d, Deh04e, Deh04g, FS04b, GH03, GR04, Hau04, IES04b, Kha03a, LMS02, LC01b, MY03, Min03, Ona02, Rui03, Sak03d, SYK03, Waz01b, Waz02c, Wün03, XZ03, YG04, wYjSlZ01, ZLS04]. Parallel [Deh03g, Esc03a, MS04b, NG02b, Pov02, Wan00c, CCY03, CYC03, Cic01, CPTZ04b, GH03, Gar01b, KSS02, KS03a, KLK02, KNJ04, KS01, Kwo03, Lew01, Rao04b, Sar02, SEB03, SARAEG04, WWW00, WW02a, Yan00a, Yan02a, wYjSlZ01, Yua03, Zan00, Zha02b, Zhu04c, Zhu04b]. parallel-series [Cic01]. parallelepiped [ALO04, KHL02]. Parameter [AB03a, SEG03, Sar04a, Sar04b, YAOY03, AAH04, AB04a, ASAB03, ASK04, AL00, Bad01, BF04, CKNU00, Deh03a, EME03, HP02b, JWL03, LL04d, NS03, NS04, PS00, SM02a, SR04a, SD01, Taw00, VR03a, VR03b, ZI04]. Parameters [ST03a, WFC04, ADARAM04, ADASM04, Bel01, EG04e, EG04a, EG04d, EG04b, GG02b, HM00, KIY00, OAB04, WHT04]. Parametric [EL04b, Kha04, EA04a, Juk04, RT03]. Parametrically [EB02a, BDGG04]. Pareto [AAMAE03, Sla04]. Parity [SYK02]. part [BN02, ÁR02, AR04c, Kay04h, Liu03b]. Partial [Pet03, AS04h, AV04b, BJD^{+}03, Deh04b, Fad04, GHL00, GM03b, HOL02, HLO02, JCL01, KP03, Keş03a, Keş04d, KLK02, KKM04, LSY03, Li03b, LTS02, Lin04a, Lin04b, Lin04c, Luo02b, Mař04b, Meh00b, TE02b, Waz00a, Waz01b, Waz02c, Waz03b, Waz03k, ZZN04]. partially [Fan03, HLL03, Oqu03]. particle [AL00, BK02, ER02, Pla03]. particles [IN04]. particulate [KE02]. partitioned [WZ04b]. Partitioning [ST04, IbS01, MR00a, bSI01]. party [LLL04, Sta00, Tse03]. Pascal [CEM04, EM03b, EM03c, EM03g]. past [SR04b, Zak04a]. patch [CWY04, WFT03]. patching [RC04c]. path [ZL04a]. Pathology [Ada01]. paths [AAH04]. patterns [AGES04, AM00, Liu04f, Ram03d, TCL02, Waz02e, Waz02h, Waz03a]. pavement [WZ01b]. PCR [WW02a]. PD [CO00]. PDE [BK03, HA04, IES04b, MS04b, SS04b, Şim04a, SS04c, XX03]. PDEs [BP01]. peaks [SZ03a]. penalty [Gon04, GL00, yN04a]. penalty/least [GL00]. penalty/least-squares [GL00]. pendulum [CL04b, EESESS03]. Penrose [Rak04, Tia04b, Wan04a, Wei00b, WW01, Wei01, WW03a, Wei03b]. Percentage [NG04]. perfectly [Zak04c]. Performance [DM04, KNJ04, SSPA01, ZZ04a, Abu00, BE02, BK02, JAK04b, Liu03e, SN01, TCS01]. perimeter [BRVI00]. period [KY03, Tan03, Tar03]. Periodic [CLW02, DH04, EEB03b, HNA04a, HL04e, HL04c, HL04d, Jia03b, LG03, LGZ04, LG04c, PP01, SW04b, XCD04a, XCD04c, AAHAD04a, BH01, Bar03, BRS04, CC02, CC03d, CC03c, CHC03, Che04d, CWY04, CLC04, Çin04a, CFS04, De 01, Dur04, EOEA00, EOM02, EOM03b, EOM03d, EOM08, ES03a, FL03, Fen03, GB03, HC02a, HC02b, HZ04b, HF03, HL04b, HL04f, JGW02, Lan04b, Li04g, LZ04e, LZ04d, LwWW03, LL04d, LG04b, LG04a, LLW04, MM02b, NO03, Pen04, RCC04, Sak03c, TZ04, Ten02, WL03, XCCC04, Yan01c, Yan02d, Yan02c, Yan03g, YC03, Yan03w, Yan04e, YC04, YL04d, YTM00, YH04, ZZ02b, Zha04c, ZC04a, d’O03b]. Periodically [Ram01a, DF02]. periodicity [GM00, XCD04e, ZHD04]. periodogram [Gha01]. Peristaltic [ESH02, EG04g, Mek04, AEE04, AG03, EEE02]. Permanence [LC02b, ZJ04]. permeability [Has03b, HSE03, HSE04]. permeable [HEM04, Hlo04]. permission [Ino02b, YIN00a, YIN00b, YIN00c].

Persistence [XC02b, XC02c, XCD04d, XCD04e, XCD04f, AT04, SC03c]. personal [Cas00]. perspective [Sah04a]. perspectives [ZG02]. Perturbation [CLX02, gWxZ04, WC01, Wei02c, ZW03a, ZW03b, ZClC03, CGVC04, Din03a, EM03h, GGRS03, HI04a, He03a, He03d, He04a, He04b, He04d, Kam02a, Kol02, LW03b, RC04a, Wan04c, Wei00a, WW03a, WWL04, XC02d, XC04]. Perturbations [AAAE03a, Dia04, NS03]. perturbative [GV04]. Perturbed [Din03b, Kho04, Aln04, AV04a, Bog04, DD02, DK04, Dem02a, Dem03c, DW02, Dur04, Jay03, JCZ02, KP02a, KP02b, KP03, KS04a, LW04b, LM04b, LK04, NR02a, NR02b, RR02a, RC03, RC04b, RC04c, Sak03e, SR02a, SR02b, SS04b, SS04c, Sol02b, Sol04, VR03a, VR04b, VR02a, VR02b, VR03b, Xen03]. Perturbing [McR01, Has00, HS00, Sol02a]. pervaded [RH04]. Petri [NV01, NK01]. Petrila [Pop04]. Petrov [HB00, TM03]. Petryshyn [AO01a]. Petryshyn-Leggett-Williams [AO01a]. Petviashvili [sLqZ03a, sLqZ03b, Waz01c]. phase [Ald04, CM04b, EM04a, HIN04, KST00, KE04b, LW04a, MAR04a, Nar02, San00, Tsi01, TS02, VCD04]. phase-lag [LW04a, Tsi01, TS02]. philosophy [Ver04]. Photon [KNU00, BMMRS04]. photovoltaic [ALDP01]. physical [ESG03, Gia03, Kay04c, Waz04g, Zay02b, Zay02a, Zay03a]. physics [Esc02, Esc03b, Sev02, Zay03c, SDR03]. phytoplankton [MSO00]. piece [Zhu04a]. piecewise [Asa04a, GLR02, ZAH02, Zay03a, Zay03d]. piecewise-linearized [GLR02]. piezoceramic [FRRSCS02]. piezoelectric [OT04]. piezothermoelastic [DWC04]. pipe [Lee04c]. pipes [AESS00b]. PKP [sLqZ03a]. places [BL01]. Planar [BA04, APS04, CG03, Din04b, MV04, Ram03b, Ram03c, Ram04c, Tin01]. plane [AM00, BGT04, DWC04, Hil04, Kwo01, Nar02, SL04a, SP03b]. planetary [KST00]. plankton [GR03]. planning [Mak04b, TLLKB03, dOVS03]. plants [AQD04]. plasma [BG04]. plate [Abd02c, CL03a, ENEA04, Elb01, EB02b, Elb03, FRRSCS02, He03b, KNJ04, Kuo04, Oqu03, RWC03, TS01, Xu02, Xu04, YX04, Zak04a]. plateaus [SZ03a]. plates [CCY03, EEE03, IH01, dJ04]. players [AE04]. plus [NCÁHCLP04]. ply [OT04]. Poincaré [EL04a, HMM03, Mog04]. Poincaré-map [HMM03]. Point [WL04d, AAH04, AO01a, AZ04, zBxF03, CT03a, DSC01, DG03, Din04a, EG04f, ES02a, GA04, HZ04a, HY04, HM04, JF04, KP02a, KA04, KR03c, KR03b, Kha04, Kok03, Kor03, Kum02, Kum03a, Kum03b, Kum03c, Kum03d, Kum03e, Kum09, LL04a, LY02a, Liu02a, LY02b, Liu03a, Liu03b, Liu04a, Liu04c, LQG04, ML04b, Ma04a, Moh03, MSJ04, NR02b, ÖA04, PPS04, RR02a, RC03, RC04b, RC04c, SB04a, SV02, Sil02, Sil03, Sil04, WSX03, Yan03v, YWY03, ZW03c]. Points [SE04, BMR03, Gin04, Kan04, LMG00, LW00, NG04, Wol04]. Pointwise [Ehr02]. Poisson [AB02b, GA04, Meh03, SKM04]. Pol [EL04b, LwWW03]. polar [Hel01]. polarization [KST00, MTK^{+}00]. POLDER [KIY00]. poles [Abd02c]. policy [CP04, MADT03, MAR04a, RAH01, Tad03, VCV01]. Political [BF03]. polluted [DLYC04]. Polygamma [RS02b]. polygon [MV04]. polyharmonic [DX04, XYD02, XD03]. polyhedral [GLVW00]. polymer [ALO03b]. polymers [AESS00a]. Polynomial [CMM02, AD04a, AS03d, BÇ02, Boy03a, BW04, EMED03, Fuy04, HM04, Has03a, Ji02, KV03, Keş03a, Keş03b, Keş04d, KS02a, KM03a, KS02b, MM03, NCÁHCLP03a, PR04b, ŞBY04, SC03d, Sta03, ST04, Yal02, gZC04, Zhu04b]. polynomially [KG03]. polynomials [Abd03b, AM02b, ÁND02, AGM02, BJ04, BŞY03, BBK^{+}03, Bİ04b, CS03b, DSC01, DRS03, DC03, DCS03, Dat03, DGR04, Dat04, EMED03, ES00, GMGC01, HIS04, HR02, HZ04c, IKS02, LOS00, LCS03, MM02a, ÖKOD04, PSS00, Pom01, RS02a, RR02b, UK03, Waz00b, Wün03, YS00a, dBGM02]. polytropic [El 03]. Population [JD03, ADAAM03, AD00, Ahm04b, DLYC04, Gho03, HA02, HM00, JK04c, Li04c, LG04b, LLW04, Mut03, PR04a, QMWAZK04, Sah04b, SYK02, WW04, XC01, Yan01a, d’O03b]. Population-level [JD03]. populations [HA02, MR04, NG04, ÖM02, PS00]. poroelastic [AAA02]. porosity [LPS02]. porous [AESS00b, AG03, AO04b, AH02, AQBT04, CX04, EA04d, ENEA04, ES03a, ESH02, Elb01, EB02b, Elb03, EB04f, EEE03, Has03b, HSE03, HEM03, HEM04, HSE04, HNA04b, LWZ00, LPS02, (Mu04, NM03, Qui04, SR04b, Waz01e, Xu04, Yan02h, Zak03b, Zak04b]. portfolio [ZN04]. posed [BGT04, WN04]. positioning [CC03f, FT00, Far00, Lun01]. positioning/inertial [FT00, Far00]. positions [EGE02]. Positive [CCP03, CWY04, HL04f, LZ04e, LZ04d, Liu02a, LL04d, Liu04a, Liu04c, Pen04, SL04b, XYD02, XD03, Yan02f, YC04, YZ04, YL04d, Afr04, zBxF03, zBxF04, BJD^{+}03, BRS04, Che04d, CLC04, Çin04b, Çin04e, Çin04d, Çin04f, DX04, DNS03, ESEA02, ESAD04, FL03, Guo03b, Guo04, HP02b, Lan04a, Lan04b, LL04a, LC04a, Li04b, LR04, LNS04, LS04b, LS04c, LWL00, Liu03g, LQG04, LOZ02, LYT04, LG04b, Ma04a, NS03, OLT04, PS04, Sak03d, SL03a, SL03b, SL04d, WL03, Waz03c, WJ04, YGL01, Yan03k, YC03, Yan04i, YWY03, Yao02, Yao03a, Yao03b, Zay03d, ZKSD02, ZY04, Zha04a]. posterior [ASM03]. posteriori [AAAE03b, AO00c, Liu00, OIO03]. Potential [DK04, Abd00, Abd02a, Cin04h, Dob00, Has04a, sLqZ03a, sLqZ03b, MA04b, Sub02a, Yan03a, de 00]. potentials [CD03b, ÇD04, Kin01]. power [AR04a, ADARAM04, KM03a, MK02, OASM04, Yoo04]. practical [JL04, LZ01a, LZ01b]. precedence [MJCM03]. preconditional [Yin03]. Preconditioned [QS04, Yua00, Zha00, Zha04b, GH04]. preconditioner [Che01b, Zha02b]. preconditioners [CW02, CMM02, qJkSlS03]. Preconditioning [KS02a, WZ03, WZ04d, Zha01, Zha02b]. predation [EOM04b, JB02]. predator [CWY04, EGAR03, EGB03b, HL04b, HL04e, HL04c, JB02, KL03, KSC02, LHL04, WL03, XC02a, XC02b, XC02c, XCD04b, XCD04d, XCD04e, XCD04a, XCD04c, XCD04f, ZJ04]. predator-prey [XCD04f]. predicting [Gon04]. prediction [BBN04, NK01, WZ02]. Predictor [Din04e, NNAK04, GG00]. predictor-corrector [GG00]. Preface [AM02a, SDR03]. preferences [BHH02]. Preissman [WWQ04]. presence [ADARAM04, CDH01, Coo04, Ezz04, (Mu04, NO03, Zak03a]. presentation [AD01, ŞBY04]. preserving [LCH02]. pressure [CY02, EE03, EEES03, EE04a, EE04b, ESEE02, FRRSCS02, Rao02, SD02b]. prey [CWY04, EGAR03, EGB03b, EOM04b, HL04b, HL04e, HL04c, JB02, KL03, KSC02, LHL04, WL03, XC02a, XC02b, XC02c, XCD04b, XCD04d, XCD04e, XCD04a, XCD04c, XCD04f, ZJ04, ZC04b]. price [BS00]. pricing [DDX02, ML03]. primal [AAAE03b]. primary [EESESS03]. primes [HLLL03]. primitive [YX04]. principal [AB04a, Rol02]. principle [He04g, TY04, Ver03b]. principles [BCI03, Ver03c]. priori [GV00]. private [HLLL03]. Probabilistic [Dua00, GK02, Gzy01, NU00b, SJM04]. Probability [WBW01, ASKT03, Oht04, SD01, Tag00a, Tag00b, Tag01a, Tag01b, Tag01d, Tag01e, Tag02a, Tag02b]. Probability-one [WBW01]. Problem [Zak04b, AEL04, AGES04, Abd01a, Abd01b, Abd02b, AB02a, Abd02d, AM03, AS04a, AS02b, AhL00, AZ02, ALO03b, AÖE04, AO01b, AO04c, AEK04, BB02c, Bae04b, Bah04, BH01, Bao02, BD03, BGVHN02, BEAB04, CHL01, Çat03, Çat04, ÇB04b, ÇB04d, CLL00, CT03a, CL04b, CW03, DDX02, Deh00, Deh03f, Deh03g, Deh04c, DH04, Dur04, DNS03, EL04a, EMY04, ERN03, ESK04b, ES03b, EK04, Fat04, FQ04, FS04b, GGB03, Gia03, Guo03a, GM03b, GM04b, Gzy01, HS00, Has04b, HAS04f, He03b, Hlo04, HP02b, HHT04, IOAB01, IPPT03, IES04b, Jay03, JW03, JY00, Kam02a, KML04, Khu04, KL00b, KHL02, KW04a, KB04b, Kor03, Kru03, KE04a, KE04b, KW04b, Lee00, LOS02, Lee04c, Li04f, Liu02a, LY02b, Liu03a, Liu03b, gLgW04, LC04c, Lon00, LLW04, ML04b, Man03, Mel00, MT04, Moh03, MD00, Moh00]. problem [MV04, Mun03, MLA01, NHCLPSR04, NS04, OSL03, Opp00, OK03b, PPS04, Pen04, Qui04, Ram03a, RKS04a, RKS04b, RWC03, San00, SM02a, Sev02, Sev04, SL04a, ŞÇ03b, Şim04b, SD02b, SD02a, ST03b, Sub02a, SY03, SW01, SL04d, TE00, TE02a, TE02b, Tag00a, Tag01a, Tag01c, TM04b, VCD04, WZ04c, WC01, WW03a, WW00a, XFL04, Yan00b, Yan02a, Yan03f, Yan03i, Yan03h, Yan03n, Yan03v, YZ04, YSS00, YS01a, YTM00, YS03, ZZ01, Zay02b, Zay02a, Zay02c, ZAH02, Zay03b, Zay03a, Zay03e, ZI04, Zay04a, ZN04, Zha04b, ZH04, dOVS03]. problems [AAEA03, Abd02c, AS04d, AD02, AM02b, ANS02, AZ04, ASN03, AAMEST04, All04b, ÁR02, AR04c, Asa04c, zBxF03, BMMRS04, BM02, Bog04, CPL00, CH02, CH03, Cao01, CY01c, CL03a, Cha04a, Che01b, CC02, CX03, CC04b, CC03f, CMOS01, CMF04, CLD03, De 02a, De 03, De 01, Deh03c, DMT02, Din02, DWC04, EGBH03, EKE02, ERN01, ES02a, EEK03a, EEYK03, FZ04b, GHW04b, GLM04, Gla04, GL00, GV00, HA03, Has00, Has04c, Has02b, Has04d, HKA04b, He03g, JF04, JCL00, JWL03, JGW02, JBMR02, JK04c, KP02a, KP03, KA04, Kad04, KR03a, KR03c, KR03b, KR04, KS03a, Kay04e, KA03c, Kha04, KL00a, KH02, KA00, KNJ04, Kum02, Kum03a, Kum03b, Kum03c, Kum03d, Kum03e, Kum09, Kuo04, LM04a, Lan04a, Lan04b, LCH02, LL04a, Les01, LW03a, Lia04, LC03b, LY02a, Liu03f, Liu03g, Liu04b]. problems [Liu04a, Liu04c, LQG04, MM04a, Ma04a, Mae01, MLG00, MdL04, MSJ04, MSO00, Mun03, NK03, NR02a, NR02b, NU00b, Noo04c, NAS04, Ode02, OOAA04, QW03, RdL01, RO01, RR02a, RC03, RC04a, RC04b, RC04c, Rui03, Sha03, SR02b, SR04a, STHN02, SV04, SP03b, SL03b, Sun04b, TCS01, TLX04, Tro04, TS02, Tsi03, TS01, VR03a, VR04b, VR02b, VR03b, VK04, Vul00, Wan00a, Wan04c, WFC04, Wan04a, Waz01g, Waz01h, Waz02f, Wei04, WN04, WJ04, WC02, Wu04a, Xen03, YG04, YL04a, Yan00a, Yan02f, Yan03l, Yan03u, Yan03w, Yan04a, YD04, YWY03, Yao02, Yao03a, Yao03b, You04, Yu04, Zay03c, Zay04b, ZWRL03, ZY00, ZL04a, ZHZ04]. procedure [Dje00, GG02a, JSD04]. procedures [Deh04c, GR04]. Proceedings [SDR03]. process [GA04]. processes [AB02b, ESG03, EP00, Gha01, Gha03, Jol00b, JCZ02, Oht04, TB04]. processing [BEV02]. processor [Pov02]. Product [HHJ01, DMT02, MLA01, SM02a, SK04]. production [QWY02, QW03]. productive [JK03a]. products [CZT04, CT04a, RR02b, Şim04a, Tia04b, Tri04, YAYA03]. Professor [Cas00]. profile [LM02]. profiles [Mur03]. program [ŞBY04]. programmed [LRMSVO01]. programming [AS04d, AD01, Cha04a, ES03b, FK03, FQ04, Isk04, JLST04, Lee01, Lew01, Liu01, Mae01, MJCM03, MT04, MD00, NN04a, yN04a, Pap00, SS04a, SZ04, TCS01, You04, YE04, ZWRL03, ZZ04c, dOVS03]. programs [DG03, Iwa01, LSJR00, yN04c]. progress [BK03, JAK04b]. project [MB04]. projected [Zhu04a]. Projection [Ras02b, Bao02, NWX03, NN04b, ZW04a]. projections [DW04, MMER03]. projective [Ras02c]. projectors [CT03b]. prolate [Boy03b]. Propagation [AR03, AS03b, Dem02b, Sei03, Sel03, enNAaA02, AAA02, BV02, ENAAM01, ENAAF04]. Properties [LL04c, ZL04a, AAMAE03, AEK04, BEV02, BG04, BRS04, Bİ04b, CD03b, Cin04h, Coo04, DMS01, DMT02, ESEA02, JBMR02, Lac03, Li03b, LC04a, Liu03h, LS04e, MTK^{+}00, NU00b, NO03, OK03a, Tan03, TB04, Zed03, vAG03]. property [ERN03, Koz03, RS04a]. prospection [Sev04]. protein [Dua00]. protocol [CJ04b, HWWM03]. protocols [LLL04, Tse03]. prototype [TCL02]. prototyping [NV01]. prototyping-based [NV01]. prox [Mou04]. prox-regular [Mou04]. proximity [Tag03a, Tag03c]. proxy [CCH04, HWW03, HW04, HC04b]. pseudo [DRS03, FK03, Pom01, YG04, YÇAM04]. pseudo-Euclidean [YÇAM04]. pseudo-hypersurfaces [YÇAM04]. pseudo-inverses [FK03]. pseudo-orthogonal [Pom01]. pseudo-parabolic [YG04]. pseudo-polynomials [DRS03]. pseudoinverses [CW04a, ST03b]. Pseudomonotone [Noo03c]. pseudospectral [BH03, GH04]. public [HW04, HL04g, Sha04b, TJC03, WL04c]. publications [KN00]. published [Pop04]. pull [ADG03]. pulsate [SS02]. Pulsatile [ES03a, AG03]. Pulse [Sab18, d’O04, MK03b, MM04b]. pulsed [LC04c]. pursuit [IOAB01]. push [ADG03]. push-to-pull [ADG03]. pyroelectric [DWC04].

quadrant [Liu03e]. quadrant-recursive [Liu03e]. Quadratic [NCÁHCLP04, ZG03, ABEM04, CG03, Dje00, EBE03, LC03a, gLgW04, MD00, TCS01, WW00b, WSX03]. quadrature [AhL00, APS04, EM03d, FS04a, Özd03b, Özd03a, ÖK03d, SJM04]. quadrature-based [AhL00]. quadrilateral [KL00b, KLK02]. Qualitative [DMS01, DMT02, Has00, HS00, SL04c, GM03a, ZC04b]. quality [SH01]. quantitative [TE02b]. quantities [HY04]. quantity [AA04a, De 02b]. quantum [Esc03c, YCC02]. quarter [BGT04]. Quartic [NAS04, EBE03, Sal00]. quartile [Mut03]. Quasi [De 03, Ino04a, BAB01, Che03, Din00, Din01, Din04d, Kap04, Li04d, LW00, LYT04, MM04a, MM04c, Moh03, Noo03a, Noo03b, Noo04b, NNAK04, SAV04b, SC03d, Tro04, ZY00, ZL04a]. quasi- [ZY00]. Quasi-clusterability [Ino04a]. quasi-elliptic [Kap04]. quasi-exact [BAB01]. quasi-fixed [LW00]. quasi-linear [Li04d, LYT04, MM04a, Moh03]. quasi-linearization [MM04c]. quasi-monomiality [Che03, SC03d]. quasi-monotone [ZL04a]. quasi-Monte [SAV04b]. quasi-solutions [Tro04]. Quasi-steady [De 03]. quasi-variational [Din00, Din04d]. quasi-variational-like [Din01, Noo04b]. quasilinear [Aas03b, Gül04b, Jia02, LOS02, Lin04a, Lin04c, Yan04i, Yan04e, Yan04h]. quasilinearization [ANS02, BH01, KAVM00]. quasivariational [sC03a, Din03b, LK04, LDKU04]. quaternionic [ÇT04b]. quaternions [EGE02, EGE03]. quenching [CY01a, CY01b]. queue [ASAI03, AS04g, CP04, MADT03, MANAM04, MAR04a, Mor04, Tar02]. queueing [CM04b, Tad01, Tad03, Tar03]. queues [AS04g]. Queuing [AS04f]. quintic [KA03c]. quorum [Tad03].

R [XYD02]. rabies [AFRH02]. Rabinowitsch [He04g]. radial [HS01, JHS^{+}04, LNW03, WHG02, Yoo04]. radiating [Rad04, RM02]. radiation [EE03, EEES03, EE04a, EE04b, ENEA04, ESEE02, ED02, KEB03, RPT04, TLLKB03, Zed02]. radiative [NU00b, Wan00a]. radii [NS03]. radius [LCH00]. Ramanujan [KL02]. ramp [KW04b]. Random [Kan04, ADASM04, AL03, Bao02, Bel01, EB04a, KC04, Tad01, Tan03]. range [EA04b, EB04b]. ranges [Hel01]. Rank [CT04a, LW04e, LW04d, TW03, Tia03, Tia04b, ZTD04, ZB04a, ZB04b]. ranked [ADASM04, AS04e, HM00, Mut03, MAD04, SM02b]. Ranking [JLS^{+}04c, JSdN04, JSLS04]. ranks [Tia04a]. Rao [Par02]. Raphson [Abb03]. Rapid [NV01, Yan03i]. rapidly [CC04b]. raster [BR01]. rate [ASAI03, CCY03, CYC03, EB04b, EG04e, EG04d, Gra03, Sar04a, Sar04b, Xen03, YAYA03]. rates [CCS04a, d’O03b]. Ratio [KC04, AAMK01, CWY04, KK04, XC02a, XC02b, XCD04d, XCD04e]. ratio-dependent [CWY04, XC02a, XC02b, XCD04d, XCD04e]. Rational [PR04a, BGVHN02, EEYK03, Pap00, Sta03, ST04, YL03a, YLEM04]. rationality [BF03, YA03]. rationalized [MMA04, RO01]. ratios [BRVI00, WZ04c]. Rayleigh [AAA03, AAHAD04b, enNAaAd04, Dem03a]. Razumikhin [LS02]. re [MANAM04, RS01a]. re-entrant [RS01a]. re-service [MANAM04]. Reachability [CFS04]. reaction [AGES04, Abd04, DXL02, Hel01, Jay03, LM04b, LPS02, RS01a, SR04a, SK02, TZ04, Thu04c, VR02a, Waz00a, ZL03]. reaction-diffusion [Hel01, RS01a, SK02, Waz00a]. reaction-mineralogy-porosity [LPS02]. reactions [Bao02, GGRS03, SV02, YAYA03]. reactive [Ram04b, Ram04c]. reactor [Dun02, MMR01]. Real [Tag03d, Ana01, CG03, Kır04, KÖ04, MM02b, Pap00, Tag01d, Tag02b, Tag03b, YL02]. real-time [YL02]. realistically [d’O04]. Realization [IYM00]. reasoning [HV01, TKK00]. reciprocity [SG01]. recombination [GGRS03]. reconstructing [GV04]. reconstruction [GV00, Gzy02, XL04a, YDW02]. reconstructions [Gzy02]. record [AY03, Jah03]. Recovering [Tag01e]. recovery [AAMEST04, Sha04b, SD02b, TJC03, WLX00]. rectangular [CC02, HA04, KH02, Sei03, WLW03]. Recurrence [Cho02b, KS02b, ÁND02, EM03a, Liu04d]. recurrences [Dik04]. Recurrent [Wei00b, HC03, Mor04]. recursion [HY04]. Recursive [Nak02, CTZ03, CPTZ04a, EA04c, EOREA00, EOAM03, EOAE04, LZ04c, Liu03e, Nak04b, Tan03, VCV01, YL03b, YL03a, YL04b, YCME04, YLEM04]. recycling [Kan00b]. Reduced [LRHRD01, Din03a, WBW01, Wu03, Wu04b, ZClC03]. Reduced-order [LRHRD01, WBW01]. Reducing [BH03, Hos04]. reduction [ADG03, Bon02, CCKS02, HA04, Mar04c, Ode02, OIO03, RC03]. reference [EEKS04]. referential [TE03]. Refinement [Yan03t, yWShX02]. reflection [BGWX03, Mat00, Yil04]. reflectionless [Kin01]. reforms [VCV01]. reformulation [Yu04]. regime [HEM03]. region [AH02, ES03b, Moh00, Sun04b, Thu04c, Yu04, ZAH02, Zay04a, Zay04b, Zhu04a]. regions [Jol00a, KST00, MPS04, RS01a]. regress [JAK04b]. regression [AYW04, CBK00, EB04a, EB04b, NN04a, TH04, Wan02]. regula [WSX03]. regular [Cic01, FRRSCS02, MV04, Mou04, WLW03, ZI04]. Regularization [PZZF02, WLX00, WS01, SS04b, SS04c, YSS00]. Regularization-based [WLX00]. regularized [Kay04d, Wan04a, YYS03]. regularizing [MS01]. regulation [YL04d]. regulators [LRMSVO01]. Reissner [YX04]. Related [SDR03, Bat04, CSA03, DW04, Has04c, KSL02, LOS00, LW04e, SK04, Sri03]. relation [ÁND02, NG02a, YIN00a]. Relations [ENR04, Sav04a]. relationships [Abd01b, Abd02d, AS04b, LCS03, YB00b]. relative [MR04, Sah04b, WC02]. relative-motion [WC02]. relativistic [EEES03, EE04b, ESEE02]. relativity [Khu03b]. relaxation [enNAaAd04, AEK04, CO00, EKE04, EEK03b, EEKS04, Jia03b, Sam04, SZ04, Yua03, Zak03a, Zak03b, Zak04b]. relaxed [YÇAM04]. releases [CY01c]. reliabilities [Sar03b]. Reliability [Sar02, SARAEG04, Ami01, Aru03, Cic01, EG04a, EG04b, Gra03, Koł01, Kur02, KS01, Mil03, Sar03a, SEB03]. reliable [DR01, Kay04e, LC03a, Waz01h, Waz02g]. REM [BT00]. remark [KS01]. Remarks [Der03]. reminiscences [Cas00]. reneging [AS04f]. renewal [AY04a, EP00]. repair [Ana03, AG04]. repairable [ST03a, SEG03]. Repairing [HL04g]. Representation [Che01c, ALDP01, CF03, RW03, WW01, WD03, Wei03a, WQ03, WW03c, Wei03b, WZ04e]. representations [Bat04, JKP03, LS04d, WC04c]. representing [AEC03]. reproducing [lLgC03]. require [ABP04]. Reroute [Mak04b]. research [Tie03]. reservoirs [Meh00a, Meh01]. residence [SD01]. residual [AY04b, Bon02, CL03a, EBA03, SZ03a]. residuals [LC03b]. resighting [HA02]. resistant [WL04c]. Resolution [Meh00b, BHH02, YL02]. resolvent [Abd00, pFjH03, LDKU04]. resonance [EESESS03, EBE03, EEB03b, KB04a, LY02b, Liu03a, Liu03b, Yan03m]. resonances [ML00]. resonantly [EB03a]. resonator [DNS03]. resource [FSLMC03]. resources [MB04, RAH01, Sla04]. respect [Bil04, Bil07]. respiration [BT00]. response [BS04, EOM04a, Nak02, Ram04a, XC02c]. Restarted [BJ03, BJS04]. restarting [NK04]. restricted [JF04, rWS04, WLW03]. restrictions [BI04a, JSdN04]. Restrictive [EEYK03, IES04a, IES04b, IR04, Ism04]. result [Tun04a, Yan01c, Yan03g]. resulting [Lun00, WW00a]. results [CH02, CH03, Che03, Che04d, CO00, De 04, GAZK03, Men02, Sad03, Tun04b, Yan02c, Yan02g, Yan03i, Zay02b, Zha04a]. retarded [AÇB03, ÇAB03, ÇAB04]. retraction [Hab04b]. retracts [MAE04a]. retrial [Mor04]. Retrieval [LCH00, MTK^{+}00]. return [JMV04b]. returns [JSD04]. Reverse [CZT04, CHL01, SW02, WZ04a]. reverse-order [SW02]. reversible [CGG01, Yan03c, Yan04b, Yan04c]. review [HP02b]. Revised [HH01]. Riccati [ETBAN04, LCXZ04, LF04]. ridge [AYW04, CQL01, MC00, Wan02]. ridge-type [CQL01, MC00]. Riemann [Mus00a, Mus00b, Mus00d, RS02b]. Riesz [ÇD04, Cin04h, Kan00a]. right [Bry02]. rigid [EGE02, EG02, EGE03, EG03a, EG03b, EG04c, EGT04, EG04f]. ring [gZC04]. rings [Ras02b, vBU02]. risk [Hür04]. RKMK [SMQ04]. RLW [DSI04]. Robin [NR02a, Ram03d, Zay02c, ZAH02, Zay03a, Zay03d, Zay04b]. robot [CS00, HkT03, HT04]. robotic [Jam01]. Robust [AQD04, Che04b, HGS04, LL03, Par03, Par04b, WL04c, CW04b, Lie04, PKW04, PKLW04, Shi02]. Robustness [Ram04b]. Rodrigues [RES04b]. rods [EEB03b]. role [ZG02]. rolling [EEB03a]. root [CL04a, Coo04, SV02]. root-finding [SV02]. Rosenau [LM02]. Rosenbrock [CCK04]. rotating [AR04d, Rad04]. rotation [AAAaZ99a, AaAaZ00a, Meh00b, YÇA04]. rotational [EG02, EGE03, EGT04]. rotations [Ana01]. rotors [EG02, EG03b]. rough [Hil04]. roughness [Rao02]. round [ML04d]. Rounding [Cao03b]. routinely [Yan01a]. RSA [HLLL03, HLL03]. RSA-based [HLL03]. rubella [JDV04]. rule [SM00b, rWS04, WQ04, Wei02a]. Ruled [AAABH04, KAK05, KAK06, Kas06, AB04a, Küç04]. rules [Özd03b, Özd03a, ÖK03d]. rulings [AAABH04, KAK05, KAK06, Kas06]. run [Kat01]. Runge [ÁR02, AR04c, EME03, EMR03, HT00, IbS01, KS03a, MS04c, OOAA04, RP00, Wu03, Wu04b, XL04c, bSI01]. rupture [DYH04].

saddle [Din04a]. safeguard [Shi02]. SAGE [LCH00]. Saharan [SP03a]. Saka [Pan08]. salesman [Lon00]. sample [ÇG04, Coo04]. samples [Mut03]. Sampling [XL04a, ADASM04, Ahm04b, AS04e, HM00, KC04, KM02, MPS04, MAD04, Yan04a]. sapling [SM02b]. satellite [CCC01, CQL01, EE03, EE04a, EE04b, Lun00, RCC04, TM00]. satellites [HIS04]. Satsuma [HF03, Kay04f]. saturated [Qui04]. Sawada [ESK04a]. scalar [ÁR02, De 02b]. scale [AAEA03, CC04b, JK03a, JSD04, JMV04b, Lay02, MIM00, PJPL04, Par04b]. scaled [Hür04]. scales [ML04b, Sak04, SL04b]. scarce [Tro04]. scattering [AA04d, Hil04, KEB03, Kru03, Kwo01, ML00]. Schauder [BFGG04]. schedule [CM04b]. schedules [MADT03]. scheduling [MB04, Pov02]. scheme [Bah03, BJD^{+}03, CHC04, CCH04, GH03, HIN04, HWW03, HHT04, JLM^{+}03, JL04, Kum03b, KE04c, LCH02, LHL03, LL04b, LC04d, ML04a, ML03, Moh04, Pei04, PTG03, Ras04f, SB04b, Thu04c, WLX00, WWQ04, Wan04d, WL04c, YCH04, wYjSjS03, ZC03, ZZN04, ZX04]. schemes [Che01a, CJT03, Deh03d, Deh04d, Dur04, GhW04a, HW04, HC04b, HL04g, OOAA04, Pet03, RK03, WWQ04, XY04, XFL04, wYjSlZ01, Zha00]. scholar [Nat00]. School [SDR03]. Schrödinger [Aas03a, Dem03b, JLM^{+}03, KS00, PGyL03, Ram02, Sub02a, YS01a, ZC03, ZaYD04]. Schryer [BAB01]. Schryer-Walker [BAB01]. Schur [Tia04a]. Schwarz [CX03, Rui03, SKM04, XL03, Yan00a, YD04]. screen [Kwo01]. SDD [MMER03]. SDS [BMR01]. seals [Yüc04]. search [GLWY04, GG02b, JT04, Kah05, Kah06, RS03, Shi04, SYY04, YK03]. seasonal [PdR01]. SeaWiFS [TRC03]. Secant [AB03b, HR01]. Second [NCÁHCLP03a, Abd03a, AN03, AO00b, ABP04, AF03, Asa04b, BBV04a, BDD00, BCI03, CCP03, CCY03, CDH01, DH04, DS03, EM04b, ESS00a, Guo02, HAS04f, Jan04, Jia02, JL03b, JL03c, Jia03a, Keş03a, Keş04d, KEE03b, KL00a, KH02, KLK02, KR01, KM03a, Kum03d, Kum03e, Kum09, LM04a, LL04a, Li03c, LS03a, LW03a, LNS04, Li04d, Li04e, LW04a, LY02a, Liu03f, LQG04, LG03, LG04a, MM04a, MK03b, Mas03, MM04c, Moh03, Özd03b, PGX00, Pen03, RKS04b, Şah04c, Sak03a, Sak03e, SW03, SW04a, SW04b, Sor02, SL03b, SL04b, Thu04c, VR04a, VR03a, VR04b, VR03b, Wan01, Waz02f, lWCH03, lWyCjT03, YY04a, YL04a, Yan02e, Yan02a, Yan02g, Yan03a, Yan04g, YY04b, ZL04b]. Second-order [NCÁHCLP03a, Abd03a, Asa04b, CDH01, DS03, Guo02, Jan04, Jia02, JL03b, Keş03a, KL00a, KH02, KLK02, LM04a, LS03a, LNS04, Li04d, Li04e, LY02a, LQG04, LG04a, PGX00, Pen03, Sak03a, Sak03e, SW03, SW04b, SL03b, SL04b, Thu04c, VR03b, Wan01, Waz02f, lWCH03, lWyCjT03, Yan02e, Yan02a, Yan04g, YY04b]. secret [YCH04]. section [CM04a]. sections [FZ04a, Ras02d]. sector [RAH01]. securities [BS00, CCY04, SMF04]. Sedenions [II00]. sediment [De 03]. seeking [Mus00c]. Segalman [HAS04f]. segmentation [Fuy04]. Seidel [Özb04]. Selecting [AAMK01]. Selection [Bon02, ZN04]. Selective [MMER03]. Self [NN04b, Sol03b, Che01b, HW04, HL04g, Ino04a, Ino04b, Lee04a, Pla03, Rad04, Sal03a, Sal03b, SS02, Sha04b, TE03, TJC03]. Self-adaptive [NN04b]. self-adjoint [Sal03a, Sal03b]. self-certified [HW04, HL04g, Sha04b, TJC03]. self-gravitating [Rad04]. Self-limitation [Sol03b]. self-organized [Pla03]. self-pairing [Lee04a]. self-pulsate [SS02]. self-referential [TE03]. selfadjoint [CX03]. selfgravitational [Rad03]. semi [ÇT04b, De 01, EG04a, EG04b, Gra03, HSE03, Kuo04, Mat01, NM03, yN04a]. semi- [De 01]. semi-Euclidean [ÇT04b]. semi-infinite [HSE03, Kuo04, Mat01, NM03]. semi-Markov [EG04a, EG04b, Gra03]. semi-penalty [yN04a]. semiconductor [GF00, GS02, SS02]. semicontinuity [MAK03, MAE04b]. Semiconvergence [Cao04b]. semidefinite [Yu04]. Semigroups [EB04d, ASK04, EB04c]. semiiterative [WW02b]. semilinear [CY01a, CY01b, PS04, Yan02b, Yan03f, wYjSlZ01, ZH04]. seminormed [AET04]. semiopen [MAK03]. semiretracts [MAK03]. semiseparation [MAE04b]. sense [NHCJLLP04]. Sensitivity [JLM04, MC00, JW03, KB04a]. sentiments [Ino02a]. separability [Kin01]. separated [AR04c]. Separation [MA04b, All04a, HIN04]. sequence [AET04, AB04b, Bil03, Bil04, Bil07, EA04c, EOREA00, EOAE04, EAA04, KŞ04b, Li03a, Mak04b, MS04d, MRŽ04, Thu04a, YL03b, YL03a, YL04b, YCME04, YLEM04]. sequences [De 04, DÖ03a, EOAM03, GÇ04, GEA04, KA03b, KY03, ÖAD03b, Özk03]. Sequential [GHW04b, HM04, BMR00, RT03, WH02, BMR03]. Series [CCS04b, AEC03, Bar04, Bat04, ÇB03a, ÇB03b, Çel04a, CCS03, Cic01, EG04e, EG04d, GA04, Hai00, He01, KM03a, KS01, LCS03, Mil03, ÖAD03a, RES04a, RS02b, RS04a, Sar02, SARAEG04, STHN02, Sri03, Thu04c, WX03b]. series- [Mil03]. series-parallel [Cic01, KS01]. series/parallel [Sar02]. server [AS04g, MADT03, MAR04a, Tad01]. service [CP04, MANAM04, MAR04a]. Set [ESI01, ADASM04, AS04e, EA00, HHK01, HM00, Li03a, Mut03, MAD04, Oht04, SS04a, SM02b, Sla04]. set-chains [HHK01]. Set-valued [ESI01]. sets [GNS02, MAK03, MAE04b, MAK04a, Sla04, SC03d, Ver00]. settled [ÇB04b, ÇB04d]. seven [EME03, ESK04a]. seven-order [ESK04a]. Several [Liu02c, Waz03i, CLD03, Luo02a, NG04, PS00, Sak03d, SL04c]. severe [NR02b]. shadow [Lun00]. Shafer [BN02]. shallow [Rao04b, WHG02]. shape [ASAB03, BRVI00, SP03b, Zay03d]. shaped [Khu03a]. shapes [Meh00a, Meh01]. shared [JLSS04, JAK04b, JAK04a, Kay04h]. sharing [Sor01b, YCH04]. Sharp [Bai00, Ras02c]. sheet [MM04c, Vaj01, VR04a]. sheets [Ram03b, Ram03c]. shift [AD04b, DRS03]. ship [EEB03a]. Shishkin [Xen03]. shock [EMY04, Vul00]. shocks [Rao02, YR01]. Shooting [NR02b, SKM04]. Showalter [SS04c]. Shrödinger [Che01a]. sic [EA00, DYH04]. SiC/SiC [DYH04]. side [CL03a, CL04b, CHL02]. sided [AÇB03]. sign [ALO03a, YWY03]. signal [NHCJLLP04, NHCLPSR04, YDW02]. signature [HWW03, HW04, HLL03, HC04b, HL04g, LH04a, Sha04b, TJC03, Wan04d]. signatures [Fan03]. Signed [Ino04b, Ino02a, Ino04a]. signing [WH02]. signomial [SZ04]. Signorini [HS00]. Similarities [Thu03, Thu04b]. simple [ADASM04, Dje00, Gho03, He03d, Ino02b, KC04, TCS01, WSX03, YIN00b, YIN00c, YCC02]. simplex [ADG03, Tin01, YXC03]. simplified [BK02]. Simulating [HT00]. Simulation [JDV04, RP00, BMR00, BMR01, CW03, HkT03, JLM^{+}03, KES04b, Kay04d, NV01, NK01, PGyL03, QBK02, Ven03, Vou03]. simulations [SKM04]. simulator [BE02]. Simultaneous [BÇ02, Boy03a, CCKS02, EESESS03, IOAB02, PR04b]. simultaneously [Zhu04b]. Sinc [KB04b]. sine [Fen04a, Kay03b, Kay04b, Ram01b, WZ01a, Waz04h]. sine-Gordon [Fen04a, Kay03b, Kay04b, Ram01b, WZ01a]. Single [CT03a, Sal03b, CH04, EBE03, MADT03, MAR04a, Moo01, SPH04, WW04, YL04d, Zhu04c]. single-degree-of-freedom [EBE03]. Single-point [CT03a]. single-step [Zhu04c]. Singular [GGRS03, KEB03, Sha03, Abd03c, AN03, ALO03a, All04b, AD04b, zBxF04, CC03a, CF03, Cao04b, CY01a, CY01b, Dar03b, DX04, ES02a, EM03h, Gin04, HY04, JJ04, KA04, KAVM00, Kam02a, KR03a, KR03b, KR04, KW04a, Kum02, Kum03a, Kum03b, Kum03c, Kum03d, Kum03e, Kum09, LL04a, LZL00, Liu03g, LOZ02, LCN04, MSJ04, OOAA04, RC04a, SW01, Wan04c, Waz02f, WW02b, WWW03, Wei04, XD03, YL04a, Yan02f, Yan03u, YZ04, ZW03b, ZW04a, dK02]. Singularities [LM04a, CCP03, Hos04, Lan04a, ZG02]. Singularly [AV04a, KP03, Aln04, Bog04, DD02, Dur04, Jay03, KP02a, KP02b, KS04a, LM04b, NR02a, NR02b, RR02a, RC03, RC04b, RC04c, SR02a, SR02b, SS04b, SS04c, VR03a, VR04b, VR02a, VR02b, VR03b, Xen03]. sinusoidal [CWZ02]. SIR [PTG03]. situ [BE02]. Sivashinsky [Waz04a]. six [EMR03]. sixth [He03g, Tun04a, Waz01g]. sixth-order [He03g, Waz01g]. size [AD00, AL00, EOM04b, HA02, JK03a, JD03]. size-structured [AD00, JD03]. Skan [Asa04a, Asa04b]. skein [GMGC01]. skein-template [GMGC01]. slab [ENAAF04]. slabs [NU00a]. sleep [BT00]. slope [ADAMM03]. slopes [Waz02h]. SLR [HIS04]. sluice [Pet02, Pop04]. small [BN02, CC02, FL02b, KST00, KBÖ00, SR04a]. smooth [BK02, HP02a, SZW03, Yu04, ZAH02, Zay03b, Zay03a, Zay03d]. smoothed [Ric00]. smoother [Nak03, Nak04a, Nak04b]. smoothing [HL04a, NHCLPSR04, SH01, ZG03]. Sobolev [MM02a, ÇD04, Dur04, SY02, dBGM02]. Sobolev-type [MM02a]. society [RAH01]. sociodynamics [TE02b]. soft [ITN00]. software [FH04]. soil [QBK02, SD02b]. solar [EE03, EEES03, EE04a, EE04b, ESEE02]. solid [enNAaAd04]. Solitary [IRS04, Kay04f, Kay04g, Waz02h, Kay04d, Waz02e, Waz03a]. Solitary-wave [Kay04g, Kay04d]. Soliton [HF03, BV02, CZ04, Kay04c, sLqZ03a, LqZ04, Waz01c]. soliton-like [CZ04, sLqZ03a, LqZ04]. Solitons [Bis04b, Bis03, Bis04a, Bis04c]. Solution [BB02b, BBV04d, BTBI03, BBI03, BBI04, BI04a, GHL00, He04e, JKP03, MLA01, Ode02, PZJ03, SV02, WJ00, Zha02a, Abd03d, AS02b, ASEM03, AD02, ALO04, AÖ04a, ASAI03, AS04g, ADG03, AGS03, Asa04a, BH02, BSJ04, Bae04a, zBxF04, BP01, BÇ02, BEAB04, Cao04a, Çat04, ÇB03b, Çel04a, ÇB04c, Çel04b, ÇB04d, CL04b, CHC03, CWY04, DSI04, Deh03f, Deh03e, Deh04e, Deh04f, Dem02a, Dem03b, Dem03c, Dem04, DG03, DF04, DWC04, Dob00, EEES03, EE04a, EGBH03, ENEA04, ENAAM01, ER02, ESEA02, ESK04d, EsEKEA03, El 03, EEAES01a, EEAES01b, EK04, Fat04, Fen04a, Gal03, GB03, Gül04a, Gül04b, GHSJ00, GZ00, HDZ04, HES05, Has04c, HAM04, HS03, HC01, HL04c, IOAB02, IR04, Jay03, JW03, KP02a, KSS02, Kay03b, KES04c, Kay04e, Kha03a, KA03c, Kha04, Khu02, KB04b, Kor03]. solution [lLgC03, Li04g, Lia03a, LY02a, LL04c, Lun01, MK02, MM03, MM04b, MS04c, Man03, MR03, MM04c, Moh03, Moo01, Mun03, NR02b, OASM04, ÖAÖ03, QBK02, Ras03b, Ras04d, RKS04b, Sab18, SKM04, STHN02, Sor03, SW01, Tad01, TZ04, VCD04, rWS04, Waz01g, Wei02a, WW02a, yWShX02, XZ03, Xu04, XCD04a, YS00a, YGL01, Yan02d, Yan02c, Yan03g, YZ04, YÖ00, YYS03, YTM00, ZK03, Zha04c, ZL04a]. Solutions [Aya04b, HI04a, Waz03j, AMI03, Abd04, Afr04, AO00b, AD04a, AS03d, BPJ03, zBxF03, BDD00, BK03, BAB01, CPL00, CHL01, CCP03, CLW02, CCS04a, ÇKB03, CHL02, CC03d, CC03c, CZLZ04, CZ04, Che04d, CLC04, CDH01, CC00a, CC00b, Çin04b, Çin04e, Çin04d, Çin04f, Çin04g, DX04, Din00, DW02, Din02, Din04d, DFF04, DG04, EEB03a, EOM02, EOM03b, EOM03d, EOM08, ES02a, ESAD04, EA00, FH04, FL03, Fen03, FL02b, FNO04, GG00, GM00, GF00, GS02, Guo00, Guo03b, Guo04, GKaM01, HZL02, Has00, HS00, HKA04b, HC02a, HC02b, HF03, HL04b, HL04f, HL04e, HL04d, IOAB01, IH01, IG03, IRS04, JGW02, Jia03b, Kay03a, Kay04c, KI04, KES04b, Kay04d, Kay04f, Kay04g, Keş03a, Keş03b, Keş04d, Kha03b, KERG04, KM03a, KS02b, KH01, KBÖ00, KW04b, Lac03, Lan04a, Lan04b, LSY03, LL04a, LCZ03, sLqZ03a]. solutions [sLqZ03b, Li04b, LR04, Li04f, LNS04, LS04b, LqZ04, LS04c, LHM04, LZ04e, LZ04d, LZ04b, LwWW03, Liu02a, Liu03g, LJ04, LL04d, Liu04a, Liu04c, LQG04, LOZ02, LG03, LYT04, LG04b, LG04a, LGZ04, LG04c, LBE00, ML04b, MdS04, Ma04a, MP03, Mel00, MdL04, MY03, Min03, Min04, MV04, NO03, NAS04, OLT04, PGX00, PS04, Pen04, RdL01, Ras03c, Ras04c, Ras04e, Sad02, Sad04b, Sha02, SW04b, SZK04, ŞÇ03b, Şim04a, SP03b, SL03a, SL03b, SL04d, SL04b, TE00, TE02a, Ten02, Tro04, Tsi03, TT04, TM04b, Tun04b, Tun04c, VM03, WL03, WL04a, Waz01a, Waz01c, Waz01d, Waz02c, Waz02d, Waz02e, Waz02h, Waz03f, Waz03e, Waz03d, Waz03g, Waz03k, Waz03i, Waz04a, Waz04b, Waz04d, WG04c, Waz04i, Waz04h, Waz04k, Waz04j, WJ04, WX04, XYD02, XD03, XCD04c, YY04a, Yal02, YG04, YL04a, Yan01c, Yan02b, Yan02f, Yan03b, Yan03f, Yan03c]. solutions [Yan03d, Yan03e, Yan03h, Yan03k, Yan03l, YC03, Yan03w, Yan04d, Yan04b, Yan04c, Yan04i, Yan04e, YC04, YWY03, Yao02, Yao03a, Yao03b, YL04d, You04, YH04, ZL01, ZZ02b, ZKSD02, ZS04b, ZY04, Zha04a, ZTD04, ZS04a, ZL03, ZaYD04, ZHD04]. Solvability [ESEBD03, LY02b, Liu03a, Liu03b, De 02a, Ism04, ZY00, ZHZ04]. solvable [LTS02]. solve [Asa04c, BV02, ÇB04b, CPTZ04a, lLgC02, MZ04b, Mor01, Sha04a]. solver [AhL00, Pov02]. Solving [AAEA03, BB02c, BB03, ETBAN04, GS00, IbS01, MK03b, MMA04, MK04, NS04, SM02a, WQ04, bSI01, dK02, AA04b, AB02a, AKKN04, AO01b, AÇB03, CY04, ÇB03a, ÇAB03, ÇAB04, CL03a, CC03e, CTZ03, DKX00, DDX02, Din04e, Dis01, Dje00, DNS03, EHM03, EKE02, EM03g, ES02b, ESAA03, ESK04c, ES03b, FS04a, GM03b, Han01, Has06, Has02b, Has04d, He03c, IES04a, JF04, JCL00, qJkSlS03, KP02b, KA04, KML04, KR03c, KEE03b, LCH02, lLgC03, LW03a, Li04f, LS03b, MS04b, NG02b, yN04b, PR04a, RK03, RKS04a, RC03, RC04b, Sal00, SB04a, SZ03a, SYK03, Sun04b, SMQ04, TM04a, VR02b, VR03b, Wan00b, WBW01, WS01, Wan04b, Waz01f, Waz01h, Waz02f, yWShX02, YW02, YW03, Zay04b, ZZ04a, Zhi04, ZW04a]. SOM [qLzWcC03]. Some [ADAAM03, Ahm04b, AB04b, CCS03, Che03, CFS02, CLD03, CO00, DG04, ESEA02, FS03, He04f, KKSY04, Kha03b, KERG04, LOS00, LZ04b, LS04d, Lin04c, Liu03h, Mus00a, Mus00b, Mus00d, NK03, NWX03, Noo04a, Noo04d, PSS00, RS02b, TS01, Wan04a, XC04, Yan02g, ZZ04b, AS04d, AD02, AAMEST04, Bad01, Bai00, Bat04, CTHK03, CZLZ04, CTHT04, Cic01, De 04, EBMA03, EB04d, ESG03, EAA04, HV08, Has02b, JVFM04, Kay03a, Kay04c, KÖ04, KM03b, LSY03, ML04c, OK03a, SM02b, SC03d, VK04, Waz01a, Yan03e]. Soret [CO01]. source [BPJ03, BF04, CT03a, CX04, Fat04]. sources [CD02, EP00]. Southern [TRC03]. space [AAHAD04b, AAAE03a, ALDP01, AET04, CC03b, ÇT04b, DW02, DG04, EMY04, EsEKEA03, Fen04a, Fu03, GB02, Guo02, Guo03a, Guo03b, Hab04b, Han01, Hau04, HIİ04, IOAB01, Küç04, lLgC03, Liu03f, MTK^{+}00, Moh03, Moh04, NM03, yN04b, yN04c, ÖKOD04, RW03, Sam04, STHN02, WQ03, Wei03b, YÇA04]. spacecraft [EEES03, ESEE02]. spaces [AAMEST04, All04a, AET04, AB04b, BS01, BRST02, Bil03, sC03a, Dar03b, Din02, EAA04, Far03, GÇ04, Guo00, Guo04, Hua03, JCZ02, KŞ04b, KM03b, LqZ04, MAK04a, MS04d, MRŽ04, ÖA04, Sor01b, Waz02c, Waz02d, Waz02e, WC01, XL04a, XC04, YÇAM04, ZClC03]. spanned [TCL02]. sparse [CMM02, Zha01, Zha02b]. sparseness [Par02]. Spatial [BR01, AAH04, Hel01, Liu03c, Liu03d, Liu03e, Moh03]. spatially [DF02]. Spatio [GR03, Ram03d]. Spatio-temporal [GR03, Ram03d]. Special [Kin01, SDR03, AaAaZ00b, BBK^{+}03, Dat03, DWC04, EGBH03, EM03e, FH03, Kır04, KÖ04, MCS00, ÖKOD04, Ras03b]. specialized [Wan04d]. specialties [DG04]. species [AT04, BB02c, CWY04, EOI03, EOM04b, KL03, LHL04, SYK02, SPH04, Ten00, Ten02, XCC02, XC02b, YL04d, ZC04b]. specific [SP03a, Waz04e]. specification [Deh03d, Deh04e]. specifications [Deh03e, Deh04a]. specified [ZX04]. specifying [AS02b]. Spectral [Abd01b, Abd02d, AS04b, BH02, ERN03, LCH00, SB03, Tan03, TB04]. spectrum [Kok03, LW04c]. speed [WLX00, XLLL03]. speed-up [XLLL03]. SPH [HL04a]. sphere [SR04b]. Spherical [Hel01, AAAaZ99b, AAHAD04a]. spherically [AAAaZ99b, AaAaZ00b]. spheroidal [Boy03b]. Spiral [Ram04c, Ram04b]. Spline [EHM03, Bah04, Ehr02, KP02a, KA04, Kha04, KS02a, Kum03e, Kum09, KE04a, MR00b, MSJ04, Moo01, Sal00]. splines [ASN03, Bel04, DSI04, HL04a, KENM03, KA03c, NAS04, RC04c]. Splitting [CWS02, CC03a, CC03b, Moh04]. splittings [CWL00, WLW03]. spread [AFRH02, GCSS04]. spring [EESESS03]. SQP [ZZ04c]. square [Wei00a]. squared [CCKS02]. squares [AAEA03, Bel04, CY02, Din02, GL00, GR04, JW03, JY00, Juk04, KL00a, Liu00, MLA01, Wan04a, WC01, WW02a, WW03a, WHT04, Yan00b]. squaring [Wei00c, WWW00]. Squeezing [FRRSCS02]. Srivastava [LS04e]. Stability [AS04h, CH04, EESESS03, Far04, Gop04, HLO02, HHZ04, LRMSVO01, LwWW03, LS03b, NS03, NO03, RG03, SS04a, Sol03a, Sun04a, THY02, YÖ00, Aln04, CD03a, Cao03a, CW04b, CLF04, CHC03, Che04a, CCK04, EA03, EBMA03, EOI03, ESS00b, ESSA00, ESSA01, Far02, GLVW00, GLR02, GR03, HC04a, HC03, HL04b, JLM04, KSC02, LZ01a, LZ01b, LSK00, LH04b, Li04g, LZ04c, Li04c, LC04b, LK04, Mar01, McR01, MG03b, QMWAZK04, Qui03, Qui04, Rad03, Sad04a, Sad04b, Sev04, Shi02, Sol02b, Sol02c, Sol04, Sou03, Sub02b, Tag00b, TM03, Tun04b, WL03, WHM04, XH04, XCC02, XC02c, XCD04b, XCD04d, XCD04f, XL04b, XL04c, YB03, YB04, Yam05a, Yam05b, YLEM04, YY04b, YYS03, wYjSlZ01, ZL01, ZC04a, ZC02, ZZC04]. stabilizable [AQD04, Ame01]. Stabilization [LC02a, EG02, EG03b, EGB03c, GB02, HGS04, LL00, LC03a, PJPL04, WZ03, dlS03]. Stabilized [KA00, Cao04a, Gon04]. stabilizes [HL04a]. stabilizing [HML^{+}02]. Stable [CLL00, HIN04, KL00b, Sal00, Din04a, GH03, KS00, KHL02, LW03a, LW04a, Moh04, Sor03, UvB01, XC04, d’O03b, dlS03]. stage [Ahm04b, Cao03b, DLYC04, JK04c, KM02, LWL00, SP03a, XC01, XCD04b, XCD04d, XCD04e, XCD04c, XCD04f]. stage-specific [SP03a]. stage-structured [DLYC04, JK04c, XC01, XCD04f]. stages [EME03, EMR03]. Stagnation [ZW04b]. stamps [Abd01b]. Stancu [YXC03]. standard [Deh03b, Ram04a, XFL04]. Starlike [OK03c, OK03a, ÖYY04]. starling [Els04]. State [EMY04, Sam04, AESS00b, ALDP01, Ana03, AG04, BD04b, Che04b, EG04a, EG04b, Koł01, LZ04d, PJ04, SWY00]. state-dependent [LZ04d]. state-space [ALDP01]. states [EGB03c, Pla03, YCC02]. static [CC03f, Lun01]. Stationary [Pla03, EB04b, Far02, Gha01, Gha03, Hlo04, NHCJLLP04, SB04a, Vul00, YY02, Yoo04]. Statistical [Gha01, Gha03, KAAD01, EP00, Jol00a, Sah04a, TB04]. statistics [ÇG04, Jah03]. steady [AESS00b, Ana03, AG04, De 03, EGB03c, Man03, SWY00]. steady-state [AESS00b]. steady-states [EGB03c]. steal [Kat01]. Steepest [Ric00]. steering [SSPA01]. Stefan [EK04, KE04b, LOS02, OSL03, San00, VCD04]. stenosed [ES03a]. Step [ÖAD03a, Özk03, CW02, Deh03b, Esc03a, KA03b, LW03a, LW04a, LK04, SZ03b, Tsi03, Ven03, Wu04b, ZG03, Zhu04c]. stepped [Yüc04]. stepsize [XL04b, XL04c]. Stieltjes [BGVHN02, Tag02a]. stiff [ÁR02, AR04c, Bao02, LCH02, Sal00, WX01]. Stirling [EM03b, EM03c, EMED03, EM04b]. stirred [ZL03]. Stochastic [WZ02, AFRH02, AS04h, CLF04, DE00, EBMA03, EGB03b, EG04c, Fad04, FQ04, HHT04, Isk04, KKM04, LF04, Mun03, Nak02, Nak03, Nak04a, Nak04b, Olu03b, Olu04, QMWAZK04, RS03, TE04, WHM04, ZZN04]. stochastically [TE03]. Stock [GP03, ML03, ySW01]. Stoke [KSS02]. Stokes [BK04, Çağ04, Cao04a, CLL00, CY02, ESS04, HAS04f, Hlo04, KL00b, KHL02, LLP02, Man03, She03, SM04c, Zed03]. Stoneley [AAA03]. stop [ySW01]. stop-loss [ySW01]. strain [DWC04]. straining [Ram04c]. strata [NK03]. strategies [CCY04, Esc03a, JDV04, ÖM02, ySW01, XLLL03, Zan00]. strategy [ADG03, WZ03, d’O04]. stratification [AS03b]. stratified [NU00a]. stream [AEG03]. streamline [SY02, SM04c]. streams [Rad04]. strength [AS03a, SM04a]. stress [AAHAD04b, APS04, DYH04, Mel01, Zak03b]. stress-rupture [DYH04]. stress-temperature [Mel01]. stresses [AAAaZ99b, AAAaZ99a, AAAAZ01, AaAaZ00b, AaAaZ00a, AZ02, ENAAF04]. stretching [AEG03, EB03d, EB04f, MM04c, Vaj01, VR04a, Zak04a]. Strict [LZ01a, LZ01b, ZG03]. stripe [SS02]. stripping [SH02]. Strong [All04a, HHT04, Bil04, Bil07, Gül04a, Gül04b, KH01, MZ04a, MAK03, Sla03, YR01]. Strongly [GEA04, CC00a, EL04a, FZ04a, MAK03, dK02]. Structural [BRS04, Coo04, Dua00, Qui03]. Structure [Sla04, Abd00, CW04a, EBE03, EB03a, GB02, KSJ02, Li04a, LHL04, XCD04b, XCD04d, XCD04e, XCD04c]. Structured [Dia04, AD00, DLYC04, JD03, JK04c, XC01, XCD04f]. structures [BR01, Liu03c, MCS04, Waz02a, Waz02b, Waz03c, Waz03a, Waz03j, Waz04e, Waz04c, Waz04f, Waz04g, Waz04h]. studies [DD02, NZP^{+}04]. Study [ES04, GKK02, EM04a, ERN01, EEE03, GGRS03, GM03a, KK04, LPS02, Man03, MT04, Tun04c, Waz03b, Waz03k, Waz04a, WG04b, Waz04i, YAOY03]. Sturm [AhL00, HS02, Has03a, Has04a, Has04b, MA04b, SH02, WJ04, Yan03i, Yan03u]. Sub [Ras02d, SS03, SP03a]. Sub-adjoint [Ras02d]. sub-meshing [SS03]. sub-Saharan [SP03a]. subclass [AKÖ04, Kad03, KA03a, KÖY03, Orh03b, ÖYY04]. subfamily [OK04]. subgrid [Lay02]. subgroups [Isı03]. subject [Deh02, Deh04a, Deh04e, EA04a, Jol00b, Ram03d]. subjected [AAHAD04a]. subliminal [LH04a]. sublinear [LS03a, Yan04c]. submatrices [LW04d]. Subproper [WLW03]. subspace [Bai00, ZZ04a]. substitute [ÇB04b]. substitution [WL04c]. substrates [YAYA03]. subtypes [Gum02]. Successive [Wei00c, WWW00, ÇAB03, ÇB04b, ÇAB04, ŞÇ03b]. suction [AO04b, Elb01, Elb03, EB04f, TS01]. Sueo [Cas00, Nat00, KN00]. Sufficient [CLC04, HA04, YC03, EA03, EOM03b, Juk04, LC01b, TZ04]. Sugisaka [Sug00]. Sugisaka-Ueno-Chandrasekhar [Sug00]. Suitable [JK04b]. sum [MK02, MK03a, ML04a, WZ04c]. summability [He01, RS04a, RS04b, Sav04a]. summable [GEA04]. Summation [Liu04d, Mak04b]. Summations [Chu03]. Summed [CCKS02]. sums [HP02b]. sup [TM03]. Super [CTHK03, CTHT04, GH01]. Super-connectivity [CTHK03]. super-edge-connectivity [CTHK03]. Super-Halley [GH01]. superlinear [Afr04]. superlinearly [KC00]. superposed [Rad03]. superposition [EP00]. support [IKS02]. supported [WHG02]. surface [AR04b, AEG03, AS04c, AO04b, Cap01, EB02b, EB03d, EB04e, EB04f, Has03b, HSE03, HEM04, HSE04, IH01, KG04, LC04c, Pet02, Pop04, QBK02, RP00, RM02, RM03, Sel03, TH04, ZS02]. Surfaces [TH04, AAABH04, AB04a, AAE04, CYC03, KAK05, KAK06, Kas06, Küç04, QS04, Ras02c, TM00]. survey [AS04d, KP02b, KP03]. surveys [HA02]. survival [LC04a, LC02b]. SVD [YY02]. Switching [EOM04b]. Sylvester [dlS03]. Symbolic [CS03b, sLqZ03b, CZLZ04, GM04b, KI04]. Symmetric [AGES04, Ana01, CEM04, CC04c, EEES03, EMED03, ESEE02, GS00, Gar01b, Kay04h, LWL00, LS04f, Moh00, ZHZ04]. symmetries [SM04b]. Symmetry [ML04d, Tan03, Ver03c, YIN00c, BP01, DK04, Ver03a, Ver03b]. symmetry-violations [Ver03c]. synchronization [Yas03]. synchronizor [Yan01b]. synthesis [FT00]. system [Abd01a, AT04, AO00a, Ahm04a, ASN03, All04b, All04c, AR04c, Ana03, AG04, Aya04b, BBV04c, BBV04d, Bae04a, zBxF04, BT00, BBI03, BBI04, CC03a, CCS04a, CC04a, Che04d, CWY04, CCC01, CM04b, Çin04d, CPTZ04a, CFS04, DXL02, De 02b, Din04c, DFF04, EESESS03, EB03b, EG02, EGB03a, EG03b, EG04e, EG04d, EKE03, ESK04d, EEB03b, GS02, HF03, HL04b, HL04f, HL04c, HL04d, Ino03, KAVM00, Kha03a, KERG04, KE02, Lew01, LZ04b, LM04b, LC02b, LHL04, MMA04, MK04, MS04c, MSO00, yN04b, Oda01, PS04, RK03, RKS04a, Ras04f, SM00a, Sar02, SEB03, Sar03b, ST03a, SEG03, SARAEG04, SV02, STHN02, SB04b, SN01, SL03a, SL04d, Tad01, Tad03, Tar03, Tun04a, VR04b, WJ00, WH00, WL03, WC04a, WLW03, WA01, XC02b, XC02c, XL04b, YC04, Yas02, Yas03, ZK03, Zha02a, ZW03a]. system [ZJ04, ZL03, ZLS04, ZW04a]. systems [ALDP01, ALP03, AD04a, AS03d, Ame01, AF03, Aru03, BBV04a, Bai00, BH01, BS01, BDS03, BMR00, BMR03, BL01, BCT04, BRS04, CLW02, CW02, Cao03b, CW04b, Cao04a, CY04, CH04, Cao04b, CKNU00, ÇB04c, CG03, Che04a, Che04b, CL03b, Cic01, CTZ03, CFS02, CW03, CLD03, DMS01, DR01, DF02, EB02a, EM03g, EOM03c, EOI03, EOM03a, ESS00b, ESSA00, ESSA01, EM03h, FT00, Far00, FSLMC03, FS04a, Fu03, Fu04, Gal03, GHW04b, GS00, Gar01b, GM04a, GL00, GKaM01, He00, HLO02, HC04a, HC01, HML^{+}02, HHZ04, HGS04, IN04, Jam01, JLMC03, KML04, KES04a, Koł01, KAAD01, KSJ02, KS01, LB01, LRHRD01, LZL00, LSK00, LM03, Li03b, LR04, LNS04, LS04b, LS04c, LL00, LL03, LC02a, LC03a, Lie04, Lin04a, Lin04b, Lin04c, LY02a, LCN04, MS04b, Men02, MR00b, Mil03, Min04]. systems [Mor01, MS01, Nak02, NCÁHCLP03b, Nak03, Nak04a, Nak04b, NG02b, NV01, Noo03a, Par03, PJPL04, PJ04, Par04a, PKW04, PKLW04, Par04b, PGyL03, Pla03, RHB04, Sad03, SW04a, Shi02, SW04b, Sol02b, Sol02c, Sol03a, Sol04, SC03c, Sun04c, SM04d, Sun04a, Ten00, Ten02, Tia03, Tin01, UvBP00, UvB01, Wan00c, Waz00a, Waz03h, WW02b, WWW03, WW00a, WZ02, WX01, yWShX02, XCC02, Yan01b, Yan03c, Yan03l, Yan03n, Yan04b, Yan04c, YB00a, wYjSlZ01, YH04, Zan00, ZZ02b, ZS04b, Zha04a, ZC04b, ZW03b, dlS03]. Szego [OK03b].

T [Pan08]. tableau [ADG03]. tables [MMP03, Olu04]. tabu [JT04]. Tachibana [Mağ04]. TAGE [MSJ04]. tail [Bry02]. tailed [Tag02b, TV04]. tangential [AAE04]. tanh [LZ04a, LZ04f, Waz04j]. tanh-function [LZ04a]. tapered [BD04c, CM04a]. tapers [TB04]. target [JM02, Mus00c, Oht04]. target-seeking [Mus00c]. tau [PR04a, Esc03a, HS03]. Taylor [Sab18, IES04a, IES04b, Ism04, Keş03b, Keş04d, MM03, MM04b, STHN02, WX03b, YS00a, Yal02]. Taylor-series [STHN02]. technique [Bah04, pFjH03, GLWY04, Has04d, He03a, HZ04b, Jay03, JLMC03, Kam02a, LLCC03, RP00, Yao02, Yao03a, Zha01]. techniques [BP01, Bay04, BM02, BS00, ÇB04a, Deh02, Deh03g, Deh04g, Deh04h, GGRS03, KP02b, RS01a, RHB04, SMF04]. technological [VK04]. technology [ALO03b, JBS04]. telecommunication [Mak04b]. telegraph [Abd04]. temperature [EB04e, EESF04, EEKS04, IH01, Mel01, Mur03, NK03, Pan08, QBK02, Yan02h]. temperature-dependent [EB04e]. temperatures [DYH04]. template [GMGC01]. Temporal [SM00b, GR03, Ram03d]. tend [TE04]. tension [HL04a, KP02a]. Tensional [Nys01]. tensor [CS03a, KSJ02, MS04a]. term [BPJ03, Bae04a, EM03a, Fat04, FSLMC03, Gül04a, Gül04b, KR01, LOS02, LF04, PB02, SY02, WG04a, Waz04g]. terminal [JM02]. terms [Ino02a, Ino02b, Kay04g, Rao02, SC04, Tag03a, Tag03c, WL04a, Waz03h, XL04b, YS00a, YIN00b]. ternary [Pla03]. terrain [Der03]. tessellations [DG02]. test [AY04b, PS00, Tan03]. Testing [EBSAG04, AY04a, AY04b, EBA03, EB03c, JBS04, NG04, dlS03]. Tests [MMP03, ADAMM03, ADARAM04, Coo04, Par02]. tether [RCC04]. tethered [CCC01, CQL01]. textile [JK04b]. TGA [GG02b]. th [CL04a, EOM02, EOM03d, EOM08, Guo00, Guo03a, HC02a, HC02b, Lac03, Mat00, Tra00, Yan03j, Yan03v]. th-order [EOM02, Yan03j]. their [ÁND02, BK04, DF04, DW04, Gia03, Ino02b, KS02b, KH01, ML04c, STB03, SS01, WX03a, WWW03, XL04a]. theorem [AO01a, Özd03b, Özd03a, ÖK03d, ÖA04, RS04b, THY02, Tra00, ZI04, ZY00]. theorems [ALO03a, CP03b, CPTZ04b, Lin04c, LS02, WJ04, WA01, XJM04]. theoretical [KST00, YCC02]. Theory [Bis03, Bis04c, AA04d, BK03, BMR00, BMR01, CLX02, DRP04, Dob00, ES03b, Esc02, Esc03b, Esc03c, Hai00, Has00, He01, He04h, Hil04, Kah06, KENM03, KAAD01, Kru03, LZL00, Liu01, MMR01, Mel00, Pet03, SYY04, TY04, Ton00, Ver00, XC02d, YAYA03, Zhi04]. therapy [TLLKB03]. Thermal [AZ02, EESF04, Kuo04, Pan08, Zak04c, AAAaZ99b, AAAaZ99a, AAAAZ01, AaAaZ00b, AaAaZ00a, enNAaAd04, ENAAF04, EEK03b, EEKS04, HEM03, RPT04, Zak03a]. Thermal-diffusion [EESF04, Pan08]. thermistor [Bah04, Çat04, KE04a, KW04b]. thermo [AAHAD04a, enNAaAd04, EESF04, Pan08]. thermoelastic [EMY04, KERG04, Lee04c, OT04]. thermoelasticity [EEKS04, Mel01, Qui03]. thermomechanical [EMY04]. thermosyphon [CO01]. thermoviscoelastic [AEK04]. thermoviscoelasticity [EKE04]. Theta [KK04]. thin [AR04b, Dem02b, He03b]. Third [FS04a, SDR03, ASN03, ABP04, CH03, FS03, GM03b, KA03c, LS04b, LS04c, MdS04, NAS04, VR02a, VR02b, Waz03b]. Third-order [FS04a, ASN03, ABP04, FS03, KA03c, LS04b, LS04c, NAS04, VR02b, Waz03b]. Thomas [He03h, Lia03a]. Three [Deh04g, EB03b, KHL02, LQG04, AL00, Deh03d, Deh04f, EG04a, EG04d, EG04b, EM03a, EOI03, EG04g, GM04b, HT04, IOAB01, KL03, LWZ00, Liu02a, LK04, LHL04, Moh03, yN04c, QS04, Sal03a, SH01, Sil04, Ver03a, Wu03, Yür04a, Zay02c, Zay04a, ZW03c]. three-dimension [yN04c]. Three-dimensional [KHL02, Deh03d, Deh04f, GM04b, IOAB01, LWZ00, QS04, SH01, Sil04, Yür04a, Zay02c]. three-joint [HT04]. three-layered [EG04g]. Three-level [Deh04g]. Three-mode [EB03b]. three-parameter [AL00]. three-point [Liu02a]. three-species [EOI03]. three-step [LK04]. three-term [EM03a]. Threshold [Vou03, Wan04d, HWW03, Oht04, WH00]. throughflow [Mur01]. Tikhonov [Wan04a, WN04]. Time [DF02, GA04, AESS00a, AFRH02, Ana03, AG04, AM00, AA04d, BRS04, CW04b, CY01b, CGG01, Che04a, CWY04, Che04b, CM04b, CFS02, CFS04, DD00, Dun02, EG03a, EGB03a, EGAR03, EGB03c, EG04e, EG04d, EP00, Gha01, Gha03, HA03, HC04a, HGS04, HC03, JM02, KAK06, Küç04, KBÖ00, KE04b, Lee04b, Les01, ML04b, MP03, Mat00, McR01, MG03b, NCÁHCLP03b, Nak03, Nak04a, Nak04b, Sak03b, Sak04, San00, Shi02, SL04b, SD01, TET02, TE04, TB04, Thu04c, XL04a, XFL04, XC02b, XCD04b, XCD04a, XCD04f, YB03, YB04, Yam05a, Yam05b, YL04d, YL02, YH04, Zak03b, Zak04b, ZS04b, ZC02]. time-delay [HGS04]. time-delays [Che04a, Che04b]. time-dependent [Dun02, Lee04b, Les01, Mat00]. time-fractional [MP03]. time-harmonic [AM00]. time-like [Küç04]. Time-periodically [DF02]. time-reversible [CGG01]. time-variant [HA03]. time-varying [HC03, Shi02, ZS04b, ZC02]. time-warped [XL04a]. Timed [NK01]. timelike [AAABH04, KAK05, Kas06]. times [enNAaAd04, EBA03, EKE04, EEK03b, Mor04, Sam04, d’O04]. Timoshenko [JY00]. titled [Kah05, Kah06]. Titus [Pop04]. Toeplitz [Gün04, KSJ02, SB03, TB02]. together [BB02c, Zay04a]. tolerance [Ami01]. tolerant [CTHT04]. Tomographic [Gzy02]. Topics [SDR03, Wan04a]. topological [All04a]. topology [MAE04b]. torsional [HkT03]. Torus [BŞY03, KU03, UKŞ03]. Total [Kee03a, Sla04, Sol02c]. totality [AS04e]. toxicants [XC01]. Traceability [HLL03]. traceable [CCH04]. tracking [AB03a, HB04]. trade [BS00]. trade-offs [BS00]. traffic [EP00]. trailing [RCC04]. trajectories [Nar02, Sor01c]. trajectory [Küç04, KG04]. transfer [AEG03, DKV04, EA04d, ED02, EB03d, EB04e, EB04f, EESF04, HEM03, JBS04, KE02, LLCC03, NU00b, Pan08, Ram00a, Ram00b, Ram00c, Ram01a, Ram03b, Wan00a, WL02, Zak04a]. transform [Aya03, Aya04a, Aya04b, EESS04, JCL01, SEK01, Tag01b, Tag01d, Tag01e, Tag02b, Tag03b, TV03, TV04, YW02, YW03]. transformation [AESS00a, Bil03, CC04a, CJ04a, EEES03, HZ04a, Has02a, Has02b, Has04d, JCL00, JWL03, JJ04, Kul03, Kuo04]. transformations [Ahn03, CCS03, Cho02b, MS04d, MRŽ04, Yil04]. transformed [XZ03]. transforms [BBV04b, CWZ02, Dat04, Tri04]. Transient [AAAaZ99b, AAAaZ99a, AAAAZ01, AaAaZ00b, AaAaZ00a, AS03c, AO04b, BS04, BGWX03, Lee04c, ZS02, AS03a, ASAI03, AS04g, CHL02, CJ04a, OT04, RM03, Rao04a, Sel03, Sel04a, Sel04b, Tar02]. transmission [BGWX03, TM04b]. transmitted [Yan01a, d’O03b]. transport [BMMRS04, CJ04a, De 03, ESH02, EG04g, GM04b, HB00, HT00, LLCC03, LNW03, ML04d, SC02, wYjSjS03]. transportation [PPS04]. transverse [Ezz04, Zak03a]. transversely [AAAAZ01, enNAaA02, AZ02]. trapezoid [DP01]. trapezoidal [AA04a]. trapped [PGyL03]. Traveling [SZ01, Fen04a, KI04, Lon00, Waz04j]. Travelling [SPH04, Dem04, Din04c, LCZ03]. treat [Ras03a]. treatment [AN03, EE03, EE04a, KW04a, TLLKB03, Waz01b, Waz02g]. trees [BRST02]. Trefftz [AD02]. trenches [Khu03a]. Triangular [Ame01, LMS02]. tridiagonal [Ahm04a, CTZ03, CPTZ04a, EM03a, EM04c, Kok03, LCN04, NG02b]. tridigonal [Kay04h]. trigonometric [CL01, Chu03]. Triple [SW02, Rad03]. truncated [AS04g, DCS03]. trust [ES03b, Sun04b, Yu04, Zhu04a]. Tseng [XY04]. tube [BD04c]. tubes [AAAaZ99a, AaAaZ00a, BD04b, Dem02b]. tumor [GB03]. tunnel [Sei03]. turbulent [De 02b]. Twin [YWY03]. twisted [TY04]. Two [CC04b, JCL01, Liu04h, ÖK03d, SY03, Waz04k, WX01, Wu04b, Yan03v, AESS00b, AAH04, AGES04, AAMK01, Ahm04b, AZ04, ASKT03, ASK04, Ald04, AH02, Ami01, AEK04, Aya03, AÇB03, BE02, Bah03, BF04, BA04, BM02, BS04, BR01, Boy03a, BW04, Bİ04b, Cao03b, CZT04, CWY04, CM04b, Deh00, Deh01, Deh03c, Deh04a, EM04a, EB02a, EB03a, EG04e, EKE04, ES02a, ESK04d, EEE03, EEK03b, Fat04, Gum02, GHSJ00, HT04, IES04b, KP02a, KA04, KR03c, KR03b, Kay04b, KM02, Kha04, Kor03, Kum02, KSC02, Kum03a, Kum03b, Kum03c, Kum03d, Kum03e, Kum09, LWT00, LL04a, LW03a, LZ04c, LW04a, LWL00, Liu04a, LW04e, ML04b, MADT03, Man03, MdL04, MSJ04, Moh04, MR04, NR02b, RS01a, Ram03d, Ram04b, Ram04c, RM03, Rao04a, RCS03, RR02a, RC03, RC04b, RC04c, Sal03b, Sam04, San00]. two [SKM04, SM04a, SV02, Sta00, Tsi03, VR03a, VR03b, XL03, XL04b, YAYA03, ZHD04, Zhu04a]. two-cell [AGES04]. two-degree-of-freedom [EB03a]. Two-dimensional [JCL01, AEK04, Aya03, BE02, Bah03, BF04, BR01, Deh00, Deh01, Deh03c, Deh04a, ESK04d, Fat04, GHSJ00, MdL04, RS01a, Ram03d, Ram04b, Ram04c, RM03, RCS03]. two-layer [BS04]. two-level [XL03]. two-link [HT04]. two-parameter [AAH04, VR03a, VR03b]. two-party [Sta00]. two-patch [CWY04]. two-phase [Ald04, San00]. two-piece [Zhu04a]. Two-point [Yan03v, AZ04, ES02a, KP02a, KA04, KR03c, Kum02, Kum03a, Kum03b, Kum03c, Kum03d, Kum03e, Kum09, ML04b, NR02b, RC03, RC04b, RC04c, SV02]. two-prey [KSC02]. Two-scale [CC04b]. two-species [CWY04]. two-stage [Ahm04b, Cao03b, KM02, LWL00]. Two-step [Wu04b, LW03a, LW04a, Tsi03]. two-variable [Ami01, BW04]. type [AZTK00, BPJ03, Bae04b, BJD^{+}03, CQL01, Dar03a, Din03b, EM03g, ESES04, ESSEF04, EK04, GKK02, GM03b, Hel01, KASK01, Kay04g, LM04a, Lia03b, Lin04a, Lin04b, LS02, MAR04a, MM02a, Min04, MC00, NHCJLLP04, PB02, Par03, Ras03b, RKS04b, Şah04c, SR02a, Sun04c, eT04, Thu04a, TS02, VR02a, WL04b, WFC04, Waz01f, Waz03e, WA01, XC02c, XX03, XCD04b, Yan03p, Yan03q, Yan03u, dBGM02, FZ04a, Waz03j]. types [FH03, LZ04b, Zay04a].

Ueno [Sug00, Cas00, KN00, Nat00]. UHF [HEM03]. UHF/UMF [HEM03]. ultra [All04a, KKM04]. ultra-discretization [KKM04]. ultrasonic [BGWX03]. ultraspherical [Rid04]. UMF [HEM03]. UMVU [Sah04b]. un-stirred [ZL03]. unbalanced [HLLL03]. unbiased [AYW04]. Unbounded [YL04a, AAHAD04a, Fu04, Guo02, KM03b, Liu03f]. unboundedness [Liu03g]. Uncertain [Liu01, Ame01, CW04b, Che04b, GM04a, KAAD01, Lie04, NCÁHCLP03b, NCÁHCLP03a, NHCJLLP04, NHCLPSR04, NCÁHCLP04, Par03, PJ04, PKLW04, Par04b, Shi02]. uncertainty [BHH02]. Unconditional [wYjSlZ01]. unconditionally [GH03, Moh04, PTG03]. unconstrained [Gzy02, Shi04, WXO02]. undeniable [LH04a]. underdetermined [MS01]. undesirable [HV08, JVFM04]. undiscounted [Oht04]. unfolding [Hab04a]. Unicity [HS02]. unidirectional [CCY03, CYC03, HNA04a]. unidirectionally [Yan01b]. Unified [ASKT03, CWZ02, EM03d, KASK01, Mat01].

Uniform [Bae04b, Bog04, CX03, Oqu03, SC03c, AEG03, BPJ03, Bae04a, Bry02, ÇG04, CCS04a, DEK04, HEM04, Kum03b, Mek04, MdOPF04, Mur03]. uniformity [De 02b]. uniformly [ÖK03d]. unifying [Liu01]. unilateral [Bae04b, Has00]. Unique [Ism04]. Uniqueness [AS02b, CL04a, DXL02, Fen03, GF00, Gzy01, HDZ04, Has04c, MdOPF04, QMWAZK04, Sev02, Yan02c, Yan03g]. unit [Coo04, MM02a]. univalent [Kad03, KÖY03]. univariate [NG04]. universal [SB04a]. unknown [BF04, Fat04, Has00, HS00]. unstable [DNS03, LRHRD01, Thu04c, WL04b, YB00a]. Unsteady [AR04d, CCY03, CYC03, ENEA04, EB03d, HNA04b, Moh03, She03, Waz01e, Yür04b, Zak03b]. unstructured [qLzWcC03]. unsymmetric [HS01, Wu04a]. unwrapping [EM04a]. update [Zhu04a]. Updating [MR04, Wei00a]. Upper [Yan03w, AS03b, CPL00, CHL01, CDH01, NO03, Tag02c, Yan03l]. upwind [Dis01]. upwind-biased [Dis01]. Urysohn [Dar03a, ESEBD03]. Use [Tro04, MCC01, Mut03, WW02b]. used [AY04a, BS00, Liu02b, Sai02]. user [WC04a]. Using [ALDP01, Asa04c, JK03a, JSLS04, MS04c, Tia04b, ADAAM03, ADASM04, Ahm04b, ABP04, Ami01, Asa04a, BTBI03, CHL02, CZLZ04, CJ04b, CQL01, EGE02, EG02, EGE03, EG03b, EGT04, EG04f, EM03e, ETBAN04, EEAES01b, FH04, Gon04, GMGC01, GKaM01, HIS04, Has02b, HZ04c, HW04, Hür04, HL04g, IbS01, JLSS04, JF04, JLS^{+}04c, JCL00, Kap04, KS03a, Kay03b, KI04, KA03c, LS04a, LC03b, LBE00, MM04b, MMA04, MM04c, MTK^{+}00, Mog04, Moh00, Nak02, NCÁHCLP03b, Nak03, NCÁHCLP03a, NHCJLLP04, Nak04a, NCÁHCLP04, NS04, OASM04, RA03, RP00, RC04c, RHB04, Ric00, Sab18, Sah04a, SJM04, SEB03, Sar03b, Sar04a, Sar04b, Sha04b, SZ04, SV02, SH01, TJC03, Wan04d, WG04b, WL02, WHT04, Xen03, YIN00a, YB00a, YB00b, YW02, YW03, bSI01, dGKG^{+}00, dlS03, JLST04]. utility [VCV01].

V [Esc02]. vacation [CM04b, MADT03, MAR04a]. vacations [MAR04a]. vaccinated [Yan01a]. Vaccination [ÖM02, GM03a, JDV04, d’O04]. vaccine [d’O03b]. vaccine-induced [d’O03b]. valent [Noo04a, OK04]. valently [AKÖ04, Kad04]. valid [Wan04d]. valid-signature [Wan04d]. Validating [Mus00e]. validity [CW02]. valuation [AA04c, GP03, SMF04]. value [AS02b, AD02, ANS02, AZ04, ASN03, zBxF03, BH01, CH02, CH03, De 02a, De 03, Deh00, Deh03g, Deh04c, DMT02, DH04, EGBH03, ERN01, ES02a, EEK03a, FZ04b, FS04b, GGB03, Guo03a, Has04c, Has04d, He03g, IES04b, JCL00, Jan04, Jay03, JGW02, JBMR02, KP02a, KA04, KR03a, KR03c, KR03b, KR04, KS03a, KA03c, Kha04, KW04a, KA00, Kum02, Kum03a, Kum03b, Kum03c, Kum03d, Kum03e, Kum09, LM04a, Lan04a, Lan04b, LCH02, LL04a, LW03a, Li04f, LC03b, LY02a, Liu02a, LY02b, Liu03f, Liu03g, Liu03a, Liu03b, Liu04b, Liu04a, Liu04c, LQG04, ML04b, Ma04a, MSJ04, MLA01, NK03, NR02b, Ode02, Pen04, RKS04a, RR02a, RC03, RC04b, RC04c, Sha03, SR02b, SV02, STHN02, ŞÇ03b, SD02a, Sor01b, SY03, SL03b, SL04d, TM04b, VR03a, VR04b, VR02b, VR03b, Waz01g, Waz01h, Waz02f, Wei04, WJ04]. value [YL04a, Yan02a, Yan02f, Yan03i, Yan03h, Yan03l, Yan03u, Yan03v, YD04, YWY03, Yao02, Yao03a, Yao03b, YZ04]. valued [AEC03, ESI01, ESSA01, Sol04]. values [AY03, ÇG04, CF03, FH04, GLVW00, Li02, Tag03b]. Vandermonde [EM03b, EM03c, EM03f, EM03e, Mor01]. vanilla [ML03]. vapor [Lee04c].

Varadharajan [CJT03]. Variable [Has03b, HSE03, HEM03, Rad04, AEE04, AS03d, AS04f, Ami01, BCI03, BW04, ÇB04b, ÇB04d, CG03, CC03d, CC03c, ED02, EB02b, EB04e, EG04g, EEK03b, HSE04, IH01, Mas03, Moh04, Moo01, NZP^{+}04, RR02b, TE04, Waz01b, Waz02c, WG04c, XL04b, XL04c, YY04a, YY04b, ZZC04]. variable-coefficient [NZP^{+}04]. variables [AAMK01, ASKT03, Amm04, Bon02, Bİ04b, YX04]. variance [AY04b, CBK00, LCH00]. variant [AÇB03, ÇAB03, ÇAB04, FS03, HA03, Waz02a]. variants [Sha04b, TJC03, Waz03a, Waz04a, Waz04b, Waz04c, Waz04f, Waz04k]. variation [AAE04, Kee03a]. Variational [He00, He03f, He03g, He03h, He04g, Aas03b, AÖE04, CF03, Din00, Din01, Din04d, Din04e, pFjH03, Gla04, He03e, KB04a, MCC01, Mou04, Noo03a, Noo03b, Noo03c, NWX03, Noo04b, NNAK04, NN04b, Noo04d, RO01, SMF04, TY04, ZZ02a, ZY00]. variational-like [Din04e]. variations [Par04b]. various [sLqZ03b, Liu01, MAE04a]. varying [ASAI03, Cap01, HC03, RH04, Shi02, ZS04b, ZC02]. vector [Bis03, CGG01, LHM04, Mak04b, RHB04, Sad04a, WX03b]. vector-network [RHB04]. vectors [WC04b]. vehicle [SWL00]. velcro [BF03]. velocity [CY02, Yen04]. ventilated [LLCC03]. venting [SD02b]. verification [Khu03b, Wan04d]. verifiers [ZX04]. versa [Tag03d]. version [GK02]. vertical [AS03c, AO04b, BS04, BBN04, ENEA04, EB02b, Elb03, Has03b, Xu04, ZS02]. vertically [d’O03b]. via [AS04h, Che04b, CEM04, ESSA01, GKaM01, IPPT03, JKP03, KR03b, ML03, MY03, Pap00, Sol04, SC03d, TV03, Ver03c, Zhi04]. vibrating [Zay02b, Zay02c, Zay03d]. Vibration [EA04a]. vibrations [AEL04, GB02]. vice [Tag03d]. VIDEs [GG00]. view [DSC01]. violations [Ver03c].

Virtual [BL01, KU01]. viscoelastic [AAHAD04a, enNAaAd04, Dem03a, HNA04a, Zak03a, Zak04a]. viscosity [AEE04, ED02, EB04e, EESF04, EG04g, HEM03, Lay02, Pan08]. Viscous [Vaj01, BEAB04]. visible [TM00]. VOCs [LLCC03]. Voigt [CYC03]. Volterra [Abd02b, Abd02a, Abd03c, AM03, AMI03, AS04b, AT04, Aln04, BBV04c, Bad01, BFGG04, BBI03, Dar03a, Dar04, EOI03, ESSEF04, GHSJ00, HHJ01, HL04d, LR04, LC02b, MM03, MS04c, PR04a, Ten00, Tin01, Waz02g, XCC02, XCD04b, XCD04a, YS00a, Yal02, Yan01a, YC04, ZJ04]. Volume [Ano02i, Ano03h, Ano03i, Ano03l, Ano03j, Ano03k, Ano04a, Ano04g, Ano04h, Ano04i, Ano04b, Ano04c, Ano04j, Ano04k, Ano04d, Ano04l, Ano04e, Ano04f, Ano04m, AO00c, AEK04, CCY03, CYC03, Ram00c]. Volumes [Ano04n, Ano03m]. Voronoi [DG02]. vortex [HSE03]. vortices [HSE04]. vorticity [CY02]. vote [WW00a]. voter [YIN00b]. voters [Ino02b, YIN00a, YIN00c]. voting [CJT03, Ino03, Sta00]. Vries [Dem02a, Dem03c, KBÖ00]. vs [He04h].

waiting [San00]. Walker [BAB01]. walks [Kan04]. wall [Mas03]. Walsh [Gla04]. Walsh-wavelet [Gla04]. Wang [Has06]. warped [XL04a]. water [FSLMC03, Rao04b, WHG02]. wave [enNAaA02, AAA02, AL03, BPJ03, Bae04a, Bae04b, BI04a, BGWX03, Dem04, ENAAM01, FZ04a, Fen04a, Gül04a, Gül04b, Gzy01, HZL02, HF03, IRS04, KSL02, KI04, Kay04d, Kay04f, Kay04g, LCZ03, LZ04a, PB02, Ram04c, RKS04b, SK04, eT04, Waz01a, WG04c, Waz04j, ZAH02, Zay03b, Zay03e, Zay04a, Zay04b, ZC03]. wave-like [WG04c]. waveform [Jia03b, Yua03]. wavefronts [SPH04]. waveguides [Bis04b, Yen04]. Wavelet [AV04b, CL01, ESAA03, Gla04, SL04a]. wavelet-Galerkin [ESAA03]. wavelets [DEK04]. waves [AAA03, AAHAD04b, enNAaAd04, AM00, AA04d, BD04b, BD04c, BN02, Cha04b, Dem02b, Din04c, Esc03c, Ram04b, Sei03, SZ01]. wavy [EEE03]. way [CJ04b, KML04]. Weak [Çağ04, Sla03, CWL00, GF00, GS02, Hua03, MAE04a, MdOPF04, Tro04, YÖ00]. Weakly [Dar03b, Ami01, Dem02b, JJ04, Wan00c]. weather [BBN04]. web [EORI01]. wedge [ED02, Has03b, HEM03, Khu03a]. wedge-shaped [Khu03a]. Weighted [Deh04h, MM02b, MAD04, WN04, CW04a, CY04, CL03a, Far03, Isk04, JW03, LC03b, MS04d, RW03, RS04a, SSPA01, SW02, Wan04a, WZ04b, Wei00b, WWW00, WW01, Wei01, Wei02a, WW03a, Wei03a, Wei03b, WWL04]. weights [JF04, JSdN04, LP03]. weights-restricted [JF04]. which [AB04a, CGVC04, TET02]. white [EG04c, NCÁHCLP04]. Whittaker [SS00]. wide [NHCJLLP04]. wide-sense [NHCJLLP04]. Wiener [Nak04b]. Williams [AO01a]. Wills [KL02]. wind [KB04b]. wind-driven [KB04b]. Wintner [Tra00]. Wintner-Leighton [Tra00]. within [Ram00c]. without [CJ04b, EB04c, HL04c, yS01b, Wan00c, WW00b, ZG03]. work [KU01, UKŞ03]. Workshop [SDR03]. worlds [BL01]. Wright [MP03]. wronkler [ÇB04b, ÇB04d].

XML [LC04d]. XVII [Esc03b]. XVIII [Esc03c].

Yang [Koz03].

Zakharov [LCZ03]. zero [CGVC04, Hab04b, RA03, Sor02, Wol04]. Zeros [dBGM02, ÁND02, HP02a, PR04b, Wol02, WSX03, WX04, Zhu04b]. zeta [Liu04d, Mus00a, Mus00b, Mus00d, CCS04b, KKSY04, LS04d, RS02b, Sri03].

References

Abdel-Aziz:2000:CAL

[AA00] Mohammedi R. Abdel-Aziz.

Convexity analysis of the largest dependent eigenvalue functions of eigensystems. Applied Mathematics and Computation, 109 (1):93–100, March 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/82/21/
28/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/82/21/28/article.
pdf.

Abbasbandy:2004:NTF

[AA04a] S. Abbasbandy and B. Asady.

The nearest trapezoidal fuzzy number to a fuzzy quantity. Applied Mathematics and Computation, 156 (2):381–386, September 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abbasbandy:2004:NMS

[AA04b] S. Abbasbandy and B. Asady.

Newton’s method for solving fuzzy nonlinear equations. Applied Mathematics and Computation, 159(2): 349–356, December 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Arciniega:2004:EDM

[AA04c] Armando Arciniega and Edward Allen. Extrapolation of difference methods in option valuation. Applied Mathematics and Computation, 153(1):165–186, May 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ates:2004:EFF

[AA04d] Ülkü (Dınlemez) Ateş and İbrahim Ethem Anar. Eigenvalues of the far field operator for time harmonic acoustic waves by an inhomogeneous medium and inverse scattering theory. Applied Mathematics and Computation, 158(3):835– 851, November 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ahmed:2002:EWP

[AAA02] S. M. Ahmed and A. M. Abd-Alla. Electromechanical wave propagation in a cylindrical poroelastic bone with cavity. Applied Mathematics and Computation, 133(2–3):257–286, December 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abd-Alla:2003:SRW

[AAA03] A. M. Abd-Alla and S. M.

Ahmed. Stoneley and Rayleigh waves in a nonhomogeneous orthotropic elastic medium under the influence of gravity. Applied Mathematics and Computation, 135(1):187–200, February 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abd-All:1999:TTSb

[AAAaZ99a] A. M. Abd-All, A. N.

Abd-alla, and N. A. Zeidan. Transient thermal stresses in a rotation non-homogeneous cylindrically orthotropic composite tubes. Applied Mathematics and Computation, 105(2–3):253–269, November 12, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/amc/sub/1999/105/
2-3/6323.pdf;   http:/
/www.elsevier.com/cgi-
bin/cas/tree/store/amc/
cas_sub/browse/browse.
cgi?year=1999&volume=105&
issue=2-3&aid=6323. See

erratum [AaAaZ00a].

Abd-All:1999:TTSa

[AAAaZ99b] A. M. Abd-All, A. N. Abdalla, and N. A. Zeidan. Transient thermal stresses in a spherically orthotropic elastic medium with spherical cavity. Applied Mathematics and Computation, 105(2–3):231–252, November 12, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/amc/sub/1999/105/
                                        2-3/6314.pdf;   http:/
                                        /www.elsevier.com/cgi-
                                        bin/cas/tree/store/amc/
                                        cas_sub/browse/browse.
                                        cgi?year=1999&volume=105&
                                        issue=2-3&aid=6314. See

erratum [AaAaZ00b].

Abd-alla:2000:ETTb

[AaAaZ00a] A. M. Abd-alla, A. N. Abdalla, and N. A. Ziedan. Erratum to: “Transient thermal stresses in a rotation non-homogeneous cylindrically orthotropic composite tubes” [Appl. Math. Comput. 105 (1999) 253– 269]. Applied Mathematics and Computation, 112 (2–3):346, June 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/86/23/
                                        36/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/86/23/36/article.

pdf. See [AAAaZ99a].

Abd-alla:2000:ETTa

[AaAaZ00b] A. M. Abd-alla, A. N. Abdalla, and N. A. Ziedan. Erratum to: “Transient thermal stresses in a spherically orthotropic elastic medium with special cavity” [Appl. Math. Comput. 105 (1999) 231–252]. Applied Mathematics and Computation, 112(2–3):345, June 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/86/23/
35/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/86/23/35/article.

pdf. See [AAAaZ99b].

Abd-Alla:2001:TTS

[AAAAZ01] A. M. Abd-Alla, A. N.

Abd-Alla, and N. A. Zeidan. Transient thermal stresses in a transversely isotropic infinite circular cylinder. Applied Mathematics and Computation, 121 (1):93–122, May 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/105/21/
27/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/105/21/27/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002659.

Abdel-All:2004:RST

[AAABH04] Nassar H. Abdel-All, Rashad A.

Abdel-Baky, and Fathi M. Hamdoon. Ruled surfaces with timelike rulings. Applied Mathematics and Computation, 147(1):241–253, January 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See errata [KAK05, KAK06, Kas06].

Abdel-All:2003:PCF

[AAAE03a] Nassar H. Abdel-All and H. N. Abd-Ellah. Perturbations of curvatures flow in a hyperbolic space. Applied Mathematics and Computation, 143(2–3):349–360, November 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Achchab:2003:PEE

[AAAE03b] B. Achchab, S. Achchab, A. Agouzal, and R. Ellaia. On a posteriori error estimator for primal, equilibrium and mixed approximation of diffusion equations. Applied Mathematics and Computation, 134 (1):83–92, January 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

All:2004:TVH

[AAE04] N. H. Abdel All and H. N. Abd-Ellah. The tangential variation on hyperruled surfaces. Applied Mathematics and Computation, 149(2): 475–492, February 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abdel-Aziz:2003:SLS

[AAEA03] Mohammedi R. Abdel-Aziz and Mahmoud M. El-Alem. Solving large-scale constrained least-squares problems. Applied Mathematics and Computation, 137(2– 3):571–587, May 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abdel-All:2004:GCT

[AAH04] Nassar H. Abdel-All and Fathi M. Hamdoon. A geometric characterisation of two-parameter spatial motions with many locally onedimensional point paths. Applied Mathematics and Computation, 153(1):19–25, May 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abd-Alla:2004:MTV

[AAHAD04a] A. M. Abd-Alla, H. A. H.

Hammad, and S. M. Abo-Dahab. Magneto-thermo-viscoelastic interactions in an unbounded body with a spherical cavity subjected to a periodic loading. Applied Mathematics and Computation, 155(1):235–248, July 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abd-Alla:2004:RWMb

[AAHAD04b] A. M. Abd-Alla, H. A. H. Hammad, and S. M. Abo-Dahab. Rayleigh waves in a magnetoelastic halfspace of orthotropic material under influence of initial stress and gravity field. Applied Mathematics and Computation, 154(2):583– 597, July 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abdel-All:2003:GPP

[AAMAE03] N. H. Abdel-All, M. A. W. Mahmoud, and H. N. Abd-Ellah. Geometrical properties of Pareto distribution. Applied Mathematics and Computation, 145(2–3): 321–339, December 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ali:2004:FCD

[AAMEST04] Sahar Mohammed Ali, N. Abdel-Mottaleb, E. M. El-Shobaky, and Wataru Takahashi. Finite co-dimensional Banach spaces and some bounded recovery problems. Applied Mathematics and Computation, 153(3):785–792, June 14, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Agarwal:2001:SBT

[AAMK01] Satish K. Agarwal, Dhaifalla K.

Al-Mutairi, and Pranesh Kumar. Selecting between two concomitant variables in ratio method of estimation. Applied Mathematics and Computation, 124(2):227–234, October 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/113/31/32/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000916.

Aassila:2003:ECS

[Aas03a] M. Aassila. Exact controllability of the Schrödinger equation. Applied Mathematics and Computation, 144(1):89–106, November 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Aassila:2003:VID

[Aas03b] Mohammed Aassila. On a variational inequality for a degenerate quasilinear hyperbolic equation. Applied Mathematics and Computation, 137(1):1–14, May 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abdou:2002:MSI

[AB02a] M. A. Abdou and A. A.

Badr. On a method for solving an integral equation in the displacement contact problem. Applied Mathematics and Computation, 127 (1):65–78, March 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/9/12/123/28/30/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300301000030.

Aggoun:2002:ADM

[AB02b] Lakhdar Aggoun and Lakdere Benkherouf. M-ary detection of Markov-modulated Poisson processes in inventory models. Applied Mathematics and Computation, 132(2–3):315–324, November 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Aggoun:2003:PTI

[AB03a] Lakhdar Aggoun and Lakdere Benkherouf. Parameter tracking for an inventory model. Applied Mathematics and Computation, 146(2–3): 393–401, December 31, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Amat:2003:HOS

[AB03b] Sergio Amat and Sonia Busquier. On a higher order Secant method. Applied Mathematics and Computation, 141(2–3):321–329, September 5, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abdel-Baky:2004:LCW

[AB04a] Rashad A. Abdel-Baky. On a line congruence which has the parameter ruled surfaces as principal ruled surfaces. Applied Mathematics and Computation, 151 (3):849–862, April 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Aydin:2004:SND

[AB04b] Cafer Aydın and Feyzı Başar. Some new difference sequence spaces. Applied Mathematics and Computation, 157(3):677–693, October 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abbasbandy:2003:INR

[Abb03] S. Abbasbandy. Improving Newton–Raphson method for nonlinear equations by modified Adomian decomposition method. Applied Mathematics and Computation, 145(2–3):887–893, December 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abbas:2004:FIF

[Abb04a] S. E. Abbas. Fuzzy βirresolute functions. Applied Mathematics and Computation, 157(2):369–380, October 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abbas:2004:FIM

[Abb04b] S. E. Abbas. Fuzzy Sirresolute mappings. Applied Mathematics and Computation, 155(2):329–343, August 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abdou:2000:FIE

[Abd00] M. A. Abdou. Fredholm integral equation with potential kernel and its structure resolvent. Applied Mathematics and Computation, 107(2–3):169–180, January 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/72/22/
                                        28/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/72/22/28/article.
                                        pdf.

Abdou:2001:IEM

[Abd01a] M. A. Abdou. Integral equation with Macdonald kernel and its application to a system of contact problem. Applied Mathematics and Computation, 118(1):83–94, February 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/96/20/24/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/96/20/
24/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399001502.

Abdou:2001:SRI

[Abd01b] M. A. Abdou. Spectral relationships for integral operators in contact problem of impressing stamps. Applied Mathematics and Computation, 118(1):95–111, February 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/96/20/25/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/96/20/
25/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399001678.

Abdou:2002:FVIb

[Abd02a] M. A. Abdou. Fredholm– Volterra integral equation and generalized potential kernel. Applied Mathematics and Computation, 131(1): 81–94, September 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abdou:2002:FVIa

[Abd02b] M. A. Abdou. Fredholm– Volterra integral equation of the first kind and contact problem. Applied Mathematics and Computation, 125(2–3):177–193, January 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        9/12/120/31/29/abstract.
                                        html;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300300001181.

Abdou:2002:FPI

[Abd02c] M. A. Abdou. Fundamental problems for infinite plate with a curvilinear hole having finite poles. Applied Mathematics and Computation, 125(1):79–91, January 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        9/12/120/27/32/abstract.
                                        html;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630030000117X.

Abdou:2002:SRI

[Abd02d] M. A. Abdou. Spectral relationships for the integral equation with Macdonald kernel and contact problem. Applied Mathematics and Computation, 125(1):93–103, January 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/120/27/33/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001193.

Abdallah:2003:ONO

[Abd03a] Sabah Hafez Abdallah. Oscillatory and non-oscillatory behaviour of second-order neutral delay differential equations. Applied Mathematics and Computation, 135(2–3):333–344, March 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abdelwahid:2003:MMA

[Abd03b] Fawzi Abdelwahid. A mathematical model of Adomian polynomials. Applied Mathematics and Computation, 141(2–3):447–453, September 5, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abdou:2003:FVIa

[Abd03c] M. A. Abdou. Fredholm– Volterra integral equation with singular kernel. Applied Mathematics and Computation, 137(2–3):231–243, May 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abdou:2003:SLN

[Abd03d] M. A. Abdou. On the solution of linear and nonlinear integral equation. Applied Mathematics and Computation, 146(2–3):857–871, December 31, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abdusalam:2004:AAS

[Abd04] H. A. Abdusalam. Analytic and approximate solutions for Nagumo telegraph reaction diffusion equation. Applied Mathematics and Computation, 157(2):515– 522, October 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Amat:2004:FCM

[ABEM04] Sergio Amat, Sonia Busquier, Dhafer El kebir, and José Molina. A fast Chebyshev’s method for quadratic equations. Applied Mathematics and Computation, 148(2): 461–474, January 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Amat:2004:DFT

[ABP04] Sergio Amat, Sonia Busquier, and Sergio Plaza. Dynamics of a family of third-order iterative methods that do not require using second derivatives. Applied Mathematics and Computation, 154 (3):735–746, July 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abuelmaatti:2000:IPB

[Abu00] Muhammad Taher Abuelma’atti.

Intermodulation performance of biased-nonlinearities. Applied Mathematics and Computation, 107(2–3):211–224, January 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/72/22/
31/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/72/22/31/article.
pdf.

Angel:2001:BC

[AC01] E. Angel and J. Casti.

The Bellman Continuum. Applied Mathematics and Computation, 120(1–3):1, May 10, 2001. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/104/21/
21/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/104/21/21/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002817.

Aykut:2003:MTS

[AÇB03] Arzu Aykut, Ercan Çelik, and Mustafa Bayram. The modified two sided approximations method and Padé approximants for solving the differential equation with variant retarded argument. Applied Mathematics and Computation, 144(2–3):475– 482, December 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ackleh:2000:MAN

[AD00] Azmy S. Ackleh and Keng Deng. A monotone approximation for a nonlinear nonautonomous sizestructured population model. Applied Mathematics and Computation, 108(2–3):103– 113, February 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/79/22/
                                        25/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/79/22/25/article.
                                        pdf.

Anderson:2001:IGP

[AD01] Russell K. Anderson and Moshe Dror. An interactive graphic presentation for multiobjective linear programming. Applied Mathematics and Computation, 123(2):229–248, September 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/110/29/32/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000710.

Abou-Dina:2002:ITM

[AD02]  M. S. Abou-Dina.  Im-

plementation of Trefftz method for the solution of some elliptic boundary value problems. Applied Mathematics and Computation, 127(1):125–147, March 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/123/28/33/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300301000637.

Akyuz-Dacioglu:2004:CPS

[AD04a]  Ayegül  Akyüz-Dacıoğlu.

Chebyshev polynomial solutions of systems of linear integral equations. Applied Mathematics and Computation, 151(1):221–232, March 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Amer:2004:CNS

[AD04b] S. M. Amer and S. Dardery.

On a class of nonlinear singular integral equations with shift on a closed contour.

Applied Mathematics and Computation, 158(3):781– 791, November 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Adamatzky:2001:PCD

[Ada01] Andrew Adamatzky. Pathology of collective doxa. Automata models. Applied Mathematics and Computation, 122(2):195–228, July 25, 2001. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/9/12/106/25/
                                        24/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/9/12/106/25/24/article.
                                        pdf;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300300000254.

Adamatzky:2003:AAM

[Ada03] Andrew Adamatzky. Affectons: automata models of emotional interactions. Applied Mathematics and Computation, 146(2–3):579–594, December 31, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abu-Dayyeh:2003:SEF

[ADAAM03] Walid A. Abu-Dayyeh, M. S. Ahmed, R. A. Ahmed, and Hassen A. Muttlak. Some estimators of a finite population mean using auxiliary information. Applied Mathematics and Computation, 139(2–3):287–298, July 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abu-Dayyeh:2003:LCB

[ADAJM03] Walid Abu-Dayyeh, Jehad Al-Jararha, and Kailash C. Madan. LRT conjecture for bivariate normal. Applied Mathematics and Computation, 145(1):73–84, December 20, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abu-Dayyeh:2003:EBS

[ADAMM03] Walid A. Abu-Dayyeh, Marwan A. Al-Momani, and Hassen A. Muttlak. Exact Bahadur slope for combining independent tests for normal and logistic distributions. Applied Mathematics and Computation, 135(2– 3):345–360, March 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abu-Dayyeh:2004:LPC

[ADARAM04] W. A. Abu-Dayyeh, Z. R.

Al-Rawi, and M. Ma. Al-Momani. Local power for combining independent tests in the presence of nuisance parameters for the normal distribution. Applied Mathematics and Computation, 147(3):687–698, January 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abu-Dayyeh:2004:LPE

[ADASM04] Walid A. Abu-Dayyeh, Sameer A. Al-Subh, and Hassen A. Muttlak. Logistic parameters estimation using simple random sampling and ranked set sampling data. Applied Mathematics and Computation, 150 (2):543–554, March 8, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Arsham:2003:AST

[ADG03] H. Arsham, T. Damij, and J. Grad. An algorithm for simplex tableau reduction: the push-to-pull solution strategy. Applied Mathematics and Computation, 137(2– 3):525–547, May 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Agiza:2004:CDN

[AE04] H. N. Agiza and A. A. Elsadany. Chaotic dynamics in nonlinear duopoly game with heterogeneous players. Applied Mathematics and Computation, 149(3): 843–860, February 22, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abul-Ez:2003:OBS

[AEC03] M. A. Abul-Ez and D. Constales. On the order of basic series representing Clifford valued functions. Applied Mathematics and Computation, 142(2–3):575–584, October 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

AbdElNaby:2004:EEF

[AEE04] Abd El Hakeem Abd El Naby, A. E. M. El Misery, and I. I. El Shamy. Effects of an endoscope and fluid with variable viscosity on peristaltic motion. Applied Mathematics and Computation, 158(2):497– 511, November 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abo-Eldahab:2003:CHT

[AEG03] Emad M. Abo-Eldahab and Ahmed F. Ghonaim. Convective heat transfer in an electrically conducting micropolar fluid at a stretching surface with uniform free stream. Applied Mathematics and Computation, 137(2– 3):323–336, May 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Aouadi:2004:REV

[AEK04] M. Aouadi and A. S. El-Karamany. The relaxation effects of volume properties in two-dimensional generalized thermoviscoelastic problem. Applied Mathematics and Computation, 151 (3):689–711, April 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abd-El-Latif:2004:PFM

[AEL04] G. M. Abd-El-Latif. On a problem of a field method and its applications to the non-linear vibrations. Applied Mathematics and Computation, 147(1):267–289, January 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abd-El-Salam:2000:CTI

[AESS00a] M. R. Abd-El-Salam and M. H. Shehata. Coordinate transformation and implicit time discretization for diffusion in glassy polymers. Applied Mathematics and Computation, 113(2–3):199–218, July 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/87/23/25/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/87/23/
25/article.pdf.

Abd-El-Salam:2000:ESS

[AESS00b] M. R. Abd-El-Salam and M. H. Shehata. Establishment of steady-state flow in mixture consisting of two electrically and nonelectrically conducting fluids in porous pipes. Applied Mathematics and Computation, 113(2–3):235–248, July 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/87/23/27/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/87/23/
27/article.pdf.

Altin:2004:SSS

[AET04] Yavuz Altin, Mikâil Et, and Binod Chandra Tripathy. The sequence space |N̄_{p}|(M, r, q, s) on seminormed spaces. Applied Mathematics and Computation, 154(2):423–430, July 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

An:2003:CSS

[AF03] Yukun An and Xianling Fan. On the coupled systems of second and fourth order elliptic equations. Applied Mathematics and Computation, 140(2–3):341–351, August 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Afrouzi:2004:EPS

[Afr04] G. A. Afrouzi. Existence of positive solutions on indefinite superlinear elliptic equations. Applied Mathematics and Computation, 157(3):841–848, October 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Allen:2002:DTD

[AFRH02] Linda J. S. Allen, David A. Flores, Ruwan K. Ratnayake, and John R. Herbold. Discrete-time deterministic and stochastic models for the spread of rabies. Applied Mathematics and Computation, 132(2–3):271– 292, November 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Afifi:2003:IPF

[AG03] N. A. S. Afifi and N. S. Gad. Interaction of peristaltic flow with pulsatile fluid through a porous medium. Applied Mathematics and Computation, 142(1):167–176, September 20, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ananda:2004:SSA

[AG04] Malwane M. A. Ananda and Jinadasa Gamage. On steady state availability of a system with lognormal repair time. Applied Mathematics and Computation, 150(2):409–416, March 8, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abdel-Gawad:2004:SPD

[AGES04] H. I. Abdel-Gawad and A. M. El-Shrae. Symmetric patterns in the Dirichlet problem for a two-cell cubic autocatalytor reaction model. Applied Mathematics and Computation, 150 (3):623–645, March 17, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Area:2002:CPO

[AGM02] I. Area, E. Godoy, and F. Marcellán. q-Coherent pairs and q-orthogonal polynomials. Applied Mathematics and Computation, 128(2– 3):191–216, May 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Arsham:2003:LCG

[AGS03] Hossein Arsham, Miro Gradisar, and Mojca Indihar Stemberger. Linearly constrained global optimization: a general solution algorithm with applications. Applied Mathematics and Computation, 134(2–3):345–361, January 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Allan:2002:FMI

[AH02] F. M. Allan and M. H. Hamdan. Fluid mechanics of the interface region between two porous layers. Applied Mathematics and Computation, 128(1):37–43, May 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Adomaitis:2000:CQB

[AhL00] Raymond A. Adomaitis and Yi hung Lin. A collocation… quadrature-based Sturm–Liouville problem solver. Applied Mathematics and Computation, 110(2– 3):205–223, April 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/83/23/
                                        27/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/83/23/27/article.
                                        pdf.

Ahmad:2004:SET

[Ahm04a] Faiz Ahmad. A system of equations with a tridiagonal coefficient matrix. Applied Mathematics and Computation, 159(2):435–438, December 6, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ahmed:2004:SEF

[Ahm04b] M. S. Ahmed. Some estimators for a finite population mean under twostage sampling using multivariate auxiliary information. Applied Mathematics and Computation, 153 (2):505–511, June 4, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ahn:2003:GGT

[Ahn03] Young-Ho Ahn. Generalized Gauss transformations. Applied Mathematics and Computation, 142(1):113– 122, September 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Al-Khaled:2004:NCM

[AKKN04] Kamel Al-Khaled, Dogan Kaya, and Muhammad Aslam Noor. Numerical comparison of methods for solving parabolic equations. Applied Mathematics and Computation, 157(3): 735–743, October 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Akbulut:2004:SVF

[AKÖ04] Sezgin Akbulut, Ekrem Kadioğlu, and Murat Özdemir. On the subclass of p-valently functions. Applied Mathematics and Computation, 147(1):89–96, January 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Alexandrov:2000:NTP

[AL00] Mikhail D. Alexandrov and Andrew A. Lacis. A new three-parameter cloud/ aerosol particle size distribution based on the generalized inverse Gaussian density function. Applied Mathematics and Computation, 116(1–2):153–165, November 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/91/20/32/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/91/20/
32/article.pdf.

Anh:2003:HAR

[AL03] V. V. Anh and N. N. Leonenko. Harmonic analysis of random fractional diffusion– wave equations. Applied Mathematics and Computation, 141(1):77–85, August 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Aldas:2004:ATP

[Ald04]  Kemal Aldas.  Applica-

tion of a two-phase flow model for hydrogen evolution in an electrochemical cell. Applied Mathematics and Computation, 154 (2):507–519, July 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Adjakou:2001:USS

[ALDP01] R. Adjakou, C. Lishou, N. Dieye, and L. Protin. Using state-space representation for the modelisation of photovoltaic systems. Applied Mathematics and Computation, 124(1):129–138, November 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/113/27/33/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000898.

Alla:2004:SUS

[All04a] M. A. Fath Alla. Strong and ultra separation axioms on L-fuzzy topological spaces. Applied Mathematics and Computation, 157(2): 573–587, October 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Allahverdiev:2004:DEP

[All04b] Bilender P. Allahverdiev.

Dissipative eigenvalue problems for a singular Dirac system. Applied Mathematics and Computation, 152 (1):127–139, April 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Allahviranloo:2004:NMF

[All04c] Tofigh Allahviranloo. Numerical methods for fuzzy system of linear equations. Applied Mathematics and Computation, 155(2):493– 502, August 6, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Alnasr:2004:SBD

[Aln04] Modi H. Alnasr. The P - stability of backward differentiation method for singularly perturbed Volterra integral equations with delay argument. Applied Mathematics and Computation, 154(3):881–887, July 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Agarwal:2003:ETO

[ALO03a] Ravi P. Agarwal, Haishen Lü, and Donal O’Regan. Existence theorems for the one-dimensional singular p– Laplacian equation with sign changing nonlinearities. Applied Mathematics and Computation, 143(1):15–38, October 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ahn:2003:EEF

[ALO03b] Min Jung Ahn, Hyun Young Lee, and Mi Ray Ohm. Error estimates for fully discrete approximation to a free boundary problem in polymer technology. Applied Mathematics and Computation, 138(2–3):227–238, June 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Akinola:2004:SAE

[ALO04] A. P. Akinola, O. P. Layeni, and M. A. Olagunju. A solution by anisotropic expansion for a composite parallelepiped deformed into cylinder. Applied Mathematics and Computation, 149 (3):599–611, February 22, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Agarwal:2003:ABN

[ALP03] Ravi P. Agarwal, Wan-Tong Li, and P. Y. H. Pang. Asymptotic behavior of nonlinear difference systems. Applied Mathematics and Computation, 140(2–3):307– 316, August 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anar:2000:FFP

[AM00] İbrahim Ethem Anar and Adil Misir. Far-field patterns for time-harmonic acoustic plane waves in an inhomogeneous medium. Applied Mathematics and Computation, 111(1):103–119, May 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/

                                        gej-ng/29/17/20/85/21/
                                        27/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/85/21/27/article.
                                        pdf.

Agarwal:2002:P

[AM02a] R. P. Agarwal and G. V. Milovanović. Preface. Applied Mathematics and Computation, 128(2–3):149, May 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Agarwal:2002:EPI

[AM02b] Ravi P. Agarwal and Gradimir V. Milovanovi. Extremal problems, inequalities, and classical orthogonal polynomials. Applied Mathematics and Computation, 128(2– 3):151–166, May 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abdou:2003:FVIb

[AM03] M. A. Abdou and Osama L. Moustafa. Fredholm– Volterra integral equation in contact problem. Applied Mathematics and Computation, 138(2–3):199–215, June 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Amat:2003:NME

[Ama03] Sergio Amat. Nonseparable multiresolution with error control. Applied Mathematics and Computation, 145 (1):117–132, December 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Amemiya:2001:TCU

[Ame01] Takashi Amemiya. Triangular configuration of uncertain systems stabilizable by means of feedback controller. Applied Mathematics and Computation, 120 (1–3):45–54, May 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/104/21/
25/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/104/21/25/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002842.

Aminzadeh:2001:BTC

[Ami01] M. S. Aminzadeh. Bootstrap tolerance and confidence limits for two-variable reliability using independent and weakly dependent observations. Applied Mathematics and Computation, 122(1):81–93, July 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/106/21/
26/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/106/21/26/article.

                                        pdf;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300300000163.

Abdou:2003:NSF

[AMI03] M. A. Abdou, Khamis I. Mohamed, and A. S. Ismail. On the numerical solutions of Fredholm–Volterra integral equation. Applied Mathematics and Computation, 146(2–3):713–728, December 31, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ammar:2004:GCF

[Amm04] E. E. Ammar. On generalized convexity of fuzzy variables maps. Applied Mathematics and Computation, 157(1):65–75, September 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abdou:2003:NTS

[AN03] M. A. Abdou and A. A. Nasr. On the numerical treatment of the singular integral equation of the second kind. Applied Mathematics and Computation, 146(2–3): 373–380, December 31, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anastassakis:2001:GMD

[Ana01] E. Anastassakis. A geometrical method for diagonalizing real, symmetric 3 × 3 matrices through Euler rotations. Applied Mathematics and Computation, 117(2–3):193–201, January 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/92/23/25/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/92/23/
25/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399001733.

Ananda:2003:CIS

[Ana03] Malwane M. A. Ananda.

Confidence intervals for steady state availability of a system with exponential operating time and lognormal repair time. Applied Mathematics and Computation, 137(2–3):499–509, May 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Alvarez-Nodarse:2002:DZD

[ÁND02] R. Álvarez-Nodarse and Jesús S. Dehesa. Distributions of zeros of discrete and continuous polynomials from their recurrence relation. Applied Mathematics and Computation, 128(2– 3):167–190, May 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2000:Ia

[Ano00a] Anonymous. Index. Applied Mathematics and Computation, 107(2–3):235, January 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/72/22/
                                        33/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/72/22/33/article.
                                        pdf.

Anonymous:2000:Ib

[Ano00b] Anonymous. Index. Applied Mathematics and Computation, 108(2–3):197–198, February 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/79/22/
                                        32/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/79/22/32/article.
                                        pdf.

Anonymous:2000:Ic

[Ano00c] Anonymous. Index. Applied Mathematics and Computation, 109(2–3):307–308, March 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/82/23/
                                        36/abstract.html;   http:
                                        //www.elsevier.nl/gej-

ng/29/17/20/82/23/36/article.
pdf.

Anonymous:2000:Id

[Ano00d] Anonymous. Index. Applied Mathematics and Computation, 110(2–3):265, April 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/83/23/
31/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/83/23/31/article.
pdf.

Anonymous:2000:Ie

[Ano00e] Anonymous. Index. Applied Mathematics and Computation, 110(2–3):267–277, April 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/83/23/
32/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/83/23/32/article.
pdf.

Anonymous:2000:If

[Ano00f] Anonymous. Index. Applied Mathematics and Computation, 111(2–3):231–232, May 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/85/23/
29/abstract.html;   http:

                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/85/23/29/article.
                                        pdf.

Anonymous:2000:Ig

[Ano00g] Anonymous. Index. Applied Mathematics and Computation, 112(2–3):347–349, June 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/86/23/
                                        37/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/86/23/37/article.
                                        pdf.

Anonymous:2000:Ih

[Ano00h] Anonymous. Index. Applied Mathematics and Computation, 113(2–3):327–328, July 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/87/23/33/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/87/23/
                                        33/article.pdf.

Anonymous:2000:Ii

[Ano00i] Anonymous. Index. Applied Mathematics and Computation, 114(2–3):329–331, September 11, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/
                                        10/9/12/89/22/36/abstract.

html;   http://www.elsevier.
nl/gej-ng/10/9/12/89/22/
36/article.pdf.

Anonymous:2000:Ij

[Ano00j] Anonymous. Index. Applied Mathematics and Computation, 115(2–3):231–232, October 27, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/90/23/32/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/90/23/
32/article.pdf.

Anonymous:2000:Ik

[Ano00k] Anonymous. Index. Applied Mathematics and Computation, 116(3):307–308, December ??, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/91/23/26/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/91/23/
26/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300300000485.

Anonymous:2001:AIa

[Ano01a] Anonymous. Author index.

Applied Mathematics and Computation, 118(2–3):349– 350, March 9, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/

                                        9/12/96/25/34/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/96/25/
                                        34/article.pdf;   http:
                                        //www.sciencedirect.com/
                                        science/article/pii/S0096300301000029.

Anonymous:2001:AIb

[Ano01b] Anonymous. Author index. Applied Mathematics and Computation, 119(2–3):349– 350, April 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/99/25/36/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/99/25/
                                        36/article.pdf;   http:
                                        //www.sciencedirect.com/
                                        science/article/pii/S0096300301000352.

Anonymous:2001:AIc

[Ano01c] Anonymous. Author index. Applied Mathematics and Computation, 120(1– 3):313–314, May 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/9/12/104/21/
                                        45/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/9/12/104/21/45/article.
                                        pdf;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300301000662.

Anonymous:2001:AId

[Ano01d] Anonymous. Author index. Applied Mathematics and Computation, 121(2– 3):397–399, June 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/105/25/
40/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/105/25/40/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300301001060.

Anonymous:2001:AIe

[Ano01e] Anonymous. Author index. Applied Mathematics and Computation, 122(3): 407–408, August 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/106/26/
31/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/106/26/31/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300301001436.

Anonymous:2001:AIf

[Ano01f] Anonymous. Author index. Applied Mathematics and Computation, 123(3): 421–422, October 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/9/12/110/32/35/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300301002028.

Anonymous:2001:AIg

[Ano01g] Anonymous. Author index. Applied Mathematics and Computation, 124(3):413–414, December 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        9/12/113/32/39/abstract.
                                        html;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300301003083.

Anonymous:2001:I

[Ano01h] Anonymous. Index. Applied Mathematics and Computation, 117(2–3):351–352, January 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/92/23/36/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/92/23/
                                        36/article.pdf;   http:
                                        //www.sciencedirect.com/
                                        science/article/pii/S009630030000103X.

Anonymous:2001:MAI

[Ano01i] Anonymous. Master author index. Applied Mathematics and Computation, 120(1– 3):315–328, May 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/9/12/104/21/
                                        46/abstract.html;   http:
                                        //www.elsevier.nl/gej-

ng/10/9/12/104/21/46/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300301000674.

Anonymous:2002:AIa

[Ano02a] Anonymous. Author index. Applied Mathematics and Computation, 125(2–3): 423–424, January 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/9/12/120/31/48/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300301003095.

Anonymous:2002:AIb

[Ano02b] Anonymous. Author index. Applied Mathematics and Computation, 126(2–3): 389–390, March 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/9/12/122/30/43/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300301003411.

Anonymous:2002:AIc

[Ano02c] Anonymous. Author index.

Applied Mathematics and Computation, 127(2–3):415– 416, April 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/123/27/48/abstract.
html;   http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300302000048.

Anonymous:2002:AId

[Ano02d] Anonymous. Author index. Applied Mathematics and Computation, 128(2– 3):415–416, May 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2002:AIe

[Ano02e] Anonymous. Author index. Applied Mathematics and Computation, 129(2– 3):551–553, July 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2002:AIf

[Ano02f] Anonymous. Author index. Applied Mathematics and Computation, 130(2–3): 629–631, August 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2002:AIg

[Ano02g] Anonymous. Author index. Applied Mathematics and Computation, 131(2–3):615– 617, September 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2002:AIh

[Ano02h] Anonymous. Author index. Applied Mathematics and Computation, 132(2–3):649– 652, November 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2002:AIV

[Ano02i] Anonymous. Author index to volume. Applied Mathematics and Computation, 133(2–3):655–657, December 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2002:EB

[Ano02j]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 133(2– 3):CO2, December 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2002:EBP

[Ano02k]  Anonymous.  Editorial

Board page. Applied Mathematics and Computation, 133(1):ii, November 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2003:AIa

[Ano03a] Anonymous. Author index. Applied Mathematics and Computation, 134(2–3): 607–609, January 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2003:AIb

[Ano03b] Anonymous. Author index. Applied Mathematics and Computation, 135 (2–3):??, March 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2003:AIc

[Ano03c] Anonymous. Author index. Applied Mathematics and Computation, 136 (2–3):??, March 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2003:AId

[Ano03d] Anonymous. Author index. Applied Mathematics and Computation, 137(2– 3):595–598, May 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2003:AIe

[Ano03e] Anonymous. Author index. Applied Mathematics and Computation, 138(2– 3):555–557, June 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2003:AIf

[Ano03f] Anonymous. Author index. Applied Mathematics and Computation, 139(2– 3):555–557, July 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2003:AIg

[Ano03g] Anonymous. Author index. Applied Mathematics and Computation, 144(2–3): 569–571, December 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2003:AIVb

[Ano03h] Anonymous. Author index to volume. Applied Mathematics and Computation, 141(2–3):631–633, September 5, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2003:AIVc

[Ano03i] Anonymous. Author index to volume. Applied Mathematics and Computation, 142(2–3):585–587, October 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2003:AIVe

[Ano03j] Anonymous. Author index to volume. Applied Mathematics and Computation, 145(2–3):895–899, December 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2003:AIVf

[Ano03k] Anonymous. Author index to volume. Applied Mathematics and Computation, 146(2–3):897–900, December 31, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2003:AIVd

[Ano03l] Anonymous. Author index to volume 143. Applied Mathematics and Computation, 143(2–3):571–573, November 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2003:AIVa

[Ano03m] Anonymous. Author Index to Volumes 121–140. Applied Mathematics and Computation, 140(2–3):549–589, August 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2003:EBa

[Ano03n]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 134 (1):ii, January 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2003:EBb

[Ano03o]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 134(2– 3):CO2, January 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2003:EBc

[Ano03p]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 135 (1):ii, February 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2003:EBd

[Ano03q]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 135 (2–3):CO2, March 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2003:EBe

[Ano03r]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 136 (1):ii, March 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2003:EBf

[Ano03s]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 136 (2–3):CO2, March 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2003:EBg

[Ano03t]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 137 (1):ii, May 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2003:EBh

[Ano03u]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 138 (1):ii, June 1, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2003:EBi

[Ano03v]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 139 (1):ii, July 1, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2003:EBj

[Ano03w]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 140 (1):ii, July 30, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2003:EBk

[Ano03x]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 141 (1):i, August 20, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2003:EBl

[Ano03y]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 142 (1):ii, September 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2003:EBm

[Ano03z]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 143 (1):ii, October 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2003:EBn

[Ano03-27]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 144 (1):ii, November 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2003:EBo

[Ano03-28]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 145 (1):ii, December 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2003:EBp

[Ano03-29]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 146 (1):ii, December 30, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2004:AIVa

[Ano04a] Anonymous. Author index to volume. Applied Mathematics and Computation, 147(3):909–913, January 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2004:AIVf

[Ano04b] Anonymous. Author index to volume. Applied Mathematics and Computation, 151(3):903–907, April 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2004:AIVg

[Ano04c] Anonymous. Author index to volume. Applied Mathematics and Computation, 152(3):905–909, May 13, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2004:AIVj

[Ano04d] Anonymous. Author index to volume. Applied Mathematics and Computation, 155(3):883–887, August 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2004:AIVl

[Ano04e] Anonymous. Author index to volume. Applied Mathematics and Computation, 157(3):889–892, October 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2004:AIVm

[Ano04f] Anonymous. Author index to volume. Applied Mathematics and Computation, 158(3):887–891, November 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2004:AIVb

[Ano04g] Anonymous. Author index to volume 148. Applied Mathematics and Computation, 148(3):893–897, January 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2004:AIVc

[Ano04h] Anonymous. Author index to volume 149. Applied Mathematics and Computation, 149(3):901–905, February 22, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2004:AIVd

[Ano04i] Anonymous. Author index to volume 150. Applied Mathematics and Computation, 150(3):889–893, March 17, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2004:AIVh

[Ano04j] Anonymous. Author index to volume 153. Applied Mathematics and Computation, 153(3):903–908, June 14, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2004:AIVi

[Ano04k] Anonymous. Author index to volume 154 (2004). Applied Mathematics and Computation, 154(3):889– 893, July 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2004:AIVk

[Ano04l] Anonymous. Author index to volume 156. Applied Mathematics and Computation, 156(3):883–887, September 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2004:AIVn

[Ano04m] Anonymous. Author index to volume 159. Applied Mathematics and Computation, 159(3):903–907, December 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2004:AIVe

[Ano04n] Anonymous. Author index to volumes 141–150. Applied Mathematics and Computation, 150(3):895– 926, March 17, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2004:EBa

[Ano04o]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 147 (1):ii, January 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2004:EBb

[Ano04p]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 148 (1):ii, January 20, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2004:EBc

[Ano04q]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 149 (1):ii, February 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2004:EBd

[Ano04r]  Anonymous.  Editorial

board. Applied Mathematics and Computation, 150 (1):i, February 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2004:EBe

[Ano04s]  Anonymous.  Editorial

board. Applied Mathematics and Computation, 151 (1):ii, March 30, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2004:EBf

[Ano04t]  Anonymous.  Editorial

board. Applied Mathematics and Computation, 152 (1):ii, April 26, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2004:EBg

[Ano04u]  Anonymous.  Editorial

board. Applied Mathematics and Computation, 153 (1):ii, May 25, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2004:EBh

[Ano04v]  Anonymous.  Editorial

board. Applied Mathematics and Computation, 154 (1):ii, June 25, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2004:EBi

[Ano04w]  Anonymous.  Editorial

board. Applied Mathematics and Computation, 155 (1):ii, July 26, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2004:EBj

[Ano04x]  Anonymous.  Editorial

board. Applied Mathematics and Computation, 156 (1):ii, August 25, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2004:EBk

[Ano04y]  Anonymous.  Editorial

board. Applied Mathematics and Computation, 157 (1):ii, September 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2004:EBl

[Ano04z]  Anonymous.  Editorial

board. Applied Mathematics and Computation, 158 (1):ii, October 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2004:EBm

[Ano04-27]  Anonymous.  Editorial

board. Applied Mathematics and Computation, 159 (1):ii, November 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ahmad:2002:GQM

[ANS02] Bashir Ahmad, Juan J. Nieto, and Naseer Shahzad. Generalized quasilinearization method for mixed boundary value problems. Applied Mathematics and Computation, 133(2–3):423– 429, December 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Agarwal:2000:CSD

[AO00a] Ravi P. Agarwal and Donal O’Regan. A coupled system of difference equations. Applied Mathematics and Computation, 114(1):39–49, August 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/89/20/23/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/89/20/
23/article.pdf.

Agarwal:2000:MNS

[AO00b] Ravi P. Agarwal and Donal O’Regan. Multiple nonnegative solutions for second order impulsive differential equations. Applied Mathematics and Computation, 114(1):51–59, August 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.nl/gej-ng/
10/9/12/89/20/24/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/89/20/
24/article.pdf.

Agouzal:2000:PEE

[AO00c] Abdellatif Agouzal and Fabienne Oudin. A posteriori error estimator for finite volume methods. Applied Mathematics and Computation, 110(2–3):239–250, April 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/83/23/
29/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/83/23/29/article.
pdf.

Agarwal:2001:GPL

[AO01a] Ravi P. Agarwal and Donal O’Regan. A generalization of the Petryshyn-Leggett-Williams fixed point theorem with applications to integral inclusions. Applied Mathematics and Computation, 123(2):263–274, September 25, 2001. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/9/12/110/29/34/abstract.
html;   http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300300000771.

Alvarez:2001:IMS

[AO01b] Sixto Jesus Alvarez and Rachid Oujja. An iterative method for solving a free boundary problem for an infinite journal bearing. Applied Mathematics and Computation, 122 (1):15–26, July 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/9/12/106/21/
                                        22/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/9/12/106/21/22/article.
                                        pdf;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300300000084.

Aksan:2004:NSB

[AÖ04a] E. N. Aksan and A. Özdeş. A numerical solution of Burgers’ equation. Applied Mathematics and Computation, 156(2):395–402, September 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Al-Odat:2004:TND

[AO04b] M. Q. Al-Odat. Transient non-Darcy mixed convection along a vertical surface in porous medium with suction or injection. Applied Mathematics and Computation, 156(3):679–694, September 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Alvarez:2004:NNA

[AO04c] Sixto J. Alvarez and Rachid Oujja. A new numerical approach of a lubrication free boundary problem. Applied Mathematics and Computation, 148(2): 393–405, January 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Aksan:2004:VAP

[AÖE04] E. N. Aksan, A. Özdeş, and A. Esen. A variational approximation to the problem of the deflection of a bar. Applied Mathematics and Computation, 150 (2):525–531, March 8, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ascenzi:2004:CQF

[APS04] Oscar Ascenzi, Lorenzo Pareschi, and Fausto Segala. Convergence of a quadrature formula for the approximation of stress intensity factor for planar cracks. Applied Mathematics and Computation, 158(3):597– 617, November 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Allan:2004:DGM

[AQBT04] Fathi M. Allan, Naji Qatanani, Imad Barghouthi, and Khaled M. Takatka. Dusty gas model of flow through naturally occurring porous media. Applied Mathematics and Computation, 148(3):809–821, January 30, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Alonso-Quesada:2004:RAC

[AQD04] S. Alonso-Quesada and M. De la Sen. Robust adaptive control of discrete nominally stabilizable plants. Applied Mathematics and Computation, 150(2):555– 583, March 8, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Alvarez:2002:ICG

[ÁR02] Jorge Álvarez and Jesús Rojo. An improved class of generalized Runge–Kutta methods for stiff problems. Part I: The scalar case. Applied Mathematics and Computation, 130(2–3):537–560, August 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abdel-Rahman:2003:PBI

[AR03] Reda G. Abdel-Rahman. Propagation of boundary of inhomogeneous heat conduction equation. Applied Mathematics and Computation, 141(2–3):231–239, September 5, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abdel-Rahman:2004:FNN

[AR04a] Gamal M. Abdel-Rahman.

Flow of a non-Newtonian power law through a conical bearing in an applied magnetic field. Applied Mathematics and Computation, 159(1):237–246, November 25, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abdel-Rahman:2004:FFT

[AR04b] Gamal M. Abdel-Rahman.

The fluid flow in the thin films between the immobile conic surface. Applied Mathematics and Computation, 153(1):59–67, May 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Alvarez:2004:ICG

[AR04c] Jorge Alvarez and Jesús Rojo. An improved class of generalized Runge–Kutta methods for stiff problems. Part II: The separated system case. Applied Mathematics and Computation, 159(3):717–758, December 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anilkumar:2004:UMC

[AR04d] D. Anilkumar and S. Roy. Unsteady mixed convection flow on a rotating cone in a rotating fluid. Applied Mathematics and Computation, 155(2):545–561, August 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Arulmozhi:2003:DMR

[Aru03]  G. Arulmozhi.  Direct

method for reliability computation of k-out-of-n: G systems. Applied Mathematics and Computation, 143(2–3):421–429, November 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Adams:2001:MME

[AS01] John C. Adams and Piotr Smolarkiewicz. Modified multigrid for 3D elliptic equations with crossderivatives. Applied Mathematics and Computation, 121(2–3):301–312, June 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/9/12/105/25/
                                        34/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/9/12/105/25/34/article.

pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000047.

Abo-Seida:2002:AEF

[AS02a] Osama M. Abo-Seida. The analyticity of the electromagnetic field in an isotropic medium. Applied Mathematics and Computation, 127(2–3):361–364, April 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/123/27/43/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300301000145.

Abo-Seida:2002:USB

[AS02b] Osama M. Abo-Seida. Uniqueness solution for the boundary value problem defined by specifying the components of the electromagnetic field. Applied Mathematics and Computation, 132(2–3):553– 558, November 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abo-Seida:2003:TBM

[AS03a] Osama Abo-Seida. The transient behaviour of the magnetic field strength at any distance above the duct. Applied Mathematics and Computation, 138(1):91–98, June 1, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abo-Seida:2003:PEF

[AS03b] Osama M. Abo-Seida. Propagation of electromagnetic field in an abnormal stratification of the upper ionosphere. Applied Mathematics and Computation, 142(2–3): 409–416, October 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abo-Seida:2003:TFV

[AS03c] Osama M. Abo-Seida. Transient fields of a vertical electric dipole on an M-layered dielectric medium. Applied Mathematics and Computation, 145(2–3):289–296, December 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Akyuz:2003:CPS

[AS03d] Ayşegül Akyüz and Mehmet Sezer. Chebyshev polynomial solutions of systems of high-order linear differential equations with variable coefficients. Applied Mathematics and Computation, 144(2–3):237–247, December 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abdou:2004:IEC

[AS04a] M. A. Abdou and F. A.

Salama. Integral equation and contact problem. Applied Mathematics and Computation, 149(3):735–746, February 22, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abdou:2004:VFI

[AS04b] M. A. Abdou and F. A.

Salama. Volterra–Fredholm integral equation of the first kind and spectral relationships. Applied Mathematics and Computation, 153 (1):141–153, May 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abo-Seida:2004:EEF

[AS04c] Osama M. Abo-Seida. Estimation of the electromagnetic field created at the Earth’s surface by an overhead line current. Applied Mathematics and Computation, 149(2):559–565, February 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abo-Sinna:2004:MOF

[AS04d] Mahmoud A. Abo-Sinna.

Multiple objective (fuzzy) dynamic programming problems: a survey and some applications. Applied Mathematics and Computation, 157(3):861–888, October 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Al-Saleh:2004:TRS

[AS04e] Mohammad Fraiwan Al-Saleh. On the totality of ranked set sampling. Applied Mathematics and Computation, 147(2):527–535, January 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Al-Seedy:2004:QFV

[AS04f] R. O. Al-Seedy. Queuing with fixed and variable channels considering balking and reneging concepts. Applied Mathematics and Computation, 156(3):755– 761, September 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Al-Seedy:2004:TSN

[AS04g] R. O. Al-Seedy. A transient solution of the nontruncated queue M/M/2 with balking, and an additional server for longer queues (Krishnamoorthi discipline). Applied Mathematics and Computation, 156 (3):763–769, September 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anabtawi:2004:SCL

[AS04h] M. J. Anabtawi and S. Sathananthan. Stability and convergence via Lyapunovlike functionals of stochastic parabolic partial differential equations. Applied Mathematics and Computation, 157(1):201–218, September 27, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Asaithambi:2004:NSF

[Asa04a] Asai Asaithambi. Numerical solution of the Falkner– Skan equation using piecewise linear functions. Applied Mathematics and Computation, 159(1):267–273, November 25, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Asaithambi:2004:SOF

[Asa04b] Asai Asaithambi. A secondorder finite-difference method for the Falkner–Skan equation. Applied Mathematics and Computation, 156 (3):779–786, September 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Asaithambi:2004:UAD

[Asa04c] Asai Asaithambi. Using automatic differentiation to solve concentration dependent diffusion problems. Applied Mathematics and Computation, 159(3):641–650, December 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Al-Saleh:2003:ESP

[ASAB03] Mohammad F. Al-Saleh and Fatima K. Al-Batainah. Estimation of the shape parameter k of the negative binomial distribution. Applied Mathematics and Computation, 143(2–3):431–441, November 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Al-Seedy:2003:NTS

[ASAI03] Ragab Omarah Al-Seedy and Fawziah M. Al-Ibraheem. New transient solution to the M/M/∞ queue with varying arrival and departure rate. Applied Mathematics and Computation, 135(2–3):425–428, March 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abo-Seida:2003:SEF

[ASEM03] Osama M. Abo-Seida and Ahmed I. El-Maghrabi. The solution of electromagnetic fields by equivalent magnetic network. Applied Mathematics and Computation, 135(1):179–185, February 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Al-Sharif:2004:GTP

[ASK04] Sh. Al-Sharif and R. Khalil.

On the generator of two parameter semigroups. Applied Mathematics and Computation, 156(2):403–414, September 6, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Al-Saqabi:2003:UPD

[ASKT03] B. N. Al-Saqabi, S. L.

Kalla, and V. K. Tuan. Unified probability density function involving a confluent hypergeometric function of two variables. Applied Mathematics and Computation, 146(1):135–152, December 30, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Al-Saleh:2003:NPE

[ASM03] M. Fraiwan Al-Saleh and Fawaz A. Masoud. A note on the posterior expected loss as a measure of accuracy in Bayesian methods. Applied Mathematics and Computation, 134(2–3): 507–514, January 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Al-Said:2003:CSM

[ASN03] E. A. Al-Said and M. A. Noor. Cubic splines method for a system of third-order boundary value problems. Applied Mathematics and Computation, 142(2–3):195– 204, October 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Al-Showaikh:2004:ODM

[AST04] F. N. M. Al-Showaikh and E. H. Twizell. Onedimensional measles dynamics. Applied Mathematics and Computation, 152(1): 169–194, April 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abdurahman:2004:PNS

[AT04] Xamxinur Abdurahman and Zhidong Teng. On the persistence of a nonautonomous n-species Lotka– Volterra cooperative system. Applied Mathematics and Computation, 152 (3):885–895, May 13, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Angelova:2004:SPD

[AV04a] Ivanka Tr. Angelova and Lubin G. Vulkov. Singularly perturbed differential equations with discontinuous coefficients and concentrated factors. Applied Mathematics and Computation, 158 (3):683–701, November 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Avudainayagam:2004:WBM

[AV04b] A. Avudainayagam and C. Vani. Wavelet based multigrid methods for linear and nonlinear elliptic partial differential equations. Applied Mathematics and Computation, 148(2):307– 320, January 26, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abu-Youssef:2003:CCD

[AY03] S. E. Abu-Youssef. On characterization of certain distributions of record values. Applied Mathematics and Computation, 145(2–3): 443–450, December 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abu-Youssef:2004:MIN

[AY04a] S. E. Abu-Youssef. Moment inequality on new renewal better than used class of life distributions with hypothesis testing application. Applied Mathematics and Computation, 149(3): 651–659, February 22, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abu-Youssef:2004:NTM

[AY04b] S. E. Abu-Youssef. Nonparametric test for monotone variance residual life class of life distributions with hypothesis testing applications. Applied Mathematics and Computation, 158 (3):817–826, November 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ayaz:2003:TDD

[Aya03]  Fatma Ayaz.  On the

two-dimensional differential transform method. Applied Mathematics and Computation, 143(2–3):361–374, November 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ayaz:2004:ADT

[Aya04a]  Fatma Ayaz.  Applica-

tions of differential transform method to differentialalgebraic equations. Applied Mathematics and Computation, 152(3):649–657, May 13, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ayaz:2004:SSD

[Aya04b]  Fatma Ayaz.  Solutions

of the system of differential equations by differential transform method. Applied Mathematics and Computation, 147(2):547–567, January 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Akdeniz:2004:MOA

[AYW04] Fikri Akdeniz, Güzin Yüksel, and Alan T. K. Wan. The moments of the operational almost unbiased ridge regression estimator. Applied Mathematics and Computation, 153(3):673–684, June 14, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ahmed:2002:TSP

[AZ02] S. M. Ahmed and N. A. Zeiden. Thermal stresses problem in non-homogeneous transversely isotropic infinite circular cylinder. Applied Mathematics and Computation, 133(2–3):337–350, December 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Akhmet:2004:CTP

[AZ04] M. Akhmet and A. Zafer.

Controllability of two-point nonlinear boundary-value problems by the numericalanalytic method. Applied Mathematics and Computation, 151(3):729–744, April 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Al-Zamel:2000:GET

[AZTK00] A. Al-Zamel, V. K. Tuan, and S. L. Kalla. Generalized elliptic-type integrals and asymptotic formulas. Applied Mathematics and Computation, 114(1):13–25, August 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/89/20/21/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/89/20/
21/article.pdf.

Barbat:2004:PDT

[BA04] Tiberiu Barbat and Nasser Ashgriz. Planar dynamics of two interacting bubbles in an acoustic field. Applied Mathematics and Computation, 157(3):775– 824, October 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Braselton:2001:SWQ

[BAB01] James P. Braselton, Martha L. Abell, and Lorraine M. Braselton. Schryer-Walker quasi-exact solutions to the Landau–Lifshitz–Gilbert equations. Applied Mathematics and Computation, 124(2):151–167, October 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.com/
gej-ng/10/9/12/113/31/
28/abstract.html;   http:
//www.sciencedirect.com/
science/article/pii/S0096300300000837.

Badr:2001:SPM

[Bad01] A. A. Badr. On some parameter methods for nonlinear Volterra integral equation. Applied Mathematics and Computation, 117(1):15–22, January 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/92/20/21/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/92/20/
21/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399001848.

Bae:2004:UDS

[Bae04a] Jeong Ja Bae. On uniform decay of the solution for a damped nonlinear coupled system of wave equations with nonlinear boundary damping and memory term. Applied Mathematics and Computation, 148(1): 207–223, January 20, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bae:2004:UDU

[Bae04b] Jeong Ja Bae. Uniform decay for the unilateral problem associated to the Kirchhoff type wave equations with nonlinear boundary damping. Applied Mathematics and Computation, 156(1):41–57, August 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bahadir:2003:FIF

[Bah03] A. Refik Bahadır. A fully implicit finite-difference scheme for two-dimensional Burgers’ equations. Applied Mathematics and Computation, 137(1):131–137, May 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bahadir:2004:ACB

[Bah04] A. R. Bahadır. Application of cubic B-spline finite element technique to the thermistor problem. Applied Mathematics and Computation, 149(2):379–387, February 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bai:2000:SEB

[Bai00] Zhong-Zhi Bai. Sharp error bounds of some Krylov subspace methods for non-Hermitian linear systems. Applied Mathematics and Computation, 109(2–3):273– 285, March 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/82/23/
33/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/82/23/33/article.
pdf.

Bao:2002:RPM

[Bao02] Weizhu Bao. The random projection method for a model problem of combustion with stiff chemical reactions. Applied Mathematics and Computation, 130(2– 3):561–571, August 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Barrio:2003:CAP

[Bar03] Roberto Barrio. Compression of almost-periodic data. Applied Mathematics and Computation, 134(2–3):431– 444, January 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Barrio:2004:AID

[Bar04] R. Barrio. Algorithms for the integration and derivation of Chebyshev series. Applied Mathematics and Computation, 150(3):707– 717, March 17, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Batir:2004:IRS

[Bat04] Necdet Batir. Integral representations of some series involving \binom{2k}{k^{−1}k^{−n}} and some related series. Applied Mathematics and Computation, 147(3):645–667, January 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bayram:2004:ACA

[Bay04] Mustafa Bayram. Application of computer algebra matrix operation techniques to the control of metabolic networks. Applied Mathematics and Computation, 152 (1):289–297, April 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Babolian:2002:OCA

[BB02a] E. Babolian and J. Biazar. On the order of convergence of Adomian method. Applied Mathematics and Computation, 130(2–3):383–387, August 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Babolian:2002:SNE

[BB02b] E. Babolian and J. Biazar. Solution of nonlinear equations by modified Adomian decomposition method. Applied Mathematics and Computation, 132 (1):167–172, October 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Babolian:2002:SPB

[BB02c] E. Babolian and J. Biazar. Solving the problem of biological species living together by Adomian decomposition method. Applied Mathematics and Computation, 129(2–3):339–343, July 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Babolian:2003:SCE

[BB03] E. Babolian and J. Biazar.

Solving concrete examples by Adomian method. Applied Mathematics and Computation, 135(1):161–167, February 15, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Benkherouf:2003:DIM

[BBA03] Lakdere Benkherouf, Amin Boumenir, and Lakhdar Aggoun. A diffusion inventory model for deteriorating items. Applied Mathematics and Computation, 138(1):21–39, June 1, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Biazar:2003:SSV

[BBI03] J. Biazar, E. Babolian, and R. Islam. Solution of a system of Volterra integral equations of the first kind by Adomian method. Applied Mathematics and Computation, 139(2–3):249–258, July 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Biazar:2004:SSO

[BBI04] J. Biazar, E. Babolian, and R. Islam. Solution of the system of ordinary differential equations by Adomian decomposition method. Applied Mathematics and Computation, 147(3):713–719, January 16, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Biazar:2003:AAC

[BBK^{+}03] J. Biazar, E. Babolian, G. Kember, A. Nouri, and R. Islam. An alternate algorithm for computing Adomian polynomials in special cases. Applied Mathematics and Computation, 138(2– 3):523–529, June 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bourchtein:2004:CVD

[BBN04] Andrei Bourchtein, Ludmila Bourchtein, and Maxim Naumov. On correctness of the vertical discretization in numerical weather prediction models. Applied Mathematics and Computation, 158(2):513–527, November 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Babolian:2004:DMA

[BBV04a] E. Babolian, J. Biazar, and A. R. Vahidi. The decomposition method applied to systems of Fredholm integral equations of the second kind. Applied Mathematics and Computation, 148(2): 443–452, January 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Babolian:2004:NCM

[BBV04b] E. Babolian, J. Biazar, and A. R. Vahidi. A new computational method for Laplace transforms by decomposition method. Applied Mathematics and Computation, 150(3):841–846, March 17, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Babolian:2004:DMS

[BBV04c] E. Babolian, J. Biazar, and A. R. Vahidi. On the decomposition method for system of linear equations and system of linear Volterra integral equations. Applied Mathematics and Computation, 147(1):19–27, January 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Babolian:2004:SSN

[BBV04d] E. Babolian, J. Biazar, and A. R. Vahidi. Solution of a system of nonlinear equations by Adomian decomposition method. Applied Mathematics and Computation, 150(3):847–854, March 17, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bayram:2002:SSP

[BÇ02] Mustafa Bayram and Ercan Çelik. Simultaneous solution of polynomial equations. Applied Mathematics and Computation, 133(2–3): 533–538, December 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Boulbrachene:2004:FEA

[BC04] M. Boulbrachene and B. Chentouf. The finite element approximation of Hamilton– Jacobi–Bellman equations: the noncoercive case. Applied Mathematics and Computation, 158(2):585–592, November 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Barteneva:2003:MAM

[BCI03] Irina V. Barteneva, Alberto Cabada, and Alexander O. Ignatyev. Maximum and anti-maximum principles for the general operator of second order with variable coefficients. Applied Mathematics and Computation, 134(1):173–184, January 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bru:2004:GMB

[BCT04] Rafael Bru, Carmen Coll, and Néstor Thome. Gramian matrices and balanced model of generalized systems. Applied Mathematics and Computation, 148(2):341–350, January 26, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Boglaev:2003:DDA

[BD03] Igor Boglaev and Vic Duoba.

Domain decomposition for an advection–diffusion problem with parabolic layers. Applied Mathematics and Computation, 146(1):27–53, December 30, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Babolian:2004:NIA

[BD04a] E. Babolian and A. Davari.

Numerical implementation of Adomian decomposition method. Applied Mathematics and Computation, 153 (1):301–305, May 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bakirta:2004:MNW

[BD04b] İkay Bakirta and Hilmi Demıray. Modulation of nonlinear waves near the marginal state of instability in fluid-filled elastic tubes. Applied Mathematics and Computation, 149(1): 83–101, February 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bakirtas:2004:AMN

[BD04c] İlkay Bakirtaş and Hilmi Demiray. Amplitude modulation of nonlinear waves in a fluid-filled tapered elastic tube. Applied Mathematics and Computation, 154 (3):747–767, July 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bainov:2000:OBS

[BDD00] D. D. Bainov, M. B. Dimitrova, and A. B. Dishliev. Oscillation of the bounded solutions of impulsive differential-difference equations of second order. Applied Mathematics and Computation, 114(1):61–68, August 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/89/20/25/abstract.

html;   http://www.elsevier.
nl/gej-ng/10/9/12/89/20/
25/article.pdf.

Bartuccelli:2004:GAO

[BDGG04] M. V. Bartuccelli, J. H. B.

Deane, G. Gentile, and S. A. Gourley. Global attraction to the origin in a parametrically driven nonlinear oscillator. Applied Mathematics and Computation, 153(1):1–11, May 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Balachandran:2003:CNE

[BDS03] K. Balachandran, J. P.

Dauer, and S. Sangeetha. Controllability of nonlinear evolution delay integrodifferential systems. Applied Mathematics and Computation, 139(1):63–84, July 1, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bahadir:2002:PCD

[BE02] A. Refik Bahadir and F. Brian Ellerby. On the performance of certain direct and iterative methods on equations arising on a two-dimensional in situ combustion simulator. Applied Mathematics and Computation, 125(2–3):347–358, January 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.

                                        elsevier.com/gej-ng/10/
                                        9/12/120/31/43/abstract.
                                        html;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300300001363.

Bulut:2004:NSV

[BEAB04] Hasan Bulut, Mahmut Ergüt, Vedat Asil, and Roza Horvath Bokor. Numerical solution of a viscous incompressible flow problem through an orifice by Adomian decomposition method. Applied Mathematics and Computation, 153 (3):733–741, June 14, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bell:2001:AML

[Bel01] Bradley M. Bell. Approximating the marginal likelihood estimate for models with random parameters. Applied Mathematics and Computation, 119(1):57–75, March 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/99/20/24/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/99/20/
                                        24/article.pdf;   http:
                                        //www.sciencedirect.com/
                                        science/article/pii/S0096300399002301.

Beliakov:2004:LSS

[Bel04]  Gleb Beliakov.  Least

squares splines with free knots: global optimization approach. Applied Mathematics and Computation, 149(3):783–798, February 22, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Benjumea:2002:NPF

[BEV02] Juan C. Benjumea, Francisco J. Echarte, and Juan Núñez Valdes. New properties of filiform Lie algebras and its computational processing. Applied Mathematics and Computation, 130(1):53–62, July 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bocso:2003:PER

[BF03] A. Bocsó and M. Farkas. Political and economic rationality leads to velcro bifurcation. Applied Mathematics and Computation, 140(2– 3):381–389, August 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Baran:2004:DUS

[BF04] Emine Can Baran and Afet Golayoğlu Fatullayev. Determination of an unknown source parameter in two-dimensional heat equation. Applied Mathematics and Computation, 159(3): 881–886, December 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Berenguer:2004:LVI

[BFGG04] M. I. Berenguer, M. A. Fortes, A. I. Garralda Guillem, and M. Ruiz Galán. Linear Volterra integrodifferential equation and Schauder bases. Applied Mathematics and Computation, 159(2):495–507, December 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Boza:2001:NMC

[BFN01] Luis Boza, Eugenio M. Fedriani, and Juan Núñez. A new method for classifying complex filiform Lie algebras. Applied Mathematics and Computation, 121(2– 3):169–175, June 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/105/25/
25/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/105/25/25/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002702.

Boza:2003:CFL

[BFN03] Luis Boza, Eugenio M. Fedriani, and Juan Núñez. Complex filiform Lie algebras of dimension 11. Applied Mathematics and Computation, 141(2–3):611–630, September 5, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bennett:2004:APD

[BG04] D. L. Bennett and S. A.

Gourley. Asymptotic properties of a delay differential equation model for the interaction of glucose with plasma and interstitial insulin. Applied Mathematics and Computation, 151 (1):189–207, March 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bu:2004:GGL

[BGT04] Charles Bu, Hongjun Gao, and Kimitoshi Tsutaya. The generalized Ginzburg– Landau equation: posed in a quarter plane. Applied Mathematics and Computation, 159(3):667–674, December 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bultheel:2002:RSM

[BGVHN02] A. Bultheel, P. Gonzalez-Vera, E. Hendriksen, and Olav Njåstad. A rational Stieltjes moment problem. Applied Mathematics and Computation, 128(2– 3):217–235, May 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Buchanan:2003:TRT

[BGWX03] James L. Buchanan, Robert P. Gilbert, Armand Wirgin, and Yongzhi Xu. Transient reflection and transmission of ultrasonic wave in cancellous bone. Applied Mathematics and Computation, 142(2–3):561–573, October 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bainov:2001:MQP

[BH01] D. D. Bainov and S. G. Hristova. The method of quasilinearization for the periodic boundary value problem for systems of impulsive differential equations. Applied Mathematics and Computation, 117(1):73–85, January 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/92/20/26/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/92/20/
26/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399001563.

Babolian:2002:MSM

[BH02] E. Babolian and M. M. Hosseini. A modified spectral method for numerical solution of ordinary differential equations with non-analytic solution. Applied Mathematics and Computation, 132(2–3):341–351, November 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Babolian:2003:RIP

[BH03] E. Babolian and M. M.

Hosseini. Reducing index, and pseudospectral methods for differential-algebraic equations. Applied Mathematics and Computation, 140(1):77–90, July 30, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ben-Haim:2002:GMC

[BHH02] Yakov Ben-Haim and Keith W.

Hipel. The graph model for conflict resolution with information-gap uncertainty in preferences. Applied Mathematics and Computation, 126(2–3):319–340, March 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/122/30/39/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001612.

Biazar:2004:SWE

[BI04a] J. Biazar and R. Islam.

Solution of wave equation by Adomian decomposition method and the restrictions of the method. Applied Mathematics and Computation, 149(3):807–814, February 22, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Buyukyazici:2004:APG

[Bİ04b] İbrahim Büyükyazıcı and Ertan İbikli. The approximation properties of generalized Bernstein polynomials of two variables. Applied Mathematics and Computation, 156(2):367–380, September 6, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bilgin:2003:MTC

[Bil03] Tunay Bilgin. Matrix transformation on certain sequence spaces. Applied Mathematics and Computation, 146(2–3):433–436, December 31, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bilgin:2004:LSC

[Bil04] Tunay Bilgin. Lacunary strong A-convergence with respect to a sequence of modulus functions. Applied Mathematics and Computation, 151(3):595–600, April 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See corrigendum [Bil07].

Bilgin:2007:CLS

[Bil07] Tunay Bilgin. Corrigendum to “Lacunary strong Aconvergence with respect to a sequence of modulus functions” [Appl. Math. Comput. 151 (3) (2004) 595– 600]. Applied Mathematics and Computation, 185 (1):755, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [Bil04].

Biswas:2003:TNK

[Bis03]  Anjan Biswas.  Theory

of non-Kerr law generalized vector solitons. Applied Mathematics and Computation, 136(2–3):443–452, March 15, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Biswas:2004:ADN

[Bis04a] Anjan Biswas. Adiabatic dynamics of non-Kerr law solitons. Applied Mathematics and Computation, 151 (1):41–52, March 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Biswas:2004:SMO

[Bis04b] Anjan Biswas. Solitons in magneto-optic waveguides. Applied Mathematics and Computation, 153(2):387– 393, June 4, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Biswas:2004:TNK

[Bis04c]  Anjan Biswas.  Theory

of non-Kerr law solitons. Applied Mathematics and Computation, 153(2):369– 385, June 4, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Babolian:2003:RAM

[BJ03] E. Babolian and Sh. Javadi. Restarted Adomian method for algebraic equations. Applied Mathematics and Computation, 146(2–3):533–541, December 31, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Babolian:2004:NMC

[BJ04] E. Babolian and Sh. Javadi. New method for calculating Adomian polynomials. Applied Mathematics and Computation, 153(1):253– 259, May 25, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Boswell:2003:PNS

[BJD^{+}03] Graeme P. Boswell, Helen Jacobs, Fordyce A. Davidson, Geoffrey M. Gadd, and Karl Ritz. A positive numerical scheme for a mixed-type partial differential equation model for fungal growth.

Applied Mathematics and Computation, 138(2–3):321– 340, June 20, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Babolian:2004:RAM

[BJS04] E. Babolian, Sh. Javadi, and H. Sadeghi. Restarted Adomian method for integral equations. Applied Mathematics and Computation, 153(2):353–359, June 4, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bonet:2002:SAE

[BK02] Javier Bonet and Sivakumar Kulasegaram. A simplified approach to enhance the performance of smooth particle hydrodynamics methods. Applied Mathematics and Computation, 126(2–3):133–155, March 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/122/30/27/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001430.

Balser:2003:RPT

[BK03] Werner Balser and Vladimir Kostov. Recent progress in the theory of formal solutions for ODE and PDE. Applied Mathematics and Computation, 141(1):113– 123, August 20, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bae:2004:FEA

[BK04] Hyeong-Ohk Bae and Do Wan Kim. Finite element approximations for the Stokes equations on curved domains, and their errors. Applied Mathematics and Computation, 148(3):823–847, January 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bellman:2001:VWM

[BL01] Kirstie L. Bellman and Christopher Landauer. Virtual worlds as meeting places for formal systems. Applied Mathematics and Computation, 120(1–3):165– 173, May 10, 2001. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/104/21/
34/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/104/21/34/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002398.

Bhaskar:2002:MIT

[BM02] T. G. Bhaskar and Farzana A. McRae. Monotone iterative techniques for nonlinear problems involving the difference of two monotone functions. Applied Mathematics and Computation, 133(1):187–192, November 25, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Belleni-Morante:2004:NAI

[BMMRS04] A. Belleni-Morante, R. Monaco, R. Riganti, and F. Salvarani. A numerical approach for inverse problems in photon transport inside an interstellar cloud. Applied Mathematics and Computation, 154(1):115– 126, June 25, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Barrett:2000:ETS

[BMR00] C. L. Barrett, H. S.

Mortveit, and C. M. Reidys. Elements of a theory of simulation II: sequential dynamical systems. Applied Mathematics and Computation, 107(2–3):121–136, January 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/72/22/
26/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/72/22/26/article.
pdf.

Barrett:2001:ETS

[BMR01] C. L. Barrett, H. S. Mortveit, and C. M. Reidys. Elements of a theory of simulation III: equivalence of SDS. Applied Mathematics and Computation, 122(3): 325–340, August 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/9/12/106/26/
                                        24/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/9/12/106/26/24/article.
                                        pdf;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300300000424.

Barrett:2003:EIS

[BMR03] C. L. Barrett, H. S. Mortveit, and C. M. Reidys. ETS IV: Sequential dynamical systems: fixed points, invertibility and equivalence. Applied Mathematics and Computation, 134(1):153– 171, January 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Boyd:2002:SHP

[BN02] John P. Boyd and Andrei Natarov. Shafer (Hermite–Padé) approximants for functions with exponentially small imaginary part with application to equatorial waves with critical latitude. Applied Mathematics and Computation, 126(1):109–117, February 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/122/27/32/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001417.

Boglaev:2004:UNM

[Bog04] Igor Boglaev. Uniform numerical methods on arbitrary meshes for singularly perturbed problems with discontinuous data. Applied Mathematics and Computation, 154(3):815–833, July 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bonham:2002:SHV

[Bon02] Charles D. Bonham. Selection of hydrocarbon variables to assess reduction of residual oil on nutrient enriched beaches. Applied Mathematics and Computation, 126(2–3):361–376, March 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/122/30/41/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001685.

Bourchtein:2002:ENM

[Bou02] Andrei Bourchtein. Ellipticity of normal mode initialization equations. Applied Mathematics and Computation, 133(1):193–211, November 25, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Boyd:2003:CPE

[Boy03a] John P. Boyd. Chebyshev polynomial expansions for simultaneous approximation of two branches of a function with application to the onedimensional Bratu equation. Applied Mathematics and Computation, 143(2–3):189– 200, November 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Boyd:2003:LMN

[Boy03b] John P. Boyd. Large mode number eigenvalues of the prolate spheroidal differential equation. Applied Mathematics and Computation, 145(2–3):881–886, December 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Barco:2001:NSS

[BP01] M. A. Barco and G. E. Prince. New symmetry solution techniques for firstorder non-linear PDEs. Applied Mathematics and Computation, 124(2):169–196, October 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/113/31/29/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000850.

Bae:2003:UDS

[BPJ03] Jeong Ja Bae, Jong Yeoul Park, and Jin Mun Jeong. On uniform decay of solutions for wave equation of Kirchhoff type with nonlinear boundary damping and memory source term. Applied Mathematics and Computation, 138(2–3):463– 478, June 20, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bogaert:2001:SAT

[BR01] J. Bogaert and R. Rousseau.

Spatial aggregation of twodimensional objects in raster data structures. Applied Mathematics and Computation, 119(2–3):117–126, April 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/99/25/20/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/99/25/
20/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399002350.

Bru:2004:SPP

[BRS04] Rafael Bru, Sergio Romero, and Elena Sánchez. Structural properties of positive periodic discrete-time linear systems: canonical forms. Applied Mathematics and Computation, 153 (3):697–719, June 14, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bastert:2002:LST

[BRST02] Oliver Bastert, Dan Rockmore, Peter F. Stadler, and Gottfried Tinhofer. Landscapes on spaces of trees. Applied Mathematics and Computation, 131(2–3):439– 459, September 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bogaert:2000:AAP

[BRVI00] J. Bogaert, R. Rousseau, P. Van Hecke, and I. Impens. Alternative area-perimeter ratios for measurement of 2D shape compactness of habitats. Applied Mathematics and Computation, 111(1):71–85, May 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/85/21/
                                        25/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/85/21/25/article.
                                        pdf.

Bryc:2002:UAR

[Bry02] W. Bryc. A uniform approximation to the right normal tail integral. Applied Mathematics and Computation, 127(2–3):365–374, April 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/123/27/44/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300301000157.

Buetow:2000:TOB

[BS00] Gerald W. Buetow and James S. Sochacki. The trade-offs between alternative finite difference techniques used to price derivative securities. Applied Mathematics and Computation, 115(2–3):177–190, October 27, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/90/23/28/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/90/23/
28/article.pdf.

Balachandran:2001:CIS

[BS01] K. Balachandran and R. Sakthivel. Controllability of integrodifferential systems in Banach spaces. Applied Mathematics and Computation, 118(1):63–71, February 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/96/20/22/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/96/20/
22/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300300000400.

Bishay:2004:TRV

[BS04] Samira T. Bishay and Adel A. S. Abo Seliem. Transient response of a vertical electric dipole above a two-layer medium. Applied Mathematics and Computation, 151 (2):401–410, April 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

binSuleiman:2001:SDD

[bSI01] Mohamed bin Suleiman and Fudziah Ismail. Solving delay differential equations using componentwise partitioning by Runge–Kutta method. Applied Mathematics and Computation, 122(3):301–323, August 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/106/26/
23/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/106/26/23/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000394.

Babolian:2004:NSF

[BSJ04] E. Babolian, H. Sadeghi, and Sh. Javadi. Numerically solution of fuzzy differential equations by Adomian method. Applied Mathematics and Computation, 149 (2):547–557, February 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bayram:2003:ACA

[BŞY03] Mustafa Bayram, Hakan Şimşek, and Necmettin Yıldırım. Automatic calculation of Alexander polynomials of (3, k)-torus knots. Applied Mathematics and Computation, 136(2–3):505– 510, March 15, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Batzel:2000:MIC

[BT00] J. J. Batzel and H. T. Tran.

Modeling instability in the control system for human respiration: applications to infant non-REM sleep. Applied Mathematics and Computation, 110(1):1–51, April 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/83/21/
21/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/83/21/21/article.
pdf.

Biazar:2003:SKM

[BTBI03] J. Biazar, M. Tango, E. Babolian, and R. Islam. Solution of the kinetic modeling of lactic acid fermentation using Adomian decomposition method. Applied Mathematics and Computation, 144(2–3):433–439, December 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bachiri:2002:FDM

[BV02] Hafid Bachiri and Luis Vázquez. Finite difference method to solve Maxwell’s equations for soliton propagation. Applied Mathematics and Computation, 126(2–3):213–229, March 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/122/30/32/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001521.

Bu:2004:ACD

[BW04] Fanbin Bu and Yimin Wei. The algorithm for computing the Drazin inverses of two-variable polynomial matrices. Applied Mathematics and Computation, 147(3): 805–836, January 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Celik:2003:OSA

[ÇAB03] Ercan Çelik, Arzu Aykut, and Mustafa Bayram. The ordinary successive approximations method and Padé approximants for solving a differential equation with variant retarded argument. Applied Mathematics and Computation, 144(1):173– 180, November 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Celik:2004:MSA

[ÇAB04] Ercan Çelık, Arzu Aykut, and Mustafa Bayram. The modified successive approximations method and Padé approximants for solving the differential equation with variant retarded argument. Applied Mathematics and Computation, 151(2):393– 400, April 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Caglar:2004:WIB

[Çağ04] Atife Çağlar. Weak imposition of boundary conditions for the Navier–Stokes equations. Applied Mathematics and Computation, 149(1): 119–145, February 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Casasus:2002:DMO

[CAH02] Luis Casasús and Waleed Al-Hayani. The decomposition method for ordinary differential equations with discontinuities. Applied Mathematics and Computation, 131(2–3):245–251, September 25, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cao:2001:NAE

[Cao01]  Yanzhao Cao.  Numeri-

cal approximations of exact controllability problems by optimal control problems for parabolic differential equations. Applied Mathematics and Computation, 119(2–3):127–145, April 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/99/25/21/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/99/25/
21/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399002519.

Cao:2003:GAS

[Cao03a] Jinde Cao. Global asymptotic stability of delayed bidirectional associative memory neural networks. Applied Mathematics and Computation, 142(2–3):333–339, October 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cao:2003:REA

[Cao03b] Zhi-Hao Cao. Rounding error analysis of two-stage iterative methods for large linear systems. Applied Mathematics and Computation, 139(2–3):371–381, July 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cao:2004:FIS

[Cao04a] Zhi-Hao Cao. Fast iterative solution of stabilized Navier–Stokes systems. Applied Mathematics and Computation, 157(1):219–241, September 27, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cao:2004:SEI

[Cao04b] Zhi-Hao Cao. Semiconvergence of extrapolated iterative method for singular linear systems. Applied Mathematics and Computation, 156(1):131–136, August 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cap:2001:EMA

[Cap01] Ferdinand F. Cap. Eigenfrequencies of membranes of arbitrary boundary and with varying surface mass density. Applied Mathematics and Computation, 124(3):319–329, December 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/113/32/29/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000849.

Casti:2000:PSU

[Cas00] John L. Casti. Professor Sueo Ueno — personal reminiscences. Applied Mathematics and Computation, 116(1–2):3–4, November 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/91/20/21/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/91/20/21/article.
pdf.

Catal:2003:NAO

[Çat03] Seval Çatal. Numerical approximation for the oxygen diffusion problem. Applied Mathematics and Computation, 145(2–3):361–369, December 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Catal:2004:NST

[Çat04] Seval (Alku) Çatal. Numerical solution of the thermistor problem. Applied Mathematics and Computation, 152 (3):743–757, May 13, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Covachev:2001:DAI

[CAY01] Valéry Covachev, Haydar Akça, and Fuat Yeniçerioğlu. Difference approximations for impulsive differential equations. Applied Mathematics and Computation, 121(2–3):383–390, June 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/105/25/
38/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/105/25/38/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S009630030000014X.

Celik:2003:AON

[ÇB03a] Ercan Çelik and Mustafa Bayram. Arbitrary order numerical method for solving differential-algebraic equation by Padé series. Applied Mathematics and Computation, 137(1):57–65, May 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Celik:2003:NSD

[ÇB03b] Ercan Çelik and Mustafa Bayram. On the numerical solution of differentialalgebraic equations by Padé series. Applied Mathematics and Computation, 137 (1):151–160, May 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Celik:2004:AGB

[ÇB04a] Ercan Çelik and Mustafa Bayram. Application of Gröbner basis techniques to enzyme kinetics. Applied Mathematics and Computation, 153(1):97–109, May 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Celik:2004:BSS

[ÇB04b] Ercan Çelık and Mustafa Bayram. The basic successive substitute approximations method and Padé approximations to solve the elasticity problem of settled of the wronkler ground with variable coefficients. Applied Mathematics and Computation, 154(2):495– 505, July 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Celik:2004:NSDa

[ÇB04c] Ercan Çelık and Mustafa Bayram. Numerical solution of differential–algebraic equation systems and applications. Applied Mathematics and Computation, 154 (2):405–413, July 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Celik:2004:NSE

[ÇB04d] Ercan Çelik and Mustafa Bayram. A numerical solution of the elasticity problem of settled of the wronkler ground with variable coefficients. Applied Mathematics and Computation, 150 (3):821–831, March 17, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chaturvedi:2000:BAD

[CBK00] Anoop Chaturvedi, M. I.

Bhatti, and Kuldeep Kumar. Bayesian analysis of disturbances variance in the linear regression model under asymmetric loss functions. Applied Mathematics and Computation, 114(2–3):149–153, September 11, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/89/22/23/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/89/22/
23/article.pdf.

Choo:2000:FDA

[CC00a] S. M. Choo and S. K.

Chung. Finite difference approximate solutions for the strongly damped extensible beam equations. Applied Mathematics and Computation, 112(1):11–32, June 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/86/21/
22/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/86/21/22/article.
pdf.

Choo:2000:FEG

[CC00b] S. M. Choo and S. K. Chung. Finite element Galerkin solutions for the nonplanar oscillatory beam equations. Applied Mathematics and Computation, 114(2–3):279–301, September 11, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/89/22/33/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/89/22/
33/article.pdf.

Chen:2001:NOM

[CC01] Yurong Chen and Dayong Cai. Nonlinear orthomin(k) methods. Applied Mathematics and Computation, 124(3):351–363, December 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/113/32/33/abstract.
html;   http://www.sciencedirect.

com/science/article/pii/
S0096300300001016.

Chen:2002:MRE

[CC02] Jinru Chen and Junzhi Cui. A multiscale rectangular element method for elliptic problems with entirely small periodic coefficients. Applied Mathematics and Computation, 130 (1):39–52, July 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cai:2003:SIM

[CC03a] Jing Cai and GuoLiang Chen. A splitting iterative method for α–β generalized inverse and singular linear system. Applied Mathematics and Computation, 145(2–3):221–232, December 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chao:2003:ISD

[CC03b] Zhu Chao and Guoliang Chen. Index splitting for the Drazin inverse of linear operator in Banach space. Applied Mathematics and Computation, 135(2–3):201–209, March 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:2003:EEA

[CC03c] Anping Chen and Jinde Cao.

Erratum to “Existence and attractivity of almost periodic solutions for cellular neural networks with distributed delays and variable coefficients”: [Appl. Math. Comput. 134 (2002) 125– 140]. Applied Mathematics and Computation, 138 (1):169, June 1, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [CC03d].

Chen:2003:EAA

[CC03d] Anping Chen and Jinde Cao. Existence and attractivity of almost periodic solutions for cellular neural networks with distributed delays and variable coefficients. Applied Mathematics and Computation, 134(1): 125–140, January 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See erratum [CC03c].

Chen:2003:IOB

[CC03e] Yurong Chen and Dayong Cai. Inexact overlapped block Broyden methods for solving nonlinear equations. Applied Mathematics and Computation, 136(2–3):215– 228, March 15, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Choi:2003:AGS

[CC03f] Eunwoo Choi and David A. Cicci. Analysis of GPS static positioning problems. Applied Mathematics and Computation, 140(1):37–51, July 30, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:2004:ADTa

[CC04a] Chao-Kuang Chen and Shih-Shin Chen. Application of the differential transformation method to a nonlinear conservative system. Applied Mathematics and Computation, 154(2):431– 441, July 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:2004:TSF

[CC04b] Jinru Chen and Junzhi Cui. Two-scale finite element method for nonselfadjoint elliptic problems with rapidly oscillatory coefficients. Applied Mathematics and Computation, 150 (2):585–601, March 8, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cui:2004:ACS

[CC04c] Lihong Cui and Zhengxing Cheng. An algorithm for constructing symmetric orthogonal multiwavelets by matrix symmetric extension. Applied Mathematics and Computation, 149(1): 227–243, February 5, 2004.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cho:2001:IOD

[CCC01] S. Cho, J. E. Cochran, and D. A. Cicci. Identification and orbit determination of a tethered satellite system. Applied Mathematics and Computation, 117(2–3):301–312, January 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/92/23/33/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/92/23/
                                        33/article.pdf;   http:
                                        //www.sciencedirect.com/
                                        science/article/pii/S0096300399001824.

Chen:2004:TPM

[CCH04] Tzer-Shyong Chen, Yu-Fang Chung, and Kuo-Hsuan Huang. A traceable proxy multisignature scheme based on the elliptic curve cryptosystem. Applied Mathematics and Computation, 159(1):137– 145, November 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cong:2004:GSR

[CCK04] Yuhao Cong, Jianing Cai, and Jiaoxun Kuang. The GPL-stability of Rosenbrock methods for delay differential equation. Applied Mathematics and Computation, 150(2):533–542, March 8, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Censor:2002:SSD

[CCKS02] Yair Censor, Nir Cohen, Tuvia Kutscher (Kotzer), and Joseph Shamir. Summed squared distance error reduction by simultaneous multiprojections and applications. Applied Mathematics and Computation, 126(2– 3):157–179, March 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/9/12/122/30/28/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001442.

Cabada:2003:PSC

[CCP03] Alberto Cabada, José Angel Cid, and Rodrigo L. Pouso. Positive solutions for a class of implicit and discontinuous second order functional differential equations with singularities. Applied Mathematics and Computation, 137(1):89–99, May 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chan:2003:SFH

[CCS03] Whei-Ching C. Chan, Kung-Yu Chen, and H. M. Srivastava. Some families of hypergeometric generating functions associated with multiple series transformations. Applied Mathematics and Computation, 144(1): 27–59, November 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cavalcanti:2004:EUD

[CCS04a] M. M. Cavalcanti, V. N. Domingos Cavalcanti, and M. L. Santos. Existence and uniform decay rates of solutions to a degenerate system with memory conditions at the boundary. Applied Mathematics and Computation, 150(2):439–465, March 8, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Choi:2004:SIZ

[CCS04b] Junesang Choi, Young Joon Cho, and H. M. Srivastava. Series involving the Zeta function and multiple Gamma functions. Applied Mathematics and Computation, 159(2):509–537, December 6, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:2003:UUFa

[CCY03] Chun-I Chen, Cha’o-Kuang Chen, and Yue-Tzu Yang. Unsteady unidirectional flow of second grade fluid between the parallel plates with different given volume flow rate conditions. Applied Mathematics and Computation, 137(2–3):437–450, May 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chan:2004:OSE

[CCY04] Pui-Kit Chan, Wai-Ki Ching, and Siu-Pang Yung. Optimal strategies in equity securities and derivatives. Applied Mathematics and Computation, 151(3):615– 643, April 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Calude:2002:EMM

[CD02] Cristian S. Calude and Monica Dumitrescu. Entropic measures, Markov information sources and complexity. Applied Mathematics and Computation, 132(2–3):369– 384, November 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cao:2003:ESD

[CD03a] Jinde Cao and Meifang Dong. Exponential stability of delayed bi-directional associative memory networks. Applied Mathematics and Computation, 135(1):105– 112, February 15, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cinar:2003:CPP

[CD03b] Inan Cinar and Hakki Duru. On continuity properties of potentials depending on λdistance. Applied Mathematics and Computation, 139(2– 3):531–534, July 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cinar:2004:HLS

[ÇD04] İnan Çinar and Hakki Duru. The Hardy–Littlewood–Sobolev inequality for (β, γ)-distance Riesz potentials. Applied Mathematics and Computation, 153(3):757–762, June 14, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cherpion:2001:CMI

[CDH01] M. Cherpion, C. De Coster, and P. Habets. A constructive monotone iterative method for second-order BVP in the presence of lower and upper solutions. Applied Mathematics and Computation, 123(1):75–91, September 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        9/12/110/27/31/abstract.
                                        html;   http://www.sciencedirect.

com/science/article/pii/
S0096300300000588.

Celik:2004:NSC

[Çel04a] Ercan Çelik. On the numerical solution of chemical differential–algebraic equations by Padé series. Applied Mathematics and Computation, 153(1):13–17, May 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Celik:2004:NSDb

[Çel04b] Ercan Çelık. On the numerical solution of differential– algebraic equations with index-2. Applied Mathematics and Computation, 156 (2):541–550, September 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cheon:2004:ESP

[CEM04] Gi-Sang Cheon and Moawwad El-Mikkawy. Extended symmetric Pascal matrices via hypergeometric functions. Applied Mathematics and Computation, 158(1):159– 168, October 25, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cao:2003:NVR

[CF03] Zhi-Hao Cao and Li-Hong Feng. A note on variational representation for singular values of matrix. Applied Mathematics and Computation, 143(2–3):559–563, November 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Coll:2002:SID

[CFS02] Carmen Coll, Màrius J. Fullana, and Elena Sánchez. Some invariants of discretetime descriptor systems. Applied Mathematics and Computation, 127(2–3):277–287, April 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/123/27/37/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300301000054.

Coll:2004:ROI

[CFS04] Carmen Coll, Màrius J. Fullana, and Elena Sánchez. Reachability and observability indices of a discretetime periodic descriptor system. Applied Mathematics and Computation, 153 (2):485–496, June 4, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chavarriga:2003:IAC

[CG03] Javier Chavarriga and Maite Grau. Invariant algebraic curves linear in one variable for planar real quadratic systems. Applied Mathematics and Computation, 138(2– 3):291–308, June 20, 2003.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Calik:2004:EVS

[ÇG04] Sinan Çalik and Mehmet Güngör. On the expected values of the sample maximum of order statistics from a discrete uniform distribution. Applied Mathematics and Computation, 157(3): 695–700, October 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chavarriga:2001:IFT

[CGG01] J. Chavarriga, I. A. García, and J. Giné. Isochronicity into a family of timereversible cubic vector fields. Applied Mathematics and Computation, 121(2–3):129– 145, June 15, 2001. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/105/25/
22/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/105/25/22/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002672.

Castro-Gonzalez:2004:CMW

[CGVC04] N. Castro-González and J. Vélez-Cerrada. Characterizations of matrices which eigenprojections at zero are equal to a fixed perturbation. Applied Mathematics and Computation, 159(3): 613–623, December 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cabada:2002:IND

[CH02] Alberto Cabada and Seppo Heikkilä. Implicit nonlinear discontinuous functional boundary value ϕ-Laplacian problems: extremality results. Applied Mathematics and Computation, 129(2– 3):537–549, July 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cabada:2003:ECR

[CH03] A. Cabada and S. Heikkilä. Extremality and comparison results for discontinuous implicit third order functional initial-boundary value problems. Applied Mathematics and Computation, 140(2–3): 391–407, August 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cao:2004:SCL

[CH04] D. Q. Cao and Ping He. Stability criteria of linear neutral systems with a single delay. Applied Mathematics and Computation, 148(1): 135–143, January 20, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chang:2004:MBG

[Cha04a] Ching-Ter Chang. On the mixed binary goal programming problems. Applied Mathematics and Computation, 159(3):759–768, December 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chapko:2004:IEM

[Cha04b] Roman Chapko. An integral equation method for the numerical analysis of gravity waves in a channel with free boundary. Applied Mathematics and Computation, 159(1):247–266, November 25, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:2003:ESA

[CHC03] Anping Chen, Lihong Huang, and Jinde Cao. Existence and stability of almost periodic solution for BAM neural networks with delays. Applied Mathematics and Computation, 137 (1):177–193, May 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:2004:MCK

[CHC04] Tzer-Shyong Chen, Kuo-Hsuan Huang, and Yu-Fang Chung. Modified cryptographic key assignment scheme for overcoming the incorrectness of the CHW scheme. Applied Mathematics and Computation, 159 (1):147–155, November 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:2001:NSN

[Che01a] Jingbo Chen. New schemes for the nonlinear Shrödinger equation. Applied Mathematics and Computation, 124(3):371–379, December 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/113/32/35/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001119.

Chen:2001:MMM

[Che01b]  Jinru Chen.  Multigrid

method and multilevel additive preconditioner for mixed element method for non-self-adjoint and indefinite problems. Applied Mathematics and Computation, 119(2–3):229–247, April 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/99/25/29/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/99/25/
29/article.pdf;   http:

//www.sciencedirect.com/
science/article/pii/S0096300399002738.

Chen:2001:RAD

[Che01c] Yong-Lin Chen. Representation and approximation for the Drazin inverse A^{(}d). Applied Mathematics and Computation, 119(2–3):147–160, April 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/99/25/22/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/99/25/
22/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399002520.

Cheikh:2003:SRQ

[Che03] Youssèf Ben Cheikh. Some results on quasi-monomiality. Applied Mathematics and Computation, 141(1):63–76, August 20, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:2004:NSC

[Che04a] Jenq-Der Chen. New stability criteria for a class of neutral systems with discrete and distributed timedelays: an LMI approach. Applied Mathematics and Computation, 150(3):719– 736, March 17, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:2004:RCU

[Che04b] Jenq-Der Chen. Robust control for uncertain neutral systems with time-delays in state and control input via LMI and GAs. Applied Mathematics and Computation, 157(2):535–548, October 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:2004:BEH

[Che04c] Wenyan Chen. The blow-up estimate for heat equations with non-linear boundary conditions. Applied Mathematics and Computation, 156(2):355–366, September 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:2004:NRP

[Che04d]  Y. Chen.  New results

on positive periodic solutions of a periodic integrodifferential competition system. Applied Mathematics and Computation, 153 (2):557–565, June 4, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cabada:2001:MMN

[CHL01] Alberto Cabada, Patrick Habets, and Susana Lois. Monotone method for the Neumann problem with lower and upper solutions in the reverse order. Applied Mathematics and Computation, 117(1):1–14, January 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/92/20/20/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/92/20/
20/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399001496.

Chen:2002:DSA

[CHL02] Cha’o-Kuang Chen, Chen-I Hung, and Zong-Yi Lee. Double side approach method to obtain solutions for transient nonlinear heat conduction using genetic algorithms. Applied Mathematics and Computation, 133(2–3):431–444, December 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Choe:2000:GCF

[Cho00] Geon Ho Choe. Generalized continued fractions. Applied Mathematics and Computation, 109(2–3):287– 299, March 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/82/23/
34/abstract.html;   http:
//www.elsevier.nl/gej-

                                        ng/29/17/20/82/23/34/article.
                                        pdf.

Cho:2002:NAN

[Cho02a] Chung-Ki Cho. A new approach for numerical identification of free boundary. Applied Mathematics and Computation, 133(1):131– 145, November 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Choe:2002:RTA

[Cho02b] Geon Ho Choe. Recurrence of transformations with absolutely continuous invariant measures. Applied Mathematics and Computation, 129(2–3):501–516, July 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chu:2003:STF

[Chu03] Wen-Chang Chu. Summations on trigonometric functions. Applied Mathematics and Computation, 141(1): 161–176, August 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cichocki:2001:LRF

[Cic01] Adam Cichocki. Limit reliability functions of some homogeneous regular seriesparallel and parallel-series systems of higher order. Applied Mathematics and Computation, 117(1):55–72, January 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/92/20/25/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/92/20/
25/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399001551.

Cinar:2004:PCX

[Çin04a]  C. Çinar.  On the pe-

riodic cycle of x_{n+1} = (a_{n} + b_{n}x_{n})/c_{n}x_{n−1}. Applied Mathematics and Computation, 150(1):1–4, February 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cinar:2004:PSDa

[Çin04b] C. Çinar. On the positive solutions of the difference equation x_{n+1} = (x_{n−1})/(1 + x_{n}x_{n−1}). Applied Mathematics and Computation, 150(1):21–24, February 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cinar:2004:DE

[Çin04c] Cengiz Çinar. On the difference equation x_{n+1} = x_{n−1}/(−1 + x_{n}x_{n−1}). Applied Mathematics and Computation, 158(3):813–816, November 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cinar:2004:PSDc

[Çin04d]  Cengiz Çinar.  On the

positive solutions of the difference equation system x_{n+1} = 1/y_{n}, y_{n+1} = y_{n}/(x_{n−1}y_{n−1}). Applied Mathematics and Computation, 158(2):303–305, November 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cinar:2004:PSDb

[Çin04e] Cengiz Çinar. On the positive solutions of the difference equation x_{n+1} = ax_{n−1}/(1 + bx_{n}x_{n−1}). Applied Mathematics and Computation, 156(2):587–590, September 6, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cinar:2004:PSDd

[Çin04f] Cengiz Çinar. On the positive solutions of the difference equation x_{n+1} = x_{n−1}/(1 + ax_{n}x_{n−1}). Applied Mathematics and Computation, 158(3):809–812, November 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cinar:2004:SDE

[Çin04g] Cengiz Çinar. On the solutions of the difference equation x_{n+1} = x_{n−1}/(−1 + ax_{n}x_{n−1}). Applied Mathematics and Computation, 158(3):793–797, November 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cinar:2004:PGR

[Cin04h] Inan Cinar. The (p, q, l)- properties of a generalized Riesz potential. Applied Mathematics and Computation, 153(3):751–755, June 14, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:2004:ADTb

[CJ04a] Cha’o-Kuang Chen and Shin-Ping Ju. Application of differential transformation to transient advective– dispersive transport equation. Applied Mathematics and Computation, 155 (1):25–38, July 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chien:2004:IAM

[CJ04b] Hung-Yu Chien and Jinn-Ke Jan. Improved authenticated multiple-key agreement protocol without using conventional one-way function. Applied Mathematics and Computation, 147(2): 491–497, January 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Carbonell:2002:NMC

[CJB02] F. Carbonell, J. C. Jimenez, and R. Biscay. A numerical method for the computation of the Lyapunov exponents of nonlinear ordinary differential equations. Applied Mathematics and Computation, 131(1):21–37, September 10, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chien:2003:CMV

[CJT03] Hung-Yu Chien, Jinn-Ke Jan, and Yuh-Min Tseng. Cryptanalysis on Mu–Varadharajan’s e-voting schemes. Applied Mathematics and Computation, 139(2–3):525–530, July 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Choe:2003:KCG

[CK03] Geon Ho Choe and Chihurn Kim. The Khintchine constants for generalized continued fractions. Applied Mathematics and Computation, 144(2–3):397–411, December 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Celik:2003:NSC

[ÇKB03] Ercan Çelik, Erdal Karaduman, and Mustafa Bayram. Numerical solutions of chemical differential-algebraic equations. Applied Mathematics and Computation, 139(2–3):259–264, July 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Casti:2000:CSF

[CKNU00] John Casti, Robert Kalaba, Harriet Natsuyama, and Sueo Ueno. Cauchy systems for Fredholm integral equations with parameter imbedding. Applied Mathematics and Computation, 108 (1):47–54, February 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/79/21/
25/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/79/21/25/article.
pdf.

Chen:2001:GTW

[CL01] Wen-Sheng Chen and Wei Lin. Galerkin trigonometric wavelet methods for the natural boundary integral equations. Applied Mathematics and Computation, 121(1):75–92, May 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/9/12/105/21/
                                        26/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/9/12/105/21/26/article.
                                        pdf;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300399002635.

Chang:2003:ADS

[CL03a] Cheng-Long Chang and Zong-Yi Lee. Applying the double side method of weighted residual for solving circle plate large deformation problems. Applied Mathematics and Computation, 141(2–3):477–490, September 5, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ching:2003:IMF

[CL03b] Wai Ki Ching and Anthony W. Loh. Iterative methods for flexible manufacturing systems. Applied Mathematics and Computation, 141(2–3):553–564, September 5, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cao:2004:UMR

[CL04a] Yanhua Cao and Jianzhou Liu. Uniqueness of matrix mth root. Applied Mathematics and Computation, 157(3):605–622, October 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chang:2004:ADS

[CL04b] Cheng-Long Chang and Zong-Yi Lee. Applying the double side method to solution nonlinear pendulum problem. Applied Mathematics and Computation, 149 (3):613–624, February 22, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:2004:AAD

[CL04c] Wenhai Chen and Zhengyi Lu. An algorithm for Adomian decomposition method. Applied Mathematics and Computation, 159(1):221– 235, November 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:2004:SCE

[CLC04] Fengde Chen, Faxin Lin, and Xiaoxin Chen. Sufficient conditions for the existence positive periodic solutions of a class of neutral delay models with feedback control. Applied Mathematics and Computation, 158 (1):45–68, October 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cui:2003:SOP

[CLD03] Bao Tong Cui, Yongqing Liu, and Feiqi Deng. Some oscillation problems for impulsive hyperbolic differential systems with several delays. Applied Mathematics and Computation, 146(2–3): 667–679, December 31, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cao:2004:MSE

[CLF04] Wanrong Cao, Mingzhu Liu, and Zhencheng Fan. MS-stability of the Euler– Maruyama method for stochastic differential delay equations. Applied Mathematics and Computation, 159 (1):127–135, November 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cha:2000:SNM

[CLL00] Youngjoon Cha, Miyoung Lee, and Sungyun Lee. Stable nonconforming methods or the Stokes problem. Applied Mathematics and Computation, 114(2–3):155–174, September 11, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/89/22/24/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/89/22/
24/article.pdf.

Cao:2002:PSH

[CLW02] Jinde Cao, Qiong Li, and Shidong Wan. Periodic solutions of the higherdimensional non-autonomous systems. Applied Mathematics and Computation, 130(2– 3):369–382, August 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:2002:PTG

[CLX02] Guo-Liang Chen, Guo-Ming Liu, and Yi-Feng Xue. Perturbation theory for the generalized Bott–Duffin inverse and its applications. Applied Mathematics and Computation, 129(1):145– 155, June 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Campo:2004:MAC

[CM04a] Antonio Campo and Biagio Morrone. Meshless approach for computing the heat liberation from annular fins of tapered cross section. Applied Mathematics and Computation, 156(1): 137–144, August 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Choudhury:2004:TPB

[CM04b] Gautam Choudhury and Kailash C. Madan. A two phase batch arrival queueing system with a vacation time under Bernoulli schedule. Applied Mathematics and Computation, 149(2): 337–349, February 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Correa:2004:CPI

[CMF04] F. J. S. A. Corrêa, Silvano D. B. Menezes, and J. Ferreira. On a class of problems involving a nonlocal operator. Applied Mathematics and Computation, 147(2): 475–489, January 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cerdan:2002:PPB

[CMM02] J. Cerdán, J. Marín, and A. Martínez. Polynomial preconditioners based on factorized sparse approximate inverses. Applied Mathematics and Computation, 133(1):171–186, November 25, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Clavero:2001:NEA

[CMOS01] C. Clavero, J. J. H. Miller, E. O’Riordan, and G. I. Shishkin. Numerical experiments for advectiondiffusion problems in a channel with a 180^{∘} bend. Applied Mathematics and Computation, 118(2–3):223–246, March 9, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/96/25/26/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/96/25/
26/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399002155.

Cvetkovic:2000:SCR

[CO00] Ljiljana Cvetković and Jadranka Obrovski. Some convergence results of PD relaxation methods. Applied Mathematics and Computation, 107(2–3):103–112, January 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/72/22/
24/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/72/22/24/article.
pdf.

Casas:2001:NAC

[CO01] Ángela Jiménez Casas and Alfonso Matías Lozano Ovejero. Numerical analysis of a closed-loop thermosyphon including the Soret effect. Applied Mathematics and Computation, 124(3):289–318, December 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/113/32/28/abstract.
html;   http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300300000758.

Cook:2004:FSP

[Coo04] Steven Cook. Finite-sample properties of modified unit root tests in the presence structural change. Applied Mathematics and Computation, 149(3):625–640, February 22, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chang:2003:CCM

[CP03a] H. C. Chang and N. Prabhu. Canonical coordinates method for equality-constrained nonlinear optimization. Applied Mathematics and Computation, 140(1):135–158, July 30, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Climent:2003:CCT

[CP03b] Joan-Josep Climent and Carmen Perea. Convergence and comparison theorems for a generalized alternating iterative method. Applied Mathematics and Computation, 143(1):1–14, October 20, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Choudhury:2004:BAQ

[CP04] Gautam Choudhury and Madhuchanda Paul. A batch arrival queue with an additional service channel under N-policy. Applied Mathematics and Computation, 156(1):115–130, August 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cabada:2000:GMU

[CPL00] Alberto Cabada, Rodrigo L.

Pouso, and Eduardo Liz. A generalization of the method of upper and lower solutions for discontinuous first order problems with nonlinear boundary conditions. Applied Mathematics and Computation, 114(2–3):135–148, September 11, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/89/22/22/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/89/22/
22/article.pdf.

Climent:2004:BRD

[CPTZ04a] Joan-Josep Climent, Carmen Perea, Leandro Tortosa, and Antonio Zamora. A BSP recursive divide and conquer algorithm to solve a tridiagonal linear system. Applied Mathematics and Computation, 159(2): 459–484, December 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Climent:2004:CTP

[CPTZ04b] Joan-Josep Climent, Carmen Perea, Leandro Tortosa, and Antonio Zamora. Convergence theorems for parallel alternating iterative methods. Applied Mathematics and Computation, 148(2):497–517, January 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cicci:2001:LTS

[CQL01] D. A. Cicci, C. Qualls, and T. A. Lovell. A look at tethered satellite identification using ridge-type estimation methods. Applied Mathematics and Computation, 119(2–3):297–316, April 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/99/25/33/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/99/25/
                                        33/article.pdf;   http:
                                        //www.sciencedirect.com/
                                        science/article/pii/S0096300399002325.

Chaudhry:2004:EHC

[CQSP04] M. Aslam Chaudhry, Asghar Qadir, H. M. Srivastava, and R. B. Paris. Extended hypergeometric and confluent hypergeometric functions. Applied Mathematics and Computation, 159(2): 589–602, December 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cho:2000:ECB

[CS00] Sung-Bae Cho and Geum-Beom Song. Evolving CAM-Brain to control a mobile robot. Applied Mathematics and Computation, 111(2– 3):147–162, May 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/85/23/
23/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/85/23/23/article.
pdf.

Cengiz:2003:DLT

[CS03a] N. Cengiz and A. A. Salimov. Diagonal lift in the tensor bundle and its applications. Applied Mathematics and Computation, 142 (2–3):309–319, October 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Choi:2003:SIA

[CS03b] H.-W. Choi and J.-G. Shin.

Symbolic implementation of the algorithm for calculating Adomian polynomials. Applied Mathematics and Computation, 146(1):257– 271, December 30, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Choi:2003:MGR

[CSA03] Junesang Choi, H. M. Srivastava, and V. S. Adamchik. Multiple Gamma and related functions. Applied Mathematics and Computation, 134(2–3):515–533, January 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:2001:MIF

[CSD01] Zhaodou Chen, Yanan Shen, and Jun Ding. The Möbius inversion and Fourier coefficients. Applied Mathematics and Computation, 117(2–3): 161–176, January 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/
                                        10/9/12/92/23/23/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/92/23/
                                        23/article.pdf;   http:
                                        //www.sciencedirect.com/
                                        science/article/pii/S0096300399001708.

Chan:2003:SPB

[CT03a] C. Y. Chan and H. Y. Tian. Single-point blow-up for a degenerate parabolic problem with a nonlinear source of local and nonlocal features. Applied Mathematics and Computation, 145(2–3): 371–390, December 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Coll:2003:OPG

[CT03b] Carmen Coll and Néstor Thome. Oblique projectors and group involutory matrices. Applied Mathematics and Computation, 140(2–3): 517–522, August 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chuai:2004:REI

[CT04a] Jianjun Chuai and Yongge Tian. Rank equalities and inequalities for Kronecker products of matrices with applications. Applied Mathematics and Computation, 150(1):129–137, February 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Coken:2004:QIC

[ÇT04b] A. Ceylan Çöken and Abide Tuna. On the quaternionic inclined curves in the semi-Euclidean space E_{2}^{4}. Applied Mathematics and Computation, 155(2): 373–389, August 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:2003:SCS

[CTHK03] Y-Chuang Chen, Jimmy J. M. Tan, Lih-Hsing Hsu, and Shin-Shin Kao. Superconnectivity and super-edge-connectivity for some interconnection networks. Applied Mathematics and Computation, 140(2–3):245–254, August 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:2004:SSF

[CTHT04] Y-Chuang Chen, Chang-Hsiung Tsai, Lih-Hsing Hsu, and Jimmy J. M. Tan. On some super fault-tolerant Hamiltonian graphs. Applied Mathematics and Computation, 148(3):729–741, January 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Climent:2003:NRD

[CTZ03] Joan-Josep Climent, Leandro Tortosa, and Antonio Zamora. A note on the recursive decoupling method for solving tridiagonal linear systems. Applied Mathematics and Computation, 140 (1):159–164, July 30, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cao:2002:VSM

[CW02] Zhi-Hao Cao and Yang Wang. On validity of m-step multisplitting preconditioners for linear systems. Applied Mathematics and Computation, 126(2–3):199–211, March 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/122/30/31/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S009630030000151X.

Cotroneo:2003:SPH

[CW03] Tommaso Cotroneo and Jan C. Willems. The simulation problem for high order linear differential systems. Applied Mathematics and Computation, 145(2–3): 821–851, December 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cai:2004:DSW

[CW04a] Jianfeng Cai and Yimin Wei. Displacement structure of weighted pseudoinverses. Applied Mathematics and Computation, 153 (2):317–335, June 4, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cao:2004:DDR

[CW04b] Jinde Cao and Jun Wang.

Delay-dependent robust stability of uncertain nonlinear systems with time delay. Applied Mathematics and Computation, 154 (1):289–297, June 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cao:2000:NWS

[CWL00] Zhi-Hao Cao, He-Bing Wu, and Zhong-Yun Liu. A note on weak splittings of matrices. Applied Mathematics and Computation, 112(2– 3):265–275, June 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/86/23/
                                        29/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/86/23/29/article.
                                        pdf.

Chen:2002:SBO

[CWS02] Xin Chen, Wei Wang, and Yongzhong Song. Splitting based on the outer inverse of matrices. Applied Mathematics and Computation, 132(2–3):353–368, November 10, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:2004:PPS

[CWY04] Shihua Chen, Feng Wang, and Todd Young. Positive periodic solution of two-species ratio-dependent predator–prey system with time delay in two-patch environment. Applied Mathematics and Computation, 150(3):737–748, March 17, 2004. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic).

Cheng:2002:FUC

[CWZ02] L. Z. Cheng, Z. Wang, and Z. H. Zhang. Fast unified computation of the multidimensional discrete sinusoidal transforms. Applied Mathematics and Computation, 132(2–3):455–487, November 10, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:2003:UCS

[CX03] Jinru Chen and Xuejun Xu. Uniform convergence and Schwarz method for the mortar element method for non-selfadjoint and indefinite problems. Applied Mathematics and Computation, 136(2–3):517–533, March 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:2004:BPM

[CX04] Youpeng Chen and Chunhong Xie. Blow-up for a porous medium equation with a localized source. Applied Mathematics and Computation, 159(1):79–93, November 25, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chan:2001:BQD

[CY01a] C. Y. Chan and Jian Yang. Beyond quenching for degenerate singular semilinear parabolic equations. Applied Mathematics and Computation, 121(2–3):185– 201, June 15, 2001. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/105/25/
27/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/105/25/27/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002775.

Chan:2001:NQI

[CY01b] C. Y. Chan and Jian Yang. No quenching in infinite time for degenerate singular semilinear parabolic equations. Applied Mathematics and Computation, 121 (1):29–35, May 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/105/21/
23/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/105/21/23/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002763.

Chan:2001:PPN

[CY01c] C. Y. Chan and S. I. Yuen. Parabolic problems with nonlinear absorptions and releases at the boundaries. Applied Mathematics and Computation, 121(2– 3):203–209, June 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/105/25/
28/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/105/25/28/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002787.

Chang:2002:ALS

[CY02] Ching L. Chang and Suh-Yuh Yang. Analysis of the L^{2} least-squares finite element method for the velocity–vorticity–pressure Stokes equations with velocity boundary conditions. Applied Mathematics and Computation, 130(1):121– 144, July 25, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cao:2004:NWF

[CY04] Zhi-Hao Cao and Xiao-Yun Yu. A note on weighted FOM and GMRES for solving nonsymmetric linear systems. Applied Mathematics and Computation, 151 (3):719–727, April 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:2003:UUFb

[CYC03] Chun-I Chen, Yue-Tzu Yang, and Cha’o-Kuang Chen. Unsteady unidirectional flow of a Voigt fluid between the parallel surfaces with different given volume flow rate conditions. Applied Mathematics and Computation, 144(2–3):249–260, December 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:2004:NMS

[CZ04] Huaitang Chen and Hongqing Zhang. New multiple soliton-like solutions to the generalized (2 + 1)- dimensional KP equation. Applied Mathematics and Computation, 157(3):765– 773, October 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:2004:NES

[CZLZ04] Yong Chen, Xuedong Zheng, Biao Li, and Hongqing Zhang. New exact solutions for some nonlinear differential equations using symbolic computation. Applied Mathematics and Computation, 149(1):277–298, February 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cao:2004:ROL

[CZT04] Chongguang Cao, Xian Zhang, and Xiaomin Tang. Reverse order law of group inverses of products of two matrices. Applied Mathematics and Computation, 158(2):489–495, November 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Darwish:2003:IEU

[Dar03a] Mohamed Abdalla Darwish. On integral equations of Urysohn–Volterra type. Applied Mathematics and Computation, 136 (1):93–98, March 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Darwish:2003:WSF

[Dar03b] Mohamed Abdalla Darwish.

Weakly singular functionalintegral equation in infinite dimensional Banach spaces. Applied Mathematics and Computation, 136(1):123– 129, March 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Darwish:2004:NFV

[Dar04] Mohamed Abdalla Darwish.

On nonlinear Fredholm– Volterra integral equations with hysteresis. Applied Mathematics and Computation, 156(2):479–484, September 6, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dattoli:2003:OMF

[Dat03] G. Dattoli. Operational methods, fractional operators and special polynomials. Applied Mathematics and Computation, 141(1): 151–159, August 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dattoli:2004:ITC

[Dat04] G. Dattoli. Integral transforms and Chebyshev-like polynomials. Applied Mathematics and Computation, 148(1):225–234, January 20, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

deBruin:2002:ZSO

[dBGM02] M. G. de Bruin, W. G. M. Groenevelt, and H. G. Meijer. Zeros of Sobolev orthogonal polynomials of Hermite type. Applied Mathematics and Computation, 132 (1):135–166, October 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dattoli:2003:NFH

[DC03] G. Dattoli and C. Cesarano. On a new family of Hermite polynomials associated to parabolic cylinder functions. Applied Mathematics and Computation, 141(1): 143–149, August 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

daCosta:2003:CED

[dCD03] N. C. A. da Costa and F. A.

Doria. Consequences of an exotic definition for P = NP. Applied Mathematics and Computation, 145(2–3): 655–665, December 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See addendum [dCD06].

daCosta:2006:ACE

[dCD06] N. C. A. da Costa and F. A. Doria. Addendum to “Consequences of an exotic formulation for P = NP” [Appl. Math. Comput. 145 (2–3) (2003) 655–665]. Applied Mathematics and Computation, 172(2):1364–1367, January 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [dCD03].

Dattoli:2003:NTP

[DCS03] Giuseppe Dattoli, Clemente Cesarano, and Dario Sacchetti. A note on truncated polynomials. Applied Mathematics and Computation, 134(2–3):595–605, January 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Davies:2000:BTD

[DD00] Penny J. Davies and Dugald B. Duncan. On the behaviour of time discretisations of the electric field integral equation. Applied Mathematics and Computation, 107(1):1–26, January ??, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/72/17/17/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/72/17/
                                        17/article.pdf.

Dash:2002:ANS

[DD02] Ranjan K. Dash and Prabir Daripa. Analytical and numerical studies of a singularly perturbed Boussinesq equation. Applied Mathematics and Computation, 126(1):1–30, February 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        9/12/122/27/27/abstract.
                                        html;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300301001667.

Deeba:2002:ASB

[DDX02] Elias Deeba, Ghassan Dibeh, and Shishen Xie. An algorithm for solving bond pricing problem. Applied Mathematics and Computation, 128(1):81–94, May 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

deCastroSantos:2000:AFN

[de 00] Luis Carlos de Castro Santos. An adjoint formulation for the non-linear potential flow equation. Applied Mathematics and Computation, 108(1):11–21, February 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/79/21/
22/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/79/21/22/article.
pdf.

DeSchepper:2001:FEM

[De 01] H. De Schepper. A finite element method for differential eigenvalue problems with mixed classical and (semi- )periodic boundary conditions. Applied Mathematics and Computation, 121 (1):1–15, May 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/105/21/
21/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/105/21/21/article.
pdf;   http://www.sciencedirect.

                                        com/science/article/pii/
                                        S009630039900260X.

DeChant:2002:CIC

[De 02a] Lawrence J. De Chant. A computationally implementable criterion for the solvability of boundary value problems on an infinite domain. Applied Mathematics and Computation, 125(1):35–47, January 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        9/12/120/27/29/abstract.
                                        html;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300300001065.

DeChant:2002:INU

[De 02b] Lawrence J. De Chant. Introduction of non-uniformity through linearization of the system governing turbulent, mixing of a scalar quantity in a 2-d mixing layer. Applied Mathematics and Computation, 130(2–3):399–413, August 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

DeChant:2003:QSA

[De 03] Lawrence J. De Chant. Quasi-steady approximations to initial value problems with application to sediment transport. Applied Mathematics and Computation, 134(2–3):535–552, January 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

DeMarchi:2004:LSS

[De 04] Stefano De Marchi. On Leja sequences: some results and applications. Applied Mathematics and Computation, 152(3):621–647, May 13, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Duan:2000:NAE

[DE00] Jinqiao Duan and Vincent J. Ervin. On nonlinear amplitude evolution under stochastic forcing. Applied Mathematics and Computation, 109(1):59–65, March 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/82/21/
25/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/82/21/25/article.
pdf.

Dehghan:2000:FDM

[Deh00]  Mehdi Dehghan.  A fi-

nite difference method for a non-local boundary value problem for two-dimensional heat equation. Applied Mathematics and Computation, 112(1):133–142, June 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/86/21/
                                        32/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/86/21/32/article.
                                        pdf.

Dehghan:2001:CNF

[Deh01] Mehdi Dehghan. Crank– Nicolson finite difference method for two-dimensional diffusion with an integral condition. Applied Mathematics and Computation, 124(1):17–27, November 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        9/12/113/27/28/abstract.
                                        html;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630030000031X.

Dehghan:2002:NTP

[Deh02] Mehdi Dehghan. Numerical techniques for a parabolic equation subject to an overspecified boundary condition. Applied Mathematics and Computation, 132(2– 3):299–313, November 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dehghan:2003:FCP

[Deh03a] Mehdi Dehghan. Finding a control parameter in onedimensional parabolic equations. Applied Mathematics and Computation, 135(2– 3):491–503, March 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dehghan:2003:FSM

[Deh03b]  Mehdi Dehghan.  Frac-

tional step methods for parabolic equations with a non-standard condition. Applied Mathematics and Computation, 142(1):177– 187, September 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dehghan:2003:ICF

[Deh03c] Mehdi Dehghan. Identifying a control function in two-dimensional parabolic inverse problems. Applied Mathematics and Computation, 143(2–3):375–391, November 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dehghan:2003:LES

[Deh03d] Mehdi Dehghan. Locally explicit schemes for threedimensional diffusion with a non-local boundary specification. Applied Mathematics and Computation, 138(2– 3):489–501, June 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dehghan:2003:NSP

[Deh03e] Mehdi Dehghan. Numerical solution of a parabolic equation with non-local boundary specifications. Applied Mathematics and Computation, 145(1):185–194, December 20, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dehghan:2003:NSO

[Deh03f] Mehdi Dehghan. Numerical solution of one-dimensional parabolic inverse problem. Applied Mathematics and Computation, 136(2–3):333– 344, March 15, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dehghan:2003:PTB

[Deh03g] Mehdi Dehghan. Parallel techniques for a boundary value problem with nonclassic boundary conditions. Applied Mathematics and Computation, 137(2–3):399– 412, May 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dehghan:2004:AAD

[Deh04a] Mehdi Dehghan. Application of the Adomian decomposition method for twodimensional parabolic equation subject to nonstandard boundary specifications. Applied Mathematics and Computation, 157(2): 549–560, October 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dehghan:2004:NCC

[Deh04b] Mehdi Dehghan. Numerical computation of a control function in a partial differential equation. Applied Mathematics and Computation, 147(2):397–408, January 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dehghan:2004:NPB

[Deh04c] Mehdi Dehghan. Numerical procedures for a boundary value problem with a non-linear boundary condition. Applied Mathematics and Computation, 147(1): 291–306, January 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dehghan:2004:NSO

[Deh04d] Mehdi Dehghan. Numerical schemes for one-dimensional parabolic equations with nonstandard initial condition. Applied Mathematics and Computation, 147(2): 321–331, January 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dehghan:2004:NSP

[Deh04e] Mehdi Dehghan. Numerical solution of a parabolic equation subject to specification of energy. Applied Mathematics and Computation, 149(1):31–45, February 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dehghan:2004:NST

[Deh04f] Mehdi Dehghan. Numerical solution of the threedimensional advection–diffusion equation. Applied Mathematics and Computation, 150(1):5–19, February 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dehghan:2004:TLT

[Deh04g] Mehdi Dehghan. Threelevel techniques for onedimensional parabolic equation with nonlinear initial condition. Applied Mathematics and Computation, 151(2):567–579, April 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dehghan:2004:WFD

[Deh04h] Mehdi Dehghan. Weighted finite difference techniques for the one-dimensional advection–diffusion equation. Applied Mathematics and Computation, 147(2):

307–319, January 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dubeau:2004:NUH

[DEK04] François Dubeau, Said Elmejdani, and Riadh Ksantini. Non-uniform Haar wavelets. Applied Mathematics and Computation, 159(3):675–693, December 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Demiray:2002:NSP

[Dem02a] Hilmi Demiray. A note on the solution of perturbed Korteweg–de Vries equation. Applied Mathematics and Computation, 132(2–3):643– 647, November 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Demiray:2002:PWN

[Dem02b] Hilmi Demiray. Propagation of weakly nonlinear waves in fluid-filled thin elastic tubes. Applied Mathematics and Computation, 133(1): 29–41, November 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Demir:2003:RBC

[Dem03a] H. Demir. Rayleigh–Benard convection of viscoelastic fluid. Applied Mathematics and Computation, 136(2–3): 251–267, March 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Demiray:2003:ASD

[Dem03b] Hilmi Demiray. An analytical solution to the dissipative nonlinear Schrödinger equation. Applied Mathematics and Computation, 145(1):179–184, December 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Demiray:2003:NAS

[Dem03c] Hilmi Demiray. A note on the analytical solution to the modified perturbed Korteweg–de Vries equation. Applied Mathematics and Computation, 134(2–3): 501–505, January 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Demiray:2004:TWS

[Dem04] Hilmi Demiray. A travelling wave solution to the KdV–Burgers equation. Applied Mathematics and Computation, 154(3):665–670, July 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dercole:2003:RCH

[Der03] Fabio Dercole. Remarks on the computation of the horizon of a digital terrain. Applied Mathematics and Computation, 146(2–3): 627–641, December 31, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Duan:2002:TPF

[DF02] Jinqiao Duan and Xin-Chu Fu. Time-periodically forced amplitude evolution in spatially extended nonlinear systems. Applied Mathematics and Computation, 127(2–3):215–219, April 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/123/27/32/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001673.

Diethelm:2004:MOF

[DF04] Kai Diethelm and Neville J.

Ford. Multi-order fractional differential equations and their numerical solution. Applied Mathematics and Computation, 154 (3):621–640, July 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Du:2004:OSD

[DFF04] Lili Du, Wei Fu, and Mingshu Fan. Oscillatory solutions of delay hyperbolic system with distributed deviating arguments. Applied Mathematics and Computation, 154(2):521–529, July 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Du:2002:GGO

[DG02] Qiang Du and Max Gunzburger. Grid generation and optimization based on centroidal Voronoi tessellations. Applied Mathematics and Computation, 133(2–3): 591–607, December 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Diasparra:2003:EAI

[DG03] M. Diasparra and H. Gzyl. Entropic approach to interior point solution of linear programs. Applied Mathematics and Computation, 143(2–3):339–347, November 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Duru:2004:SSS

[DG04] Hakki Duru and Aytekin Gulle. Some specialties of the solutions of the differential equations in Banach space. Applied Mathematics and Computation, 157(3): 667–675, October 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

deGaris:2000:BAB

[dGKG^{+}00] Hugo de Garis, Michael Korkin, Felix Gers, Eiji Nawa, and Michael Hough. Building an artificial brain using an FPGA based CAM-Brain Machine. Applied Mathematics and Computation, 111(2–3):163–192, May 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/85/23/
24/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/85/23/24/article.
pdf.

Dattoli:2004:CMO

[DGR04] G. Dattoli, B. Germano, and P. E. Ricci. Comments on monomiality, ordinary polynomials and associated bi-orthogonal functions. Applied Mathematics and Computation, 154 (1):219–227, June 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ding:2004:PBV

[DH04] Wei Ding and Maoan Han.

Periodic boundary value problem for the second order impulsive functional differential equations. Applied Mathematics and Computation, 155(3):709–726, August 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Diao:2004:SPD

[Dia04] Huaian Diao. Structured perturbations of Drazin inverse. Applied Mathematics and Computation, 158(2): 419–432, November 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dikici:2004:GRF

[Dik04] Ramazan Dikici. General recurrences in finite p-groups. Applied Mathematics and Computation, 158(2):445– 458, November 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ding:2000:EAS

[Din00] Xie Ping Ding. Existence and algorithm of solutions for generalized mixed implicit quasi-variational inequalities. Applied Mathematics and Computation, 113(1):67–80, July 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/
                                        10/9/12/87/21/25/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/87/21/
                                        25/article.pdf.

Ding:2001:GQV

[Din01] Xie Ping Ding. Generalized quasi-variational-like inclusions with nonconvex functionals. Applied Mathematics and Computation, 122(3):267–282, August 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/106/26/
21/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/106/26/21/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000278.

Ding:2002:ESE

[Din02] J. Ding. On the existence of solutions to equality constrained least-squares problems in infinite dimensional Hilbert spaces. Applied Mathematics and Computation, 131(2–3):573–581, September 25, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ding:2003:PRM

[Din03a] J. Ding. On the perturbation of the reduced minimum modulus of bounded linear operators. Applied Mathematics and Computation, 140(1):69–75, July 30, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ding:2003:PIT

[Din03b] Xie Ping Ding. Perturbed Ishikawa type iterative algorithm for generalized quasivariational inclusions. Applied Mathematics and Computation, 141(2–3):359–373, September 5, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ding:2004:IBAb

[Din04a] Changming Ding. Intertwined basins of attraction generated by the stable manifold of a saddle point. Applied Mathematics and Computation, 153 (3):779–783, June 14, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ding:2004:IBAa

[Din04b] Changming Ding. On the intertwined basins of attraction for planar flows. Applied Mathematics and Computation, 148(3):801–805, January 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ding:2004:ETW

[Din04c] Wei Ding. Existence of travelling waves in a lattice dynamical system. Applied Mathematics and Computation, 158(2):529–536, November 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ding:2004:ASC

[Din04d] Xie Ping Ding. Algorithms of solutions for completely generalized mixed implicit quasi-variational inclusions. Applied Mathematics and Computation, 148(1):47–66, January 20, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ding:2004:PCI

[Din04e] Xie Ping Ding. Predictor– corrector iterative algorithms for solving generalized mixed variational-like inequalities. Applied Mathematics and Computation, 152(3):855–865, May 13, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Diskin:2001:EMM

[Dis01] Boris Diskin. Efficient multigrid methods for solving upwind-biased discretizations of the convection equation. Applied Mathematics and Computation, 123(3): 343–379, October 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/9/12/110/32/30/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000734.

deLimaSantos:2004:BCM

[dJ04] Mauro de Lima Santos and F. Junior. A boundary condition with memory for Kirchhoff plates equations. Applied Mathematics and Computation, 148(2):475– 496, January 26, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Djerdjour:2000:SPS

[Dje00] Mohamed Djerdjour. A simple procedure for solving a continuous quadratic mathematical model. Applied Mathematics and Computation, 113(2–3):161–174, July 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/87/23/23/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/87/23/
23/article.pdf.

deKlerk:2002:SSS

[dK02] Johan H. de Klerk. Solving strongly singular integral equations by L_{p} approximation methods. Applied Mathematics and Computation, 127(2–3):311–326, April 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/123/27/39/abstract.
html;   http://www.sciencedirect.

com/science/article/pii/
S0096300301000091.

Davison:2004:PSG

[DK04] A. H. Davison and A. H.

Kara. Potential symmetry generators and associated conservation laws of perturbed nonlinear equations. Applied Mathematics and Computation, 156(1): 271–285, August 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dogan:2004:NIH

[DKV04] Abdulkadir Dogan, Yuksel Kaplan, and T. Nejat Veziroğlu. Numerical investigation of heat and mass transfer in a metal hydride bed. Applied Mathematics and Computation, 150(1): 169–180, February 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Deeba:2000:ASN

[DKX00] E. Deeba, S. A. Khuri, and Shishen Xie. An algorithm for solving a nonlinear integro-differential equation. Applied Mathematics and Computation, 115(2–3):123–131, October 27, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/90/23/24/abstract.

                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/90/23/
                                        24/article.pdf.

delaSen:2003:ASN

[dlS03] M. de la Sen. Adaptive stabilization of non-necessarily inversely stable first-order systems by using estimates modification based on testing the Sylvester determinant. Applied Mathematics and Computation, 141(2–3): 261–280, September 5, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Duan:2004:EDS

[DLYC04] Lixia Duan, Qishao Lu, Zaizhong Yang, and Lansun Chen. Effects of diffusion on a stage-structured population in a polluted environment. Applied Mathematics and Computation, 154 (2):347–359, July 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Digalakis:2004:PCM

[DM04] J. Digalakis and K. Margaritis. Performance comparison of memetic algorithms. Applied Mathematics and Computation, 158(1):237– 252, October 25, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Deo:2001:QPN

[DMS01] S. G. Deo, K. N. Murty, and Michael D. Shaw. Qualitative properties of nonlinear difference systems. Applied Mathematics and Computation, 123(1):27–36, September 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/110/27/28/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000539.

Deo:2002:QPA

[DMT02] S. G. Deo, K. N. Murty, and J. Turner. Qualitative properties of adjoint Kronecker product boundary value problems. Applied Mathematics and Computation, 133(2–3):287–295, December 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Duszczyk:2003:NMS

[DNS03] Bernard Duszczyk, Michael P.

Newell, and Stephen J. Sugden. Numerical methods for solving the eigenvalue problem for a positive branch confocal unstable resonator. Applied Mathematics and Computation, 140(2–3):427– 443, August 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dikici:2003:AFS

[DÖ03a] Ramazan Dikici and Engin Özkan. An application of Fibonacci sequences in groups. Applied Mathematics and Computation, 136(2– 3):323–331, March 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

dOnofrio:2003:GSV

[d’O03b] Alberto d’Onofrio. Globally stable vaccine-induced eradication of horizontally and vertically transmitted infectious diseases with periodic contact rates and diseasedependent demographic factors in the population. Applied Mathematics and Computation, 140(2–3):537–547, August 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

dOnofrio:2004:MPV

[d’O04] Alberto d’Onofrio. Mixed pulse vaccination strategy in epidemic model with realistically distributed infectious and latent times. Applied Mathematics and Computation, 151(1):181– 187, March 30, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dobner:2000:MES

[Dob00] H.-J. Dobner. A method for estimating the solution of integral equations encountered in potential theory. Applied Mathematics and Computation, 109(2–3): 199–204, March 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/82/23/
28/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/82/23/28/article.
pdf.

Dogan:2004:GFE

[Dog04]  Abdulkadir Dogan.  A

Galerkin finite element approach to Burgers’ equation. Applied Mathematics and Computation, 157(2): 331–346, October 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

deOliveira:2003:GPP

[dOVS03] Fabiane de Oliveira, Neida Maria Patias Volpi, and Carlos Roberto Sanquetta. Goal programming in a planning problem. Applied Mathematics and Computation, 140 (1):165–178, July 30, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dedic:2001:ETF

[DP01] Lj. Dedić and M. Matićand J.

Pečarić. On Euler trapezoid formulae. Applied Mathematics and Computation, 123(1):37–62, September 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/110/27/29/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000540.

Djerdjour:2001:BBA

[DR01] Mohamed Djerdjour and Kamel Rekab. A branch and bound algorithm for designing reliable systems at a minimum cost. Applied Mathematics and Computation, 118(2–3):247–259, March 9, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/96/25/27/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/96/25/
27/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399002179.

Dattoli:2004:CTB

[DRP04] G. Dattoli, P. E. Ricci, and P. Pacciani. Comments on the theory of Bessel functions with more than one index. Applied Mathematics and Computation, 150 (3):603–610, March 17, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dattoli:2003:GSO

[DRS03] G. Dattoli, P. E. Ricci, and D. Sacchetti. Generalized shift operators and pseudopolynomials of fractional order. Applied Mathematics and Computation, 141(1): 215–224, August 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dzurina:2003:OCS

[DS03] J. Džurina and I. P.

Stavroulakis. Oscillation criteria for second-order delay differential equations. Applied Mathematics and Computation, 140(2–3):445–453, August 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dattoli:2001:LLP

[DSC01] G. Dattoli, H. M. Srivastava, and C. Cesarano. The Laguerre and Legendre polynomials from an operational point of view. Applied Mathematics and Computation, 124(1):117–127, November 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/113/27/32/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000825.

Dag:2004:ACB

[DSI04] İdris Dağ, Bülent Saka, and Dursun Irk. Application of cubic B-splines for numerical solution of the RLW equation. Applied Mathematics and Computation, 159(2): 373–389, December 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Duan:2000:PSD

[Dua00]  Jinqiao Duan.  Proba-

bilistic structural dynamics of protein folding. Applied Mathematics and Computation, 113(1):97–100, July 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/87/21/27/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/87/21/
                                        27/article.pdf.

Dung:2002:TDB

[Dun02] Le Dung. On a timedependent bio-reactor model with chemotaxis. Applied Mathematics and Computation, 131(2–3):531–558, September 25, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Duru:2004:DSS

[Dur04]  Hakki Duru.  Difference

schemes for the singularly perturbed Sobolev periodic boundary problem. Applied Mathematics and Computation, 149(1):187–201, February 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ding:2002:BPS

[DW02] Jiu Ding and Yimin Wei.

Bounds for perturbed solutions of linear operator equations in Hilbert space. Applied Mathematics and Computation, 132(2–3):293–298, November 10, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Djordjevic:2004:OEP

[DW04] Dragan S. Djordjević and Yimin Wei. Operators with equal projections related to their generalized inverses. Applied Mathematics and Computation, 155(3):655– 664, August 16, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ding:2004:ASS

[DWC04] H. J. Ding, H. M. Wang, and W. Q. Chen. Analytical solution of a special non-homogeneous pyroelectric hollow cylinder for piezothermoelastic axisymmetric plane strain dynamic problems. Applied Mathematics and Computation, 151(2):423–441, April 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Debnath:2004:EPE

[DX04] Lokenath Debnath and Xingye Xu. The existence of positive entire solutions to singular nonlinear polyharmonic equations in R^{n}. Applied Mathematics and Computation, 151 (3):679–688, April 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Davidson:2002:EUL

[DXL02] F. A. Davidson, R. Xu, and J. Liu. Existence and uniqueness of limit cycles in an enzyme-catalysed reaction system. Applied Mathematics and Computation, 127(2–3):165–179, April 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        9/12/123/27/28/abstract.
                                        html;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300301000650.

DiCarlo:2004:FMS

[DYH04] J. A. DiCarlo, H. M. Yun, and J. B. Hurst. Fracture mechanisms for SiC fibers and SiC/SiC composites under stress-rupture conditions at high temperatures. Applied Mathematics and Computation, 152 (2):473–481, May 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Elzohny:2000:ABS

[EA00] Habiba A. Elzohny and Sabah Hafez Abd Allah. On analyse [sic] the behaviour of the solutions on a bounded set. Applied Mathematics and Computation, 112(1): 1–9, June 1, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/86/21/
21/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/86/21/21/article.
pdf.

El-Afifi:2003:NSC

[EA03] M. M. El-Afifi. Necessary and sufficient conditions for boundedness and stability of N-order difference equation. Applied Mathematics and Computation, 141(2–3): 427–445, September 5, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Eissa:2004:VCC

[EA04a] M. Eissa and Y. A. Amer.

Vibration control of a cantilever beam subject to both external and parametric excitation. Applied Mathematics and Computation, 152 (3):611–619, May 13, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Adawy:2004:JEM

[EA04b] T. M. El-Adawy. The joint essential maximal numerical range. Applied Mathematics and Computation, 148(3): 793–799, January 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Afifi:2004:RS

[EA04c] M. M. El-Afifi. On the recursive sequence x_{n+1} = \frac{α+βx_{n}+γx_{n−1}}{Bx_{n}+Cx_{n−1}} . Applied Mathematics and Computation, 147(3):617–628, January 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Amin:2004:DDE

[EA04d] M. F. El-Amin. Double dispersion effects on natural convection heat and mass transfer in non-Darcy porous medium. Applied Mathematics and Computation, 156 (1):1–17, August 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Afifi:2003:DE

[EAA03] M. M. El-Afifi and A. M. Ahmed. On the difference equation x_{n+1} = \frac{a+αx_{n}+αx_{n−1}+···+αx_{n−k+2}}{x_{n−k+1}} . Applied Mathematics and Computation, 144(2–3):537–542, December 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Et:2004:SGS

[EAA04] Mikâil Et, Hıfsı Altinok, and Yavuz Altin. On some generalized sequence spaces. Applied Mathematics and Computation, 154 (1):167–173, June 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Bassiouny:2002:PEN

[EB02a] A. F. El-Bassiouny. Parametrically excited non-linear systems: a comparison of two methods. Applied Mathematics and Computation, 132(2–3):385–410, November 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Elbashbeshy:2002:MCA

[EB02b] E. M. A. Elbashbeshy and M. A. Bazid. The mixed convection along a vertical plate with variable surface heat flux embedded in porous medium. Applied Mathematics and Computation, 125(2–3):317–324, January 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        9/12/120/31/41/abstract.
                                        html;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630030000134X.

El-Bassiouny:2003:MIR

[EB03a] A. F. El-Bassiouny. Modal interaction of resonantly forced oscillations of twodegree-of-freedom structure. Applied Mathematics and Computation, 134(2–3):217– 242, January 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Bassiouny:2003:TMI

[EB03b] A. F. El-Bassiouny. Threemode interaction in harmonically excited system with cubic nonlinearities. Applied Mathematics and Computation, 139(2–3):201–230, July 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Bassiouny:2003:TEA

[EB03c] Ahmed H. El-Bassiouny. On testing exponentiality against IFRA alternatives. Applied Mathematics and Computation, 146(2–3):445– 453, December 31, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Elbashbeshy:2003:HTU

[EB03d] E. M. A. Elbashbeshy and M. A. A. Bazid. Heat transfer over an unsteady stretching surface with internal heat generation. Applied Mathematics and Computation, 138(2–3):239–245, June 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Bassiouny:2004:AIL

[EB04a] Ahmed H. El-Bassiouny.

Asymptotic inference for LSE in multivariate continuous regression models with long-memory random fields. Applied Mathematics and Computation, 151 (1):251–262, March 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Bassiouny:2004:RCN

[EB04b] Ahmed H. El-Bassiouny. On the rate of convergence to the normal law for LSE in multivariate continuous regression model with longrange dependence stationary errors. Applied Mathematics and Computation, 149(2): 369–378, February 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Borai:2004:EES

[EB04c] Mahmoud M. El-Borai.

Evolution equations without semigroups. Applied Mathematics and Computation, 149(3):815–821, February 22, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Borai:2004:SSN

[EB04d] Mahmoud M. El-Borai. Semigroups and some nonlinear fractional differential equations. Applied Mathematics and Computation, 149(3):823–831, February 22, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Elbashbeshy:2004:ETD

[EB04e] E. M. A. Elbashbeshy and M. A. A. Bazid. The effect of temperature-dependent viscosity on heat transfer over a continuous moving surface with variable internal heat generation. Applied Mathematics and Computation, 153(3):721–731, June 14, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Elbashbeshy:2004:HTP

[EB04f] E. M. A. Elbashbeshy and M. A. A. Bazid. Heat transfer in a porous medium over a stretching surface with internal heat generation and suction or injection. Applied Mathematics and Computation, 158(3):799– 807, November 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Bassiouny:2003:GFA

[EBA03] Ahmed H. El-Bassiouny and Ibrahim A. Alwasel. A goodness of fit approach to testing mean residual times. Applied Mathematics and Computation, 143(2–3):301–307, November 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Bassiouny:2003:DSD

[EBE03] A. F. El-Bassiouny and M. Eissa. Dynamics of a single-degree-of-freedom structure with quadratic, cubic and quartic nonlinearities to a harmonic resonance. Applied Mathematics and Computation, 139(1):1–21, July 1, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Borai:2003:ASS

[EBMA03] Mahmoud M. El-Borai, Osama L. Moustafa, and Hamdy M. Ahmed. Asymptotic stability of some stochastic evolution equations. Applied Mathematics and Computation, 144(2–3): 273–286, December 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Bassiouny:2004:TEA

[EBSAG04] Ahmed H. El-Bassiouny, Ammar M. Sarhan, and M. Al-Garian. Testing exponentiality against NBUFR (NWUFR). Applied Mathematics and Computation, 149(2):351–358, February 12, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Eom:2002:GAB

[ECL02] Tae-Dok Eom, Changkyu Choi, and Ju-Jang Lee. Generalized asymmetrical bidirectional associative memory for multiple association. Applied Mathematics and Computation, 127(2–3):221– 233, April 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/123/27/33/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001703.

Elbashbeshy:2002:ERF

[ED02] E. M. A. Elbashbeshy and M. F. Dimian. Effect of radiation on the flow and heat transfer over a wedge with variable viscosity. Applied Mathematics and Computation, 132(2–3):445–454, November 10, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Enna:2003:ATE

[EE03] Ahmed A. El-Enna. Analytical treatment of the Earth oblateness and solar radiation pressure effects on an artificial satellite I. The equations of motion. Applied Mathematics and Computation, 138(2– 3):443–453, June 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Enna:2004:ATE

[EE04a] Ahmed. A. El-Enna. Analytical treatment of the earth oblateness and solar radiation pressure effects on an artificial satellite: II — The solution. Applied Mathematics and Computation, 151 (1):121–145, March 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Enna:2004:EMJ

[EE04b] Ahmed A. El-Enna. The equations of motion of the joint effects of the relativistic and solar radiation pressure on an earth artificial satellite. Applied Mathematics and Computation, 149(2): 359–368, February 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Elshehawey:2001:ESE

[EEAES01a] E. F. Elshehawey, Elsayed M. E. Elbarbary, N. A. S. Afifi, and Mostafa El-Shahed. An exact solution of the endolymph equation. Applied Mathematics and Computation, 124(3):331–335, December 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/113/32/30/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S009630030000093X.

Elshehawey:2001:SEE

[EEAES01b] E. F. Elshehawey, Elsayed M. E. Elbarbary, N. A. S. Afifi, and Mostafa El-Shahed. On the solution of the endolymph equation using fractional calculus. Applied Mathematics and Computation, 124(3):337–341, December 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/113/32/31/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000941.

Eissa:2003:ANS

[EEB03a] M. Eissa and A. F. El-Bassiouny. Analytical and numerical solutions of a non-linear ship rolling motion. Applied Mathematics and Computation, 134(2–3): 243–270, January 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Elnagger:2003:PNP

[EEB03b] A. M. Elnagger and A. F.

El-Bassiouny. Periodic and non-periodic combination resonance in kinematically excited system of rods. Applied Mathematics and Computation, 139(2–3):179–199, July 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

ElHakeem:2002:EEG

[EEE02] Abd El Hakeem, Abd El Naby, and A. E. M. El Misiery. Effects of an endoscope and generalized Newtonian fluid on peristaltic motion. Applied Mathematics and Computation, 128 (1):19–35, May 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Elshehawey:2003:EIM

[EEE03] E. F. Elshehawey, Elsayed M. E. Elbarbary, and Nasser S. Elgazery. Effect of inclined magnetic field on magneto fluid flow through a porous medium between two inclined wavy porous plates (numerical study). Applied Mathematics and Computation, 135(1):85–103, February 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Enna:2003:JER

[EEES03] A. A. El-Enna and M. I. El-Saftawy. The joint effect of relativistic and direct solar radiation pressure on an axially symmetric spacecraft: II. The solution and the elements of transformation. Applied Mathematics and Computation, 138(2– 3):247–265, June 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Elbarbary:2003:CFD

[EEK03a] Elsayed M. E. Elbarbary and M. El-Kady. Chebyshev finite difference approximation for the boundary value problems. Applied Mathematics and Computation, 139(2–3):513–523, July 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ezzat:2003:MTR

[EEK03b] Magdy A. Ezzat and Ahmed S. El-Karamany. Magnetothermoelasticity with two relaxation times in conducting medium with variable electrical and thermal conductivity. Applied Mathematics and Computation, 142(2–3):

449–467, October 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ezzat:2004:DME

[EEKS04] Magdy A. Ezzat, Ahmed S.

El-Karamany, and Angail A. Samaan. The dependence of the modulus of elasticity on reference temperature in generalized thermoelasticity with thermal relaxation. Applied Mathematics and Computation, 147(1): 169–189, January 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Eissa:2003:SPS

[EESESS03] M. Eissa, S. A. El-Serafi, M. El-Sheikh, and M. Sayed. Stability and primary simultaneous resonance of harmonically excited non-linear spring pendulum system. Applied Mathematics and Computation, 145(2–3):421– 442, December 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Eldabe:2004:TDD

[EESF04] N. T. Eldabe, A. G. El-Saka, and Ashraf Fouad. Thermaldiffusion and diffusionthermo effects on mixed freeforced convection and mass transfer boundary layer flow for non-Newtonian fluid with temperature dependent viscosity. Applied Mathematics and Computation, 152(3):867–883, May 13, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See comment [Pan08].

Eldabe:2004:EFH

[EESS04] N. T. Eldabe, M. El-Shahed, and M. Shawkey. An extension of the finite Hankel transform. Applied Mathematics and Computation, 151(3):713–717, April 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Elbarbary:2003:RCR

[EEYK03] Elsayed M. E. Elbarbary, Salah M. Elsayed, Ismail K. Youssef, and Ahmed M. M. Khodier. Restrictive Chebyshev rational approximation and applications to heatconduction problems. Applied Mathematics and Computation, 136(2–3):395–403, March 15, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Gohary:2002:GSR

[EG02] Awad El-Gohary. Global stabilization of a rotational motion of a rigid body using rotors system. Applied Mathematics and Computation, 133(2–3):297–307, December 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Gohary:2003:OCA

[EG03a] Awad El-Gohary. Optimal control of an angular motion of a rigid body during infinite and finite time intervals. Applied Mathematics and Computation, 141(2–3): 541–551, September 5, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Gohary:2003:OSR

[EG03b] Awad El-Gohary. Optimal stabilization of a rigid body motion using rotors system. Applied Mathematics and Computation, 136(2– 3):229–239, March 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Gohary:2004:BEPb

[EG04a] A. El-Gohary. Bayesian estimations of parameters in a three state reliability semi-Markov models. Applied Mathematics and Computation, 154(1):53–67, June 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Gohary:2004:EPT

[EG04b] A. El-Gohary. Estimations of parameters in a three state reliability semi-Markov model. Applied Mathematics and Computation, 154 (2):389–403, July 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Gohary:2004:CRB

[EG04c] A. El-Gohary. On the control of a rigid body motion affected by stochastic white Gaussian noises. Applied Mathematics and Computation, 151(1):69–80, March 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Gohary:2004:BEPc

[EG04d] Awad El-Gohary. Bayes estimation of parameters in a three non-independent component series system with time dependent failure rate. Applied Mathematics and Computation, 158(1):121– 132, October 25, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Gohary:2004:BEPa

[EG04e] Awad El-Gohary. Bayesian estimation of the parameters in two non-independent component series system with dependent time failure rate. Applied Mathematics and Computation, 154 (1):41–51, June 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Gohary:2004:ORB

[EG04f] Awad El-Gohary. On the orientation of a rigid body using point masses. Applied Mathematics and Computation, 151(1):163–179, March 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Elshehawey:2004:PTT

[EG04g] E. F. Elshehawey and Z. M.

Gharsseldien. Peristaltic transport of three-layered flow with variable viscosity. Applied Mathematics and Computation, 153 (2):417–432, June 4, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Gohary:2003:OCN

[EGAR03] Awad El-Gohary and A. S.

Al-Ruzaiza. Optimal control of non-homogeneous prey– predator models during infinite and finite time intervals. Applied Mathematics and Computation, 146(2–3): 495–508, December 31, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Gohary:2003:OCL

[EGB03a] Awad El-Gohary and F. Bukhari.

Optimal control of Lorenz system during different time intervals. Applied Mathematics and Computation, 144(2–3):337–351, December 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Gohary:2003:OCS

[EGB03b] Awad El-Gohary and Fawzy A. Bukhari. Optimal control of stochastic prey–predator models. Applied Mathematics and Computation, 146(2–3):403–415, December 31, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Gohary:2003:OSS

[EGB03c] Awad El-Gohary and Fawzy A. Bukhari. Optimal stabilization of steady-states of the genital herpes epidemic during infinite and finite time intervals. Applied Mathematics and Computation, 137(1):33–47, May 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Gamel:2003:NMS

[EGBH03] Mohamed El-Gamel, S. H. Behiry, and H. Hashish. Numerical method for the solution of special nonlinear fourth-order boundary value problems. Applied Mathematics and Computation, 145(2–3):717–734, December 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Gohary:2002:ECE

[EGE02] Awad El-Gohary and Ebrahim R.

Elazab. Exponential control of equilibrium positions of a rigid body using quaternions. Applied Mathematics and Computation, 132 (1):173–186, October 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Gohary:2003:ECR

[EGE03] Awad El-Gohary and Ebrahim R.

Elazab. Exponential control of a rotational motion of a rigid body using quaternions. Applied Mathematics and Computation, 137(2– 3):195–207, May 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Gohary:2004:OCR

[EGT04] A. I. El-Gohary and T. S.

Tawfik. Optimal control of the rotational motion of a rigid body using moving masses. Applied Mathematics and Computation, 153 (2):453–465, June 4, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Hawary:2003:SCM

[EHM03] H. M. El-Hawary and S. M.

Mahmoud. Spline collocation methods for solving delay-differential equations. Applied Mathematics and Computation, 146(2–3): 359–372, December 31, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ehrich:2002:PEB

[Ehr02] S. Ehrich. Pointwise error bounds for orthogonal cardinal spline approximation. Applied Mathematics and Computation, 128(2– 3):237–260, May 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Esen:2004:NSS

[EK04] A. Esen and S. Kutluay. A numerical solution of the Stefan problem with a Neumann-type boundary condition by enthalpy method. Applied Mathematics and Computation, 148(2):321–329, January 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Kady:2002:CEM

[EKE02] M. El-Kady and Elsayed M. E. Elbarbary. A Chebyshev expansion method for solving nonlinear optimal control problems. Applied Mathematics and Computation, 129(2–3):171–182, July 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Kady:2003:OCA

[EKE03] M. El-Kady and ElSayed M. E. ElBarbary. Optimal control approach for system of ordinary differential equations. Applied Mathematics and Computation, 135(2– 3):277–285, March 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Karamany:2004:BIE

[EKE04] Ahmed S. El-Karamany and Magdy A. Ezzat. Boundary integral equation formulation for the generalized thermoviscoelasticity with two relaxation times. Applied Mathematics and Computation, 151(2):347– 362, April 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Elshaer:2003:ESE

[El 03] Amal M. El shaer. Exact solution of equations of motion a polytropic gas. Applied Mathematics and Computation, 141(2–3):597–609, September 5, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Latif:2004:PML

[EL04a] G. M. Abd El-Latif. On a problem of modified Lindstedt–Poincaré for certain strongly non-linear oscillators. Applied Mathematics and Computation, 152 (3):821–836, May 13, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Latif:2004:PEF

[EL04b] G. M. Abd El-Latif. Parametric excitation for forcing van der Pol oscillator. Applied Mathematics and Computation, 147(1): 255–265, January 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Elbashbeshy:2001:LMC

[Elb01] E. M. A. Elbashbeshy. Laminar mixed convection over horizontal flat plate embedded in a non-Darcian porous medium with suction and injection. Applied Mathematics and Computation, 121(2– 3):123–128, June 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/105/25/
21/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/105/25/21/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S009630039900274X.

Elbashbeshy:2003:MCA

[Elb03] E. M. A. Elbashbeshy. The mixed convection along a vertical plate embedded in non-Darcian porous medium with suction and injection. Applied Mathematics and Computation, 136(1):139– 149, March 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Elshahed:2004:BFC

[Els04] Moustafa Elshahed. Blood flow in capillary under starling hypothesis. Applied Mathematics and Computation, 149(2):431–439, February 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Mikkawy:2003:NTT

[EM03a] Moawwad El-Mikkawy. A note on a three-term recurrence for a tridiagonal matrix. Applied Mathematics and Computation, 139(2– 3):503–511, July 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Mikkawy:2003:CBPa

[EM03b] Moawwad El-Mikkawy. On a connection between the Pascal, Vandermonde and Stirling matrices. I. Applied Mathematics and Computation, 145(1):23–32, December 20, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Mikkawy:2003:CBPb

[EM03c] Moawwad El-Mikkawy. On a connection between the Pascal, Vandermonde and Stirling matrices. II. Applied Mathematics and Computation, 146(2–3):759–769, December 31, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Mikkawy:2003:UAN

[EM03d] Moawwad El-Mikkawy. A unified approach to Newton– Cotes quadrature formulae. Applied Mathematics and Computation, 138(2–3):403– 413, June 20, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Mikkawy:2003:VIU

[EM03e] Moawwad El-Mikkawy. Vandermonde interpolation using special associated matrices. Applied Mathematics and Computation, 141(2–3): 589–595, September 5, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Mikkawy:2003:EIG

[EM03f] Moawwad E. A. El-Mikkawy. Explicit inverse of a generalized Vandermonde matrix. Applied Mathematics and Computation, 146(2–3): 643–651, December 31, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Mikkawy:2003:SLS

[EM03g] Moawwad E. A. El-Mikkawy.

On solving linear systems of the Pascal type. Applied Mathematics and Computation, 136(1):195–202, March 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Etchechoury:2003:NSP

[EM03h] M. Etchechoury and C. Muravchik. Nonstandard singular perturbation systems and higher index differentialalgebraic systems. Applied Mathematics and Computation, 134(2–3):323–344, January 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Egidi:2004:CST

[EM04a] N. Egidi and P. Maponi. A comparative study of two fast phase unwrapping algorithms. Applied Mathematics and Computation, 148(3):599–629, January 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Mikkawy:2004:NSM

[EM04b] Moawwad E. A. El-Mikkawy.

A note on the Stirling matrix of the second kind. Applied Mathematics and Computation, 151(1):147– 151, March 30, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Mikkawy:2004:IGT

[EM04c] Moawwad E. A. El-Mikkawy. On the inverse of a general tridiagonal matrix. Applied Mathematics and Computation, 150(3):669–679, March 17, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Mikkawy:2003:GFP

[EME03] M. E. A. El-Mikkawy and M. M. M. Eisa. A general four-parameter non-FSAL embedded Runge–Kutta algorithm of orders 6 and 4 in seven stages. Applied Mathematics and Computation, 143(2–3):259–267, November 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Mikkawy:2003:CBS

[EMED03] Moawwad El-Mikkawy and Beih El-Desouky. On a connection between symmetric polynomials, generalized Stirling numbers and the Newton general divided difference interpolation polynomial. Applied Mathematics and Computation, 138(2– 3):375–385, June 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Mikkawy:2003:NON

[EMR03] Moawwad El-Mikkawy and El-Desouky Rahmo. A new optimized non-FSAL embedded Runge–Kutta– Nystrom algorithm of orders 6 and 4 in six stages. Applied Mathematics and Computation, 145(1):33–43, December 20, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Maghraby:2004:SSA

[EMY04] Nasser M. El-Maghraby and Hamdy M. Yossef. State space approach to generalized thermoelastic problem with thermomechanical shock. Applied Mathematics and Computation, 156(2): 577–586, September 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ekart:2001:NGA

[EN01] Anikó Ekárt and S. Z.

Németh. A noncontinuous generalization of the arithmetic-geometric mean. Applied Mathematics and Computation, 124(2):261– 279, October 25, 2001. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/9/12/113/31/36/abstract.

                                        html;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300300000989.

El-Naggar:2004:PTS

[ENAAF04] A. M. El-Naggar, A. M. Abd-Alla, and M. A. Fahmy. The propagation of thermal stresses in an infinite elastic slab. Applied Mathematics and Computation, 157(2): 307–312, October 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Naggar:2001:ASE

[ENAAM01] A. M. El-Naggar, A. M. Abd-Alla, and S. R. Mahmoud. Analytical solution of electro-mechanical wave propagation in long bones. Applied Mathematics and Computation, 119(1):77–98, March 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/99/20/25/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/99/20/
                                        25/article.pdf;   http:
                                        //www.sciencedirect.com/
                                        science/article/pii/S0096300399002313.

El-Naby:2004:FDS

[ENEA04] M. A. Abd El-Naby, Elsayed M. E. Elbarbary, and Nader Y. AbdElazem. Finite difference solution of radiation effects on MHD unsteady free-convection flow over vertical porous plate. Applied Mathematics and Computation, 151(2):327– 346, April 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abd-alla:2002:MLW

[enNAaA02] Abo el-nour N. Abd-alla and Ibrahim A. A. Abbas. Magnetoelastic longitudinal wave propagation in a transversely isotropic circular cylinder. Applied Mathematics and Computation, 127(2–3):347–360, April 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/123/27/42/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300301000121.

Abd-alla:2004:RWMa

[enNAaAd04] Abo el-nour N. Abd-alla and S. M. Abo-dahab. Rayleigh waves in magneto-thermo-viscoelastic solid with thermal relaxation times. Applied Mathematics and Computation, 149(3):861–877, February 22, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Echarte:2004:RAI

[ENR04] F. J. Echarte, J. Núñez, and F. Ramírez. Relations among invariants of complex filiform Lie algebras. Applied Mathematics and Computation, 147(2): 365–376, January 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Owaidy:2004:GAR

[EOAE04] H. M. El-Owaidy, A. M. Ahmed, and Z. Elsady. Global attractivity of the recursive sequence x_{n+1} = (α − βx_{n−1})/(γ + x_{n}). Applied Mathematics and Computation, 151(3):827–833, April 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Owaidy:2003:RS

[EOAM03] H. M. El-Owaidy, A. M. Ahmed, and M. S. Mousa. On the recursive sequences x_{n+1} = \frac{−αx_{n−1}}{β±x_{n}}. Applied Mathematics and Computation, 145(2–3):747–753, December 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Owaidy:2004:ABD

[EOAM04] H. M. El-Owaidy, A. M. Ahmed, and M. S. Mousa. On asymptotic behaviour of the difference equation x_{n+1} = α + \frac{x_{n−k}}{x_{n}}. Applied Mathematics and Computation, 147(1):163–167, January 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Owaidy:2000:NPC

[EOEA00] H. M. El-Owaidy and M. M.

El-Afifi. A note on the periodic cycle of X_{n+2} = (1 + X_{n+1})/(X_{n}). Applied Mathematics and Computation, 109(2–3):301–306, March 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/82/23/
35/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/82/23/35/article.
pdf.

El-Owaidy:2002:MMB

[EOI02] H. M. El-Owaidy and M. Ismail. A mathematical model of bilingualism. Applied Mathematics and Computation, 131(2–3):415–432, September 25, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Owaidy:2003:GST

[EOI03] H. M. El-Owaidy and M. Ismail. Global stability for three-species Lotka–Volterra systems with delay. Applied Mathematics and Computation, 135(2–3):301–306, March 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Owaidy:2002:PSO

[EOM02] H. El-Owaidy and H. Y. Mohamed. On the periodic solutions for an nth-order difference equations. Applied Mathematics and Computation, 131(2–3):461–467, September 25, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Owaidy:2003:LON

[EOM03a] H. El-Owaidy and H. Y. Mohamed. Linearized oscillation for non-linear systems of delay differential equations. Applied Mathematics and Computation, 142(1): 17–21, September 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Owaidy:2003:NSC

[EOM03b] H. El-Owaidy and H. Y. Mohamed. The necessary and sufficient conditions of existence of periodic solutions of nonautonomous difference equations. Applied Mathematics and Computation, 136(2–3):345–351, March 15, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Owaidy:2003:GAS

[EOM03c] H. El-Owaidy and H. Y. Mohamed. On the global attractivity of systems of nonlinear difference equations. Applied Mathematics and Computation, 135(2– 3):377–382, March 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Owaidy:2003:PSO

[EOM03d] H. El-Owaidy and H. Y. Mohamed. On the periodic solutions for an nth order difference equations. Applied Mathematics and Computation, 135(2–3):383–390, March 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See erratum [EOM08].

El-Owaidy:2004:GAB

[EOM04a] H. M. El-Owaidy and A. A.

Moniem. Global asymptotic behavior of a chemostat model with delayed response in growth. Applied Mathematics and Computation, 147(1):147–161, January 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Owaidy:2004:SEP

[EOM04b] H. M. El-Owaidy and A. A.

Moniem. Switching effect of predation on global large size prey species exhibiting group defence. Applied Mathematics and Computation, 151(2):491–499, April 5, 2004. CODEN AMHCBQ.

ISSN 0096-3003 (print),

1873-5649 (electronic).

El-Owaidy:2008:EPS

[EOM08] H. El-Owaidy and H. Y. Mohamed. Erratum to “On the periodic solutions for an n-th order difference equations” [Appl. Math. Comput. 135 (2003) 383–390]. Applied Mathematics and Computation, 205(1):507, November 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [EOM03d].

El-Owaidy:2000:RSX

[EOREA00] H. M. El-Owaidy, A. A. Ragab, and M. M. El-Afifi. On the recursive sequence X_{n+1} = A/x_{n}^{p} + B/x_{n−1}^{q} + C/x_{n+2}^{s}. Applied Mathematics and Computation, 112(2– 3):277–290, June 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/86/23/
                                        30/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/86/23/30/article.
                                        pdf.

El-Owaidy:2001:MAF

[EORI01] H. M. El-Owaidy, A. A. Ragab, and M. Ismail. Mathematical analysis of a food-web model. Applied Mathematics and Computation, 121(2–3):155–167, June 15, 2001. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/105/25/
24/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/105/25/24/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002696.

Elsayed:2000:SDT

[EP00] Khaled M. Faud Elsayed and Harry G. Perros. The superposition of discretetime Markov renewal processes with an application to statistical multiplexing of bursty traffic sources. Applied Mathematics and Computation, 115(1):43–62, October 6, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/90/21/24/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/90/21/
24/article.pdf.

El-Rahman:2002:ESP

[ER02] Reda G. Abd El-Rahman.

Exact solution of the particle– cluster dynamic equation. Applied Mathematics and Computation, 132(1):105– 114, October 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Raheem:2003:MAC

[ER03] Zaki F. A. El-Raheem. Modification of the application of a contraction mapping method on a class of fractional differential equation. Applied Mathematics and Computation, 137(2–3):371– 374, May 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Reheem:2001:NSN

[ERN01] Zaki F. A. El-Reheem and A. H. Nasser. On the numerical study of nonlinear initial-boundary value problems or initial-value problems. Applied Mathematics and Computation, 119(1):1– 19, March 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/99/20/20/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/99/20/
                                        20/article.pdf;   http:
                                        //www.sciencedirect.com/
                                        science/article/pii/S0096300399002167.

El-Raheem:2003:SPD

[ERN03] Zaki F. A. El-Raheem and A. H. Nasser. On the spectral property of a Dirichlet problem with explosive factor. Applied Mathematics and Computation, 138(2– 3):355–374, June 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Sayed:2000:LPA

[ES00] Ahmed M. A. El-Sayed.

Laguerre polynomials of arbitrary (fractional) orders. Applied Mathematics and Computation, 109(1): 1–9, March 1, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/82/21/
21/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/82/21/21/article.
pdf.

El-Sayed:2002:IMC

[ES02a] Salah M. El-Sayed. Integral methods for computing solutions of a class of singular two-point boundary value problems. Applied Mathematics and Computation, 130(2–3):235–241, August 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Sayed:2002:MDM

[ES02b] Salah M. El-Sayed. The modified decomposition method for solving nonlinear algebraic equations. Applied Mathematics and Computation, 132(2–3):589–597, November 10, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Shahed:2003:PFB

[ES03a] Moustafa El-Shahed. Pulsatile flow of blood through a stenosed porous medium under periodic body acceleration. Applied Mathematics and Computation, 138(2– 3):479–488, June 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Sobky:2003:GCT

[ES03b]  Bothina El-Sobky.  A

global convergence theory for an active-trust-region algorithm for solving the general nonlinear programming problem. Applied Mathematics and Computation, 144 (1):127–157, November 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Sayied:2004:SGC

[ES04] H. K. El-Sayied. Study on generalized convex fuzzy bodies. Applied Mathematics and Computation, 152 (1):1–8, April 26, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Sayed:2003:CAD

[ESAA03] Salah M. El-Sayed and Mohammedi R. Abdel-Aziz. A comparison of Adomian’s decomposition method and wavelet-Galerkin method for solving integro-differential equations. Applied Mathematics and Computation, 136(1):151–159, March 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Sayed:2004:PDS

[ESAD04] Salah M. El-Sayed and Asmaa M. Al-Dbiban. On positive definite solutions of the nonlinear matrix equation X + A^{∗}X^{−n}A = I. Applied Mathematics and Computation, 151(2):533– 541, April 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Escultura:2002:FTG

[Esc02]  E. E. Escultura.  The

flux theory of gravitation V: the mathematics of the new physics. Applied Mathematics and Computation, 130 (1):145–169, July 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Escalante:2003:PSS

[Esc03a] René Escalante. Parallel strategies for the step by step Tau method. Applied Mathematics and Computation, 137(2–3):277–292, May 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Escultura:2003:FTGa

[Esc03b] E. E. Escultura. The flux theory of gravitation XVII. The new mathematics and physics. Applied Mathematics and Computation, 138 (1):127–149, June 1, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Escultura:2003:FTGb

[Esc03c] E. E. Escultura. The flux theory of gravitation XVIII: macro and quantum gravity, cosmo waves and applications. Applied Mathematics and Computation, 139(1):23–36, July 1, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Sayed:2002:SPE

[ESEA02] Salah M. El-Sayed and Mahmoud El-Alem. Some properties for the existence of a positive definite solution of matrix equation X + A^{∗}X^{−2^{m}}A = I. Applied Mathematics and Computation, 128(1):99–108, May 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Sayed:2003:SUI

[ESEBD03] W. G. El-Sayed, A. A. El-Bary, and M. A. Darwish. Solvability of Urysohn integral equation. Applied Mathematics and Computation, 145(2–3):487–493, December 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Saftawy:2002:RDS

[ESEE02] M. I. El-Saftawy and A. A.

El-Enna. The relativistic and direct solar radiation pressure effects on an axially symmetric spacecraft. I. The Hamiltonian and the canonical equations of motion. Applied Mathematics and Computation, 132(2–3):505–514, November 10, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-seoud:2003:ASH

[EsEKEA03] A. A. M. Abou El-seoud, M. M. El-Kady, and M. A. El-Ameen. On approximate solution of Hammerstein integral equations in the space L_{p}(p ≥ 1). Applied Mathematics and Computation, 140(1):91–104, July 30, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Sayed:2004:FIE

[ESES04] Wagdy G. El-Sayed and Ahmed M. A. El-Sayed. On the functional integral equations of mixed type and integro-differential equations of fractional orders. Applied Mathematics and Computation, 154 (2):461–467, July 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Sayed:2003:FCS

[ESG03] Ahmed M. A. El-Sayed and Fatma M. Gaafar. Fractional calculus and some intermediate physical processes. Applied Mathematics and Computation, 144(1):117– 126, November 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Shehawey:2002:PTM

[ESH02] E. F. El-Shehawey and Saleh Z. A. Husseny. Peristaltic transport of a magnetofluid with porous boundaries. Applied Mathematics and Computation, 129(2– 3):421–440, July 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Sayed:2001:SVI

[ESI01] Ahmed M. A. El-Sayed and A.-G. Ibrahim. Setvalued integral equations of fractional-orders. Applied Mathematics and Computation, 118(1):113–121, February 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/96/20/26/abstract.
                                        html;   http://www.elsevier.

nl/gej-ng/10/9/12/96/20/
26/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399000879.

El-Sayed:2004:AAS

[ESK04a] Salah M. El-Sayed and Doğan Kaya. An application of the ADM to sevenorder Sawada–Kotara equations. Applied Mathematics and Computation, 157(1): 93–101, September 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Sayed:2004:CNM

[ESK04b] Salah M. El-Sayed and Doğan Kaya. Comparing numerical methods for Helmholtz equation model problem. Applied Mathematics and Computation, 150 (3):763–773, March 17, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Sayed:2004:DMS

[ESK04c] Salah M. El-Sayed and Doğan Kaya. The decomposition method for solving (2 + 1)-dimensional Boussinesq equation and (3 + 1)- dimensional KP equation. Applied Mathematics and Computation, 157(2):523– 534, October 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Sayed:2004:NSS

[ESK04d] Salah M. El-Sayed and Doğan Kaya. On the numerical solution of the system of two-dimensional Burgers’ equations by the decomposition method. Applied Mathematics and Computation, 158(1):101–109, October 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Sheikh:2000:OCS

[ESS00a] M. M. A. El-Sheikh and R. Sallam. Oscillation criteria for second order functional differential equations. Applied Mathematics and Computation, 115(2–3):113–121, October 27, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/90/23/23/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/90/23/
23/article.pdf.

El-Sheikh:2000:SNSa

[ESS00b] M. M. A. El-Sheikh and A. A. Soliman. On stability of nonlinear systems of functional differential equations. Applied Mathematics and Computation, 107(2– 3):81–93, January 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/72/22/
22/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/72/22/22/article.
pdf.

El-Shahed:2004:GNS

[ESS04] Moustafa El-Shahed and Ahmed Salem. On the generalized Navier–Stokes equations. Applied Mathematics and Computation, 156(1): 287–293, August 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Sheikh:2000:SNSb

[ESSA00] M. M. A. El-Sheikh, A. A.

Soliman, and M. H. Abd Alla. On stability of nonlinear systems of ordinary differential equations. Applied Mathematics and Computation, 113(2–3):175–198, July 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/87/23/24/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/87/23/
24/article.pdf.

El-Sheikh:2001:SND

[ESSA01] M. M. A. El-Sheikh, A. A.

Soliman, and M. H. Abd Alla. On stability of nonlinear differential systems via cone-valued Liapunov function method. Applied Mathematics and Computation, 119(2–3):265–281, April 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/99/25/31/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/99/25/
                                        31/article.pdf;   http:
                                        //www.sciencedirect.com/
                                        science/article/pii/S0096300399002349.

El-Sayed:2004:NOF

[ESSEF04] Ahmed M. A. El-Sayed, Nagwa Sherif, and Ibrahim Abou El-Farag. A nonlinear operator functional equation of Volterra type. Applied Mathematics and Computation, 148(3):665–679, January 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tatar:2004:BWE

[eT04] Nasser eddine Tatar. Blow up for the wave equation with a nonlinear dissipation of cubic convolution type in R^{N}. Applied Mathematics and Computation, 148(3): 759–771, January 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Tawil:2004:SRD

[ETBAN04] Magdy A. El-Tawil, Ahmed A. Bahnasawi, and Ahmed Abdel-Naby. Solving Riccati differential equation using Adomian’s decomposition method. Applied Mathematics and Computation, 157(2):503–514, October 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ezzat:2004:FCE

[Ezz04] Magdy A. Ezzat. Free convection effects on extracellular fluid in the presence of a transverse magnetic field. Applied Mathematics and Computation, 151 (2):455–482, April 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Fadrani:2004:SPD

[Fad04] D. Rostamy Vamas Fadrani.

A stochastic partial differential equation for computational algorithms. Applied Mathematics and Computation, 159(2):429–434, December 6, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Fan:2003:ILC

[Fan03] Chun-I Fan. Improved lowcomputation partially blind signatures. Applied Mathematics and Computation, 145(2–3):853–867, December 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Faruqi:2000:NLM

[Far00] Farhan A. Faruqi. Nonlinear mathematical model for integrated global positioning/inertial navigation systems. Applied Mathematics and Computation, 115(2–3):191–212, October 27, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/90/23/29/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/90/23/
                                        29/article.pdf.

Farkas:2002:SSA

[Far02] M. Farkas. On the stability of stationary age distributions. Applied Mathematics and Computation, 131 (1):107–123, September 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Faragallah:2003:IWO

[Far03] M. Faragallah. Interpolation of weighted Orlicz spaces. Applied Mathematics and Computation, 145(2–3): 613–622, December 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Farkas:2004:SCN

[Far04] J. Z. Farkas. Stability conditions for the non-linear McKendrick equations. Applied Mathematics and Computation, 156(3):771–777, September 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Fatullayev:2004:NSI

[Fat04] Afet Golayoğlu Fatullayev.

Numerical solution of the inverse problem of determining an unknown source term in a two-dimensional heat equation. Applied Mathematics and Computation, 152(3):659–666, May 13, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Feng:2003:EUA

[Fen03] Chunhua Feng. On the existence and uniqueness of almost periodic solutions for delay Logistic equations. Applied Mathematics and Computation, 136(2–3):487– 494, March 15, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Feng:2004:ASG

[Fen04a] Zhaosheng Feng. An approximate sine-Gordon equation and its traveling wave solution in (n + 1)-dimensional space. Applied Mathematics and Computation, 152 (2):597–610, May 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Feng:2004:CEA

[Fen04b] Zhaosheng Feng. Comment on “On the extended applications of Homogeneous Balance Method”. Applied Mathematics and Computation, 158(2):593–596, November 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [Sen01].

Fan:2003:AEM

[FH03] Engui Fan and Y. C. Hon. Applications of extended tanh method to ‘special’ types of nonlinear equations. Applied Mathematics and Computation, 141(2–3): 351–358, September 5, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Faghihi:2004:NSC

[FH04] F. Faghihi and K. Hadad. Numerical solutions of coupled differential equations and initial values using Maple software. Applied Mathematics and Computation, 155(2):563–572, August 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Fujita:2001:ODM

[FI01] Toshiharu Fujita and Seiichi Iwamoto. An optimistic decision-making in fuzzy environment. Applied Mathematics and Computation, 120(1–3):123–137, May 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/104/21/
31/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/104/21/31/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002386.

Fan:2003:DPP

[FK03] Y. Fan and R. Kalaba.

Dynamic programming and pseudo-inverses. Applied Mathematics and Computation, 139(2–3):323–342, July 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ferm:2002:NEB

[FL02a] Lars Ferm and Per Lötstedt.

On numerical errors in the boundary conditions of the Euler equations. Applied Mathematics and Computation, 128(1):129–140, May 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ford:2002:CSS

[FL02b] Neville J. Ford and Sjoerd M. Verduyn Lunel. Characterising small solutions in delay differential equations through numerical approximations. Applied Mathematics and Computation, 131(2–3):253–270, September 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Fan:2003:EPP

[FL03] Guihong Fan and Yongkun Li. Existence of positive periodic solutions for a periodic logistic equation. Applied Mathematics and Computation, 139(2–3):311–321, July 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Franco:2004:ESF

[FNO04] Daniel Franco, Juan J. Nieto, and Donal O’Regan. Existence of solutions for first order ordinary differential equations with nonlinear boundary conditions. Applied Mathematics and Computation, 153(3):793– 802, June 14, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Feng:2004:SAG

[FQ04] Enmin Feng and Weiyi Qian. A stochastic approach to global optimization of nonlinear programming problem with many equality constraints. Applied Mathematics and Computation, 155 (1):111–120, July 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Filshtinsky:2002:SCP

[FRRSCS02] M. L. Filshtinsky, R. Rodriguez-Ramos, O. Sanchez-Casals, and Florencia Serrania. Squeezing of composite piezoceramic plate under the loading of regular normal pressure. Applied Mathematics and Computation, 129(2– 3):407–419, July 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Frontini:2003:SVN

[FS03] M. Frontini and E. Sormani. Some variant of Newton’s method with thirdorder convergence. Applied Mathematics and Computation, 140(2–3):419–426, August 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Frontini:2004:TOM

[FS04a] M. Frontini and E. Sormani. Third-order methods from quadrature formulae for solving systems of nonlinear equations. Applied Mathematics and Computation, 149(3):771–782, February 22, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Fu:2004:OCI

[FS04b] Xilin Fu and LieJune Shiau. Oscillation criteria for impulsive parabolic boundary value problem with delay. Applied Mathematics and Computation, 153(2):587– 599, June 4, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Faye:2003:LTF

[FSLMC03] Roger Marcelin Faye, Salam Sawadogo, Claude Lishou, and Félix Mora-Camino. Long-term fuzzy management of water resource systems. Applied Mathematics and Computation, 137(2– 3):459–475, May 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Faruqi:2000:EKF

[FT00] Farhan A. Faruqi and Kenneth J. Turner. Extended Kalman filter synthesis for integrated global positioning/inertial navigation systems. Applied Mathematics and Computation, 115(2–3):213–227, October 27, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/90/23/30/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/90/23/
30/article.pdf.

Fu:2003:CNF

[Fu03] Xianlong Fu. Controllability of neutral functional differential systems in abstract space. Applied Mathematics and Computation, 141(2–3): 281–296, September 5, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Fu:2004:CAN

[Fu04] Xianlong Fu. Controllability of abstract neutral functional differential systems with unbounded delay. Applied Mathematics and Computation, 151 (2):299–314, April 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Fuyong:2004:OCS

[Fuy04]  Lin Fuyong.  Orthogo-

nal continuous segmentation polynomial. Applied Mathematics and Computation, 154(3):599–607, July 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Fan:2004:KSN

[FZ04a] Xiaoming Fan and Shengfan Zhou. Kernel sections for non-autonomous strongly damped wave equations of non-degenerate Kirchhofftype. Applied Mathematics and Computation, 158(1): 253–266, October 25, 2004.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Fu:2004:FOI

[FZ04b] Xilin Fu and Liqin Zhang. Forced oscillation for impulsive hyperbolic boundary value problems with delay. Applied Mathematics and Computation, 158(3):761– 780, November 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ghazal:2004:TSP

[GA04] M. A. Ghazal and A. Mitwalli Aly. Time series with Poisson point process. Applied Mathematics and Computation, 150(1):149–157, February 27, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Galligani:2003:NAM

[Gal03] Emanuele Galligani. The Newton-arithmetic mean method for the solution of systems of nonlinear equations. Applied Mathematics and Computation, 134 (1):9–34, January 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Garey:2001:EAM

[Gar01a] L. E. Garey. An efficient algorithm for a model with a bidiagonal coefficient matrix. Applied Mathematics and Computation, 122(3): 373–383, August 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/106/26/
28/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/106/26/28/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000527.

Garey:2001:PNA

[Gar01b]  L. E. Garey.  A par-

allel numerical algorithm for near symmetric and banded systems. Applied Mathematics and Computation, 119(1):99–108, March 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/99/20/26/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/99/20/
26/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S009630039900226X.

Galue:2003:FRG

[GAZK03] L. Galué, A. Al-Zamel, and S. L. Kalla. Further results on generalized hypergeometric functions. Applied Mathematics and Computation, 136 (1):17–25, March 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gorain:2002:ECB

[GB02] Ganesh C. Gorain and Sujit K. Bose. Exact controllability and boundary stabilization of flexural vibrations of an internally damped flexible space structure. Applied Mathematics and Computation, 126(2–3):341–360, March 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        9/12/122/30/40/abstract.
                                        html;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300300001120.

Gao:2003:APS

[GB03] Hongjun Gao and Charles Bu. Almost periodic solution for a model of tumor growth. Applied Mathematics and Computation, 140 (1):127–133, July 30, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gu:2001:BGM

[GC01] Gui-Ding Gu and Zhi-Hao Cao. A block GMRES method augmented with eigenvectors. Applied Mathematics and Computation, 121(2–3):271–289, June 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/105/25/
32/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/105/25/32/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002945.

Gokhan:2004:DSS

[GÇ04] A. Gökhan and R. Çolak.

The double sequences spaces c_{2}^{P} (p) and c_{2}^{P B}(p). Applied Mathematics and Computation, 157(2):491–501, October 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ghosh:2004:MSC

[GCSS04] Mini Ghosh, Peeyush Chandra, Prawal Sinha, and J. B. Shukla. Modelling the spread of carrier-dependent infectious diseases with environmental effect. Applied Mathematics and Computation, 152(2):385–402, May 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gungor:2004:SSS

[GEA04] Mehmet Güngör, Mikâil Et, and Yavuz Altin. Strongly (V_{σ}, λ, q)-summable sequences defined by Orlicz functions. Applied Mathematics and Computation, 157(2):561– 571, October 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Guan:2000:GEU

[GF00] Ping Guan and Jishan Fan. On the global existence and uniqueness of weak solutions to the nonstationary semiconductor equations. Applied Mathematics and Computation, 114(2–3):125–133, September 11, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/89/22/21/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/89/22/
                                        21/article.pdf.

Garey:2000:BPC

[GG00] L. E. Garey and C. J. Gladwin. Best predictorcorrector methods for first order VIDEs with solutions damped at infinity. Applied Mathematics and Computation, 113(1):81–95, July 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/87/21/26/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/87/21/
                                        26/article.pdf.

Golan:2002:GMI

[GG02a] Amos Golan and Henryk Gzyl. A generalized maxentropic inversion procedure for noisy data. Applied Mathematics and Computation, 127(2–3):249–260, April 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/123/27/35/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001727.

Gune:2002:DSM

[GG02b] Mustafa Güne and Semin Güne. A direct search method for determination of DAEM kinetic parameters from nonisothermal TGA data (note). Applied Mathematics and Computation, 130(2–3):619–628, August 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gao:2003:NBV

[GGB03] Hongjun Gao, Xiaohua Gu, and Charles Bu. A Neumann boundary value problem for a generalized Ginzburg– Landau equation. Applied Mathematics and Computation, 134(2–3):553–560, January 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Galli:2003:SPT

[GGRS03] M. Galli, M. Groppi, R. Riganti, and G. Spiga. Singular perturbation techniques in the study of a diatomic gas with reactions of dissociation and recombination. Applied Mathematics and Computation, 146(2–3):509–531, December 31, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gutierrez:2001:ANM

[GH01] J. M. Gutiérrez and M. A. Hernández. An acceleration of Newton’s method: Super-Halley method. Applied Mathematics and Computation, 117(2–3):223–239, January 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/92/23/27/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/92/23/
                                        27/article.pdf;   http:
                                        //www.sciencedirect.com/
                                        science/article/pii/S0096300399001757.

Gao:2003:USP

[GH03] Jiaquan Gao and Guixia He. An unconditionally stable parallel difference scheme for parabolic equations. Applied Mathematics and Computation, 135(2–3):391–398, March 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ghoreishi:2004:PIP

[GH04] F. Ghoreishi and S. Mohammad Hosseini. A preconditioned implementation of pseudospectral methods on arbitrary grids. Applied Mathematics and Computation, 148(1):15–34, January 20, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ghazal:2001:SAB

[Gha01] M. A. Ghazal. Statistical analysis of broanded periodogram for continuous time stationary processes. Applied Mathematics and Computation, 124(3):343–349, December 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/113/32/32/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000990.

Ghazal:2003:SAS

[Gha03]  M. A. Ghazal.  Statis-

tical analysis for stationary time processes with irregular observations. Applied Mathematics and Computation, 134(2–3):363–370, January 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ghezzi:2000:BMM

[Ghe00] Luca L. Ghezzi. Bond management and max-min optimal control. Applied Mathematics and Computation, 112(1):33–40, June 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/86/21/
23/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/86/21/23/article.
pdf.

Gunzburger:2000:SEP

[GHL00] Max D. Gunzburger, Matthias Heinkenschloss, and Hyesuk Kwon Lee. Solution of elliptic partial differential equations by an optimization-based domain decomposition method. Applied Mathematics and Computation, 113(2–3):111–139, July 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/87/23/21/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/87/23/
21/article.pdf.

Ghosh:2003:ELF

[Gho03] Mini Ghosh. Effect of liming on a fish population in an acidified lake: a simple mathematical model. Applied Mathematics and Computation, 135(2–3):553–560, March 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Guoqiang:2000:EMI

[GHSJ00] Han Guoqiang, Ken Hayami, Kokichi Sugihara, and Wang Jiong. Extrapolation method of iterated collocation solution for two-dimensional nonlinear Volterra integral equations. Applied Mathematics and Computation, 112(1):49–61, June 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/86/21/
25/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/86/21/25/article.
pdf.

Gao:2004:AAA

[GhW04a] Feng Gao and Ren hong Wang. Algebraic approximants to exp(z) and applications in construction of difference schemes of first order ODE. Applied Mathematics and Computation, 149(2): 469–474, February 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gao:2004:SSL

[GHW04b] Ziyou Gao, Guoping He, and Fang Wu. Sequential systems of linear equations algorithm for nonlinear optimization problems— -general constrained problems. Applied Mathematics and Computation, 147(1): 211–226, January 5, 2004.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Giannantoni:2003:PIC

[Gia03] Corrado Giannantoni. The problem of the initial conditions and their physical meaning in linear differential equations of fractional order. Applied Mathematics and Computation, 141 (1):87–102, August 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gine:2004:AIC

[Gin04] Jaume Giné. Analytic integrability and characterization of centers for generalized nilpotent singular points. Applied Mathematics and Computation, 148(3): 849–868, January 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gorka:2002:PVM

[GK02] A. Gorka and M. Kostreva. Probabilistic version of the method of feasible directions. Applied Mathematics and Computation, 130(2–3): 253–264, August 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Guyer:2001:NMC

[GKaM01] Tolga Güyer, Onur Kiymaz, and Göksal Bilgici andŞeref Mirasyedioğlu. A new method for computing the solutions of differential equation systems using generalized inverse via Maple. Applied Mathematics and Computation, 121(2– 3):291–299, June 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/105/25/
33/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/105/25/33/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000035.

Garg:2002:SCG

[GKK02] Mridula Garg, Vimal Katta, and Shyam L. Kalla. Study of a class of generalized elliptic type integrals. Applied Mathematics and Computation, 131(2–3):607–613, September 25, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gunzburger:2000:PLS

[GL00] Max D. Gunzburger and Hyung-Chun Lee. A penalty/least-squares method for optimal control problems for first-order elliptic systems. Applied Mathematics and Computation, 107(1):57–75, January ??, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/72/17/19/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/72/17/
                                        19/article.pdf.

Ghorbanzadeh:2001:BAD

[GL01] Dariush Ghorbanzadeh and Rachid Lounes. Bayesian analysis for detecting a change in exponential family. Applied Mathematics and Computation, 124(1): 1–15, November 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.elsevier.com/gej-
                                        ng/10/9/12/113/27/27/abstract.
                                        html;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300300000291.

Glabisz:2004:DWW

[Gla04] W. Glabisz. Direct Walshwavelet packet method for variational problems. Applied Mathematics and Computation, 159(3):769–781, December 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2004:CAC

[gLgW04] Xin guo Liu and Wei guo Wang. On a class of alternating coefficient matrices quadratic eigenvalue problem. Applied Mathematics and Computation, 158 (3):619–636, November 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Garcia:2004:NIO

[GLM04] Amelia García, David J.

López, and Pablo Martín. On the numerical integration of oscillatory problems with multistep codes. Applied Mathematics and Computation, 153(2):497– 503, June 4, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Garcia-Lopez:2002:SPL

[GLR02] C. M. García-López and J. I. Ramos. S-stability of piecewise-linearized and linearized θ-methods. Applied Mathematics and Computation, 132(2–3):617–631, November 10, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gao:2000:BLV

[GLVW00] Tangan Gao, T. Y. Li, Jan Verschelde, and Mengnien Wu. Balancing the lifting values to improve the numerical stability of polyhedral homotopy continuation methods. Applied Mathematics and Computation, 114(2–3):233–247, September 11, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.

                                        elsevier.nl/gej-ng/10/
                                        9/12/89/22/29/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/89/22/
                                        29/article.pdf.

Gao:2004:CEA

[GLWY04] Ziyou Gao, W. H. K. Lam, S. C. Wong, and H. Yang. The convergence of equilibrium algorithms with nonmonotone line search technique. Applied Mathematics and Computation, 148 (1):1–13, January 20, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gopalsamy:2000:CSA

[GM00] K. Gopalsamy and S. Mohamad. Canonical solutions and almost periodicity in a discrete logistic equation. Applied Mathematics and Computation, 113(2–3):305– 323, July 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/87/23/31/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/87/23/
                                        31/article.pdf.

Gumel:2003:QSV

[GM03a] A. B. Gumel and S. M. Moghadas. A qualitative study of a vaccination model with non-linear incidence. Applied Mathematics and Computation, 143(2–3):409– 419, November 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Guyer:2003:CAA

[GM03b] Tolga Güyer and Şeref Mirasyedioğlu. A computational approach for the analytical solving of a Dirichlettype problem for third order partial differential equations. Applied Mathematics and Computation, 137 (1):139–150, May 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gonzalez:2004:LED

[GM04a] C. Sánchez González and T. M. García Muñoz. Linear estimation for discrete systems with uncertain observations: an application to the correction of declared incomes in inquiry. Applied Mathematics and Computation, 156(1):211–233, August 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Guyer:2004:SCA

[GM04b] Tolga Güyer and Şeref Mirasyedioğlu. A symbolic computation approach to a three-dimensional inverse problem for the transport equation. Applied Mathematics and Computation, 150(1):181–193, February 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gouesbet:2001:CEK

[GMGC01] G. Gouesbet and S. Meunier-Guttin-Cluzel. Computer evaluation of Kauffman polynomials by using Gauss codes, with a skein-template algorithm. Applied Mathematics and Computation, 122(2):229–252, July 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/106/25/
25/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/106/25/25/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000308.

Gewali:2002:GAF

[GNS02] L. Gewali, S. Ntafos, and A. K. Singh. Geometric approach for finding HPDcredible sets with applications. Applied Mathematics and Computation, 125(2–3): 195–207, January 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/9/12/120/31/30/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S009630030000120X.

Gonzalez:2004:PIF

[Gon04] Hernán A. González. On predicting incompressible flows by using a stabilized finite difference method with penalty. Applied Mathematics and Computation, 156 (2):439–454, September 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gopalsamy:2004:SAN

[Gop04]  K. Gopalsamy.  Stabil-

ity of artificial neural networks with impulses. Applied Mathematics and Computation, 154(3):783–813, July 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ghezzi:2003:SVA

[GP03] Luca L. Ghezzi and Carlo Piccardi. Stock valuation along a Markov chain. Applied Mathematics and Computation, 141(2–3):385–393, September 5, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gourley:2003:STD

[GR03] Stephen A. Gourley and Shigui Ruan. Spatiotemporal delays in a nutrientplankton model on a finite domain: linear stability and bifurcations. Applied Mathematics and Computation, 145(2–3):391–412, December 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Guo:2004:LSG

[GR04] Hui Guo and Hongxing Rui. Least-squares Galerkin procedures for parabolic integro-differential equations. Applied Mathematics and Computation, 150 (3):749–762, March 17, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Grabski:2003:ROS

[Gra03] Franciszek Grabski. The reliability of an object with semi-Markov failure rate. Applied Mathematics and Computation, 135(1):1–16, February 15, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Garey:2000:SBN

[GS00] L. E. Garey and R. E. Shaw. Solving banded and near symmetric systems. Applied Mathematics and Computation, 115(2–3):133–143, October 27, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/90/23/25/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/90/23/
25/article.pdf.

Guan:2002:EWS

[GS02] Ping Guan and Fuqing Sun.

Existence of weak solutions to a system of avalanchesemiconductor equations. Applied Mathematics and Computation, 133(2–3):615– 622, December 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gulle:2004:NSL

[Gül04a] Aytekin Gülle. Numerical solution of linear wave equation with strong dissipative term. Applied Mathematics and Computation, 159(2): 341–348, December 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gulle:2004:NSQ

[Gül04b] Aytekin Gülle. On the numerical solution of quasilinear wave equation with strong dissipative term. Applied Mathematics and Computation, 151(2):581–588, April 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gumel:2002:CNM

[Gum02] A. B. Gumel. A competitive numerical method for a chemotherapy model of two HIV subtypes. Applied Mathematics and Computation, 131(2–3):329–337, September 25, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gungor:2004:LBN

[Gün04] A. Dilek Güngör. Lower bounds for the norms of Cauchy–Toeplitz and Cauchy–Hankel matrices. Applied Mathematics and Computation, 157(3):599– 604, October 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Guo:2000:MNS

[Guo00] Dajun Guo. Minimal nonnegative solutions for nth order integro-differential equations in Banach spaces. Applied Mathematics and Computation, 113(1):55–65, July 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/87/21/24/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/87/21/
24/article.pdf.

Guo:2002:CSO

[Guo02]  Dajun Guo.  A class

of second-order impulsive integro-differential equations on unbounded domain in a Banach space. Applied Mathematics and Computation, 125(1):59–77, January 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/120/27/31/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001156.

Guo:2003:BVP

[Guo03a] Dajun Guo. A boundary value problem for nth order integro-differential equations in a Banach space. Applied Mathematics and Computation, 136(2–3):571–592, March 15, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Guo:2003:MPS

[Guo03b] Dajun Guo. Multiple positive solutions for first order nonlinear impulsive integrodifferential equations in a Banach space. Applied Mathematics and Computation, 143(2–3):233–249, November 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Guo:2004:MPS

[Guo04] Dajun Guo. Multiple positive solutions for a class of nonlinear integro-differential equations in Banach spaces. Applied Mathematics and Computation, 154(2):469– 485, July 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gzyl:2000:LRP

[GV00] Henryk Gzyl and Yurayh Velasquez. Linear reconstruction problems with convex constraints: influence of the a priori data. Applied Mathematics and Computation, 109(2–3):189–198, March 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/82/23/
                                        27/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/82/23/27/article.
                                        pdf.

Gzyl:2004:PAR

[GV04] Henryk Gzyl and Minaya Villasana. A perturbative approach for reconstructing diffusion coefficients. Applied Mathematics and Computation, 154 (1):1–15, June 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:2004:PAG

[gWxZ04] Wei guo Wang and Jin xi Zhao. Perturbation analysis for the generalized Cholesky factorization. Applied Mathematics and Computation, 147(2):601– 606, January 12, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Guo:2004:AOB

[GXL04] Mengshu Guo, Xiaoping Xue, and Ronglu Li. The asymptotic and oscillatory behavior of neutral difference inclusions. Applied Mathematics and Computation, 150(1):115–128, February 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gupta:2000:HAM

[GZ00] Murli M. Gupta and Jun Zhang. High accuracy multigrid solution of the 3D convection-diffusion equation. Applied Mathematics and Computation, 113(2– 3):249–274, July 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.nl/gej-ng/
10/9/12/87/23/28/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/87/23/
28/article.pdf.

Zhang:2004:GBI

[gZC04] Sheng gui Zhang and Sui Sun Cheng. Gröbner basis for an ideal of a polynomial ring over an algebraic extension over a field and its applications. Applied Mathematics and Computation, 153(1):27–58, May 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gzyl:2001:IPA

[Gzy01]  Henryk Gzyl.  Inverse

problem for the acoustic wave equation: a probabilistic approach to approximations and uniqueness. Applied Mathematics and Computation, 122 (2):179–194, July 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/9/12/106/25/
                                        23/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/9/12/106/25/23/article.
                                        pdf;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300300000229.

Gzyl:2002:TRM

[Gzy02] Henryk Gzyl. Tomographic reconstruction by maximum entropy in the mean: unconstrained reconstructions. Applied Mathematics and Computation, 129(2–3):157– 169, July 10, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

He:2002:EPS

[HA02] Huawen He and Malwane M. A. Ananda. Estimation of population size in closed animal populations from mark-resighting surveys. Applied Mathematics and Computation, 125(2–3): 387–398, January 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/9/12/120/31/46/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001454.

Hadizadeh:2003:NCN

[HA03] M. Hadizadeh and A. H.

Amiraslani. Numerical computation of the nonlinear feedback operators for the nonquadratic time-variant optimal control problems. Applied Mathematics and Computation, 144(1):159– 168, November 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hosseini:2004:SCR

[HA04] S. M. Hosseini and N. Aliev.

Sufficient conditions for the reduction of a BVP for PDE with non-local and global boundary conditions to Fredholm integral equations (on a rectangular domain). Applied Mathematics and Computation, 147(3): 669–685, January 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Habiba:2004:HCU

[Hab04a] El Zohny Habiba. Homology and chaotic unfolding of chaos manifolds. Applied Mathematics and Computation, 155(2):507–513, August 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Habiba:2004:ZDS

[Hab04b] El Zohny Habiba. On zero dimensional space and its retraction. Applied Mathematics and Computation, 155 (2):515–519, August 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Haidar:2000:ETD

[Hai00] Nassar H. S. Haidar. On an existence theory for the double series inverse. Applied Mathematics and Computation, 112(1):125–132, June 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/86/21/
                                        31/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/86/21/31/article.
                                        pdf.

Hayat:2004:SNL

[HAM04] T. Hayat, Naseer Ahmad, and F. M. Mahomed. On solution of non-linear differential equation arising in gliding motion of bacteria. Applied Mathematics and Computation, 148(3): 743–752, January 30, 2004.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See comment [HES05].

Han:2001:MMC

[Han01] Danfu Han. The majorant method and convergence for solving nondifferentiable equations in Banach space. Applied Mathematics and Computation, 118(1):73–82, February 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/96/20/23/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/96/20/
23/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399001836.

Hartfiel:2003:BI

[Har03] D. J. Hartfiel. Boxed iterations. Applied Mathematics and Computation, 143 (1):61–71, October 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hasanov:2000:QBSa

[Has00] Alemdar Hasanov. Qualitative behaviour of solutions of unilateral elliptic problems with perturbing the unknown boundary I. The theory. Applied Mathematics and Computation, 109(2–3): 249–260, March 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/82/23/
                                        31/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/82/23/31/article.
                                        pdf.

Hassan:2002:DAD

[Has02a] I. H. Abdel-Halim Hassan. Different applications for the differential transformation in the differential equations. Applied Mathematics and Computation, 129(2– 3):183–201, July 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hassan:2002:SSE

[Has02b] I. H. Abdel-Halim Hassan. On solving some eigenvalue problems by using a differential transformation. Applied Mathematics and Computation, 127(1):1–22, March 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        9/12/123/28/27/abstract.
                                        html;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300300001235.

Hasanov:2003:IPM

[Has03a] Alemdar Hasanov. An inverse polynomial method for the identification of the leading coefficient in the Sturm–Liouville operator from boundary measurements. Applied Mathematics and Computation, 140(2–3): 501–515, August 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hassanien:2003:VPEa

[Has03b]  I. A. Hassanien.  Vari-

able permeability effects on mixed convection along a vertical wedge embedded in a porous medium with variable surface heat flux. Applied Mathematics and Computation, 138(1):41–59, June 1, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hasanov:2004:DLC

[Has04a] Alemdar Hasanov. The determination of the leading coefficient in the monotone potential Sturm–Liouville operator from boundary measurements. Applied Mathematics and Computation, 152(1):141–162, April 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hasanov:2004:EAM

[Has04b] Alemdar Hasanov. Error analysis of a multisingular inverse coefficient problem for the Sturm–Liouville operator based on boundary measurement. Applied Mathematics and Computation, 150(2):493–524, March 8, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hasanov:2004:EUC

[Has04c] Alemdar Hasanov. On existence, uniqueness and convergence of approximate solution of boundary value problems related to the nonlinear operator Au: = −(k((u′)^{2})u′)′ + g(u). Applied Mathematics and Computation, 153(3):659–672, June 14, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hassan:2004:DTT

[Has04d] I. H. Abdel-Halim Hassan. Differential transformation technique for solving higherorder initial value problems. Applied Mathematics and Computation, 154 (2):299–311, July 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hassan:2004:DDD

[Has04e]  S. Z. Hassan.  On de-

layed dynamical duopoly. Applied Mathematics and Computation, 151(1):275– 286, March 30, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hayat:2004:SSP

[HAS04f] T. Hayat, S. Asghar, and A. M. Siddiqui. Stokes’ second problem for a Johnson– Segalman fluid. Applied Mathematics and Computation, 148(3):697–706, January 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hashim:2006:CNA

[Has06] Ishak Hashim. Comments on “A new algorithm for solving classical Blasius equation” by L. Wang. Applied Mathematics and Computation, 176(2):700–703, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [Wan04b].

Hausenblas:2004:NSA

[Hau04] E. Hausenblas. A note on space approximation of parabolic evolution equations. Applied Mathematics and Computation, 157(2): 381–392, October 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hossain:2000:OPG

[HB00] M. Akram Hossain and Michael E. Barber. Optimized Petrov–Galerkin model for advective-dispersive transport. Applied Mathematics and Computation, 115(1):1–10, October 6, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/90/21/21/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/90/21/21/article.
pdf.

Haid:2004:LCI

[HB04] Markus Haid and Jan Breitenbach. Low cost inertial orientation tracking with Kalman filter. Applied Mathematics and Computation, 153(2):567–575, June 4, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hu:2001:NSD

[HC01] Guang-Da Hu and Baruch Cahlon. The numerical solution of discretedelay systems. Applied Mathematics and Computation, 124(3):403–411, December 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/113/32/38/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001168.

He:2002:DPS

[HC02a] Guangming He and Jinde Cao. Discussion of periodic solutions for pth order delayed NDEs. Applied Mathematics and Computation, 129(2–3):391–405, July 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

He:2002:FIP

[HC02b] Guangming He and Jinde Cao. Further investigate on periodic solutions of pth order delayed NDEs. Applied Mathematics and Computation, 132(2–3):231–248, November 10, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Huang:2003:GAS

[HC03] He Huang and Jinde Cao.

On global asymptotic stability of recurrent neural networks with time-varying delays. Applied Mathematics and Computation, 142 (1):143–154, September 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

He:2004:ASC

[HC04a] Ping He and D. Q. Cao.

Algebraic stability criteria of linear neutral systems with multiple time delays. Applied Mathematics and Computation, 155(3):643– 653, August 16, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hwang:2004:NMP

[HC04b] Shin-Jia Hwang and Chiu-Chin Chen. New multiproxy multi-signature schemes. Applied Mathematics and Computation, 147(1):57–67, January 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Han:2004:UNS

[HDZ04] Shuxia Han, Zhiwen Duan, and Li Zhou. The uniqueness of nonnegative C^{1}- solution for nonlinear differential equations. Applied Mathematics and Computation, 153(2):433–444, June 4, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

He:2000:VIM

[He00] Ji-Huan He. Variational iteration method for autonomous ordinary differential systems. Applied Mathematics and Computation, 114(2–3):115–123, September 11, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/89/22/20/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/89/22/
                                        20/article.pdf.

He:2001:STM

[He01] Jiaxing He. On summability theory and method of Fourier series. Applied Mathematics and Computation, 117(2–3):151–159, January 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/92/23/22/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/92/23/
22/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399001575.

He:2003:HPM

[He03a] Ji-Huan He. Homotopy perturbation method: a new nonlinear analytical technique. Applied Mathematics and Computation, 135(1): 73–79, February 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

He:2003:LKE

[He03b] Ji-Huan He. A Lagrangian for von Karman equations of large deflection problem of thin circular plate. Applied Mathematics and Computation, 143(2–3):543–549, November 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

He:2003:NIM

[He03c] Ji-Huan He. A new iteration method for solving algebraic equations. Applied Mathematics and Computation, 135(1):81–84, February 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

He:2003:SPA

[He03d]  Ji-Huan He.  A simple

perturbation approach to Blasius equation. Applied Mathematics and Computation, 140(2–3):217–222, August 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

He:2003:VAB

[He03e] Ji-Huan He. A variational approach to the Burridge– Knopoff equation. Applied Mathematics and Computation, 144(1):1–2, November 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

He:2003:VAL

[He03f] Ji-Huan He. Variational approach to the Lane–Emden equation. Applied Mathematics and Computation, 143(2–3):539–541, November 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

He:2003:VAS

[He03g] Ji-Huan He. Variational approach to the sixth-order boundary value problems. Applied Mathematics and Computation, 143(2–3):537– 538, November 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

He:2003:VAT

[He03h] Ji-Huan He. Variational approach to the Thomas–Fermi equation. Applied Mathematics and Computation, 143(2–3):533–535, November 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

He:2004:AHP

[He04a] J.-H. He. Asymptotology by homotopy perturbation method. Applied Mathematics and Computation, 156 (3):591–596, September 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

He:2004:CHP

[He04b] Ji-Huan He. Comparison of homotopy perturbation method and homotopy analysis method. Applied Mathematics and Computation, 156(2):527–539, September 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

He:2004:CIA

[He04c] Ji-Huan He. He Chengtian’s inequality and its applications. Applied Mathematics and Computation, 151 (3):887–891, April 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

He:2004:HPM

[He04d] Ji-Huan He. The homotopy perturbation method for nonlinear oscillators with discontinuities. Applied Mathematics and Computation, 151(1):287–292, March 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

He:2004:SNE

[He04e] Ji-Huan He. Solution of nonlinear equations by an ancient Chinese algorithm. Applied Mathematics and Computation, 151(1):293– 297, March 30, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

He:2004:SIF

[He04f] Ji-Huan He. Some interpolation formulas in Chinese ancient mathematics. Applied Mathematics and Computation, 152(2):367– 371, May 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

He:2004:VPN

[He04g] Ji-Huan He. Variational principle for non-Newtonian lubrication: Rabinowitsch fluid model. Applied Mathematics and Computation, 157(1):281–286, September 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

He:2004:GAV

[He04h] Ji-Huan He. Zu–Geng’s axiom vs Cavalieri’s theory. Applied Mathematics and Computation, 152 (1):9–15, April 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Held:2001:SSR

[Hel01] H. Held. Spherical spatial ranges of non-polar chemicals for reaction-diffusion type dynamics. Applied Mathematics and Computation, 124(1):29–43, November 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/113/27/29/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000321.

Hassanien:2003:VVT

[HEM03] I. A. Hassanien, A. H. Essawy, and N. M. Moursy.

Variable viscosity and thermal conductivity effects on combined heat and mass transfer in mixed convection over a UHF/UMF wedge in porous media: the entire regime. Applied Mathematics and Computation, 145(2–3):667–682, December 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hassanien:2004:NCF

[HEM04] I. A. Hassanien, A. H. Essawy, and N. M. Moursy. Natural convection flow of micropolar fluid from a permeable uniform heat flux surface in porous medium. Applied Mathematics and Computation, 152(2):323– 335, May 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Haroon:2005:CSN

[HES05] Mohammed Haroon and Moustafa El-Shahed. Comment on: “On solution of non-linear differential equation arising in gliding motion of bacteria”, Applied Mathematics and Computation 148 (2004) 743–752. Applied Mathematics and Computation, 171(2):675– 676, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [HAM04].

Hon:2003:SSD

[HF03] Y. C. Hon and E. Fan. Soliton solutions and doubly periodic wave solutions for a new generalized Hirota– Satsuma coupled system. Applied Mathematics and Computation, 146(2–3):813– 827, December 31, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hua:2004:RSC

[HGS04] Changchun Hua, Xinping Guan, and Peng Shi. Robust stabilization of a class of nonlinear time-delay systems. Applied Mathematics and Computation, 155(3): 737–752, August 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Han:2001:RFF

[HH01] Qiaoming Han and Jiye Han.

Revised filled function methods for global optimization. Applied Mathematics and Computation, 119(2–3):217– 228, April 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/99/25/28/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/99/25/
28/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399002660.

Han:2001:PIM

[HHJ01] Guoqiang Han, Ken Hayami, and Wang Jiong. Product integration method for Volterra integral equation with noncompact kernel. Applied Mathematics and Computation, 121(2–3):363– 371, June 15, 2001. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/105/25/
36/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/105/25/36/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000060.

Hosaka:2001:CMS

[HHK01] Masanori Hosaka, Masayuki Horiguchi, and Masami Kurano. Controlled Markov set-chains under average criteria. Applied Mathematics and Computation, 120(1– 3):195–209, May 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/104/21/
36/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/104/21/36/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002416.

Huang:2004:SSS

[HHT04] Shu-Shiang Huang, Yenkun Huang, and Che-Yuan Tsai. Strong scheme for a stochastic Goursat problem. Applied Mathematics and Computation, 150(2):351–363, March 8, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hu:2004:SLN

[HHZ04] Guang-Da Hu, Guang-Di Hu, and Xingfu Zou. Stability of linear neutral systems with multiple delays: boundary criteria. Applied Mathematics and Computation, 148(3):707–715, January 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hasanov:2004:SPE

[HI04a] V. I. Hasanov and I. G.

Ivanov. Solutions and perturbation estimates for the matrix equations X ± A^{∗}X^{−n}A = Q. Applied Mathematics and Computation, 156(2):513–525, September 6, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hizarci:2004:CGI

[Hİ04b] S. Hizarci and A. S. İpek.

Changes in geometry: invariant areas in Euclid geometry. Applied Mathematics and Computation, 153 (2):395–401, June 4, 2004.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hizarci:2004:INE

[HIİ04] Seyfullah Hizarci, Cemalettin Işik, and A. Sabri İpek. On indicator of normals in E^{n}-Euclid space. Applied Mathematics and Computation, 154(2):443–447, July 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hillion:2004:MAR

[Hil04] Pierre Hillion. Mathematical aspects of rough plane scattering theory. Applied Mathematics and Computation, 147(3):863–880, January 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hanada:2004:SFD

[HIN04] Takao Hanada, Naoyuki Ishimura, and Masaaki Nakamura. Stable finite difference scheme for a model equation of phase separation. Applied Mathematics and Computation, 151 (1):95–104, March 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hanna:2004:UCP

[HIS04] Yousry S. Hanna, Makram Ibrahim, and S. W. Samwel.

On using Chebyshev polynomials for fitting SLR data of artificial satellites. Applied Mathematics and Computation, 158(3):655–666, November 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hayat:2004:MFO

[HKA04a] T. Hayat, Masood Khan, and S. Asghar. Magnetohydrodynamic flow of an Oldroyd 6-constant fluid. Applied Mathematics and Computation, 155(2):417– 425, August 6, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hayat:2004:ESF

[HKA04b] T. Hayat, Masood Khan, and M. Ayub. Exact solutions of flow problems of an Oldroyd-B fluid. Applied Mathematics and Computation, 151(1):105–119, March 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hou:2003:FCS

[HkT03] Xuezhang Hou and Sze kai Tsui. A feedback control and a simulation of a torsional elastic robot arm. Applied Mathematics and Computation, 142(2–3):389–407, October 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hicks:2000:LGD

[HL00] D. L. Hicks and L. M. Liebrock. Lanczos’ generalized derivative: Insights and applications. Applied Mathematics and Computation, 112(1):63–73, June 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/86/21/
                                        26/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/86/21/26/article.
                                        pdf.

He:2003:OHF

[HL03] Mengxing He and Anping Liu. The oscillation of hyperbolic functional differential equations. Applied Mathematics and Computation, 142(2–3):205–224, October 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hicks:2004:CSB

[HL04a] D. L. Hicks and L. M. Liebrock. Conservative smoothing with B-splines stabilizes SPH material dynamics in both tension and compression. Applied Mathematics and Computation, 150(1):213–234, February 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Huo:2004:EGS

[HL04b] Hai-Feng Huo and Wan-Tong Li. Existence and global stability of periodic solutions of a discrete predator–prey system with delays. Applied Mathematics and Computation, 153 (2):337–351, June 4, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Huo:2004:PSDb

[HL04c] Hai-Feng Huo and Wan-Tong Li. Periodic solution of a delayed predator–prey system without dominating instantaneous negative feedback. Applied Mathematics and Computation, 156 (3):871–882, September 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Huo:2004:PSP

[HL04d] Hai-Feng Huo and Wan-Tong Li. Periodic solutions of a periodic Lotka– Volterra system with delays. Applied Mathematics and Computation, 156(3):787– 803, September 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Huo:2004:PSDa

[HL04e] Hai-Feng Huo and Wan-Tong Li. Periodic solutions of delayed Leslie– Gower predator–prey models. Applied Mathematics and Computation, 155(3): 591–605, August 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Huo:2004:PPS

[HL04f] Hai-Feng Huo and Wan-Tong Li. Positive periodic solutions of a class of delay differential system with feedback control. Applied Mathematics and Computation, 148(1):35–46, January 20, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hwang:2004:REL

[HL04g] Shin-Jia Hwang and Yun-Hua Lee. Repairing ElGamallike multi-signature schemes using self-certified public keys. Applied Mathematics and Computation, 156 (1):73–83, August 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hwang:2003:TRB

[HLL03] Min-Shiang Hwang, Cheng-Chi Lee, and Yan-Chi Lai. Traceability on RSA-based partially signature with low computation. Applied Mathematics and Computation, 145(2–3):465–468, December 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hong:2003:BBP

[HLLL03] Hyun-Soo Hong, Ho-Kyu Lee, Hyang-Sook Lee, and Hee-Jung Lee. The better bound of private key in RSA with unbalanced primes. Applied Mathematics and Computation, 139(2– 3):351–362, July 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

He:2002:SLS

[HLO02] Mengxing He, Anping Liu, and Zhuoling Ou. Stability for large systems of partial functional differential equations: iterative analysis method. Applied Mathematics and Computation, 132(2–3):489–503, November 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hlomuka:2004:LNS

[Hlo04] Joe Hlomuka. The linearized non-stationary problem for the permeable boundary Navier–Stokes flows. Applied Mathematics and Computation, 158(3):717–727, November 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

He:2004:GAD

[HLY04] Wan-Sheng He, Wan-Tong Li, and Xin-Xue Yan. Global attractivity of the difference equation x_{n+1} = α + (x_{n−k}/x_{n}). Applied Mathematics and Computation, 151(3):879–885, April 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hossain:2000:MPP

[HM00] S. S. Hossain and H. A. Muttlak. MVLUE of population parameters based on ranked set sampling. Applied Mathematics and Computation, 108(2–3):167–176, February 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/79/22/
                                        30/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/79/22/30/article.
                                        pdf.

Hart:2004:SDH

[HM04] Andrew G. Hart and Servet Martínez. Sequential dynamics of high order polynomial automata networks: an application to the Erlang fixed-point equations. Applied Mathematics and Computation, 151(2):507– 522, April 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hu:2002:NAD

[HML^{+}02] Wei Hu, Shengwei Mei, Qiang Lu, Tielong Shen, and Akihiko Yokoyama. Nonlinear adaptive decentralized stabilizing control of multimachine systems. Applied Mathematics and Computation, 133(2–3):519–532, December 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Houben:2003:APM

[HMM03] S. H. M. J. Houben, J. M. L.

Maubach, and R. M. M. Mattheij. An accelerated Poincaré-map method for autonomous oscillators. Applied Mathematics and Computation, 140(2–3):191–216, August 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hayat:2004:PUF

[HNA04a] T. Hayat, S. Nadeem, and S. Asghar. Periodic unidirectional flows of a viscoelastic fluid with the fractional Maxwell model. Applied Mathematics and Computation, 151(1):153–161, March 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hayat:2004:HEU

[HNA04b] T. Hayat, R. Naz, and S. Asghar. Hall effects on unsteady duct flow of a non-Newtonian fluid in a porous medium. Applied Mathematics and Computation, 157 (1):103–114, September 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

He:2002:CMP

[HOL02] Mengxing He, Zhuoling Ou, and Anping Liu. Comparison method of partial functional differential equations and its application. Applied Mathematics and Computation, 125(2–3):271–286, January 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/120/31/36/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001296.

Hosseini:2004:RIM

[Hos04]  M. M. Hosseini.  Re-

ducing index method for differential-algebraic equations with constraint singularities. Applied Mathematics and Computation, 153 (1):205–214, May 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

He:2002:GCM

[HP02a] W. He and N. Prabhu. A globally convergent method for finding zeros of smooth functions. Applied Mathematics and Computation, 133(2–3):327–335, December 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Holmstrom:2002:RPE

[HP02b] Kenneth Holmström and Jöran Petersson. A review of the parameter estimation problem of fitting positive exponential sums to empirical data. Applied Mathematics and Computation, 126(1):31–61, February 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/122/27/28/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001387.

Hernandez:2001:SMD

[HR01] M. A. Hernández and M. J.

Rubio. The Secant method and divided differences Hölder continuous. Applied Mathematics and Computation, 124(2):139–149, October 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/

                                        9/12/113/31/27/abstract.
                                        html;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300300000795.

He:2002:IEO

[HR02] M. X. He and P. E. Ricci. Information entropy of orthogonal polynomials. Applied Mathematics and Computation, 128(2–3):261–274, May 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hasanov:2000:QBSb

[HS00] Alemdar Hasanov and Zahir Seyidmamedov. Qualitative behaviour of solutions of Signorini problem with perturbing the unknown boundary. Applied Mathematics and Computation, 109(2–3): 261–271, March 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/82/23/
                                        32/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/82/23/32/article.
                                        pdf.

Hon:2001:UCR

[HS01] Y. C. Hon and R. Schaback. On unsymmetric collocation by radial basis functions. Applied Mathematics and Computation, 119(2–3):177– 186, April 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/99/25/24/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/99/25/
24/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399002556.

Hasanov:2002:DLC

[HS02] Alemdar Hasanov and Zahir Seyidmamedov. Determination of leading coefficients in Sturm–Liouville operator from boundary measurements. II. Unicity and an engineering approach. Applied Mathematics and Computation, 125(1):23–34, January 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/120/27/28/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001053.

Hosseini:2003:NSC

[HS03] S. M. Hosseini and S. Shahmorad. Numerical solution of a class of Integro-Differential equations by the Tau Method with an error estimation. Applied Mathematics and Computation, 136(2–3):559–570, March 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hassanien:2003:VPEb

[HSE03] I. A. Hassanien, A. A. Salama, and A. M. Elaiw. Variable permeability effect on vortex instability of mixed convection flow in a semi-infinite porous medium bounded by a horizontal surface. Applied Mathematics and Computation, 146(2–3): 829–847, December 31, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hassanien:2004:OLV

[HSE04] I. A. Hassanien, A. A. Salama, and A. M. Elaiw. The onset of longitudinal vortices in mixed convection flow over an inclined surface in a porous medium with variable permeability. Applied Mathematics and Computation, 154(2):313– 333, July 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hossain:2000:SAD

[HT00] Md. Akram Hossain and M. Raihan Taha. Simulating advective-dispersive transport by finite elements: Criteria for accuracy of an explicit Runge–Kutta method. Applied Mathematics and Computation, 112(2–3):309– 316, June 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/86/23/
32/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/86/23/32/article.
pdf.

Hou:2004:ACT

[HT04] Xuezhang Hou and Sze-Kai Tsui. Analysis and control of a two-link and three-joint elastic robot arm. Applied Mathematics and Computation, 152(3):759– 777, May 13, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Huang:2003:NMU

[Hua03] Zhengda Huang. Newton method under weak Lipschitz continuous derivative in Banach spaces. Applied Mathematics and Computation, 140(1):115–126, July 30, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hurlimann:2004:MOR

[Hür04] Werner Hürlimann. Measuring operational risk using a mean scaled individual risk model. Applied Mathematics and Computation, 152 (2):425–447, May 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Helman:2001:AAR

[HV01] Paul Helman and Robert Veroff. The application of automated reasoning to formal models of combinatorial optimization. Applied Mathematics and Computation, 120(1–3):175–194, May 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/104/21/
35/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/104/21/35/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002490.

Hadi-Vencheh:2008:CIO

[HV08]  Abdollah  Hadi-Vencheh.

Comment on “Inputs/outputs estimation in DEA when some factors are undesirable” [Applied Mathematics and Computation 156 (2004) 19–32]. Applied Mathematics and Computation, 202(2): 893–894, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [JVFM04].

Hsu:2004:EPS

[HW04] Chien-Lung Hsu and Tzong-Sun Wu. Efficient proxy signature schemes using selfcertified public keys. Applied Mathematics and Computation, 152(3):807–820, May 13, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hsu:2003:ITP

[HWW03] Chien-Lung Hsu, Tzong-Sun Wu, and Tzong-Chen Wu. Improvement of threshold proxy signature scheme. Applied Mathematics and Computation, 136(2–3):315–321, March 15, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hsu:2003:IMA

[HWWM03] Chien-Lung Hsu, Tzong-Sun Wu, Tzong-Chen Wu, and Chris Mitchell. Improvement of modified authenticated key agreement protocol. Applied Mathematics and Computation, 142(2–3): 305–308, October 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Haibo:2004:LRF

[HY04] Chen Haibo and Liu Yirong.

Linear recursion formulas of quantities of singular point and applications. Applied Mathematics and Computation, 148(1):163–171, January 20, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hadizadeh:2004:CAI

[HZ04a] M. Hadizadeh and A. R.

Zokayi. Concerning the action of an invertible point transformation on a class of ordinary differential equations. Applied Mathematics and Computation, 147(2): 347–353, January 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

He:2004:MIT

[HZ04b] Zhimin He and Xianming Zhang. Monotone iterative technique for first order impulsive difference equations with periodic boundary conditions. Applied Mathematics and Computation, 156 (3):605–620, September 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hou:2004:AMF

[HZ04c] Zai-En Hou and Ke-Cun Zhang. Approximation of multivariate function by using new multivariate Bernstein α-polynomials. Applied Mathematics and Computation, 154(2):335–345, July 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Han:2002:AMM

[HZL02] Bo Han, Xiaoyan Zhou, and Jiaqi Liu. Adaptive multigrid method for numerical solutions of elastic wave equation. Applied Mathematics and Computation, 133(2–3):609–614, December 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hu:2003:BAM

[HZZ03] Xi-Yan Hu, Lei Zhang, and Qin Zhang. The best approximation of matrices under inequality constraints. Applied Mathematics and Computation, 137(2–3):487– 497, May 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ismail:2001:SDD

[IbS01] Fudziah Ismail and Mohammed bin Suleiman. Solving delay differential equations using intervalwise partitioning by Runge–Kutta method. Applied Mathematics and Computation, 121 (1):37–53, May 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/105/21/
24/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/105/21/24/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002611.

Ismail:2004:RTAa

[IES04a] Hassan N. A. Ismail, Elsayed M. E. Elbarbary, and Ghada S. E. Salem. Restrictive Taylor’s approximation for solving convection– diffusion equation. Applied Mathematics and Computation, 147(2):355–363, January 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ismail:2004:RTAb

[IES04b] Hassan N. A. Ismail, Elsayed M. E. M. Elbarbary, and Ghada S. E. Salem. Restrictive Taylor’s approximation for two dimensions initial boundary value problem for parabolic PDE. Applied Mathematics and Computation, 147(3): 607–615, January 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ibrahim:2003:ESC

[IG03] A. G. Ibrahim and A. G. Gomaa. Extremal solutions of classes of multivalued differential equations. Applied Mathematics and Computation, 136(2–3):297–314, March 15, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ibrahim:2001:LNS

[IH01] F. S. Ibrahim and I. A. Hassanien. Local nonsimilarity solutions for mixed convection boundary layer flow of a micropolar fluid on horizontal flat plates with variable surface temperature. Applied Mathematics and Computation, 122 (2):133–153, July 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/106/25/
21/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/106/25/21/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000175.

Imaeda:2000:SAA

[II00] K. Imaeda and M. Imaeda.

Sedenions: algebra and analysis. Applied Mathematics and Computation, 115(2–3): 77–88, October 27, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.nl/gej-ng/
10/9/12/90/23/20/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/90/23/
20/article.pdf.

Ifantis:2002:SMO

[IKS02] E. K. Ifantis, C. G. Kokologiannaki, and P. D. Siafarikas. On the support of the measure of orthogonality of a class of orthogonal polynomials. Applied Mathematics and Computation, 128(2– 3):275–288, May 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Itiki:2004:CAG

[IN04] C. Itiki and J. José Neto. Complete automation of the generalized inverse method for constrained mechanical systems of particles. Applied Mathematics and Computation, 152(2):561–580, May 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Inohara:2000:MDD

[Ino00a] Takehiro Inohara. Meetings in deadlock and decision makers with interperception. Applied Mathematics and Computation, 109(2– 3):121–133, March 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/82/23/
                                        22/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/82/23/22/article.
                                        pdf.

Inohara:2000:CCG

[Ino00b] Takehiro Inohara. On consistent coalitions in group decision making with flexible decision makers. Applied Mathematics and Computation, 109(2–3):101–119, March 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/82/23/
                                        21/abstract.html;   http:

//www.elsevier.nl/gej-
ng/29/17/20/82/23/21/article.
pdf.

Inohara:2002:CCS

[Ino02a] Takehiro Inohara. Characterization of clusterability of signed graph in terms of Newcomb’s balance of sentiments. Applied Mathematics and Computation, 133(1): 93–104, November 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Inohara:2002:GCC

[Ino02b] Takehiro Inohara. Generalizations of the concept of core of simple games and their characterization in terms of permission of voters. Applied Mathematics and Computation, 132(1): 47–62, October 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Inohara:2003:CGI

[Ino03] Takehiro Inohara. Clusterability of groups and information exchange in group decision making with approval voting system. Applied Mathematics and Computation, 136(1):1–15, March 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Inohara:2004:QCS

[Ino04a] Takehiro Inohara. Quasiclusterability of signed graphs with negative self evaluation. Applied Mathematics and Computation, 158(1): 201–215, October 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Inohara:2004:SGN

[Ino04b] Takehiro Inohara. Signed graphs with negative self evaluation and clusterability of graphs. Applied Mathematics and Computation, 158(2):477–487, November 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ibidapo-Obe:2001:GSP

[IOAB01] O. Ibidapo-Obe, O. S. Asaolu, and A. B. Badiru. Generalized solutions of the pursuit problem in three-dimensional Euclidean space. Applied Mathematics and Computation, 119 (1):35–45, March 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.nl/gej-ng/
10/9/12/99/20/22/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/99/20/
22/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399002271.

Ibidapo-Obe:2002:NMN

[IOAB02] O. Ibidapo-Obe, O. S.

Asaolu, and A. B. Badiru. A new method for the numerical solution of simultaneous nonlinear equations. Applied Mathematics and Computation, 125(1):133–140, January 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/120/27/35/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001259.

Inverardi:2003:HMP

[IPPT03] Pierluigi Novi Inverardi, Giorgio Pontuale, Alberto Petri, and Aldo Tagliani. Hausdorff moment problem via fractional moments. Applied Mathematics and Computation, 144(1):61–74, November 20, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ismail:2004:RPA

[IR04] Hassan N. A. Ismail and Aziza A. Abd Rabboh. A restrictive Padé approximation for the solution of the generalized Fisher and Burger–Fisher equations. Applied Mathematics and Computation, 154 (1):203–210, June 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ismail:2004:ADM

[IRR04] Hassan N. A. Ismail, Kamal Raslan, and Aziza A. Abd Rabboh. Adomian decomposition method for Burger’s–Huxley and Burger’s–Fisher equations. Applied Mathematics and Computation, 159(1):291– 301, November 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ismail:2004:SWS

[IRS04] Hassan N. A. Ismail, Kamal R. Raslan, and Ghada S. E. Salem. Solitary wave solutions for the general KDV equation by Adomian decomposition method. Applied Mathematics and Computation, 154(1):17–29, June 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Isik:2003:JMF

[Isı03] Ahmet Isık. Jacobi modular forms on subgroups of the modular group. Applied Mathematics and Computation, 144(2–3):529–535, December 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Iskander:2004:FWA

[Isk04] Maged George Iskander. A fuzzy weighted additive approach for stochastic fuzzy goal programming. Applied Mathematics and Computation, 154(2):543–553, July 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ismail:2004:USR

[Ism04] Hassan N. A. Ismail. Unique solvability of restrictive Padé and restrictive Taylor’s approximations. Applied Mathematics and Computation, 152(1):89–97, April 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Inohara:2000:CIS

[ITN00] Takehiro Inohara, Shingo Takahashi, and Bunpei Nakano. Credibility of information in ‘soft’ games with interperception of emotions. Applied Mathematics and Computation, 115(1):23–41, October 6, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/90/21/23/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/90/21/
23/article.pdf.

Iwamoto:2001:CDF

[Iwa01] Seiichi Iwamoto. A class of dual fuzzy dynamic programs. Applied Mathematics and Computation, 120 (1–3):91–108, May 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/104/21/
29/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/104/21/29/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002908.

Ichikawa:2000:RMB

[IYM00] Michinori Ichikawa, Hitoshi Yamada, and Gen Matsumoto. Realization model for brain computing. Applied Mathematics and Computation, 111(2– 3):193–202, May 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/85/23/
25/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/85/23/25/article.
pdf.

Jaheen:2003:BAR

[Jah03] Zeinhum F. Jaheen. A Bayesian analysis of record statistics from the Gompertz model. Applied Mathematics and Computation, 145(2–3): 307–320, December 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jahanshahloo:2004:MMC

[JAK04a] G. R. Jahanshahloo, A. R.

Amirteimoori, and S. Kordrostami. Measuring the multi-component efficiency with shared inputs and outputs in data envelopment analysis. Applied Mathematics and Computation, 155 (1):283–293, July 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jahanshahloo:2004:MCP

[JAK04b] G. R. Jahanshahloo, A. R.

Amirteimoori, and S. Kordrostami. Multi-component performance, progress and regress measurement and shared inputs and outputs in DEA for panel data: an application in commercial bank branches. Applied Mathematics and Computation, 151(1):1–16, March 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jamshidi:2001:ACC

[Jam01] Mohammad Jamshidi. Autonomous control of complex systems: robotic applications. Applied Mathematics and Computation, 120 (1–3):15–29, May 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/104/21/
23/abstract.html;   http:

                                        //www.elsevier.nl/gej-
                                        ng/10/9/12/104/21/23/article.
                                        pdf;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300399002854.

Jankowski:2004:MMS

[Jan04] Tadeusz Jankowski. Monotone method for secondorder delayed differential equations with boundary value conditions. Applied Mathematics and Computation, 149(2):589–598, February 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jayakumar:2003:INS

[Jay03] J. Jayakumar. Improvement of numerical solution by boundary value technique for singularly perturbed one dimensional reaction diffusion problem. Applied Mathematics and Computation, 142(2–3):417–447, October 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jang:2002:NPP

[JB02] S. R.-J. Jang and J. Baglama. A nutrient-prey-predator model with intratrophic predation. Applied Mathematics and Computation, 129(2– 3):517–536, July 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jimenez:2002:DPL

[JBMR02] J. C. Jimenez, R. Biscay, C. Mora, and L. M. Rodriguez. Dynamic properties of the local linearization method for initialvalue problems. Applied Mathematics and Computation, 126(1):63–81, February 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/122/27/29/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001004.

Jayaraman:2004:TOT

[JBS04] V. Jayaraman, M. I. Bhatti, and H. Saber. Towards optimal testing of an hypothesis based on dynamic technology transfer model. Applied Mathematics and Computation, 147(1):115– 129, January 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jang:2000:SIV

[JCL00] Ming-Jyi Jang, Chieh-Li Chen, and Yung-Chin Liy. On solving the initialvalue problems using the differential transformation method. Applied Mathematics and Computation, 115(2–3):145–160, October 27, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/90/23/26/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/90/23/
26/article.pdf.

Jang:2001:TDD

[JCL01] Ming-Jyi Jang, Chieh-Li Chen, and Yung-Chin Liu. Two-dimensional differential transform for partial differential equations. Applied Mathematics and Computation, 121(2–3):261–270, June 15, 2001. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/105/25/
31/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/105/25/31/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002933.

Jung:2002:IPM

[JCZ02] Jong Soo Jung, Yeol Je Cho, and Haiyun Zhou. Iterative processes with mixed errors for nonlinear equations with perturbed m-accretive operators in Banach spaces. Applied Mathematics and Computation, 133(2–3):389–406, December 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jang:2003:PLE

[JD03] S. R.-J. Jang and S. L. Diamond. Population-level effects of density dependence in a size-structured fishery model. Applied Mathematics and Computation, 139 (1):133–155, July 1, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jazbec:2004:SMR

[JDV04] Anamarija Jazbec, Marko Delimar, and Vanja Slavić Vrzić. Simulation model of rubella — the effects of vaccination strategies. Applied Mathematics and Computation, 153(1):75–96, May 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jahanshahloo:2004:FWR

[JF04] G. R. Jahanshahloo and A. A. Foroughi. Finding a weights-restricted efficient (extreme) point and using it for solving MOLP problems. Applied Mathematics and Computation, 150(1): 203–211, February 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jiang:2002:MMC

[JGW02] Daqing Jiang, Wenjie Gao, and Aying Wan. A monotone method for constructing extremal solutions to fourth-order periodic boundary value problems. Applied Mathematics and Computation, 132(2–3):411–421, November 10, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jahanshahloo:2004:MDI

[JHS^{+}04] G. R. Jahanshahloo, F. Hosseinzadeh, N. Shoja, G. Tohidi, and S. Razavyan. A method for detecting influential observation in radial DEA models. Applied Mathematics and Computation, 147(2):415–421, January 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ji:2002:FAD

[Ji02]  Jun Ji.  A finite algo-

rithm for the Drazin inverse of a polynomial matrix. Applied Mathematics and Computation, 130(2–3): 243–251, August 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jianchu:2002:OCS

[Jia02]  Jiang Jianchu.  Oscilla-

tory criteria for second-order quasilinear neutral delay difference equations. Applied Mathematics and Computation, 125(2–3):287–293, January 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/120/31/37/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001302.

Jiang:2003:OSOb

[Jia03a] Jianchu Jiang. Oscillation of second order nonlinear neutral delay difference equations. Applied Mathematics and Computation, 146(2–3): 791–801, December 31, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jiang:2003:PWR

[Jia03b] Yao-Lin Jiang. Periodic waveform relaxation solutions of nonlinear dynamic equations. Applied Mathematics and Computation, 135(2–3):219–226, March 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Johnson:2002:CCG

[JJ02] Alan W. Johnson and Sheldon H. Jacobson. A class of convergent generalized hill climbing algorithms. Applied Mathematics and Computation, 125(2–3):359–373, January 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/120/31/44/abstract.

                                        html;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300300001375.

Johnston:2004:MNL

[JJ04] Barbara M. Johnston and Peter R. Johnston. A modified non-linear transformation method for evaluating weakly singular boundary integrals. Applied Mathematics and Computation, 148(2):519–535, January 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jahanshahloo:2003:UIO

[JK03a] G. R. Jahanshahloo and M. Khodabakhshi. Using input–output orientation model for determining most productive scale size in DEA. Applied Mathematics and Computation, 146(2–3): 849–855, December 31, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jang:2003:HOG

[JK03b] Youngho Jang and Dae San Kim. On higher order generalized Bernoulli numbers. Applied Mathematics and Computation, 137(2– 3):387–398, May 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jahanshahloo:2004:DAI

[JK04a] G. R. Jahanshahloo and M. Khodabakhshi. Determining assurance interval for non-Archimedean element in the improving outputs model in DEA. Applied Mathematics and Computation, 151(2):501–506, April 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jahanshahloo:2004:SCI

[JK04b] G. R. Jahanshahloo and M. Khodabakhshi. Suitable combination of inputs for improving outputs in DEA with determining input congestion: Considering textile industry of China. Applied Mathematics and Computation, 151(1):263– 273, March 30, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jing:2004:OHP

[JK04c] Wang Jing and Wang Ke. The optimal harvesting problems of a stagestructured population. Applied Mathematics and Computation, 148(1):235–247, January 20, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jones:2003:SDA

[JKP03] J. Jones, N. P. Karampetakis, and A. C. Pugh. Solution of discrete ARMArepresentations via MAPLE.

Applied Mathematics and Computation, 139(2–3):437– 489, July 15, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jang:2004:KCB

[JKP04] Leechae Jang, Taekyun Kim, and Dal-Won Park. Kummer congruence for the Bernoulli numbers of higher order. Applied Mathematics and Computation, 151(2):589– 593, April 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jiang:2003:OCF

[JL03a] Jianchu Jiang and Xiaoping Li. Oscillation criteria for first order nonlinear delay difference equations. Applied Mathematics and Computation, 141(2–3): 339–349, September 5, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jiang:2003:OCS

[JL03b] Jianchu Jiang and Xiaoping Li. Oscillation criteria for second-order linear difference equations. Applied Mathematics and Computation, 145(2–3):591–601, December 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jiang:2003:OSOa

[JL03c] Jianchu Jiang and Xiaoping Li. Oscillation of second order nonlinear neutral differential equations. Applied Mathematics and Computation, 135(2–3):531–540, March 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Juang:2004:PAM

[JL04] Wen-Shenq Juang and Horng-Twu Liaw. A practical anonymous multi-authority e-cash scheme. Applied Mathematics and Computation, 147(3):699–711, January 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jimenez:2003:NSS

[JLM^{+}03] Salvador Jiménez, Ignacio M. Llorente, Ana M. Mancho, Víctor M. Pérez-García, and Luis Vázquez. A numerical scheme for the simulation of blow-up in the nonlinear Schrödinger equation. Applied Mathematics and Computation, 134(2–3): 271–291, January 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jahanshahloo:2004:SSA

[JLM04] G. R. Jahanshahloo, F. Hosseinzadeh Lofti, and M. Moradi.

Sensitivity and stability analysis in DEA with interval data. Applied Mathematics and Computation, 156 (2):463–477, September 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jiang:2003:NIT

[JLMC03] Yao-Lin Jiang, Yao-Wu Liu, Kenneth K. Mei, and Richard M. M. Chen. A new iterative technique for large and dense linear systems from the MEI method in electromagnetics. Applied Mathematics and Computation, 139(1):157–163, July 1, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jahanshahloo:2004:IEI

[JLS^{+}04a] G. R. Jahanshahloo, F. Hosseinzadeh Lotfi, N. Shoja, G. Tohidi, and S. Razavyan. Input estimation and identification of extra inputs in inverse DEA models. Applied Mathematics and Computation, 156(2): 427–437, September 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jahanshahloo:2004:OED

[JLS^{+}04b] G. R. Jahanshahloo, F. Hosseinzadeh Lotfi, N. Shoja, G. Tohidi, and S. Razavyan. The outputs estimation of a DMU according to improvement of its efficiency. Applied Mathematics and Computation, 147(2): 409–413, January 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jahanshahloo:2004:RUN

[JLS^{+}04c] G. R. Jahanshahloo, F. Hosseinzadeh Lotfi, N. Shoja, G. Tohidi, and S. Razavyan. Ranking using l_{1}- norm in data envelopment analysis. Applied Mathematics and Computation, 153 (1):215–224, May 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jahanshahloo:2004:AAE

[JLSS04] G. R. Jahanshahloo, F. Hosseinzadeh Lotfi, N. Shoja, and M. Sanei. An alternative approach for equitable allocation of shared costs by using DEA. Applied Mathematics and Computation, 153(1):267–274, May 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jahanshahloo:2004:MFE

[JLST04] G. R. Jahanshahloo, F. Hosseinzadeh Lotfi, N. Shoja, and G. Tohidi. A method for finding efficient DMUs in DEA Using 0–1 linear programming. Applied Mathematics and Computation, 159(1):37–45, November 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ji:2002:OBC

[JM02] Guangcao Ji and Clyde Martin. Optimal boundary control of the heat equation with target function at terminal time. Applied Mathematics and Computation, 127(2–3):335–345, April 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/123/27/41/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S009630030100011X.

Jahanshahloo:2004:FEA

[JMV04a] G. R. Jahanshahloo, R. Kazemi Matin, and A. Hadi Vencheh. On FDH efficiency analysis with interval data. Applied Mathematics and Computation, 159(1):47–55, November 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jahanshahloo:2004:RSF

[JMV04b] G. R. Jahanshahloo, R. Kazemi Matin, and A. Hadi Vencheh. On return to scale of fully efficient DMUs in data envelopment analysis under interval data. Applied Mathematics and Computation, 154(1):31–40, June 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jolayemi:2000:MSD

[Jol00a] Joel K. Jolayemi. A model for the statistical design of multivariate control charts with multiple control regions. Applied Mathematics and Computation, 109 (1):73–91, March 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/82/21/
27/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/82/21/27/article.
pdf.

Jolayemi:2000:ODM

[Jol00b]  Joel K. Jolayemi.  An

optimal design of multiattribute control charts for processes subject to a multiplicity of assignable causes. Applied Mathematics and Computation, 114(2–3):187–203, September 11, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/89/22/26/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/89/22/
26/article.pdf.

Jahanshahloo:2004:ERS

[JSD04] Gholam Reza Jahanshahloo and Majid Soleimani-Damaneh. Estimating returns to scale in data envelopment analysis: A new procedure. Applied Mathematics and Computation, 150(1):89–98, February 27, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jahanshahloo:2004:MED

[JSdN04] G. R. Jahanshahloo, M. Soleimanidamaneh, and E. Nasrabadi. Measure of efficiency in DEA with fuzzy input–output levels: a methodology for assessing, ranking and imposing of weights restrictions. Applied Mathematics and Computation, 156(1): 175–187, August 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jahanshahloo:2004:UGL

[JSLS04] G. R. Jahanshahloo, M. Sanei, F. Hosseinzadeh Lotfi, and N. Shoja. Using the gradient line for ranking DMUs in DEA. Applied Mathematics and Computation, 151 (1):209–219, March 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ji:2004:GOT

[JT04] Mingjun Ji and Huanwen Tang. Global optimizations and tabu search based on memory. Applied Mathematics and Computation, 159 (2):449–457, December 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Juki:2004:NSC

[Juk04] Dragan Juki. A necessary and sufficient criteria for the existence of the least squares estimate for a 3-parametric exponential function. Applied Mathematics and Computation, 147 (1):1–17, January 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jahanshahloo:2004:IOE

[JVFM04] G. R. Jahanshahloo, A. Hadi Vencheh, A. A. Foroughi, and R. Kazemi Matin. Inputs/outputs estimation in DEA when some factors are undesirable. Applied Mathematics and Computation, 156(1):19–32, August 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See comment [HV08].

Ji:2003:NSS

[JW03] Jun Ji and Yimin Wei. A note on the sensitivity of the solution of the weighted linear least squares problem. Applied Mathematics and Computation, 145(2–3): 481–485, December 25, 2003.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jang:2003:ADT

[JWL03] Ming-Jyi Jang, Jiun-Shen Wang, and Yung-Chin Liu. Applying differential transformation method to parameter identification problems. Applied Mathematics and Computation, 139(2– 3):491–502, July 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jou:2000:LSF

[JY00] Jang Jou and Suh-Yuh Yang. Least-squares finite element approximations to the Timoshenko beam problem. Applied Mathematics and Computation, 115(1):63–75, October 6, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/90/21/25/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/90/21/
                                        25/article.pdf.

Junk:2004:AAF

[JY04] Michael Junk and Zhaoxia Yang. Asymptotic analysis of finite difference methods. Applied Mathematics and Computation, 158(1): 267–301, October 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jia:2004:HCM

[JYC04] Gao Jia, Xiao-Ping Yang, and Jinde Cao. The Hölder continuities of the maps in W^{1,p}(Ω, H^{n}). Applied Mathematics and Computation, 149(2):507–518, February 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kotamraju:2000:SCM

[KA00] Gnana Ramdev Kotamraju and Maruthi Ram Akella. Stabilized continuation methods for boundary value problems. Applied Mathematics and Computation, 112(2–3):317–332, June 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/86/23/
33/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/86/23/33/article.
pdf.

Kamali:2003:SCC

[KA03a] Muhammet Kamali and Sezgin Akbulut. On a subclass of certain convex functions with negative coefficients. Applied Mathematics and Computation, 145(2–3): 341–350, December 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Karaduman:2003:GSF

[KA03b] Erdal Karaduman and Hüseyin Aydın. General 2-step Fibonacci sequences in nilpotent groups of exponent p and nilpotency class 4. Applied Mathematics and Computation, 141(2–3):491–497, September 5, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Khan:2003:NST

[KA03c] Arshad Khan and Tariq Aziz. The numerical solution of third-order boundaryvalue problems using quintic splines. Applied Mathematics and Computation, 137(2– 3):253–260, May 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kadalbajoo:2004:CSS

[KA04] Mohan K. Kadalbajoo and Vivek K. Aggarwal. Cubic spline for solving singular two-point boundary value problems. Applied Mathematics and Computation, 156(1):249–259, August 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Koltchinskii:2001:SLC

[KAAD01] V. Koltchinskii, C. T. Abdallah, M. Ariola, and P. Dorato. Statistical learning control of uncertain systems: theory and algorithms. Applied Mathematics and Computation, 120 (1–3):31–43, May 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/104/21/
24/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/104/21/24/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002830.

Kadioglu:2003:SUF

[Kad03] Ekrem Kadioğlu. On subclass of univalent functions with negative coefficients. Applied Mathematics and Computation, 146(2–3):351– 358, December 31, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kadioglu:2004:MPA

[Kad04] Ekrem Kadıoğlu. On majorization problems associated with p-valently functions of complex order. Applied Mathematics and Computation, 153(1):261–265, May 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kahya:2005:CTF

[Kah05] Emin Kahya. Comment on titled “On Fibonacci search method with k-Lucas numbers”. Applied Mathematics and Computation, 162(3): 1321–1324, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300304001900.

See [YK03].

Kahya:2006:CTI

[Kah06] Emin Kahya. Comment on titled “An improvement on Fibonacci search method in optimization theory”. Applied Mathematics and Computation, 173(2):753–756, February 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [SYY04].

Kasap:2005:ERS

[KAK05] Emin Kasap, İsmail Aydemir, and Nuri Kuruoğlu. Erratum to: “Ruled surfaces with timelike rulings” [Appl. Math. Comput. 147 (2004) 241–248]. Applied Mathematics and Computation, 168(2):1461–1468, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [AAABH04, KAK06, Kas06].

Kasap:2006:ERSa

[KAK06] Emin Kasap, İsmail Aydemir, and Nuri Kuruoğlu. Erratum to: “Ruled surfaces with time like rulings” [Appl. Math. Comput. 147 (2004) 241–253]. Applied Mathematics and Computation, 174(2):1660–1667, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [AAABH04, KAK05, Kas06].

Kamar:2002:MIT

[Kam02a] Ahmed R. Abd-Ellateef Kamar. Monotone iterative technique for singular perturbation problem. Applied Mathematics and Computation, 131(2–3):559–571, September 25, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kampke:2002:GLI

[Käm02b] Thomas Kämpke. The geometry of linear infeasibility. Applied Mathematics and Computation, 129(2– 3):317–337, July 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kananthai:2000:CDK

[Kan00a]  A. Kananthai.  On the

convolutions of the diamond kernel of Marcel Riesz. Applied Mathematics and Computation, 114(1):95– 101, August 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/89/20/28/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/89/20/
28/article.pdf.

Kandil:2000:MCC

[Kan00b] F. M. Kandil. Model of competition in the chemostat with instantaneous recycling. Applied Mathematics and Computation, 108 (1):1–10, February 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/79/21/
21/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/79/21/21/article.
pdf.

Kang:2004:RWF

[Kan04] Mihyun Kang. Random walks on a finite graph with congestion points. Applied Mathematics and Computation, 153(2):601–610, June 4, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kaplan:2004:CDE

[Kap04] Abdullah Kaplan. Construction of differential equations using quasi-elliptic functions. Applied Mathematics and Computation, 152 (1):195–198, April 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Karaduman:2004:AFN

[Kar04] Erdal Karaduman. An application of Fibonacci numbers in matrices. Applied Mathematics and Computation, 147(3):903–908, January 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kasap:2006:ERSb

[Kas06] Emin Kasap. Erratum to:

“Ruled surfaces with timelike rulings” [Appl. Math. Comput. 147 (2004) 241– 253]. Applied Mathematics and Computation, 179 (1):402–405, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [AAABH04, KAK05, KAK06].

Kalla:2001:UFG

[KASK01] S. L. Kalla, B. N. Al-Saqabi, and H. G. Khajah. A unified form of gammatype distributions. Applied Mathematics and Computation, 118(2–3):175–187, March 9, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/96/25/24/abstract.

html;   http://www.elsevier.
nl/gej-ng/10/9/12/96/25/
24/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399002131.

Katsunori:2001:MOE

[Kat01] Ano Katsunori. Modified offensive earned-run average with steal effect for baseball. Applied Mathematics and Computation, 120(1– 3):279–288, May 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/104/21/
42/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/104/21/42/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002805.

Kamar:2000:GQS

[KAVM00] Ahmed R. Abd-Ellateef Kamar, Gamal M. Attia, K. Vajravelu, and M. Mosaad. Generalized quasilinearization for singular system of differential equations. Applied Mathematics and Computation, 114(1):69–74, August 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/89/20/26/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/89/20/
26/article.pdf.

Kaya:2003:ENS

[Kay03a] Doğan Kaya. An explicit and numerical solutions of some fifth-order KdV equation by decomposition method. Applied Mathematics and Computation, 144(2–3):353–363, December 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kaya:2003:NSS

[Kay03b] Doğan Kaya. A numerical solution of the sine-Gordon equation using the modified decomposition method. Applied Mathematics and Computation, 143(2–3):309–317, November 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kaya:2004:ADMa

[Kay04a] Doğan Kaya. An application of the decomposition method for the KdVB equation. Applied Mathematics and Computation, 152 (1):279–288, April 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kaya:2004:AMD

[Kay04b] Doğan Kaya. An application of the modified decomposition method for two dimensional sine-Gordon equation. Applied Mathematics and Computation, 159 (1):1–9, November 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kaya:2004:ENS

[Kay04c] Doğan Kaya. Exact and numerical soliton solutions of some nonlinear physical models. Applied Mathematics and Computation, 152 (2):551–560, May 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kaya:2004:NSS

[Kay04d] Doğan Kaya. A numerical simulation of solitary-wave solutions of the generalized regularized long-wave equation. Applied Mathematics and Computation, 149(3): 833–841, February 22, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kaya:2004:RMN

[Kay04e] Doğan Kaya. A reliable method for the numerical solution of the kinetics problems. Applied Mathematics and Computation, 156(1): 261–270, August 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kaya:2004:SWSa

[Kay04f]  Doğan Kaya.  Solitary

wave solutions for a generalized Hirota–Satsuma coupled KdV equation. Applied Mathematics and Computation, 147(1):69–78, January 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kaya:2004:SWSb

[Kay04g]  Doğan Kaya.  Solitary-

wave solutions for compound KdV-type and compound KdV–Burgers-type equations with nonlinear terms of any order. Applied Mathematics and Computation, 152(3):709–720, May 13, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kaya:2004:STE

[Kay04h] Doğan Kaya. The symmetric tridigonal eigenproblem on a shared memory multiprocessor: Part I. Applied Mathematics and Computation, 156(1):189–209, August 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Keeling:2004:VAM

[KB04a] Stephen L. Keeling and Roland Bammer. A variational approach to magnetic resonance coil sensitivity estimation. Applied Mathematics and Computation, 158(2):359–388, November 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Koonprasert:2004:BMS

[KB04b] Sanoe Koonprasert and Kenneth L. Bowers. Block matrix Sinc–Galerkin solution of the wind-driven current problem. Applied Mathematics and Computation, 155(3):607–635, August 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kutluay:2000:STS

[KBÖ00] S. Kutluay, A. R. Bahadir, and A. Özde. A small time solutions for the Korteweg– de Vries equation. Applied Mathematics and Computation, 107(2–3):203–210, January 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/72/22/
                                        30/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/72/22/30/article.
                                        pdf.

Kostreva:2000:SCM

[KC00] M. M. Kostreva and X. Chen. A superlinearly convergent method of feasible directions. Applied Mathematics and Computation, 116(3):231–244, December ??, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/91/23/20/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/91/23/
20/article.pdf.

Kadilar:2004:RES

[KC04] Cem Kadilar and Hulya Cingi. Ratio estimators in simple random sampling. Applied Mathematics and Computation, 151(3):893– 902, April 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Khan:2002:MMH

[KE02] A. R. Khan and A. Elkamel. Mathematical model for heat transfer mechanism for particulate system. Applied Mathematics and Computation, 129(2– 3):295–316, July 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kutluay:2004:BSF

[KE04a] S. Kutluay and A. Esen.

A B-spline finite element method for the thermistor problem with the modified electrical conductivity. Applied Mathematics and Computation, 156(3):621– 632, September 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kutluay:2004:IMF

[KE04b] S. Kutluay and A. Esen. An isotherm migration formulation for one-phase Stefan problem with a time dependent Neumann condition. Applied Mathematics and Computation, 150(1): 59–67, February 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kutluay:2004:LNS

[KE04c] S. Kutluay and A. Esen. A linearized numerical scheme for Burgers-like equations. Applied Mathematics and Computation, 156(2):295– 305, September 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Knisevskaja:2003:SIE

[KEB03] L. Knisevskaja, V. Engelson, and K.-F. Berggren. Singular integral equations method for computations of the scattering characteristics of a model heart exposed to electromagnetic radiation. Applied Mathematics and Computation, 138(2– 3):545–553, June 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Keeling:2003:TVB

[Kee03a] Stephen L. Keeling. Total variation based convex filters for medical imaging. Applied Mathematics and Computation, 139 (1):101–119, July 1, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Khalifa:2003:CEM

[KEE03b] Ahmed K. Khalifa, Elsayed M. E. Elbarbary, and Mohamed A. Abd Elrazek. Chebyshev expansion method for solving second and fourth-order elliptic equations. Applied Mathematics and Computation, 135(2–3):307–318, March 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kano:2003:BSC

[KENM03] Hiroyuki Kano, Magnus Egerstedt, Hiroaki Nakata, and Clyde F. Martin. Bsplines and control theory. Applied Mathematics and Computation, 145(2–3):263– 288, December 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Khalifa:2004:SNE

[KERG04] M. E. Khalifa, Reda G. Abd El-Rahman, and Marwa I. Ghonamy. Some new exact solutions for linear thermoelastic system. Applied Mathematics and Computation, 147(3):773–787, January 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kesan:2003:CPS

[Keş03a]  Cenk Keşan.  Cheby-

shev polynomial solutions of second-order linear partial differential equations. Applied Mathematics and Computation, 134(1):109– 124, January 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kesan:2003:TPS

[Keş03b] Cenk Keşan. Taylor polynomial solutions of linear differential equations. Applied Mathematics and Computation, 142(1):155–165, September 20, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kaya:2004:ADMb

[KES04a] Doğan Kaya and Salah M. El-Sayed. Adomian’s decomposition method applied to systems of nonlinear algebraic equations. Applied Mathematics and Computation, 154(2):487–493, July 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kaya:2004:NSE

[KES04b] Doğan Kaya and Salah M.

El-Sayed. A numerical simulation and explicit solutions of the generalized Burgers– Fisher equation. Applied Mathematics and Computation, 152(2):403–413, May 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kaya:2004:NSK

[KES04c] Doğan Kaya and Salah M.

El-Sayed. A numerical solution of the Klein–Gordon equation and convergence of the decomposition method. Applied Mathematics and Computation, 156(2):341– 353, September 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kesan:2004:TPS

[Keş04d] Cenk Keşan. Taylor polynomial solutions of second order linear partial differential equations. Applied Mathematics and Computation, 152(1):29–41, April 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kulkarni:2003:IDP

[KG03] Rekha P. Kulkarni and N. Gnaneshwar. Iterated discrete polynomially based Galerkin methods. Applied Mathematics and Computation, 146(1):153–165, December 30, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kucuk:2004:IBT

[KG04] Ahmet Küçük and Osman Gürsoy. On the invariants of Bertrand trajectory surface offsets. Applied Mathematics and Computation, 151 (3):763–773, April 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Koksal:2001:ODE

[KH01] M. Köksal and S. Herdem. Ordinary differential equations with strong nonlinearities and their numerical solutions with abrupt changes. Applied Mathematics and Computation, 119(2–3):249– 264, April 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/99/25/30/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/99/25/
30/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399002337.

Kim:2002:NRM

[KH02] Yongdeok Kim and Hoon Huh. New rectangular mixed finite element method for second-order elliptic problems. Applied Mathematics and Computation, 127(2–3):375–385, April 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/123/27/45/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300301000170.

Khalifa:2003:EAE

[Kha03a] M. E. Khalifa. Existence of almost everywhere solution for nonlinear hyperbolic– parabolic system. Applied Mathematics and Computation, 145(2–3):569–577, December 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Khalifa:2003:SAS

[Kha03b] M. E. Khalifa. Some analytical solutions for the advection–dispersion equation. Applied Mathematics and Computation, 139(2– 3):299–310, July 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Khan:2004:PCS

[Kha04] Arshad Khan. Parametric cubic spline solution of two point boundary value problems. Applied Mathematics and Computation, 154 (1):175–182, June 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kim:2002:TDS

[KHL02] Yongdeok Kim, Hoon Huh, and Sungyun Lee. Threedimensional stable nonconforming parallelepiped elements for the Stokes problem. Applied Mathematics and Computation, 130(2–3): 573–586, August 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Khodier:2004:PPA

[Kho04] Ahmed M. M. Khodier. Perturbed Padé approximation with high accuracy. Applied Mathematics and Computation, 148(3):753–757, January 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Khuri:2002:SCL

[Khu02] S. A. Khuri. On the solution of coupled H-like equations of Chandrasekhar. Applied Mathematics and Computation, 133(2–3):479–485, December 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Khuri:2003:BCC

[Khu03a] S. A. Khuri. Biorthogonality condition for creeping flow in wedge-shaped trenches. Applied Mathematics and Computation, 134(1):141– 145, January 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Khuri:2003:NOV

[Khu03b] S. A. Khuri. Numerical order verification of the asymptotic expansion of a nonlinear differential equation arising in general relativity. Applied Mathematics and Computation, 134(1): 147–151, January 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Khuri:2004:NAB

[Khu04]  S. A. Khuri.  A new

approach to Bratu’s problem. Applied Mathematics and Computation, 147(1): 131–136, January 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kaya:2004:ENT

[KI04] Doğan Kaya and Ibrahim E.

Inan. Exact and numerical traveling wave solutions for nonlinear coupled equations using symbolic computation. Applied Mathematics and Computation, 151 (3):775–787, April 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kincanon:2001:SMF

[Kin01]  Eric Kincanon.  Spe-

cial measure function separability and reflectionless potentials. Applied Mathematics and Computation, 123(3):409–412, October 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/110/32/33/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000862.

Kirmaci:2004:IDM

[Kır04] Uğur S. Kırmacı. Inequalities for differentiable mappings and applications to special means of real numbers and to midpoint formula. Applied Mathematics and Computation, 147(1): 137–146, January 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kawata:2000:EAO

[KIY00] Yoshiyuki Kawata, Toshiaki Izumiya, and Akihiro Yamazaki. The estimation of aerosol optical parameters from ADEOS/ POLDER data. Applied Mathematics and Computation, 116(1–2):197–215, November 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.

elsevier.nl/gej-ng/10/
9/12/91/20/35/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/91/20/
35/article.pdf.

Kultur:2004:SRT

[KK04] M. Nuri Kültür and Abdullah Kaplan. A study on ratio of the Θ–Theta functions. Applied Mathematics and Computation, 158(2): 353–358, November 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Konno:2004:SPD

[KKM04] Norio Konno, Toshihiko Kunimatsu, and Xia Ma. From stochastic partial difference equations to stochastic cellular automata through the ultra-discretization. Applied Mathematics and Computation, 155(3):727–735, August 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kanemitsu:2004:SIA

[KKSY04] S. Kanemitsu, H. Kumagai, H. M. Srivastava, and M. Yoshimoto. Some integral and asymptotic formulas associated with the Hurwitz Zeta function. Applied Mathematics and Computation, 154(3):641–664, July 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kim:2000:LSM

[KL00a] Yongdeok Kim and Sungyun Lee. Least-squares mixed method for second-order elliptic problems. Applied Mathematics and Computation, 115(2–3):89–100, October 27, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/90/23/21/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/90/23/
21/article.pdf.

Kim:2000:SNQ

[KL00b] Yongdeok Kim and Sungyun Lee. Stable nonconforming quadrilateral finite elements for the Stokes problem. Applied Mathematics and Computation, 115(2–3):101–112, October 27, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/90/23/22/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/90/23/
22/article.pdf.

Knopfmacher:2002:MER

[KL02] A. Knopfmacher and D. S. Lubinsky. Mathematica evidence that Ramanujan kills Baker–Gammel–Wills. Applied Mathematics and Computation, 128(2–3):289–302, May 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kim:2003:CTS

[KL03] Kwang Ik Kim and Zhigui Lin. Coexistence in the three species predator–prey model with diffusion. Applied Mathematics and Computation, 145(2–3):701–716, December 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kim:2002:PIG

[KLK02] Yongdeok Kim, Sungyun Lee, and Seki Kim. A parallel iterative Galerkin method based on nonconforming quadrilateral elements for second-order partial differential equations. Applied Mathematics and Computation, 127(2–3):387– 404, April 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/123/27/46/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300301000182.

Khan:2002:ATS

[KM02] A. Khan and H. A. Muttlak. Adjusted two-stage adaptive cluster sampling. Applied Mathematics and Computation, 126(1):83–95, February 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        9/12/122/27/30/abstract.
                                        html;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300300001478.

Kiymaz:2003:AAE

[KM03a] Onur Kıymaz and Şeref Mirasyedioğlu. An algorithmic approach to exact power series solutions of second order linear homogeneous differential equations with polynomial coefficients. Applied Mathematics and Computation, 139 (1):165–178, July 1, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kubiaczyk:2003:SCD

[KM03b] I. Kubiaczyk and P. Majcher. On some continuous and discrete equations in Banach spaces on unbounded intervals. Applied Mathematics and Computation, 136(2–3):463–473, March 15, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kang:2004:BAA

[KML04] Tong Kang, Changfeng Ma, and Guoping Liang. Hbased A − ϕ approaches of approximating eddy current problem by way of solving different systems inside and outside the conductor. Applied Mathematics and Computation, 155(1):1– 24, July 26, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kurita:2000:FDP

[KN00] Kazuo Kurita and Harriet H.

Natsuyama. Five decades of publications by Sueo Ueno. Applied Mathematics and Computation, 116(1–2):11– 38, November 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/91/20/23/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/91/20/
23/article.pdf.

Kulshreshtha:2004:PPM

[KNJ04] Kshitij Kulshreshtha, Neela Nataraj, and Michael Jung. Performance of a parallel mixed finite element implementation for fourth order clamped anisotropic plate bending problems in distributed memory environments. Applied Mathematics and Computation, 155(3): 753–777, August 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kalaba:2000:PDI

[KNU00] Robert E. Kalaba, Harriet H. Natsuyama, and Sueo Ueno. Photon diffusion and invariant imbedding. Applied Mathematics and Computation, 116(1–2):49–59, November 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/91/20/25/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/91/20/
                                        25/article.pdf.

Kirmaci:2004:SID

[KÖ04] U. S. Kirmaci and M. E. Özdemir. On some inequalities for differentiable mappings and applications to special means of real numbers and to midpoint formula. Applied Mathematics and Computation, 153 (2):361–368, June 4, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kokologiannaki:2003:APS

[Kok03] Chrysi G. Kokologiannaki. Absence of the point spectrum in a class of tridiagonal operators. Applied Mathematics and Computation, 136(1):131–138, March 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kolowrocki:2001:LRF

[Koł01] Krzysztof Kołowrocki. On limit reliability functions of large multi-state systems with ageing components. Applied Mathematics and Computation, 121(2– 3):313–361, June 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/105/25/
35/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/105/25/35/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000059.

Koliha:2002:EBG

[Kol02] J. J. Koliha. Error bounds for a general perturbation of the Drazin inverse. Applied Mathematics and Computation, 126(2–3):181–185, March 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/122/30/29/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001491.

Korman:2003:ACG

[Kor03] Philip Korman. An accurate computation of the global solution curve for the Gelfand problem through a two point approximation. Applied Mathematics and Computation, 139(2–3):363– 369, July 15, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kowalenko:2002:EAB

[Kow02] Victor Kowalenko. Exactification of the asymptotics for Bessel and Hankel functions. Applied Mathematics and Computation, 133(2–3): 487–518, December 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Karpuzogullari:2003:SHU

[KÖY03] Sibel Yalçin Karpuzoğullari, Metin Öztürk, and Mümin Yamankaradeniz. A subclass of harmonic univalent functions with negative coefficients. Applied Mathematics and Computation, 142(2–3): 469–476, October 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kozitsky:2003:LEF

[Koz03] Yuri Kozitsky. Laguerre entire functions and the Lee– Yang property. Applied Mathematics and Computation, 141(1):103–112, August 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kadalbajoo:2002:NSS

[KP02a] Mohan K. Kadalbajoo and Kailash C. Patidar. Numerical solution of singularly perturbed two-point boundary value problems by spline in tension. Applied Mathematics and Computation, 131(2–3):299–320, September 25, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kadalbajoo:2002:SNT

[KP02b] Mohan K. Kadalbajoo and Kailash C. Patidar. A survey of numerical techniques for solving singularly perturbed ordinary differential equations. Applied Mathematics and Computation, 130(2–3): 457–510, August 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kadalbajoo:2003:SPP

[KP03] Mohan K. Kadalbajoo and Kailash C. Patidar. Singularly perturbed problems in partial differential equations: a survey. Applied Mathematics and Computation, 134(2–3):371–429, January 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kirane:2001:ONS

[KR01] Mokhtar Kirane and Yuri V.

Rogovchenko. On oscillation of nonlinear second order differential equation with damping term. Applied Mathematics and Computation, 117(2–3):177–192, January 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/92/23/24/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/92/23/
24/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399001721.

Kanth:2003:MIB

[KR03a] A. S. V. Ravi Kanth and Y. N. Reddy. The method of inner boundary condition for singular boundary value problems. Applied Mathematics and Computation, 139(2–3):429–436, July 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kanth:2003:NMSb

[KR03b] A. S. V. Ravi Kanth and Y. N. Reddy. A numerical method for singular two point boundary value problems via Chebyshev economizition. Applied Mathematics and Computation, 146(2–3):691–700, December 31, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kanth:2003:NMSa

[KR03c] A. S. V. Ravi Kanth and Y. N. Reddy. A numerical method for solving two-point boundary value problems over infinite intervals. Applied Mathematics and Computation, 144(2–3): 483–494, December 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kanth:2004:HOF

[KR04] A. S. V. Ravi Kanth and Y. N. Reddy. Higher order finite difference method for a class of singular boundary value problems. Applied Mathematics and Computation, 155(1):249–258, July 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Krutitskii:2003:MIB

[Kru03] P. A. Krutitskii. Method of interior boundaries for the impedance problem in scattering theory. Applied Mathematics and Computation, 135(1):147–160, February 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kalogiratou:2000:SEF

[KS00] Zacharoula Kalogiratou and T. E. Simos. A P -stable exponentially fitted method for the numerical integration of the Schrödinger equation. Applied Mathematics and Computation, 112 (1):99–112, June 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/86/21/
29/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/86/21/29/article.
pdf.

Kwiatuszewska-Sarnecka:2001:RLR

[KS01] Bożena Kwiatuszewska-Sarnecka. A remark on of limit reliability function of large series-parallel systems with assisting components. Applied Mathematics and Computation, 122(2):155–177, July 25, 2001. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/106/25/
22/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/106/25/22/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000187.

Kim:2002:PLP

[KS02a] Sang Dong Kim and Byeong Chun Shin. Preconditioning C^{1} Lagrange polynomial spline collocation method of elliptic equations by finite element method. Applied Mathematics and Computation, 132(2–3):599–616, November 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Koepf:2002:RET

[KS02b] Wolfram Koepf and Dieter Schmersau. Recurrence equations and their classical orthogonal polynomial solutions. Applied Mathematics and Computation, 128(2– 3):303–327, May 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Katti:2003:PMC

[KS03a] C. P. Katti and D. K. Srivastava. On a parallel meshchopping algorithm for a class of initial value problems using fourth-order explicit Runge–Kutta method. Applied Mathematics and Computation, 143(2–3):565– 570, November 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kim:2003:BNA

[KS03b] Min-Soo Kim and Jin-Woo Son. Bernoulli numbers in p-adic analysis. Applied Mathematics and Computation, 146(1):289–297, December 30, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kadalbajoo:2004:NAS

[KS04a] M. K. Kadalbajoo and K. K. Sharma. Numerical analysis of singularly perturbed delay differential equations with layer behavior. Applied Mathematics and Computation, 157(1):11–28, September 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Karakaya:2004:LIS

[KŞ04b] Vatan Karakaya and Necip Şimşek. On lacunary invariant sequence spaces defined by a sequence of modulus functions. Applied Mathematics and Computation, 156(3):597–603, September 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kumar:2002:HBS

[KSC02] S. Kumar, S. K. Srivastava, and P. Chingakham. Hopf bifurcation and stability analysis in a harvested onepredator–two-prey model. Applied Mathematics and Computation, 129(1):107– 118, June 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kou:2002:NFA

[KSJ02] Kit-Ian Kou, Vai-Kuong Sin, and Xiao-Qing Jin. A note on the fast algorithm for block Toeplitz systems with tensor structure. Applied Mathematics and Computation, 126(2–3):187–197, March 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/122/30/30/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001508.

Kananthai:2002:ORW

[KSL02] A. Kananthai, S. Suantai, and V. Longani. On the operator ⊕^{k} related to the wave equation and Laplacian. Applied Mathematics and Computation, 132(2– 3):219–229, November 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kubiaczyk:2003:NOC

[KSM03] I. Kubiaczyk, S. H. Saker, and J. Morchalo. New oscillation criteria for first order nonlinear neutral delay differential equations. Applied Mathematics and Computation, 142(2–3):225–242, October 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Katti:2002:HEP

[KSS02] C. P. Katti, D. K. Srivastava, and S. Sivaloganathan.

Highly efficient parallel algorithm for finite difference solution to Navier– Stoke’s equation on a hypercube. Applied Mathematics and Computation, 130(2–3): 311–316, August 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kawabata:2000:NAC

[KST00] Kiyoshi Kawabata, Makoto Sato, and Larry D. Travis. A new approach to computing theoretical polarization curves along the low latitude regions of a planetary disk observed at small phase angles. Applied Mathematics and Computation, 116(1–2):115–132, November 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/91/20/30/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/91/20/
30/article.pdf.

Kalaba:2001:ADC

[KU01] Robert Kalaba and Firdaus Udwadia. Analytical dynamics with constraint forces that do work in virtual displacements. Applied Mathematics and Computation, 121(2–3):211– 217, June 15, 2001. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/105/25/
29/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/105/25/29/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S009630039900291X.

Kopuzlu:2003:MMG

[KU03] Abdullah Kopuzlu and Tamer Ugur. A method on the matrix of graph of torus knot k(2, q) and its Maple application. Applied Mathematics and Computation, 141(2–3):331–338, September 5, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kucuk:2004:DTL

[Küç04] Ahmet Küçük. On the developable time-like trajectory ruled surfaces in a Lorentz 3-space R_{1}^{3}. Applied Mathematics and Computation, 157(2):483–489, October 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kulhanek:2003:ETM

[Kul03]  J. Kulhánek.  An en-

hancement of transformation method. Applied Mathematics and Computation, 139(1):55–61, July 1, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kumar:2002:FOF

[Kum02] Manoj Kumar. A fourthorder finite difference method for a class of singular twopoint boundary value problems. Applied Mathematics and Computation, 133(2–3): 539–545, December 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kumar:2003:DMS

[Kum03a] Manoj Kumar. A difference method for singular twopoint boundary value problems. Applied Mathematics and Computation, 146(2–3): 879–884, December 31, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kumar:2003:DSB

[Kum03b] Manoj Kumar. A difference scheme based on nonuniform mesh for singular two-point boundary value problems. Applied Mathematics and Computation, 136(2–3):281–288, March 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kumar:2003:NFD

[Kum03c] Manoj Kumar. A new finite difference method for a class of singular two-point boundary value problems. Applied Mathematics and Computation, 143(2–3):551–557, November 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kumar:2003:SOF

[Kum03d]  Manoj Kumar.  A sec-

ond order finite difference method and its convergence for a class of singular twopoint boundary value problems. Applied Mathematics and Computation, 146(2–3): 873–878, December 31, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kumar:2003:SOS

[Kum03e] Manoj Kumar. A second order spline finite difference method for singular twopoint boundary value problems. Applied Mathematics and Computation, 142(2–3): 283–290, October 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See erratum [Kum09].

Kumar:2009:ESO

[Kum09] Manoj Kumar. Erratum to:

“A second order spline finite difference method for singular two-point boundary value problems” [Appl. Math. Comput. 142 (2003) 283–290]. Applied Mathematics and Computation, 209(2):430, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [Kum03e].

Kuo:2004:TBL

[Kuo04]  Bor-Lih Kuo.  Thermal

boundary-layer problems in a semi-infinite flat plate by the differential transformation method. Applied Mathematics and Computation, 150(2):303–320, March 8, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kurowicka:2002:DAA

[Kur02] Dorota Kurowicka. Domains of attraction of asymptotic reliability functions. Applied Mathematics and Computation, 128(1):1–18, May 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Karampetakis:2003:DCG

[KV03] N. P. Karampetakis and S. Vologiannidis. DFT calculation of the generalized and Drazin inverse of a polynomial matrix. Applied Mathematics and Computation, 143(2–3):501–521, November 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Koch:2004:ANT

[KW04a] Othmar Koch and Ewa Weinmüller. Analytical and numerical treatment of a singular initial value problem in avalanche modeling. Applied Mathematics and Computation, 148(2): 561–570, January 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kutluay:2004:NST

[KW04b] S. Kutluay and A. S. Wood.

Numerical solutions of the thermistor problem with a ramp electrical conductivity. Applied Mathematics and Computation, 148(1): 145–162, January 20, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kwon:2001:ISP

[Kwo01] Ohin Kwon. Inverse scattering for a plane screen. Applied Mathematics and Computation, 122(1):59– 69, July 10, 2001. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/106/21/
24/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/106/21/24/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000126.

Kwon:2003:EPI

[Kwo03]  Hee-Dae Kwon.  Effi-

cient parallel implementations of finite element methods based on the conjugate gradient method. Applied Mathematics and Computation, 145(2–3):869–880, December 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Karaduman:2003:PFS

[KY03] Erdal Karaduman and Uğur Yavuz. On the period of Fibonacci sequences in nilpotent groups. Applied Mathematics and Computation, 142(2–3):321–332, October 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kim:2004:AAA

[KY04] Hyun-Sung Kim and Kee-Young Yoo. AOP arithmetic architectures over GF(2^{m}). Applied Mathematics and Computation, 158(1):7–18, October 25, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lackova:2003:APS

[Lac03] Dáša Lacková. The asymptotic properties of the solutions of the nth order functional neutral differential equations. Applied Mathematics and Computation, 146(2–3):385–392, December 31, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lan:2004:MPSa

[Lan04a] K. Q. Lan. Multiple positive solutions of conjugate boundary value problems with singularities. Applied Mathematics and Computation, 147(2):461–474, January 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lan:2004:MPSb

[Lan04b] K. Q. Lan. Multiple positive solutions of Hammerstein integral equations and applications to periodic boundary value problems. Applied Mathematics and Computation, 154 (2):531–542, July 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Layton:2002:CBS

[Lay02] W. Layton. A connection between subgrid scale eddy viscosity and mixed methods. Applied Mathematics and Computation, 133(1):147– 157, November 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Landauer:2001:NAC

[LB01] Christopher Landauer and Kirstie L. Bellman. New architectures for constructed complex systems. Applied Mathematics and Computation, 120(1–3):149–163, May 10, 2001. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/104/21/
33/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/104/21/33/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002404.

Lundberg:2000:LAO

[LBE00] John B. Lundberg, S. Bettadpur, and R. J. Eanes. Long arc orbit determination solutions using Encke’s method. Applied Mathematics and Computation, 113(1):43–53, July 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.nl/gej-ng/
10/9/12/87/21/23/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/87/21/
23/article.pdf.

Li:2001:CEO

[LC01a] Wei Nian Li and Bao Tong Cui. A class of even order neutral differential inequalities and its applications. Applied Mathematics and Computation, 122 (1):95–106, July 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/106/21/
27/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/106/21/27/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000242.

Li:2001:NSC

[LC01b] Wei Nian Li and Bao Tong Cui. Necessary and sufficient conditions for oscillation of neutral delay parabolic differential equations. Applied Mathematics and Computation, 121(2– 3):147–153, June 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/105/25/
23/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/105/25/23/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002684.

Liaw:2002:SNS

[LC02a] Der-Cherng Liaw and Chun-Hone Chen. Stabilization of nonlinear systems in compound critical cases. Applied Mathematics and Computation, 130(2–3):317–360, August 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2002:PEB

[LC02b] Shengqiang Liu and Lansun Chen. Permanence, extinction and balancing survival in nonautonomous Lotka– Volterra system with delays. Applied Mathematics and Computation, 129(2– 3):481–499, July 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liaw:2003:LEQ

[LC03a] Der-Cherng Liaw and Chun-Hone Chen. The linearexponential-quadratic-Gaussian control for discrete systems with application to reliable stabilization. Applied Mathematics and Computation, 137(2–3):303–321, May 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lin:2003:EBE

[LC03b] Jin-Mu Lin and Cha’o-Kuang Chen. Error bounds estimate of nonlinear boundary value problems using method of weighted residuals with genetic algorithms. Applied Mathematics and Computation, 144(2–3):261– 271, December 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2004:APP

[LC04a] Wan-Tong Li and Sui Sun Cheng. Asymptotic properties of the positive equilibrium of a discrete survival model. Applied Mathematics and Computation, 157(1): 29–38, September 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liang:2004:GAS

[LC04b] Jinling Liang and Jinde Cao.

Global asymptotic stability of bi-directional associative memory networks with distributed delays. Applied Mathematics and Computation, 152(2):415–424, May 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2004:NAH

[LC04c] Kuo-Chi Liu and Han-Taw Chen. Numerical analysis for the hyperbolic heat conduction problem under a pulsed surface disturbance. Applied Mathematics and Computation, 159(3):887–901, December 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lu:2004:XMS

[LC04d] Eric Jui-Lin Lu and Rai-Fu Chen. An XML multisignature scheme. Applied Mathematics and Computation, 149(1):1–14, February 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lacis:2000:RAN

[LCH00] A. A. Lacis, B. E. Carlson, and J. E. Hansen. Retrieval of atmospheric NO_{2}, O_{3}, aerosol optical depth, effective radius and variance information from SAGE II multi-spectral extinction measurements. Applied Mathematics and Computation, 116(1–2):133–151, November 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/91/20/31/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/91/20/
31/article.pdf.

Lee:2002:MGP

[LCH02] Hung-Chang Lee, Cha’o-Kuang Chen, and Chen-I Hung. A modified grouppreserving scheme for solving the initial value problems of stiff ordinary differential equations. Applied Mathematics and Computation, 133(2–3):445–459, December 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lu:2004:EAS

[LCN04] Lin-Zhang Lu, Wai-Ki Ching, and Michael K. Ng. Exact algorithms for singular tridiagonal systems with applications to Markov chains. Applied Mathematics and Computation, 159 (1):275–289, November 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lin:2003:CCF

[LCS03] Shy-Der Lin, I-Chun Chen, and H. M. Srivastava. Certain classes of finite-series relationships and generating functions involving the generalized Bessel polynomials. Applied Mathematics and Computation, 137(2– 3):261–275, May 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2004:GRE

[LCXZ04] Biao Li, Yong Chen, Hengnong Xuan, and Hongqing Zhang. Generalized Riccati equation expansion method and its application to the (3 + 1)- dimensional Jumbo–Miwa equation. Applied Mathematics and Computation, 152(2):581–595, May 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2003:ETW

[LCZ03] Biao Li, Yong Chen, and Hongqing Zhang. Exact travelling wave solutions for a generalized Zakharov– Kuznetsov equation. Applied Mathematics and Computation, 146(2–3):653–666, December 31, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2004:GMQ

[LDKU04] Z. Liu, L. Debnath, S. M. Kang, and J. S. Ume. Generalized mixed quasivariational inclusions and generalized mixed resolvent equations for fuzzy mappings. Applied Mathematics and Computation, 149(3):879– 891, February 22, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lee:2000:OBD

[Lee00] Hyesuk Kwon Lee. An optimization-based domain decomposition method for a nonlinear problem. Applied Mathematics and Computation, 113(1):23–42, July 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/9/12/87/21/22/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/9/12/87/21/22/article.
                                        pdf.

Lee:2001:FML

[Lee01]  E. Stanley Lee.  Fuzzy

multiple level programming. Applied Mathematics and Computation, 120(1–3):79– 90, May 10, 2001. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/104/21/
28/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/104/21/28/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002891.

Lee:2004:SPM

[Lee04a] Hyang-Sook Lee. A selfpairing map and its applications to cryptography. Applied Mathematics and Computation, 151(3):671– 678, April 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lee:2004:HNM

[Lee04b] Zong-Yi Lee. Hybrid numerical method applied to 3-D multilayers hollow cylinder with time-dependent boundary conditions. Applied Mathematics and Computation, 150(1):25–43, February 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lee:2004:TTP

[Lee04c] Zong-Yi Lee. Transient thermoelastic problem on the vapor flow of heat pipe. Applied Mathematics and Computation, 154(1):127– 151, June 25, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lesnic:2001:CAI

[Les01] D. Lesnic. A computational algebraic investigation of the decomposition method for time-dependent problems. Applied Mathematics and Computation, 119(2–3):197– 206, April 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/99/25/26/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/99/25/
26/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S009630039900257X.

Lew:2001:NDP

[Lew01] Art Lew. Nondeterministic dynamic programming on a parallel coprocessing system. Applied Mathematics and Computation, 120(1– 3):139–147, May 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/104/21/
32/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/104/21/32/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002829.

Luo:2004:GDR

[LF04] Chengxin Luo and Enmin Feng. Generalized differential Riccati equation and indefinite stochastic LQ control with cross term. Applied Mathematics and Computation, 155(1):121–135, July 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lu:2003:PSK

[LG03] Shiping Lu and Weigao Ge.

Periodic solutions for a kind of second order differential equation with multiple deviating arguments. Applied Mathematics and Computation, 146(1):195–209, December 30, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lu:2004:EPS

[LG04a] Shiping Lu and Weigao Ge.

Existence of periodic solutions for a kind of secondorder neutral functional differential equation. Applied Mathematics and Computation, 157(2):433–448, October 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lu:2004:EPP

[LG04b] Shiping Lu and Weigao Ge.

Existence of positive periodic solutions for neutral population model with multiple delays. Applied Mathematics and Computation, 153(3):885–902, June 14, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lu:2004:PSNb

[LG04c] Shiping Lu and Weigao Ge. Periodic solutions of neutral differential equation with multiple deviating arguments. Applied Mathematics and Computation, 156 (3):705–717, September 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lu:2004:PSNa

[LGZ04] Shiping Lu, Weigao Ge, and Zuxiou Zheng. Periodic solutions to neutral differential equation with deviating arguments. Applied Mathematics and Computation, 152(1):17–27, April 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lee:2004:CUS

[LH04a] Narn-Yih Lee and Pei-Hsiu Ho. Convertible undeniable signature with subliminal channels. Applied Mathematics and Computation, 158(1):169–175, October 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2004:ESG

[LH04b] Xuemei Li and Lihong Huang. Exponential stability and global stability of cellular neural networks. Applied Mathematics and Computation, 147(3):843– 853, January 16, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lee:2003:NKA

[LHL03] Cheng-Chi Lee, Min-Shiang Hwang, and Li-Hua Li. A new key authentication scheme based on discrete logarithms. Applied Mathematics and Computation, 139(2–3):343–349, July 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Luo:2004:OBC

[LHL04] Zhixue Luo, Ze-Rong He, and Wan-Tong Li. Optimal birth control for predator– prey system of three species with age-structure. Applied Mathematics and Computation, 155(3):665–685, August 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2004:OSC

[LHM04] Wei Nian Li, Maoan Han, and Fan Wei Meng. Hoscillation of solutions of certain vector hyperbolic differential equations with deviating arguments. Applied Mathematics and Computation, 158(3):637–653, November 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2002:ECC

[Li02]  Jian-Lin Li.  Evaluation

of certain complex integrals involving absolute values. Applied Mathematics and Computation, 131(1): 15–19, September 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2003:OCM

[Li03a] Jun Li. Order continuous of monotone set function and convergence of measurable functions sequence. Applied Mathematics and Computation, 135(2–3):211–218, March 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2003:FOP

[Li03b] Wei Nian Li. Forced oscillation properties for certain systems of partial functional differential equations. Applied Mathematics and Computation, 143(2–3):223–232, November 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2003:BOC

[Li03c] Yuanqing Li. Bifurcation on the oscillation of a class of second order ordinary differential equations. Applied Mathematics and Computation, 142(1):51–62, September 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2004:DSG

[Li04a] Susu Li. Displacement structure of the generalized inverse A_{T,S}^{(2)}. Applied Mathematics and Computation, 156(1):33–40, August 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2004:CEPa

[Li04b] Wan-Tong Li. Classifications and existence of positive solutions of higherorder nonlinear neutral difference equations. Applied Mathematics and Computation, 152(2):351–366, May 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2004:GSO

[Li04c] Wan-Tong Li. Global stability and oscillation in nonlinear difference equations of population dynamics. Applied Mathematics and Computation, 157(1):115–125, September 27, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2004:IOC

[Li04d] Wan-Tong Li. Interval oscillation criteria for secondorder quasi-linear nonhomogeneous differential equations with damping. Applied Mathematics and Computation, 147(3):753–763, January 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2004:IOS

[Li04e] Wan-Tong Li. Interval oscillation of second-order halflinear functional differential equations. Applied Mathematics and Computation, 155(2):451–468, August 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2004:CMF

[Li04f] Xin Li. Convergence of the method of fundamental solutions for solving the boundary value problem of modified Helmholtz equation. Applied Mathematics and Computation, 159(1):113– 125, November 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2004:ESP

[Li04g] Yongkun Li. Existence and stability of periodic solution for BAM neural networks with distributed delays. Applied Mathematics and Computation, 159(3): 847–862, December 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liao:2003:EAS

[Lia03a]  Shijun Liao.  An ex-

plicit analytic solution to the Thomas–Fermi equation. Applied Mathematics and Computation, 144(2–3): 495–506, December 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liao:2003:NAA

[Lia03b] Shijun Liao. A new analytic algorithm of Lane–Emden type equations. Applied Mathematics and Computation, 142(1):1–16, September 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liao:2004:HAM

[Lia04] Shijun Liao. On the homotopy analysis method for nonlinear problems. Applied Mathematics and Computation, 147(2):499–513, January 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lien:2004:EMD

[Lie04] Chang-Hua Lien. An efficient method to design robust observer-based control of uncertain linear systems. Applied Mathematics and Computation, 158 (1):29–44, October 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lin:2004:NOS

[Lin04a] Wen-Xian Lin. A note on oscillation for systems of high order quasilinear partial differential equations of neutral type. Applied Mathematics and Computation, 156(2): 563–576, September 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lin:2004:OSH

[Lin04b]  Wen-Xian Lin.  Oscilla-

tion for systems of higherorder neutral type delay partial differential equations. Applied Mathematics and Computation, 156(1):107– 114, August 25, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lin:2004:SOT

[Lin04c] Wen-Xian Lin. Some oscillation theorems for systems of even order quasilinear partial differential equations. Applied Mathematics and Computation, 152 (2):337–349, May 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2000:EPE

[Liu00] Jinn-Liang Liu. Exact a posteriori error analysis of the least squares finite element method. Applied Mathematics and Computation, 116(3):297–305, December ??, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/91/23/25/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/91/23/
25/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399001538.

Liu:2001:UPU

[Liu01] Baoding Liu. Uncertain programming: a unifying optimization theory in various uncertain environments. Applied Mathematics and Computation, 120(1–3):227– 234, May 10, 2001. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/104/21/
38/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/104/21/38/article.

                                        pdf;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300399002428.

Liu:2002:PSN

[Liu02a] Bing Liu. Positive solutions of a nonlinear threepoint boundary value problem. Applied Mathematics and Computation, 132(1): 11–28, October 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2002:CFF

[Liu02b] Xian Liu. A computable filled function used for global minimization. Applied Mathematics and Computation, 126(2–3):271–278, March 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        9/12/122/30/36/abstract.
                                        html;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300300001570.

Liu:2002:SFF

[Liu02c] Xian Liu. Several filled functions with mitigators. Applied Mathematics and Computation, 133(2–3):375–387, December 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2003:SMPa

[Liu03a] Bing Liu. Solvability of multi-point boundary value problem at resonance (II). Applied Mathematics and Computation, 136(2–3):353– 377, March 15, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2003:SMPb

[Liu03b] Bing Liu. Solvability of multi-point boundary value problem at resonance. Part IV. Applied Mathematics and Computation, 143(2– 3):275–299, November 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2003:ASS

[Liu03c] Xian Liu. On algebraic structures of spatial ordering and location code. Applied Mathematics and Computation, 136(1):99– 122, March 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2003:OSO

[Liu03d] Xian Liu. On operations of spatial ordering and location code. Applied Mathematics and Computation, 138(2– 3):503–522, June 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2003:PQR

[Liu03e] Xian Liu. On performance of quadrant-recursive spatial orders. Applied Mathematics and Computation, 141(2–3): 499–527, September 5, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2003:BVP

[Liu03f]  Yansheng Liu.  Bound-

ary value problems for second order differential equations on unbounded domains in a Banach space. Applied Mathematics and Computation, 135(2–3):569–583, March 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2003:EUP

[Liu03g] Yansheng Liu. Existence and unboundedness of positive solutions for singular boundary value problems on halfline. Applied Mathematics and Computation, 144(2–3): 543–556, December 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2003:SPC

[Liu03h] Zhong-Yun Liu. Some properties of centrosymmetric matrices. Applied Mathematics and Computation, 141(2–3):297–306, September 5, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2004:PSF

[Liu04a]  B. Liu.  Positive solu-

tions of fourth-order two point boundary value problems. Applied Mathematics and Computation, 148(2): 407–420, January 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2004:NNB

[Liu04b] Bing Liu. A note on a nonlocal boundary value problems. Applied Mathematics and Computation, 154 (3):871–880, July 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2004:PSN

[Liu04c] Bing Liu. Positive solutions of a nonlinear fourpoint boundary value problems. Applied Mathematics and Computation, 155 (1):179–203, July 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2004:SRF

[Liu04d] Guodong Liu. Summation and recurrence formula involving the central factorial numbers and zeta function. Applied Mathematics and Computation, 149(1): 175–186, February 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2004:BAF

[Liu04e] Xian Liu. The barrier attribute of filled functions. Applied Mathematics and Computation, 149(3):641– 649, February 22, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2004:FAP

[Liu04f] Xian Liu. Four alternative patterns of the Hilbert curve. Applied Mathematics and Computation, 147(3): 741–752, January 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2004:IFM

[Liu04g] Xian Liu. The impelling function method applied to global optimization. Applied Mathematics and Computation, 151(3):745–754, April 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2004:TNC

[Liu04h] Xian Liu. Two new classes of filled functions. Applied Mathematics and Computation, 149(2):577–588, February 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2004:NDF

[Liu04i] Zhong-Yun Liu. A note on the determinant formulas computation of generalized inverse matrix Padé approximation. Applied Mathematics and Computation, 150 (3):865–873, March 17, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2004:NSD

[LJ04] Xi-Qiang Liu and Song Jiang. New solutions of the 3 + 1 dimensional Jimbo– Miwa equation. Applied Mathematics and Computation, 158(1):177–184, October 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2004:CSP

[LK04] Zeqing Liu and Shin Min Kang. Convergence and stability of perturbed threestep iterative algorithm for completely generalized nonlinear quasivariational inequalities. Applied Mathematics and Computation, 149(1):245–258, February 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liang:2000:ASD

[LL00] Yew-Wen Liang and Der-Cherng Liaw. On asymptotic stabilization of driftless systems. Applied Mathematics and Computation, 114(2–3):303–314, September 11, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/89/22/34/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/89/22/
34/article.pdf.

Liang:2003:RCN

[LL03] Yew-Wen Liang and Der-Cherng Liaw. Robust control of non-linear affine systems. Applied Mathematics and Computation, 137(2– 3):337–347, May 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lee:2004:MPS

[LL04a] Eun Kyoung Lee and Yong-Hoon Lee. Multiple positive solutions of singular two point boundary value problems for second order impulsive differential equations. Applied Mathematics and Computation, 158(3):745– 759, November 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lee:2004:FIM

[LL04b] Narn-Yih Lee and Ming-Feng Lee. Further improvement on the modified authenticated key agreement scheme. Applied Mathematics and Computation, 157 (3):729–733, October 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2004:PAS

[LL04c] M. Z. Liu and Dongsong Li.

Properties of analytic solution and numerical solution of multi-pantograph equation. Applied Mathematics and Computation, 155(3): 853–871, August 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2004:PPS

[LL04d] Ping Liu and Yongkun Li.

Positive periodic solutions of infinite delay functional differential equations depending on a parameter. Applied Mathematics and Computation, 150(1):159–168, February 27, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2003:HMC

[LLB03] Ting Li, Zhenjiang Lin, and Fengshan Bai. Heuristic methods for computing the minimal multi-homogeneous Bézout number. Applied Mathematics and Computation, 146(1):237–256, December 30, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liao:2003:TFT

[LLCC03] Chung-Min Liao, Huang-Min Liang, Jein-Wen Chen, and Jui-Sheng Chen. A transfer function technique to describe odor causing VOCs transport in a ventilated airspace with mixing/adsorption heterogeneity. Applied Mathematics and Computation, 140(2–3): 255–277, August 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2002:HSE

[lLgC02] Chun li Li and Ming gen Cui. How to solve the equation AuBu + Cu = f. Applied Mathematics and Computation, 133(2–3):643–653, December 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2003:ESS

[lLgC03] Chun li Li and Ming gen Cui. The exact solution for solving a class nonlinear operator equations in the reproducing kernel space. Applied Mathematics and Computation, 143(2–3):393–399, November 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lee:2004:MPA

[LLL04] Young-Ran Lee, Hyang-Sook Lee, and Ho-Kyu Lee. Multi-party authenticated key agreement protocols from multi-linear forms. Applied Mathematics and Computation, 159(2):317– 331, December 6, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Layton:2002:DCM

[LLP02] W. Layton, H. K. Lee, and J. Peterson. A defectcorrection method for the incompressible Navier–Stokes equations. Applied Mathematics and Computation, 129(1):1–19, June 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Luo:2004:OHC

[LLW04] Zhixue Luo, Wan-Tong Li, and Miansen Wang. Optimal harvesting control problem for linear periodic agedependent population dynamics. Applied Mathematics and Computation, 151 (3):789–800, April 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2002:BAP

[LM02] Liping Liu and Ming Mei. A better asymptotic profile of Rosenau–Burgers equation.

Applied Mathematics and Computation, 131(1):147– 170, September 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2003:FOC

[LM03] Wei Nian Li and Fan Wei Meng. Forced oscillation for certain systems of hyperbolic differential equations. Applied Mathematics and Computation, 141(2–3): 313–320, September 5, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Labbas:2004:SBV

[LM04a] Rabah Labbas and Stéphane Maingot. Singularities in boundary value problems for an abstract second-order differential equation of elliptic type. Applied Mathematics and Computation, 148(3): 645–663, January 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Linss:2004:FEA

[LM04b] Torsten Linß and Niall Madden. A finite element analysis of a coupled system of singularly perturbed reaction–diffusion equations. Applied Mathematics and Computation, 148(3):869– 880, January 30, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Labro:2000:AHM

[LMG00] H. Labro, B. Maheu, and A. Garo. Applicability of the Homotopy Method to the determination of fixed points in chemical kinetics models. Applied Mathematics and Computation, 109 (1):31–49, March 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/82/21/
23/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/82/21/23/article.
pdf.

Labbas:2002:PET

[LMS02] Rabah Labbas, Ahmed Medeghri, and Boubaker-Khaled Sadallah. On a parabolic equation in a triangular domain. Applied Mathematics and Computation, 130(2–3):511–523, August 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2004:EPSa

[LNS04] Wan-Tong Li, Ming-Fei Niu, and Jian-Ping Sun. Existence of positive solutions of BVPs for second-order nonlinear difference systems. Applied Mathematics and Computation, 152(3):779– 798, May 13, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lorentz:2003:CDT

[LNW03] R. A. Lorentz, F. J. Narcowich, and J. D. Ward. Collocation discretizations of the transport equation with radial basis functions. Applied Mathematics and Computation, 145(1):97–116, December 20, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Longani:2000:AAT

[Lon00] Vites Longani. Another approach for the traveling salesman problem. Applied Mathematics and Computation, 114(2–3):249–253, September 11, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/
                                        10/9/12/89/22/30/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/89/22/
                                        30/article.pdf.

Lee:2000:SGF

[LOS00] Poh-Aun Lee, Seng-Huat Ong, and H. M. Srivastava. Some generating functions for the Laguerre and related polynomials. Applied Mathematics and Computation, 108(2–3):129–138, February 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/79/22/
27/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/79/22/27/article.
pdf.

Lee:2002:FDA

[LOS02] H. Y. Lee, M. R. Ohm, and J. Y. Shin. Fully discrete approximation for a quasilinear Stefan problem with forcing term. Applied Mathematics and Computation, 133(2–3): 461–478, December 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lu:2002:MPS

[LOZ02] Haishen Lü, Donal O’Regan, and Chengkui Zhong. Multiple positive solutions for the one-dimensional singular p-Laplacian. Applied Mathematics and Computation, 133(2–3):407–422, December 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Landaburu:2003:MDE

[LP03] E. Landaburu and L. Pardo.

Minimum (h, ϕ)-divergences estimators with weights. Applied Mathematics and Computation, 140(1):15–28, July 30, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Logan:2002:NSR

[LPS02] J. David Logan, Mark R. Petersen, and Thomas S. Shores. Numerical study of reaction-mineralogy-porosity changes in porous media. Applied Mathematics and Computation, 127(2–3):149– 164, April 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        9/12/123/27/27/abstract.
                                        html;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300301000522.

Liu:2004:TPS

[LQG04] Xiujun Liu, Jiqing Qiu, and Yanping Guo. Three positive solutions for second-order m-point boundary value problems. Applied Mathematics and Computation, 156(3):733–742, September 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2003:OCG

[LQL03] Wenxiang Liu, Maochang Qin, and Yongkun Li. Oscillatory criteria of general nonlinear hyperbolic equations with continuous deviating arguments. Applied Mathematics and Computation, 137(2–3):451–458, May 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2004:FEF

[LqZ04] De-Sheng Li and Hong qing Zhang. A further extended tanh-function method and new soliton-like solutions to the integrable Broer–Kaup (BK) equations in (2 + 1) dimensional spaces. Applied Mathematics and Computation, 147(2):537–545, January 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2004:CEPb

[LR04] Wan-Tong Li and Youssef N.

Raffoul. Classification and existence of positive solutions of systems of Volterra nonlinear difference equations. Applied Mathematics and Computation, 155 (2):469–478, August 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Leyva-Ramos:2001:ROC

[LRHRD01] J. Leyva-Ramos, A. Hernandez-Rodriguez, and E. D. Denman. Reduced-order controllers for unstable multivariable linear systems that can be balanced. Applied Mathematics and Computation, 120(1–3):55–63, May 10, 2001. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/104/21/
26/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/104/21/26/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002866.

Leyva-Ramos:2001:SAC

[LRMSVO01] Jesus Leyva-Ramos, Jorge Alberto Morales-Saldaña, and Omar Vital-Ochoa. μstability analysis for currentprogrammed regulators. Applied Mathematics and Computation, 120(1–3):3–14, May 10, 2001. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/104/21/
22/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/104/21/22/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002878.

Luo:2002:NRT

[LS02] Zhiguo Luo and Jianhua Shen. New Razumikhin type theorems for impulsive functional differential equations. Applied Mathematics and Computation, 125(2–3):375–386, January 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/120/31/45/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001399.

Li:2003:OSO

[LS03a] Wan-Tong Li and S. H.

Saker. Oscillation of secondorder sublinear neutral delay difference equations. Applied Mathematics and Computation, 146(2–3):543–551, December 31, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2003:SNM

[LS03b] Hong Liu and Yongzhong Song. Stability of numerical methods for solving linear index-3 DAEs. Applied Mathematics and Computation, 134(1):35–50, January 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lee:2004:EEG

[LS04a] S. H. Lee and B. K. Soni.

The enhancement of an elliptic grid using appropriate control functions. Applied Mathematics and Computation, 159(3):809–821, December 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2004:EPSb

[LS04b] Wan-Tong Li and Jian-Ping Sun. Existence of positive solutions of BVPs for thirdorder discrete nonlinear difference systems. Applied Mathematics and Computation, 157(1):53–64, September 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2004:MPS

[LS04c] Wan-Tong Li and Jian-Ping Sun. Multiple positive solutions of BVPs for thirdorder discrete difference systems. Applied Mathematics and Computation, 149(2): 389–398, February 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lin:2004:SFH

[LS04d] Shy-Der Lin and H. M. Srivastava. Some families of the Hurwitz–Lerch Zeta functions and associated fractional derivative and other integral representations. Applied Mathematics and Computation, 154 (3):725–733, July 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2004:CPD

[LS04e] Jin-Lin Liu and H. M. Srivastava. Certain properties of the Dziok–Srivastava operator. Applied Mathematics and Computation, 159(2): 485–493, December 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2004:HAC

[LS04f] Xian Liu and Günther F.

Schrack. A heuristic approach for constructing symmetric Gray codes. Applied Mathematics and Computation, 155(1):55–63, July 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lakshmikantham:2000:NHA

[LSJR00] V. Lakshmikantham, S. K.

Sen, M. K. Jain, and A. Ramful. O(n^{3}) noniterative heuristic algorithm for linear programs with error-free implementation. Applied Mathematics and Computation, 110(1):53– 81, April 1, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/83/21/
22/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/83/21/22/article.
pdf.

Li:2000:NSD

[LSK00] Weiye Li, Ferenc Szidarovszky, and Yang Kuang. Notes on the stability of dynamic economic systems. Applied Mathematics and Computation, 108(2–3):85– 89, February 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/79/22/
23/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/79/22/23/article.
pdf.

Lee:2003:ESS

[LSY03] Chung-Yi Lee, H. M. Srivastava, and Wen-Chyuan Yueh. Explicit solutions of some linear ordinary and partial fractional differintegral equations. Applied Mathematics and Computation, 144(1):11–25, November 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lin:2002:CCO

[LTS02] Shy-Der Lin, Shih-Tong Tu, and H. M. Srivastava. Certain classes of ordinary and partial differential equations solvable by means of fractional calculus. Applied Mathematics and Computation, 131(2–3):223–233, September 25, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lundberg:2000:MSO

[Lun00] John B. Lundberg. Mitigation of satellite orbit errors resulting from the numerical integration across shadow boundaries. Applied Mathematics and Computation, 112(2–3):193–211, June 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/86/23/
24/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/86/23/24/article.
pdf.

Lundberg:2001:AAG

[Lun01] John B. Lundberg. Alternative algorithms for the GPS static positioning solution. Applied Mathematics and Computation, 119(1):21–34, March 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/99/20/21/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/99/20/
21/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399002192.

Luo:2002:OLO

[Luo02a] Jiaowan Luo. Oscillation and linearized oscillation of a logistic equation with several delays. Applied Mathematics and Computation, 131(2–3):469–476, September 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Luo:2002:OHP

[Luo02b] Jiaowan Luo. Oscillation of hyperbolic partial differential equations with impulses. Applied Mathematics and Computation, 133(2–3): 309–318, December 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2000:ECQ

[LW00] Xiaoying Liu and Congxin Wu. Existence of coupled quasi-fixed points for mixed monotone operators and its application to the discontinuous integral equations. Applied Mathematics and Computation, 112(2– 3):171–180, June 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/86/23/
                                        22/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/86/23/22/article.
                                        pdf.

Li:2003:TSE

[LW03a] Qinghong Li and Xinyuan Wu. A two-step explicit P -stable method for solving second order initial value problems. Applied Mathematics and Computation, 138(2–3):435–442, June 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2003:NPB

[LW03b] Xiezhang Li and Yimin Wei.

A note on the perturbation bound of the Drazin inverse. Applied Mathematics and Computation, 140(2–3): 329–340, August 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2004:TSE

[LW04a] Qinghong Li and Xinyuan Wu. A two-step explicit P - stable method of high phaselag order for second order IVPs. Applied Mathematics and Computation, 151 (1):17–26, March 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2004:EDI

[LW04b] Xiezhang Li and Yimin Wei.

An expression of the Drazin inverse of a perturbed matrix. Applied Mathematics and Computation, 153 (1):187–198, May 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2004:IMD

[LW04c] Xiezhang Li and Yimin Wei.

Iterative methods for the Drazin inverse of a matrix with a complex spectrum. Applied Mathematics and Computation, 147(3): 855–862, January 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2004:RES

[LW04d] Yonghui Liu and Musheng Wei. Rank equalities for submatrices in generalized inverse M_{T,S}^{(2)} of M. Applied Mathematics and Computation, 152(2):499–504, May 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2004:RER

[LW04e] Yonghui Liu and Musheng Wei. Rank equalities related to the generalized inverses A_{T,S}^{(2)}, B_{T_{1},S_{1}}^{(2)} of two matrices A and B. Applied Mathematics and Computation, 159(1):19–28, November 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wu:2003:OCSa

[lWCH03] Xiu li Wu, Si-Yang Chen, and Ji Hong. Oscillation of a class of second-order nonlinear ODE with impulses. Applied Mathematics and Computation, 138(2– 3):181–188, June 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2000:TSI

[LWL00] Zhong-Yun Liu, He-Bing Wu, and Lu Lin. The twostage iterative methods for symmetric positive definite matrices. Applied Mathematics and Computation, 114(1):1–12, August 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.nl/gej-ng/
10/9/12/89/20/20/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/89/20/
20/article.pdf.

Lee:2000:OAN

[LWT00] W. R. Lee, S. Wang, and K. L. Teo. An optimization approach to numerical integration in two dimensions. Applied Mathematics and Computation, 109(2– 3):205–223, March 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/82/23/
29/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/82/23/29/article.
pdf.

Liao:2003:SBP

[LwWW03] Xiaofeng Liao, Kwok wo Wong, and Zhongfu Wu. Stability of bifurcating periodic solutions for van der Pol equation with continuous distributed delay. Applied Mathematics and Computation, 146(2–3):313–334, December 31, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wu:2003:OCSb

[lWyCjT03] Xiu li Wu, Si yang Chen, and Hong ji Tang. Oscillation of a class of secondorder delay differential equation with impulses. Applied Mathematics and Computation, 145(2–3):561–567, December 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2000:MFA

[LWZ00] Hong Li, Luther White, and M. Zaman. A model for fluid accumulation and flow in a three-dimensional porous medium. Applied Mathematics and Computation, 108 (2–3):177–196, February 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/79/22/
                                        31/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/79/22/31/article.
                                        pdf.

Li:2003:CEI

[LX03] Guocheng Li and Xiaoping Xue. Controllability of evolution inclusions with nonlocal conditions. Applied Mathematics and Computation, 141(2–3):375–384, September 5, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2002:ESP

[LY02a] Bing Liu and Jianshe Yu. Existence of solution for m-point boundary value problems of secondorder differential systems with impulses. Applied Mathematics and Computation, 125(2–3):155–175, January 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/120/31/28/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001107.

Liu:2002:SMP

[LY02b] Bing Liu and Jianshe Yu.

Solvability of multi-point boundary value problem at resonance (III). Applied Mathematics and Computation, 129(1):119–143, June 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lu:2004:EEE

[LYT04] Qishao Lu, Zuodong Yang, and E. H. Twizell. Existence of entire explosive positive solutions of quasi-linear elliptic equations. Applied Mathematics and Computation, 148(2):359–372, January 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lakshmikantham:2001:SPSa

[LZ01a] V. Lakshmikantham and Y. Zhang. Strict practical stability of delay differential equation. Applied Mathematics and Computation, 118(2–3):275–285, March 9, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/96/25/29/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/96/25/
29/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399002222.

Lakshmikantham:2001:SPSb

[LZ01b] V. Lakshmikantham and Y. Zhang. Strict practical stability of delay differential equation. Applied Mathematics and Computation, 122(3):341–351, August 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/106/26/
25/abstract.html;   http:
//www.elsevier.nl/gej-

ng/10/9/12/106/26/25/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000497.

Li:2004:NET

[LZ04a] De-Sheng Li and Hong-Qing Zhang. A new extended tanh-function method and its application to the dispersive long wave equations in (2+1) dimensions. Applied Mathematics and Computation, 147(3):789–797, January 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2004:SNT

[LZ04b] De-Sheng Li and Hong-Qing Zhang. Some new types of multisoliton solutions for the (2 + 1)-dimensional higherorder Broer–Kaup system. Applied Mathematics and Computation, 152(3):847– 853, May 13, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2004:GAS

[LZ04c] Xianyi Li and Deming Zhu.

Global asymptotic stability for two recursive difference equations. Applied Mathematics and Computation, 150(2):481–492, March 8, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2004:PPSb

[LZ04d] Yongkun Li and Lifei Zhu. Positive periodic solutions for a class of higherdimensional state-dependent delay functional differential equations with feedback control. Applied Mathematics and Computation, 159(3): 783–795, December 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2004:PPSa

[LZ04e] Yongkun Li and Lifei Zhu. Positive periodic solutions of nonlinear functional differential equations. Applied Mathematics and Computation, 156(2):329–339, September 6, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lu:2004:AFE

[LZ04f] Zhuosheng Lü and Hongqing Zhang. Applications of a further extended tanh method. Applied Mathematics and Computation, 159(2):401– 406, December 6, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2000:BTL

[LZL00] Yuanqing Li, Xinzheng Zhang, and Yongqing Liu. Basic theory of linear singular discrete systems with delay. Applied Mathematics and Computation, 108 (1):33–46, February 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/79/21/
24/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/79/21/24/article.
pdf.

Ma:2001:CME

[Ma01]  Naiyang Ma.  Complete

multinomial expansions. Applied Mathematics and Computation, 124(3):365–370, December 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/113/32/34/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001028.

Ma:2004:MPS

[Ma04a] Ruyun Ma. Multiple positive solutions for nonlinear m-point boundary value problems. Applied Mathematics and Computation, 148(1):249–262, January 20, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mohammed:2004:SSL

[MA04b] A. S. Mohammed and H. A.

Atia. Separation of the Sturm–Liouville differential operator with an operator potential. Applied Mathematics and Computation, 156(2):387–394, September 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Muttlak:2004:WMR

[MAD04] Hassen A. Muttlak and Walid Abu-Dayyeh. Weighted modified ranked set sampling methods. Applied Mathematics and Computation, 151(3):645–657, April 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Madan:2003:TSQ

[MADT03] Kailash C. Madan, Walid Abu-Dayyeh, and Firas Taiyyan. A two server queue with Bernoulli schedules and a single vacation policy. Applied Mathematics and Computation, 145(1): 59–71, December 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Maeda:2001:FLP

[Mae01] Takashi Maeda. Fuzzy linear programming problems as bi-criteria optimization problems. Applied Mathematics and Computation, 120(1–3):109–121, May 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/104/21/
30/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/104/21/30/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002374.

Mahmoud:2004:VWF

[MAE04a] F. S. Mahmoud and A. A.

Abd-Ellatif. On various weak fuzzy pairwise retracts. Applied Mathematics and Computation, 155 (1):39–54, July 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mahmoud:2004:FTF

[MAE04b] F. S. Mahmoud, M. A. Fath Alla, and S. M. Abd Ellah. Fuzzy topology on fuzzy sets: fuzzy semicontinuity and fuzzy semiseparation axioms. Applied Mathematics and Computation, 153 (1):127–140, May 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Magden:2004:ATO

[Mağ04] Abdullah Mağden. On applications of the Tachibana operator. Applied Mathematics and Computation, 147 (1):45–55, January 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mahmoud:2003:SSS

[MAK03] F. S. Mahmoud, M. A. Fath Alla, and M. M. Khalaf. F_{μ}-strongly semiopen sets, F_{μ}-strong semicontinuity and F_{μ}-strongly semiretracts. Applied Mathematics and Computation, 137(2– 3):209–230, May 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mahmoud:2004:OSC

[MAK04a] F. S. Mahmoud, M. A. Fath Alla, and M. M. Khalaf. F − γ-open sets and F − γcontinuity in fuzzy bitopological spaces. Applied Mathematics and Computation, 153(1):117–126, May 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Makai:2004:RSP

[Mak04b] Márton Makai. Reroute sequence planning in telecommunication networks and compact vector summation. Applied Mathematics and Computation, 150(3):785– 801, March 17, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mancera:2003:SNS

[Man03]  P. F. A. Mancera.  A

study of a numerical solution of the steady two dimensions Navier–Stokes equations in a constricted channel problem by a compact fourth order method. Applied Mathematics and Computation, 146(2–3):771–790, December 31, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Madan:2004:QOR

[MANAM04] Kailash C. Madan, Amjad D. Al-Nasser, and Abedel-Qader Al-Masri. On \binom{M^{[x]}/G_{1}}{G_{2}/1} queue with optional re-service. Applied Mathematics and Computation, 152(1):71–88, April 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Maruster:2001:SGL

[Mar01] St. Maruster. The stability of gradient-like methods. Applied Mathematics and Computation, 117(1): 103–115, January 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.nl/gej-ng/
10/9/12/92/20/28/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/92/20/
28/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399001435.

Madan:2004:SSQ

[MAR04a] Kailash C. Madan and Adel Z. Abu Al-Rub. On a single server queue with optional phase type server vacations based on exhaustive deterministic service and a single vacation policy. Applied Mathematics and Computation, 149(3): 723–734, February 22, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Marik:2004:IAO

[Mař04b] Robert Mařík. Integral averages and oscillation criteria for half-linear partial differential equation. Applied Mathematics and Computation, 150(1):69–87, February 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Markakis:2004:RNL

[Mar04c]  M. P. Markakis.  On

the reduction of non-linear oscillator-equations to Abel forms. Applied Mathematics and Computation, 157(2): 357–368, October 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Massoudi:2003:BLF

[Mas03] Mehrdad Massoudi. Boundary layer flow of a second grade fluid with variable heat flux at the wall. Applied Mathematics and Computation, 143(2–3):201–212, November 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Matsumoto:2000:OTD

[Mat00] Masamichi Matsumoto. The nth order time-dependent reflection function for a finite homogeneous atmosphere. Applied Mathematics and Computation, 116(1– 2):61–77, November 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.nl/gej-ng/
10/9/12/91/20/26/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/91/20/
26/article.pdf.

Matsumoto:2001:UAS

[Mat01] Masamichi Matsumoto. The unified approach in a semiinfinite atmosphere. Applied Mathematics and Computation, 120(1–3):289–300, May 10, 2001. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/104/21/
43/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/104/21/43/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002465.

Milatovic:2004:AMM

[MB04] Milan Milatovic and Adedeji B. Badiru. Applied mathematics modeling of intelligent mapping and scheduling of interdependent and multi-functional project resources. Applied Mathematics and Computation, 149(3):703–721, February 22, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Moran:2000:SRT

[MC00] Kevin M. Moran and David A. Cicci. Sensitivity of ridgetype estimation methods to condition number. Applied Mathematics and Computation, 112(1):143–159, June 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/86/21/
33/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/86/21/33/article.
pdf.

Marcozzi:2001:UBC

[MCC01] Michael D. Marcozzi, Seungmook Choi, and C. S. Chen. On the use of boundary conditions for variational formulations arising in financial mathematics. Applied Mathematics and Computation, 124(2):197–214, October 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/113/31/30/abstract.

html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000874.

McRae:2001:PLF

[McR01] F. A. McRae. Perturbing Lyapunov functions and stability criteria for initial time difference. Applied Mathematics and Computation, 117(2–3):313–320, January 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/92/23/34/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/92/23/
34/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399002064.

Matsumoto:2000:ISI

[MCS00] Gen Matsumoto, John L.

Casti, and Masanori Sugisaka. Introduction to the special issue on artificial brain, brain computing, and brainware. Applied Mathematics and Computation, 111(2–3):133–135, May 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/85/23/
21/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/85/23/21/article.
pdf.

Magden:2004:HLA

[MCS04] A. Mağden, N. Cengiz, and A. A. Salimov. Horizontal lift of affinor structures and its applications. Applied Mathematics and Computation, 156(2):455–461, September 6, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mohd:2000:CEM

[MD00] Ismail Bin Mohd and Yosza Dasril. Constraint exploration method for quadratic programming problem. Applied Mathematics and Computation, 112(2–3):161–170, June 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/86/23/
                                        21/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/86/23/21/article.
                                        pdf.

Mesquita:2004:OMS

[MdL04] Maximilian S. Mesquita and Marcelo J. S. de Lemos. Optimal multigrid solutions of two-dimensional convection– conduction problems. Applied Mathematics and Computation, 152(3):725–742, May 13, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Menezes:2004:EUU

[MdOPF04] Silvano D. B. Menezes, Eliane A. de Oliveira, Ducival C. Pereira, and Jorge Ferreira. Existence, uniqueness and uniform decay for the nonlinear beam degenerate equation with weak damping. Applied Mathematics and Computation, 154(2):555–565, July 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ma:2004:ISB

[MdS04] To Fu Ma and Jair da Silva.

Iterative solutions for a beam equation with nonlinear boundary conditions of third order. Applied Mathematics and Computation, 159(1):11–18, November 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mefire:2002:MFE

[Mef02] S. M. Mefire. Mixed finite element and boundary element approximation in 3D magnetostatics for computation of the magnetic induction. Applied Mathematics and Computation, 125(2–3):399–421, January 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/120/31/47/abstract.
html;   http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300300001429.

Mehrazin:2000:OSR

[Meh00a] H. Mehrazin. Optimum shapes of reservoirs. Applied Mathematics and Computation, 110(1):83–97, April 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/83/21/
                                        23/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/83/21/23/article.
                                        pdf.

Mehrazin:2000:RPD

[Meh00b] Hashem Mehrazin. Resolution of partial differential equations by rotation. Applied Mathematics and Computation, 110(2–3):225– 237, April 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/83/23/
                                        28/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/83/23/28/article.
                                        pdf.

Mehrazin:2001:OSR

[Meh01] H. Mehrazin. Optimum shapes of reservoirs. Applied Mathematics and Computation, 117(1):23–33, January 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/92/20/22/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/92/20/
22/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399001484.

Mehrazin:2003:LEP

[Meh03] Hashem Mehrazin. Laplace equation and Poisson integral. Applied Mathematics and Computation, 145(2–3): 451–463, December 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mekheimer:2004:PFB

[Mek04] Kh. S. Mekheimer. Peristaltic flow of blood under effect of a magnetic field in a non-uniform channels. Applied Mathematics and Computation, 153 (3):763–777, June 14, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Melnik:2000:GSD

[Mel00] R. V. N. Melnik. Generalised solutions, discrete models and energy estimates for a 2D problem of coupled field theory. Applied Mathematics and Computation, 107(1):27–55, January ??, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.

                                        elsevier.nl/gej-ng/10/
                                        9/12/72/17/18/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/72/17/
                                        18/article.pdf.

Melnik:2001:DMC

[Mel01] Roderick V. N. Melnik. Discrete models of coupled dynamic thermoelasticity for stress-temperature formulations. Applied Mathematics and Computation, 122 (1):107–132, July 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/9/12/106/21/
                                        28/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/9/12/106/21/28/article.
                                        pdf;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300300000266.

Meng:2002:ORL

[Men02] Fanwei Meng. Oscillation results for linear Hamiltonian systems. Applied Mathematics and Computation, 131(2–3):357–372, September 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Moghadas:2003:DNA

[MG03a] S. M. Moghadas and A. B. Gumel. Dynamical and numerical analyses of a generalized food-chain model. Applied Mathematics and Computation, 142(1):35–49, September 20, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mohamad:2003:ESC

[MG03b] S. Mohamad and K. Gopalsamy. Exponential stability of continuous-time and discrete-time cellular neural networks with delays. Applied Mathematics and Computation, 135(1):17–38, February 15, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Milczek:2003:CLR

[Mil03] Beata Milczek. On the class of limit reliability functions of homogeneous series- “k-out-of-n” systems. Applied Mathematics and Computation, 137(1):161–176, May 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Miyakawa:2000:LSN

[MIM00] N. Miyakawa, M. Ichikawa, and G. Matsumoto. Largescale neural network method for brain computing. Applied Mathematics and Computation, 111(2–3):203–208, May 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/

                                        gej-ng/29/17/20/85/23/
                                        26/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/85/23/26/article.
                                        pdf.

Minchev:2003:OSNb

[Min03] Emil Minchev. On the oscillations of solutions of nonlinear parabolic equations. Applied Mathematics and Computation, 136(2–3):453–462, March 15, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Minchev:2004:FOS

[Min04] Emil Minchev. Forced oscillations of solutions of systems of hyperbolic equations of neutral type. Applied Mathematics and Computation, 155(2):427–438, August 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Merino:2003:FCP

[MJCM03] G. G. Merino, D. D. Jones, D. L. Clements, and D. Miller. Fuzzy compromise programming with precedence order in the criteria. Applied Mathematics and Computation, 134(1): 185–205, January 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ma:2002:GAS

[MK02] N.-Y. Ma and R. P. King.

A general analytical solution for calculating n-fold convolution power of exponentialsum distribution functions. Applied Mathematics and Computation, 133(1):83–91, November 25, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ma:2003:FCG

[MK03a] N.-Y. Ma and R. P. King.

The n-fold convolution of generalized exponentialsum distribution functions. Applied Mathematics and Computation, 142(1):23–33, September 20, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Maleknejad:2003:SSK

[MK03b] K. Maleknejad and M. Tavassoli Kajani. Solving second kind integral equations by Galerkin methods with hybrid Legendre and Block– Pulse functions. Applied Mathematics and Computation, 145(2–3):623–629, December 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Maleknejad:2004:SLIb

[MK04] K. Maleknejad and M. Tavassoli Kajani. Solving linear integro-differential equation system by Galerkin methods with hybrid functions. Applied Mathematics and Computation, 159(3): 603–612, December 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mishchenko:2000:MMD

[ML00] Michael I. Mishchenko and Andrew A. Lacis. Manifestations of morphologydependent resonances in Mie scattering matrices. Applied Mathematics and Computation, 116(1–2):167–179, November 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/91/20/33/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/91/20/
33/article.pdf.

McCartin:2003:AEP

[ML03] Brian J. McCartin and Suzanne M. Labadie. Accurate and efficient pricing of vanilla stock options via the Crandall–Douglas scheme. Applied Mathematics and Computation, 143 (1):39–60, October 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ma:2004:NAS

[ML04a] N.-Y. Ma and Feng Liu. A novel analytical scheme to compute the n-fold convolution of exponential-sum distribution functions. Applied Mathematics and Computation, 158(1):225–235, October 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ma:2004:EST

[ML04b] Ruyun Ma and Hua Luo.

Existence of solutions for a two-point boundary value problem on time scales. Applied Mathematics and Computation, 150(1):139–147, February 27, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Meng:2004:SNI

[ML04c] Fan Wei Meng and Wei Nian Li. On some new integral inequalities and their applications. Applied Mathematics and Computation, 148(2): 381–392, January 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Momoniat:2004:SAE

[ML04d] E. Momoniat and H. Laurie. Symmetry analysis of an equation for constrained optimal transport, modelling the distribution of ants round a nest. Applied Mathematics and Computation, 158(1):111–120, October 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Murty:2001:SKP

[MLA01] K. N. Murty, Vellanki N. Lakshmi, and S. Ajita. Solution of Kronecker product initial value problem involving method of least squares. Applied Mathematics and Computation, 123(1):123–132, September 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/110/27/34/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000667.

Martin:2000:IFM

[MLG00] Pablo Martín, David J. López, and Amelia García. Implementation of Falkner method for problems of the form y′′ = f(x, y). Applied Mathematics and Computation, 109(2–3):183–187, March 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/82/23/
26/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/82/23/26/article.
pdf.

Marcellan:2002:STO

[MM02a] F. Marcellán and L. Moral.

Sobolev-type orthogonal polynomials on the unit circle. Applied Mathematics and Computation, 128(2– 3):329–363, May 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mastroianni:2002:WIP

[MM02b] Giuseppe Mastroianni and Gradimir V. Milovanovi. Weighted integration of periodic functions on the real line. Applied Mathematics and Computation, 128(2– 3):365–378, May 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Maleknejad:2003:TPS

[MM03] K. Maleknejad and Y. Mahmoudi. Taylor polynomial solution of high-order nonlinear Volterra–Fredholm integro-differential equations. Applied Mathematics and Computation, 145(2–3): 641–653, December 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ma:2004:DOD

[MM04a] Zhongtai Ma and Dongtai Ma. Discontinuous oblique derivative problems for second order quasi-linear mixed equations in general domains. Applied Mathematics and Computation, 153 (2):577–585, June 4, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Maleknejad:2004:NSL

[MM04b] K. Maleknejad and Y. Mahmoudi. Numerical solution of linear Fredholm integral equation by using hybrid Taylor and Block– Pulse functions. Applied Mathematics and Computation, 149(3):799–806, February 22, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See erratum [Sab18].

Massoudi:2004:NSF

[MM04c] M. Massoudi and C. E. Maneschy. Numerical solution to the flow of a second grade fluid over a stretching sheet using the method of quasi-linearization. Applied Mathematics and Computation, 149(1):165–173, February 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Maleknejad:2004:SLIa

[MMA04] K. Maleknejad, F. Mirzaee, and S. Abbasbandy. Solving linear integro-differential equations system by using rationalized Haar functions method. Applied Mathematics and Computation, 155 (2):317–328, August 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Monsalve:2003:SAP

[MMER03] M. Monsalve, J. Moreno, R. Escalante, and M. Raydan. Selective alternating projections to find the nearest SDD ^{+} matrix. Applied Mathematics and Computation, 145(2–3):205–220, December 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Menendez:2003:TBD

[MMP03] M. L. Menéndez, D. Morales, and L. Pardo. Tests based on divergences for and against ordered alternatives in cubic contingency tables. Applied Mathematics and Computation, 134(2–3):207–216, January 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Madbouly:2001:AMH

[MMR01] N. M. Madbouly, D. F.

McGhee, and G. F. Roach. Adomian’s method for Hammerstein integral equations arising from chemical reactor theory. Applied Mathematics and Computation, 117(2–3):241–249, January 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/92/23/28/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/92/23/
                                        28/article.pdf;   http:
                                        //www.sciencedirect.com/
                                        science/article/pii/S0096300399001770.

Mastrangelo:2000:NGD

[MMT00] Michèle Mastrangelo, Victor Mastrangelo, and Jean-Marie Teuler. Non-gaussian distributions. Applied Mathematics and Computation, 109(2–3):225–247, March 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/82/23/
                                        30/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/82/23/30/article.
                                        pdf.

Moghadas:2004:AEM

[Mog04] S. M. Moghadas. Analysis of an epidemic model with bistable equilibria using the Poincaré index. Applied Mathematics and Computation, 149(3):689–702, February 22, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mohd:2000:IRA

[Moh00] Ismail Bin Mohd. Identification of region of attraction for global optimization problem using interval symmetric operator. Applied Mathematics and Computation, 110(2–3):121–131, April 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/83/23/
23/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/83/23/23/article.
pdf.

Mohanty:2003:ATS

[Moh03] R. K. Mohanty. An accurate three spatial grid-point discretization of O(k^{2} + h^{4}) for the numerical solution of one-space dimensional unsteady quasi-linear biharmonic problem of second kind. Applied Mathematics and Computation, 140(1):1– 14, July 30, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mohanty:2004:OSM

[Moh04] R. K. Mohanty. An operator splitting method for an unconditionally stable difference scheme for a linear hyperbolic equation with variable coefficients in two space dimensions. Applied Mathematics and Computation, 152(3):799–806, May 13, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Moon:2001:ESC

[Moo01] B. S. Moon. An explicit solution for the cubic spline interpolation for functions of a single variable. Applied Mathematics and Computation, 117(2–3):251–255, January 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/92/23/29/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/92/23/
                                        29/article.pdf;   http:
                                        //www.sciencedirect.com/
                                        science/article/pii/S0096300399001782.

Morhac:2001:IEF

[Mor01] Miroslav Morháč. An iterative error-free algorithm to solve Vandermonde systems. Applied Mathematics and Computation, 117(1):45–54, January 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/92/20/24/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/92/20/
                                        24/article.pdf;   http:
                                        //www.sciencedirect.com/
                                        science/article/pii/S009630039900154X.

Moreno:2004:RQR

[Mor04] P. Moreno. An M/G/1 retrial queue with recurrent customers and general retrial times. Applied Mathematics and Computation, 159(3):651–666, December 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Moudafi:2004:AAP

[Mou04]  A. Moudafi.  An algo-

rithmic approach to proxregular variational inequalities. Applied Mathematics and Computation, 155(3): 845–852, August 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mainardi:2003:WFS

[MP03] Francesco Mainardi and Gianni Pagnini. The Wright functions as solutions of the time-fractional diffusion equation. Applied Mathematics and Computation, 141(1):51–62, August 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Morales:2004:BCR

[MPS04] Domingo Morales, Leandro Pardo, and Laureano Santamaría. Bootstrap confidence regions in multinomial sampling. Applied Mathematics and Computation, 155 (2):295–315, August 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

McGrath:2000:IIC

[MR00a] Joseph F. McGrath and Rajiv Rampalli. Implicit integration with coordinate partitioning. Applied Mathematics and Computation, 111(1):7–31, May 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/85/21/
22/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/85/21/22/article.
pdf.

Micula:2000:INS

[MR00b] Gh. Micula and A. Revnic. An implicit numerical spline method for systems for ODEs. Applied Mathematics and Computation, 111 (1):121–132, May 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/85/21/
28/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/85/21/28/article.
pdf.

Marzban:2003:NSC

[MR03] H. R. Marzban and M. Razzaghi. Numerical solution of the controlled Duffing oscillator by hybrid functions. Applied Mathematics and Computation, 140(2–3): 179–190, August 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Moustafa:2004:UAR

[MR04] H. M. Moustafa and S. G.

Ramadan. Updating and asymptotic relative efficiency of a non-linear discriminant function estimated from a mixture of two Gompertz populations. Applied Mathematics and Computation, 155(1):205– 219, July 26, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Malkowsky:2004:MTBb

[MRŽ04] Eberhard Malkowsky, Vladimir Rakočević, and Snežana Živković. Matrix transformations between the sequence spaces w_{0}^{p}(Λ), v_{0}^{p}(Λ), c_{0}^{p}(Λ) (1 < p < ∞) and certain BK spaces. Applied Mathematics and Computation, 147(2):377–396, January 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Morigi:2001:RCL

[MS01] S. Morigi and F. Sgallari. A regularizing L-curve Lanczos method for underdetermined linear systems. Applied Mathematics and Computation, 121(1):55–73, May 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/9/12/105/21/
                                        25/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/9/12/105/21/25/article.
                                        pdf;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300399002623.

Magden:2004:GCL

[MS04a] A. Mağden and A. A. Salimov. Geodesics for complete lifts of affine connections in tensor bundles. Applied Mathematics and Computation, 151(3):863– 868, April 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Maleknejad:2004:PAS

[MS04b] K. Maleknejad and H. Safdari. Parallel algorithm for solving linear systems arising from PDE and integral equations. Applied Mathematics and Computation, 151(2):443–453, April 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Maleknejad:2004:URK

[MS04c] K. Maleknejad and M. Shahrezaee. Using Runge–Kutta method for numerical solution of the system of Volterra integral equation. Applied Mathematics and Computation, 149(2):399–410, February 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Malkowsky:2004:MTBa

[MS04d] E. Malkowsky and E. Savas.

Matrix transformations between sequence spaces of generalized weighted means. Applied Mathematics and Computation, 147(2):333– 345, January 12, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mohanty:2004:ACS

[MSJ04] R. K. Mohanty, P. L.

Sachdev, and Navnit Jha. An O(h^{4}) accurate cubic spline TAGE method for nonlinear singular two point boundary value problems. Applied Mathematics and Computation, 158(3):853– 868, November 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mukai:2000:IPA

[MSO00] Sonoyo Mukai, Itaru Sano, and Yasuhiko Okada. Inverse problems in the atmosphere-ocean system: estimation of aerosol characteristics and phytoplankton distribution. Applied Mathematics and Computation, 116(1–2):93–101, November 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.

                                        elsevier.nl/gej-ng/10/
                                        9/12/91/20/28/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/91/20/
                                        28/article.pdf.

Mihiotis:2004:MPS

[MT04] A. Mihiotis and I. Tsakiris. A mathematical programming study of advertising allocation problem. Applied Mathematics and Computation, 148(2):373–379, January 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Masuda:2000:RAO

[MTK^{+}00] Kazuhiko Masuda, Tsutomu Takashima, Yoshiyuki Kawata, Akihiro Yamazaki, and Masayuki Sasaki. Retrieval of aerosol optical properties over the ocean using multispectral polarization measurements from space. Applied Mathematics and Computation, 116(1–2):103–114, November 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/91/20/29/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/91/20/
                                        29/article.pdf.

Muradoglu:2004:NMC

[(Mu04] Zahir Seyidmamedov (Muradoğlu). A numerical method for coupled filtration equations in porous media in presence of the inverse compaction factor. Applied Mathematics and Computation, 155(2):439– 449, August 6, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Munk:2003:MCA

[Mun03] Claus Munk. The Markov chain approximation approach for numerical solution of stochastic control problems: experiences from Merton’s problem. Applied Mathematics and Computation, 136(1):47–77, March 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Murty:2001:ETM

[Mur01] Yedidi N. Murty. Effect of throughflow on magnetoconvection in micropolar fluids. Applied Mathematics and Computation, 123(2):249–261, September 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/110/29/33/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000746.

Murty:2003:ANU

[Mur03] Y. Narasimha Murty. Analysis of non-uniform temperature profiles on Bénard convection in micropolar fluids. Applied Mathematics and Computation, 134(2–3): 473–486, January 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Muses:2000:CSN

[Mus00a] C. Musès. Corrigendum to “Some new considerations on the Bernoulli numbers, the factorial function, and Riemann’s zeta function” [Applied Mathematics and Computation 113 (2000) 1–21]. Applied Mathematics and Computation, 115(2–3):229, October 27, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/90/23/31/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/90/23/
31/article.pdf.  See

[Mus00d, Mus00b].

Muses:2000:ESN

[Mus00b]  C. Musès.  Erratum to

“Some new considerations on the Bernoulli numbers, the factorial function, and Riemann’s zeta function” [Applied Mathematics and Computation 113 (2000) 1–21]. Applied Mathematics and Computation, 113(2–3):325–326, July 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/87/23/32/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/87/23/
32/article.pdf.  See

[Mus00d, Mus00a].

Muses:2000:NCD

[Mus00c] C. Musès. New considerations of denumerability, nondenumerability, and targetseeking algorithms. Applied Mathematics and Computation, 107(2–3):225–233, January 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/72/22/
32/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/72/22/32/article.
pdf.

Muses:2000:SNC

[Mus00d] C. Musès. Some new considerations on the Bernoulli numbers, the factorial function, and Riemann’s zeta function. Applied Mathematics and Computation, 113(1):1–21, July 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.nl/gej-ng/
10/9/12/87/21/21/abstract.
html;   http://www.elsevier.

                                        nl/gej-ng/10/9/12/87/21/
                                        21/article.pdf. See erra-

tum [Mus00b] and corrigen-

dum [Mus00a].

Muses:2000:VKC

[Mus00e] C. Musès. Validating Kepler’s conjecture: A new approach. Applied Mathematics and Computation, 110 (1):99–104, April 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/83/21/
                                        24/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/83/21/24/article.
                                        pdf.

Muttlak:2003:IUQ

[Mut03] Hassen A. Muttlak. Investigating the use of quartile ranked set samples for estimating the population mean. Applied Mathematics and Computation, 146(2–3): 437–443, December 31, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Molina:2004:PBP

[MV04] R. Molina and A. Vigueras. The planar n-body problem: regular polygon solutions. Applied Mathematics and Computation, 156(2): 321–327, September 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mosally:2002:EHB

[MWAF02] F. Mosally, A. S. Wood, and A. Al-Fhaid. An exponential heat balance integral method. Applied Mathematics and Computation, 130 (1):87–100, July 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Minchev:2003:OSNa

[MY03] Emil Minchev and Norio Yoshida. Oscillation of solutions of nonlinear parabolic equations via comparison method. Applied Mathematics and Computation, 134 (2–3):561–566, January 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ma:2004:ESG

[MZ04a] Qiaozhen Ma and Chengkui Zhong. Existence of strong global attractors for hyperbolic equation with linear memory. Applied Mathematics and Computation, 157(3):745–758, October 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Minggen:2004:HSN

[MZ04b] Cui Minggen and Chen Zhong. How to solve nonlinear operator equation A(v^{2}) + Cv = f. Applied Mathematics and Computation, 153(2):403–416, June 4, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Nakamori:2002:REI

[Nak02] Seiichi Nakamori. Recursive estimation of impulse response function using covariance information in linear continuous stochastic systems. Applied Mathematics and Computation, 131(2–3):339–347, September 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Nakamori:2003:NDF

[Nak03] Seiichi Nakamori. New design of fixed-interval smoother using covariance information in linear stochastic continuous-time systems. Applied Mathematics and Computation, 144(2–3):557– 567, December 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Nakamori:2004:NDF

[Nak04a] Seiichi Nakamori. New design of fixed-lag smoother using covariance information in linear discrete-time stochastic systems. Applied Mathematics and Computation, 156(2):415–426, September 6, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Nakamori:2004:NDR

[Nak04b]  Seiichi Nakamori.  New

design of recursive Wiener fixed-lag smoother in linear discrete-time stochastic systems. Applied Mathematics and Computation, 159(2): 303–315, December 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Narayaninsamy:2001:NFI

[Nar01] Tony Narayaninsamy. On the node fractional iterates. Applied Mathematics and Computation, 117(1): 87–102, January 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.nl/gej-ng/
10/9/12/92/20/27/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/92/20/
27/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S009630039900168X.

Narayaninsamy:2002:NAT

[Nar02] Tony Narayaninsamy. On the numerical approximation of trajectories in the phase plane. Applied Mathematics and Computation, 126(1):119–132, February 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/122/27/33/abstract.
html;   http://www.sciencedirect.

                                        com/science/article/pii/
                                        S009630030000148X.

Noor:2004:QSS

[NAS04] Muhammad Aslam Noor and Eisa E. Al-Said. Quartic splines solutions of thirdorder obstacle problems. Applied Mathematics and Computation, 153(2):307– 316, June 4, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Natsuyama:2000:SUG

[Nat00] Harriet H. Natsuyama. Sueo Ueno: gentleman scholar. Applied Mathematics and Computation, 116(1–2):5– 10, November 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/91/20/22/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/91/20/
                                        22/article.pdf.

Nakamori:2003:SOP

[NCÁHCLP03a] S. Nakamori, R. Caballero-Águila, A. Hermoso-Carazo, and J. Linares-Pérez. Secondorder polynomial estimators from uncertain observations using covariance information. Applied Mathematics and Computation, 143(2–3):319–338, November 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Nakamori:2003:NDE

[NCÁHCLP03b] Seiichi Nakamori, Raquel Caballero-Águila, Aurora Hermoso-Carazo, and Josefa Linares-Pérez. New design of estimators using covariance information with uncertain observations in linear discrete-time systems. Applied Mathematics and Computation, 135(2–3):429– 441, March 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Nakamori:2004:QEU

[NCÁHCLP04] S. Nakamori, R. Caballero-Águila, A. Hermoso-Carazo, and J. Linares-Pérez. Quadratic estimation from uncertain observations with white plus coloured noises using covariance information. Applied Mathematics and Computation, 155(1):65–79, July 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Natsuyama:2000:GEI

[NCK^{+}00] Harriet Natsuyama, John Casti, Kiyoshi Kawabata, Sonoyo Mukai, and Masanori Sugisaka. Guest editors’ introduction. Applied Mathematics and Computation, 116(1–2):1–2, November 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/9/12/91/20/20/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/9/12/91/20/20/article.
                                        pdf.

Nakajima:2002:DCM

[NG02a] Y. Nakajima and Y. P. Gunji. The dynamically changing model of exchange as interaction between cone-relation and equivalent-relation. Applied Mathematics and Computation, 126(2–3):299–318, March 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        9/12/122/30/38/abstract.
                                        html;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300300001594.

Nemani:2002:PAS

[NG02b] S. S. Nemani and L. E. Garey. Parallel algorithms for solving tridiagonal and near-circulant systems. Applied Mathematics and Computation, 130(2–3):285–294, August 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Nagar:2004:PPT

[NG04] Daya K. Nagar and Arjun K. Gupta. Percentage points for testing homogeneity of several univariate Gaussian populations. Applied Mathematics and Computation, 156(2): 551–561, September 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Nakamori:2004:CTF

[NHCJLLP04] S. Nakamori, A. Hermoso-Carazo, J. Jiménez-López, and J. Linares-Pérez. Chandrasekhartype filter for a wide-sense stationary signal from uncertain observations using covariance information. Applied Mathematics and Computation, 151(2):315–325, April 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Nakamori:2004:FIS

[NHCLPSR04] S. Nakamori, A. Hermoso-Carazo, J. Linares-Pérez, and M. I. Sánchez-Rodríguez. Fixed-interval smoothing problem from uncertain observations with correlated signal and noise. Applied Mathematics and Computation, 154(1):239–255, June 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Nketsa:2001:TPN

[NK01] Alexandre Nketsa and Nabil Ben Khalifa. Timed Petri nets and prediction to improve the Chandy-Misra conservative-distributed simulation. Applied Mathematics and Computation, 120(1– 3):235–254, May 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/104/21/
39/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/104/21/39/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002441.

Nakhi:2003:SBV

[NK03] Y. Ben Nakhi and S. L. Kalla. Some boundary value problems of temperature fields in oil strata. Applied Mathematics and Computation, 146(1):105– 119, December 30, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Najafi:2004:NRM

[NK04] H. Saberi Najafi and E. Khaleghi. A new restarting method in the Arnoldi algorithm for computing the eigenvalues of a nonsymmetric matrix. Applied Mathematics and Computation, 156 (1):59–71, August 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Natale:2003:EHC

[NM03] M. F. Natale and E. A. Santillan Marcus. The effect of heat convection on drying of porous semi-infinite space with a heat flux condition on the fixed face x = 0. Applied Mathematics and Computation, 137 (1):109–129, May 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Nasrabadi:2004:MPA

[NN04a] Mohammad Mehdi Nasrabadi and Ebrahim Nasrabadi. A mathematical-programming approach to fuzzy linear regression analysis. Applied Mathematics and Computation, 155(3):873–881, August 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Noor:2004:SAP

[NN04b] Muhammad Aslam Noor and Khalida Inayat Noor. Self-adaptive projection algorithms for general variational inequalities. Applied Mathematics and Computation, 151(3):659–670, April 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Noor:2004:PCM

[NNAK04] Muhammad Aslam Noor, Khalida Inayat Noor, and Kamel Al-Khaled. Predictor– corrector methods for general mixed quasi variational inequalities. Applied Mathematics and Computation, 157(3):643–652, October 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Njoku:2003:SPP

[NO03] Franic Ikechukwu Njoku and Pierpaolo Omari. Stability properties of periodic solutions of a Duffing equation in the presence of lower and upper solutions. Applied Mathematics and Computation, 135(2–3):471–490, March 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Noor:2003:IDS

[Noo03a] Muhammad Aslam Noor. Implicit dynamical systems and quasi variational inequalities. Applied Mathematics and Computation, 134(1):69–81, January 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Noor:2003:MQV

[Noo03b] Muhammad Aslam Noor. Mixed quasi variational inequalities. Applied Mathematics and Computation, 146(2–3):553–578, December 31, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Noor:2003:PGM

[Noo03c] Muhammad Aslam Noor.

Pseudomonotone general mixed variational inequalities. Applied Mathematics and Computation, 141(2–3): 529–540, September 5, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Noor:2004:SCV

[Noo04a] Khalida Inayat Noor. Some classes of p-valent analytic functions defined by certain integral operator. Applied Mathematics and Computation, 157(3):835–840, October 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Noor:2004:GMQ

[Noo04b] Muhammad Aslam Noor.

Generalized mixed quasivariational-like inequalities. Applied Mathematics and Computation, 156(1):145– 158, August 25, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Noor:2004:CNE

[Noo04c] Muhammad Aslam Noor.

On a class of nonconvex equilibrium problems. Applied Mathematics and Computation, 157(3):653–666, October 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Noor:2004:SDG

[Noo04d] Muhammad Aslam Noor. Some developments in general variational inequalities. Applied Mathematics and Computation, 152(1):199– 277, April 26, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Natesan:2002:ANM

[NR02a] S. Natesan and N. Ramanujam. An asymptoticnumerical method for singularly perturbed Robin problems — I. Applied Mathematics and Computation, 126(1):97–107, February 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/122/27/31/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001405.

Natesan:2002:SMS

[NR02b] S. Natesan and N. Ramanujam. “Shooting method” for the solution of singularly perturbed two-point boundary-value problems having less severe boundary layer. Applied Mathematics and Computation, 133(2–3): 623–641, December 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ngoc:2003:SRP

[NS03] Pham Huu Anh Ngoc and Nguyen Khoa Son. Stability radii of positive linear difference equations under affine parameter perturbations. Applied Mathematics and Computation, 134(2–3): 577–594, January 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Nyarko:2004:SPI

[NS04] Emmanuel Karlo Nyarko and Rudolf Scitovski. Solving the parameter identification problem of mathematical models using genetic algorithms. Applied Mathematics and Computation, 153(3):651–658, June 14, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Natsuyama:2000:III

[NU00a] Harriet Natsuyama and Sueo Ueno. Invariant imbedding and internal intensities in stratified slabs. Applied Mathematics and Computation, 107(2–3):113–120, January 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/72/22/
25/abstract.html;   http:

//www.elsevier.nl/gej-
ng/29/17/20/72/22/25/article.
pdf.

Natsuyama:2000:PMP

[NU00b] Harriet H. Natsuyama and Sueo Ueno. The probabilistic method for problems of radiative transfer: Markov properties. Applied Mathematics and Computation, 108(1):55–65, February 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/79/21/
26/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/79/21/26/article.
pdf.

Nketsa:2001:RMP

[NV01] A. Nketsa and R. Valette. Rapid and modular prototypingbased Petri nets and distributed simulation for manufacturing systems. Applied Mathematics and Computation, 120(1–3):265–278, May 10, 2001. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/104/21/
41/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/104/21/41/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S009630039900243X.

Noor:2003:SNP

[NWX03] Muhammad Aslam Noor, Yiju Wang, and Naihua Xiu. Some new projection methods for variational inequalities. Applied Mathematics and Computation, 137(2– 3):423–435, May 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Nystrom:2001:TCF

[Nys01] J. F. Nystrom. Tensional computation: further musings on the computational cosmography. Applied Mathematics and Computation, 120(1–3):211–225, May 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/104/21/
37/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/104/21/37/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002489.

Nakoulima:2004:ANS

[NZP^{+}04] O. Nakoulima, N. Zahibo, E. Pelinovsky, T. Talipova, A. Slunyaev, and A. Kurkin. Analytical and numerical studies of the variablecoefficient Gardner equation. Applied Mathematics and Computation, 152 (2):449–471, May 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ozdemir:2004:FPT

[ÖA04] Murat Özdemir and Sezgin Akbulut. A fixed point theorem for multivalued maps in hyperconvex spaces. Applied Mathematics and Computation, 157(3): 637–642, October 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Olunloyo:2004:NMM

[OAB04] V. O. S. Olunloyo, A. M. Ajofoyinbo, and A. B. Badiru. Neurofuzzy mathematical model for monitoring flow parameters of natural gas. Applied Mathematics and Computation, 149(3): 747–770, February 22, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ozkan:2003:SFSa

[ÖAD03a] Engin Özkan, Hüseyin Aydın, and Ramazan Dikici. 3-step Fibonacci series modulo m. Applied Mathematics and Computation, 143(1): 165–172, October 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ozkan:2003:AFS

[ÖAD03b] Engin Özkan, Hüseyin Aydın, and Ramazan Dikici.

Applications of Fibonacci sequences in a finite nilpotent group. Applied Mathematics and Computation, 141(2–3): 565–578, September 5, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ozi:2003:FEA

[ÖAÖ03] T. Özi, E. N. Aksan, and A. Özde. A finite element approach for solution of Burgers’ equation. Applied Mathematics and Computation, 139(2–3):417–428, July 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Osman:2004:MLN

[OASAE04] M. S. Osman, M. A. Abo-Sinna, A. H. Amer, and O. E. Emam. A multi-level non-linear multi-objective decision-making under fuzziness. Applied Mathematics and Computation, 153 (1):239–252, May 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Osman:2004:SOP

[OASM04] M. S. Osman, M. A. Abo-Sinna, and A. A. Mousa. A solution to the optimal power flow using genetic algorithm. Applied Mathematics and Computation, 155 (2):391–405, August 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Odanaka:2001:ESD

[Oda01]  Toshio Odanaka.  Envi-

ronment system and dynamic management decision. Applied Mathematics and Computation, 120(1– 3):255–263, May 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/9/12/104/21/
                                        40/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/9/12/104/21/40/article.
                                        pdf;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300399002453.

Odekunle:2002:SBV

[Ode02] M. R. Odekunle. Solution of boundary value problems by Lanczos–Legendre reduction method. Applied Mathematics and Computation, 131(2–3):321–327, September 25, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ohtsubo:2004:OTP

[Oht04]  Yoshio Ohtsubo.  Opti-

mal threshold probability in undiscounted Markov decision processes with a target set. Applied Mathematics and Computation, 149(2): 519–532, February 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Odekunle:2003:PEE

[OIO03] M. R. Odekunle, M. A.

Ibiejugba, and P. Onumanyi. A posteriori error estimator for Lanczos–Chebyshev reduction method. Applied Mathematics and Computation, 137(2–3):245–252, May 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Orhan:2003:FCS

[OK03a] Halit Orhan and Muhammet Kamali. Fractional calculus and some properties of certain starlike functions with negative coefficients. Applied Mathematics and Computation, 136(2– 3):269–279, March 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Orhan:2003:FSP

[OK03b] Halit Orhan and Muhammet Kamali. On the Fekete–Szegő problem. Applied Mathematics and Computation, 144(1):181–186, November 20, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Orhan:2003:SCC

[OK03c] Halit Orhan and Muhammet Kamali. Starlike, convex and close-to convex functions of complex order. Applied Mathematics and Computation, 135(2–3):251–262, March 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ozdemir:2003:TNT

[ÖK03d] M. Emin Özdemir and Uğur S. Kırmacı. Two new theorem on mappings uniformly continuous and convex with applications to quadrature rules and means. Applied Mathematics and Computation, 143(2–3):269– 274, November 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Orhan:2004:GSV

[OK04] Halit Orhan and Hükmi Kiziltunç. A generalization on subfamily of p-valent functions with negative coefficients. Applied Mathematics and Computation, 155 (2):521–530, August 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ozdemir:2004:AMF

[ÖKOD04] M. Emin Özdemir, Uğur S. Kırmacı, Rahim Ocak, and Ali Dönmez. On automorphic and modular forms in the space of the homogeneous polynomials with degree 2ℓ and applications to the special matrix. Applied Mathematics and Computation, 152(3):897–904, May 13, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ouyang:2004:CEP

[OLT04] Zigen Ouyang, Yongkun Li, and Qingan Tang. Classifications and existence of positive solutions of higherorder nonlinear neutral differential equations. Applied Mathematics and Computation, 148(1):105–120, January 20, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Oluyede:2003:IBK

[Olu03a] Broderick O. Oluyede. Inequalities and bounds for kernel length-biased density estimation. Applied Mathematics and Computation, 135(2–3):541–551, March 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Oluyede:2003:SID

[Olu03b] Broderick O. Oluyede. On stochastic inequalities and dependence orderings. Applied Mathematics and Computation, 146(2–3):601–610, December 31, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Oluyede:2004:LDS

[Olu04] Broderick O. Oluyede. On local dependence and stochastic inequalities with applications to contingency tables. Applied Mathematics and Computation, 151 (3):801–813, April 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ogren:2002:VSE

[ÖM02] Petter Ögren and Clyde F. Martin. Vaccination strategies for epidemics in highly mobile populations. Applied Mathematics and Computation, 127(2–3):261–276, April 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/123/27/36/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300301000042.

Onah:2002:ABG

[Ona02] Stephen E. Onah. Asymptotic behaviour of the Galerkin and the finite element collocation methods for a parabolic equation. Applied Mathematics and Computation, 127(2–3):207–213, April 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/123/27/31/abstract.

html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001661.

Ozdemir:2004:ILC

[ÖÖ04] Halim Özdemır and Ahmet Yaşar Özban. On idempotency of linear combinations of idempotent matrices. Applied Mathematics and Computation, 159(2): 439–448, December 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Odekunle:2004:CIR

[OOAA04] M. R. Odekunle, N. D. Oye, S. O. Adee, and R. A. Ademiluyi. A class of inverse Runge–Kutta schemes for the numerical integration of singular problems. Applied Mathematics and Computation, 158(1):149–158, October 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Oppenheimer:2000:CDP

[Opp00] Seth F. Oppenheimer. A convection-diffusion problem in a network. Applied Mathematics and Computation, 112(2–3):223–240, June 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/86/23/
26/abstract.html;   http:

                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/86/23/26/article.
                                        pdf.

Oquendo:2003:UDP

[Oqu03] Higidio Portillo Oquendo. Uniform decay for a plate equation partially dissipative. Applied Mathematics and Computation, 146(1): 121–134, December 30, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Orhan:2003:NCA

[Orh03a] Halit Orhan. A new class of analytic functions with negative coefficients. Applied Mathematics and Computation, 138(2–3):531–543, June 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Orhan:2003:SAF

[Orh03b] Halit Orhan. On a subclass of analytic functions with negative coefficients. Applied Mathematics and Computation, 142(2–3):243–253, October 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ohm:2003:EAM

[OSL03] M. R. Ohm, J. Y. Shin, and H. Y. Lee. Error analysis of a mixed finite element approximation of a linear Stefan problem. Applied Mathematics and Computation, 143(2–3):485–500, November 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ootao:2004:CTT

[OT04] Yoshihiro Ootao and Yoshinobu Tanigawa. Control of transient thermoelastic displacement of an angle-ply laminated cylindrical panel bonded to a piezoelectric layer. Applied Mathematics and Computation, 148(1): 263–286, January 20, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ozturk:2004:CSH

[ÖYY04] Metın Öztürk, Sıbel Yalçin, and Mümın Yamankaradenız. Convex subclass of harmonic starlike functions. Applied Mathematics and Computation, 154(2):449–459, July 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ozban:2004:ICC

[Özb04] Ahmet Yaşar Özban. Improved convergence criteria for Jacobi and Gauss–Seidel iterations. Applied Mathematics and Computation, 152(3):693–700, May 13, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ozdemir:2003:TMB

[Özd03a] M. Emin Özdemir. A theorem on mappings with bounded derivatives with applications to quadrature rules and means. Applied Mathematics and Computation, 138(2–3):425–434, June 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ozdemir:2003:TFS

[Özd03b] M. Emin Özdemir. A theorem on the first and second local modulus of continuity and applications to quadrature rules and means. Applied Mathematics and Computation, 141(2–3):579–588, September 5, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ozdemir:2003:CCH

[Özd03c] Murat Özdemir. A criterion for constantness of harmonic functions on exterior domains. Applied Mathematics and Computation, 143(2–3):251–257, November 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ozkan:2003:SFSb

[Özk03] Engin Özkan. 3-step Fibonacci sequences in nilpotent groups. Applied Mathematics and Computation, 144(2–3):517–527, December 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Pantokratoras:2008:CTD

[Pan08] Asterios Pantokratoras. Comment on “Thermal-diffusion and diffusion-thermo effects on mixed free-forced convection and mass transfer boundary layer flow for non-Newtonian fluid with temperature dependent viscosity”, by N. T. Eldabe, A. G. El-Saka and Ashraf Fouad [Applied Mathematics and Computation 152 (2004) 867–883]. Applied Mathematics and Computation, 196(2):959–961, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [EESF04].

Papamarkos:2000:ARR

[Pap00] Nikos Papamarkos. On the approximation of real rational functions via mixedinteger linear programming. Applied Mathematics and Computation, 112(1):113– 124, June 1, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/86/21/
30/abstract.html;   http:
//www.elsevier.nl/gej-

                                        ng/29/17/20/86/21/30/article.
                                        pdf.

Pardo:2002:GFT

[Par02] M. C. Pardo. Goodnessof-fit tests based on Rao’s divergence under sparseness assumptions. Applied Mathematics and Computation, 130(2–3):265–283, August 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Park:2003:RGC

[Par03] Ju H. Park. Robust guaranteed cost control for uncertain linear differential systems of neutral type. Applied Mathematics and Computation, 140(2–3):523–535, August 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Park:2004:DOB

[Par04a] Ju H. Park. On the design of observer-based controller of linear neutral delay-differential systems. Applied Mathematics and Computation, 150(1):195– 202, February 27, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Park:2004:RNF

[Par04b] Ju H. Park. Robust nonfragile control for uncertain discrete-delay largescale systems with a class of controller gain variations. Applied Mathematics and Computation, 149(1):147– 164, February 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Park:2002:CWE

[PB02] Jong Yeoul Park and Jeong Ja Bae. On coupled wave equation of Kirchhoff type with nonlinear boundary damping and memory term. Applied Mathematics and Computation, 129(1):87– 105, June 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Pourabbas:2001:MEI

[PdR01] Elaheh Pourabbas, Alberto d’Onofrio, and Maurizio Rafanelli. A method to estimate the incidence of communicable diseases under seasonal fluctuations with application to cholera. Applied Mathematics and Computation, 118(2–3):161–174, March 9, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/96/25/23/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/96/25/
23/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S009630039900212X.

Peinado:2004:CLK

[Pei04] A. Peinado. Cryptanalysis of LHL-key authentication scheme. Applied Mathematics and Computation, 152 (3):721–724, May 13, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Peng:2003:OCS

[Pen03]  Mingshu Peng.  Oscilla-

tion criteria for secondorder impulsive delay difference equations. Applied Mathematics and Computation, 146(1):227–235, December 30, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Peng:2004:PSF

[Pen04] Shiguo Peng. Positive solutions for first order periodic boundary value problem. Applied Mathematics and Computation, 158(2): 345–351, November 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Petrila:2002:MMF

[Pet02] Titus Petrila. Mathematical model for the free surface flow under a sluice gate. Applied Mathematics and Computation, 125(1):49–58, January 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/120/27/30/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001090. See

contents [Pop04].

Petropoulou:2003:PDE

[Pet03] Eugenia N. Petropoulou.

Partial difference equations arising in numerical schemes and game theory. Applied Mathematics and Computation, 141(1):185–196, August 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Fang:2003:MOR

[pFjH03] Ya ping Fang and Nan jing Huang. H-Monotone operator and resolvent operator technique for variational inclusions. Applied Mathematics and Computation, 145(2–3):795–803, December 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Peng:2000:NCO

[PGX00] Mingshu Peng, Weigao Ge, and Qianli Xu. New criteria for the oscillation and existence of monotone solutions of second-order nonlinear difference equations. Applied Mathematics and Computation, 114(1):103– 114, August 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/89/20/29/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/89/20/
29/article.pdf.

Perez-Garcia:2003:NMS

[PGyL03] Víctor M. Pérez-García and Xiao yan Liu. Numerical methods for the simulation of trapped nonlinear Schrödinger systems. Applied Mathematics and Computation, 144(2–3):215–235, December 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Park:2004:DNG

[PJ04] Ju H. Park and Ho Y. Jung. On the design of nonfragile guaranteed cost controller for a class of uncertain dynamic systems with state delays. Applied Mathematics and Computation, 150(1): 245–257, February 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Park:2004:DDO

[PJPL04] Ju H. Park, Ho Y. Jung, Jung I. Park, and Suk G. Lee. Decentralized dynamic output feedback controller design for guaranteed cost stabilization of large-scale discrete-delay systems. Applied Mathematics and Computation, 156(2):307–320, September 6, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Park:2004:RFD

[PKLW04] Ju H. Park, O. Kwon, S. Lee, and S. Won. On robust H_{∞} filter design for uncertain neural systems: LMI optimization approach. Applied Mathematics and Computation, 159(3):625–639, December 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Park:2004:LAR

[PKW04] Ju H. Park, O. Kwon, and S. Won. LMI approach to robust H_{∞} filtering for neutral delay differential systems. Applied Mathematics and Computation, 150(1): 235–244, February 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Platkowski:2003:SSO

[Pla03] Tadeusz Platkowski. Stationary self-organized states in many-particle systems with ternary interactions. Applied Mathematics and Computation, 146(2–3):701– 712, December 31, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Pommeret:2001:OPO

[Pom01] Denys Pommeret. Orthogonal and pseudo-orthogonal multi-dimensional Appell polynomials. Applied Mathematics and Computation, 117(2–3):285–299, January 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/9/12/92/23/32/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/9/12/92/23/32/article.
                                        pdf;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300399001812.

Pop:2004:CPM

[Pop04] Ioan Pop. Comments on the paper “Mathematical model for the free surface flow under a sluice gate” by Titus Petrila, published in Applied Mathematics and Computation 125 (2002) 49–

58. Applied Mathematics and Computation, 148(3): 807–808, January 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [Pet02].

Povitsky:2002:PAS

[Pov02]  Alex Povitsky.  Parallel

ADI solver based on processor scheduling. Applied Mathematics and Computation, 133(1):43–81, November 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Philos:2001:PFO

[PP01] Ch. G. Philos and I. K.

Purnaras. Periodic first order linear neutral delay differential equations. Applied Mathematics and Computation, 117(2–3):203–222, January 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/92/23/26/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/92/23/
26/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399001745.

Papamanthou:2004:CEE

[PPS04] Charalampos Papamanthou, Konstantinos Paparrizos, and Nikolaos Samaras. Computational experience with exterior point algorithms for the transportation problem. Applied Mathematics and Computation, 158(2): 459–475, November 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Parand:2004:RCT

[PR04a] K. Parand and M. Razzaghi. Rational Chebyshev tau method for solving Volterra’s population model. Applied Mathematics and Computation, 149(3):

893–900, February 22, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Petkovic:2004:GCF

[PR04b] M. S. Petković and L. Rančić. On the guaranteed convergence of the fourth order simultaneous method for polynomial zeros. Applied Mathematics and Computation, 155(2):531–543, August 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Prabhu:2003:GGD

[Pra03] N. Prabhu. Gauge groups and data classification. Applied Mathematics and Computation, 138(2–3):267–289, June 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Pandey:2000:BTH

[PS00] Alok Pandey and Ashok K. Singh. A Bayes test of homogeneity of several means for one parameter exponential populations. Applied Mathematics and Computation, 108(1):23–32, February 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/79/21/
                                        23/abstract.html;   http:
                                        //www.elsevier.nl/gej-

ng/29/17/20/79/21/23/article.
pdf.

Peng:2004:EEL

[PS04] Yahong Peng and Yongli Song. Existence of entire large positive solutions of a semilinear elliptic system. Applied Mathematics and Computation, 155(3): 687–698, August 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Pittaluga:2000:SGF

[PSS00] Giovanna Pittaluga, Laura Sacripante, and H. M. Srivastava. Some generating functions of the Laguerre and modified Laguerre polynomials. Applied Mathematics and Computation, 113(2–3):141–160, July 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/87/23/22/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/87/23/
22/article.pdf.

Piyawong:2003:UCF

[PTG03] W. Piyawong, E. H. Twizell, and A. B. Gumel. An unconditionally convergent finitedifference scheme for the SIR model. Applied Mathematics and Computation, 146(2–3): 611–625, December 31, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Papageorgiou:2002:MCN

[PY02] Nikolaos S. Papageorgiou and Nikolaos Yannakakis. Minimax control of nonlinear evolution equations. Applied Mathematics and Computation, 131(2–3):271–297, September 25, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ping:2003:SKE

[PZJ03] He Ping, Chen Zheng, and Fu Jun. Solution of KdV equation by computer algebra. Applied Mathematics and Computation, 136(2–3): 511–515, March 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Piccolomini:2002:RMD

[PZZF02] E. Loli Piccolomini, F. Zama, G. Zanghirati, and A. Formiconi. Regularization methods in dynamic MRI. Applied Mathematics and Computation, 132(2–3):325–339, November 10, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Qin:2002:NSC

[QBK02] Zhihao Qin, Pedro Berliner, and Arnon Karnieli. Numerical solution of a complete surface energy balance model for simulation of heat fluxes and surface temperature under bare soil environment. Applied Mathematics and Computation, 130 (1):171–200, July 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jin:2003:CBP

[qJkSlS03] Xiao qing Jin, Vai kuong Sin, and Li li Song. Circulant-block preconditioners for solving ordinary differential equations. Applied Mathematics and Computation, 140(2–3):409–418, August 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lu:2003:NMB

[qLzWcC03] Hong qiang Lu, Yi zhao Wu, and Song can Chen. A new method based on SOM network to generate coarse meshes for overlapping unstructured multigrid algorithm. Applied Mathematics and Computation, 140(2– 3):353–360, August 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Qi-Min:2004:EUE

[QMWAZK04] Zhang Qi-Min, Liu Wen-An, and Nie Zan-Kan. Existence, uniqueness and exponential stability for stochastic age-dependent population. Applied Mathematics and Computation, 154 (1):183–201, June 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Qatanani:2004:PCG

[QS04] Naji Qatanani and Monika Schulz. Preconditioned conjugate gradient method for three-dimensional nonconvex enclosure geometries with diffuse and grey surfaces. Applied Mathematics and Computation, 159(3): 797–807, December 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Quattromini:2003:HFF

[Qua03] M. Quattromini. Hermite filtering and form factors. Applied Mathematics and Computation, 141(1):131– 142, August 20, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Quintanilla:2003:CSS

[Qui03] Ramón Quintanilla. Convergence and structural stability in thermoelasticity. Applied Mathematics and Computation, 135(2–3):287–300, March 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Quintanilla:2004:ESP

[Qui04] R. Quintanilla. Exponential stability in porous media problem saturated by multiple immiscible fluids. Applied Mathematics and Computation, 150(3):661– 668, March 17, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Qu:2003:MMO

[QW03] Chaochun Qu and Ping Wang. Mathematical model and optimization in global production problems. Applied Mathematics and Computation, 145(1):85–95, December 20, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Qu:2002:MMO

[QWY02] Chaochun Qu, Ping Wang, and Huakang Yang. Mathematical model and optimization in production investment. Applied Mathematics and Computation, 130(2–3): 389–398, August 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Qu:2000:AME

[QY00a] Ruibin Qu and Jieping Ye. Approximation of minimum energy curves. Applied Mathematics and Computation, 108(2–3):153–166, February 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/79/22/
29/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/79/22/29/article.
pdf.

Qu:2000:EAM

[QY00b] Ruibin Qu and Jieping Ye. Efficient approximation of minimum energy curves with interpolatory constraints. Applied Mathematics and Computation, 109(2–3):151– 166, March 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/82/23/
24/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/82/23/24/article.
pdf.

Radwan:2003:IGJ

[RA03] Ahmed E. Radwan and Ramadan M. Ali. Instability of a gas jet of zero inertia dispersed in a liquid using Lagrangian’s nonlinear differential equation. Applied Mathematics and Computation, 138(2–3):171– 180, June 20, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Radwan:2003:MSS

[Rad03] Ahmed E. Radwan. Magnetoselfgravitational stability of triple superposed fluids layers of different densities. Applied Mathematics and Computation, 141(2–3):401–413, September 5, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Radwan:2004:VSS

[Rad04] Ahmed E. Radwan. Variable streams self-gravitating instability of radiating rotating gas cloud. Applied Mathematics and Computation, 148(2):331–339, January 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rashid:2001:ODE

[RAH01] Kamal M. Rashid, Marwan M. Awartani, and M. H. Hamdan. Optimal distribution of energy resources and policy implementation in an n-sector society. Applied Mathematics and Computation, 118(2–3):343–348, March 9, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/96/25/33/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/96/25/
33/article.pdf;   http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0096300399001423.

Rakha:2004:MPG

[Rak04] Medhat A. Rakha. On the Moore–Penrose generalized inverse matrix. Applied Mathematics and Computation, 158(1):185–200, October 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ramos:2000:HMTa

[Ram00a] J. I. Ramos. Heat and mass transfer in annular liquid jets: I. Formulation. Applied Mathematics and Computation, 110(2–3):133– 164, April 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/83/23/
                                        24/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/83/23/24/article.
                                        pdf.

Ramos:2000:HMTb

[Ram00b] J. I. Ramos. Heat and mass transfer in annular liquid jets: II. g-jitter. Applied Mathematics and Computation, 110(2–3):165– 183, April 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/83/23/

25/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/83/23/25/article.
pdf.

Ramos:2000:HMTc

[Ram00c] J. I. Ramos. Heat and mass transfer in annular liquid jets: III. Combustion within the volume enclosed by the jet. Applied Mathematics and Computation, 110(2– 3):185–204, April 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/83/23/
26/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/83/23/26/article.
pdf.

Ramos:2001:PFH

[Ram01a]  J. I. Ramos.  Periodi-

cally forced Hopf bifurcation in annular liquid jets with mass transfer. Applied Mathematics and Computation, 123(3):301–342, October 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/110/32/29/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000722.

Ramos:2001:SGE

[Ram01b] J. I. Ramos. The sine-Gordon equation in the finite line. Applied Mathematics and Computation, 124(1): 45–93, November 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.elsevier.com/gej-
                                        ng/10/9/12/113/27/30/abstract.
                                        html;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300300000801.

Ramos:2002:LIM

[Ram02]  J. I. Ramos.  Linearly

implicit methods for the nonlinear Schrödinger equation in nonhomogeneous media. Applied Mathematics and Computation, 133(1): 1–28, November 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ramadan:2003:AMI

[Ram03a] Mohamed A. Ramadan. An algorithm for the multiinput complex eigenvalue assignment problem. Applied Mathematics and Computation, 140(2–3):455–473, August 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ramos:2003:CDP

[Ram03b]  J. I. Ramos.  Complex

dynamics of planar liquid sheets induced by heat and mass transfer. Applied Mathematics and Computation, 145(1):151–177, December 20, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ramos:2003:ODI

[Ram03c] J. I. Ramos. Oscillatory dynamics of inviscid planar liquid sheets. Applied Mathematics and Computation, 143(1):109–144, October 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ramos:2003:STP

[Ram03d]  J. I. Ramos.  Spatio-

temporal patterns in twodimensional excitable media subject to Robin boundary conditions. Applied Mathematics and Computation, 146(1):55–72, December 30, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ramos:2004:NSE

[Ram04a] J. I. Ramos. Non-standard, explicit integration algorithms based on linearization for nonlinear dynamic response analysis. Applied Mathematics and Computation, 159(3):695–715, December 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ramos:2004:RSW

[Ram04b]  J. I. Ramos.  Robust-

ness of spiral waves in two-dimensional reactive– diffusive media. Applied Mathematics and Computation, 148(3):681–695, January 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ramos:2004:SWB

[Ram04c]  J. I. Ramos.  Spiral

wave break-up and planar front formation in two-dimensional reactive– diffusive media with straining. Applied Mathematics and Computation, 154 (3):697–711, July 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rao:2002:CBP

[Rao02] Prasada Rao. Contribution of Boussinesq pressure and bottom roughness terms for open channel flows with shocks. Applied Mathematics and Computation, 133(2–3):581–590, December 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rao:2004:MDF

[Rao04a]  Prasada Rao.  A mov-

ing domain formulation for modeling two dimensional open channel transient flows. Applied Mathematics and Computation, 154(3):769– 781, July 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rao:2004:PHM

[Rao04b] Prasada Rao. A parallel hydrodynamic model for shallow water equations. Applied Mathematics and Computation, 150(1):291–302, February 27, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rashid:2002:NFG

[Ras02a] LaiLa E. M. Rashid. Nonfinitely generated function algebras. Applied Mathematics and Computation, 130(1):1–4, July 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rashid:2002:PGR

[Ras02b] Laila E. M. Rashid. Projection geometry over rings. Applied Mathematics and Computation, 130(1):5–10, July 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rashid:2002:SCB

[Ras02c] Laila E. M. Rashid. A sharp Castelnuovo bound for the normalization of certain projective surfaces. Applied Mathematics and Computation, 131(1):141– 146, September 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rashid:2002:SAI

[Ras02d] Laila E. M. Rashid. Subadjoint ideals and hyperplane sections. Applied Mathematics and Computation, 131(1):133–139, September 10, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rashed:2003:EMT

[Ras03a] M. T. Rashed. An expansion method to treat integral equations. Applied Mathematics and Computation, 135(1):65–72, February 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rashed:2003:NSS

[Ras03b] M. T. Rashed. Numerical solution of a special type of integro-differential equations. Applied Mathematics and Computation, 143 (1):73–88, October 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rashed:2003:NSI

[Ras03c] M. T. Rashed. Numerical solutions of the integral equations of the first kind. Applied Mathematics and Computation, 145(2–3): 413–420, December 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rashed:2004:ACD

[Ras04a] M. T. Rashed. An algorithm to compute the derivatives of the function. Applied Mathematics and Computation, 156(2):493–497, September 6, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rashed:2004:LICb

[Ras04b] M. T. Rashed. Lagrange interpolation to compute the derivatives of a function. Applied Mathematics and Computation, 156(2): 499–505, September 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rashed:2004:LICa

[Ras04c] M. T. Rashed. Lagrange interpolation to compute the numerical solutions of differential, integral and integrodifferential equations. Applied Mathematics and Computation, 151(3):869–878, April 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rashed:2004:NSFa

[Ras04d] M. T. Rashed. Numerical solution of functional differential, integral and integrodifferential equations. Applied Mathematics and Computation, 156(2):485–492, September 6, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rashed:2004:NSFb

[Ras04e] M. T. Rashed. Numerical solutions of functional integral equations. Applied Mathematics and Computation, 156(2):507–512, September 6, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rasulov:2004:FDS

[Ras04f] Mahir Rasulov. The finite differences scheme for the first order system of nonlinear differential equations in a class of discontinuous functions. Applied Mathematics and Computation, 154 (3):671–681, July 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Reddy:2003:MRO

[RC03] Y. N. Reddy and P. Pramod Chakravarthy. Method of reduction of order for solving singularly perturbed twopoint boundary value problems. Applied Mathematics and Computation, 136 (1):27–45, March 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Reddy:2004:EFF

[RC04a] Y. N. Reddy and P. Pramod Chakravarthy. An exponentially fitted finite difference method for singular perturbation problems. Applied Mathematics and Computation, 154(1):83–101, June 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Reddy:2004:IVA

[RC04b] Y. N. Reddy and P. Pramod Chakravarthy. An initialvalue approach for solving singularly perturbed twopoint boundary value problems. Applied Mathematics and Computation, 155 (1):95–110, July 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Reddy:2004:NPM

[RC04c] Y. N. Reddy and P. Pramod Chakravarthy. Numerical patching method for singularly perturbed two-point boundary value problems using cubic splines. Applied Mathematics and Computation, 149(2):441–468, February 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rossi:2004:EPM

[RCC04] E. V. Rossi, D. A. Cicci, and J. E. Cochran, Jr. Existence of periodic motions of a tether trailing satellite. Applied Mathematics and Computation, 155 (1):269–281, July 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rasulov:2003:FDM

[RCS03] M. Rasulov, E. Coskun, and B. Sinsoysal. A finite differences method for a two-dimensional nonlinear hyperbolic equation in a class of discontinuous functions. Applied Mathematics and Computation, 140(2–3): 279–295, August 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rabi:2001:OCA

[RdL01] J. A. Rabi and M. J. S. de Lemos. Optimization of convergence acceleration in multigrid numerical solutions of conductiveconvective problems. Applied Mathematics and Computation, 124(2):215–226, October 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/113/31/31/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000886.

Rakha:2004:ABH

[RES04a] Medhat A. Rakha and Essam S. El-Sedy. Application of basic hypergeometric series. Applied Mathematics and Computation, 148(3): 717–723, January 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rida:2004:FCG

[RES04b] Saad Zagloul Rida and Ahmed M. A. El-Sayed. Fractional calculus and generalized Rodrigues formula. Applied Mathematics and Computation, 147(1):29–43, January 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rivera-Gallego:2003:SAN

[RG03] Wilson Rivera-Gallego. Stability analysis of numerical boundary conditions in domain decomposition algorithms. Applied Mathematics and Computation, 137(2– 3):375–385, May 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Radwan:2004:MIK

[RH04] Ahmed E. Radwan and Hoda A. Hashem. Magnetohydrodynamic instability of Kelvin–Helmholtz with annulus cylindrical interfaces pervaded by helical varying magnetic fields. Applied Mathematics and Computation, 147(1):191–210, January 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Richard:2004:CAA

[RHB04] M. J. Richard, M. Z. Huang, and M. Bouazara. Computer aided analysis and optimal design of mechanical systems using vectornetwork techniques. Applied Mathematics and Computation, 157(1):175–200, September 27, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Richardson:2000:SDU

[Ric00] Walter B. Richardson, Jr. Steepest descent using smoothed gradients. Applied Mathematics and Computation, 112(2–3):241–254, June 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/86/23/
27/abstract.html;   http:
//www.elsevier.nl/gej-

ng/29/17/20/86/23/27/article.
pdf.

Ricci:2003:MAO

[Ric03] Paolo E. Ricci. Memory of Alessandro Ossicini (1921– 1999). Applied Mathematics and Computation, 141(1): 225–230, August 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rida:2004:GUG

[Rid04] Saad Zagloul Rida. On the generalized ultraspherical or Gegenbauer functions of fractional orders. Applied Mathematics and Computation, 151(2):543– 565, April 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rasulov:2003:FDS

[RK03] Mahir Rasulov and Turhan Karaguler. Finite difference schemes for solving system equations of gas dynamic in a class of discontinuous functions. Applied Mathematics and Computation, 143(1): 145–164, October 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rasulov:2004:FDM

[RKS04a] Mahir Rasulov, Turhan Karaguler, and Bahaddin Sinsoysal. Finite difference method for solving boundary initial value problem of a system hyperbolic equations in a class of discontinuous functions. Applied Mathematics and Computation, 149(1):47–63, February 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rasulov:2004:NSC

[RKS04b] Mahir Rasulov, Turhan Karaguler, and Bahaddin Sinsoysal. Numerical solution of Cauchy problem for second order nonlinear wave equation with changeable type in a class of discontinues functions. Applied Mathematics and Computation, 147(2):423–437, January 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rao:2002:IRB

[RM02] Prasada Rao and Miguel A. Medina, Jr. An improved radiating boundary equation for free surface flows. Applied Mathematics and Computation, 132(1):73–86, October 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rao:2003:EMG

[RM03] Prasada Rao and Miguel A. Medina, Jr. Evaluation of V and W multiple grid cycles for modeling one and two-dimensional transient free surface flows. Applied Mathematics and Computation, 138(1):151–167, June 1, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Razzaghi:2001:ARH

[RO01] M. Razzaghi and Y. Ordokhani. An application of rationalized Haar functions for variational problems. Applied Mathematics and Computation, 122(3): 353–364, August 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/106/26/
26/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/106/26/26/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000503.

Rolle:2002:NAL

[Rol02] Jean-Daniel Rolle. Notes about the last principal component. Applied Mathematics and Computation, 126(2–3):231–241, March 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/122/30/33/abstract.
html;   http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300300001533.

Ronveaux:2003:FHD

[Ron03] A. Ronveaux. Factorization of the Heun’s differential operator. Applied Mathematics and Computation, 141(1): 177–184, August 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rao:2000:SFS

[RP00] Prasada Rao and B. S. Pani. Simulation of free surface flows using a Runge–Kutta technique. Applied Mathematics and Computation, 114(1):27–38, August 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/
                                        10/9/12/89/20/22/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/89/20/
                                        22/article.pdf.

Raptis:2004:ETR

[RPT04] A. Raptis, C. Perdikis, and H. S. Takhar. Effect of thermal radiation on MHD flow. Applied Mathematics and Computation, 153 (3):645–649, June 14, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Reddy:2002:NIM

[RR02a] Y. N. Reddy and K. Anantha Reddy. Numerical integration method for general singularly perturbed two point boundary value problems. Applied Mathematics and Computation, 133(2–3): 351–373, December 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ronveaux:2002:EMP

[RR02b] A. Ronveaux and L. Rebillard. Expansion of multivariable polynomials in products of orthogonal polynomials in one variable. Applied Mathematics and Computation, 128(2–3):387–414, May 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ramos:2001:DDT

[RS01a] J. I. Ramos and E. Soler.

Domain decomposition techniques for reaction-diffusion equations in two-dimensional regions with re-entrant corners. Applied Mathematics and Computation, 118(2– 3):189–221, March 9, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.nl/gej-ng/
10/9/12/96/25/25/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/96/25/
25/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399002143.

Reidys:2001:NFL

[RS01b] Christian M. Reidys and Peter F. Stadler. Neutrality in fitness landscapes. Applied Mathematics and Computation, 117(2–3):321–350, January 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/92/23/35/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/92/23/
                                        35/article.pdf;   http:
                                        //www.sciencedirect.com/
                                        science/article/pii/S0096300399001666.

Rassias:2002:CCB

[RS02a] Themistocles M. Rassias and H. M. Srivastava. A certain class of biorthogonal polynomials associated with the Laguerre polynomials. Applied Mathematics and Computation, 128(2– 3):379–385, May 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rassias:2002:SCI

[RS02b] Themistocles M. Rassias and H. M. Srivastava. Some classes of infinite series associated with the Riemann Zeta and Polygamma functions and generalized harmonic numbers. Applied Mathematics and Computation, 131(2–3):593–605, September 25, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Raphael:2003:DSA

[RS03] B. Raphael and I. F. C.

Smith. A direct stochastic algorithm for global search. Applied Mathematics and Computation, 146(2–3):729– 758, December 31, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rhoades:2004:LPA

[RS04a] B. E. Rhoades and Ekrem Savaş. Local property of absolute weighted mean summability of Fourier series. Applied Mathematics and Computation, 153 (2):445–452, June 4, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rhoades:2004:SFT

[RS04b] B. E. Rhoades and Ekrem Savaş. A summability factor theorem and applications. Applied Mathematics and Computation, 153 (1):155–163, May 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rekab:2003:SDE

[RT03] Kamel Rekab and Mohamed Tahir. A sequential design for estimating a nonlinear parametric function. Applied Mathematics and Computation, 138(1):113– 120, June 1, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rui:2003:MSM

[Rui03] Hongxing Rui. Multiplicative Schwarz methods for parabolic problems. Applied Mathematics and Computation, 136(2–3):593–610, March 15, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rakocevic:2003:RAW

[RW03] Vladimir Rakočević and Yimin Wei. The representation and approximation of the W -weighted Drazin inverse of linear operators in Hilbert space. Applied Mathematics and Computation, 141(2–3):455–470, September 5, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Romanov:2003:IPL

[RWC03] Vladimir G. Romanov, Cheng-I Weng, and Tei-Chen Chen. An inverse problem for a layered elastic plate. Applied Mathematics and Computation, 137(2– 3):349–369, May 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:2004:CRS

[rWS04] Guo rong Wang and Jie Sun. A Cramer rule for solution of the general restricted matrix equation. Applied Mathematics and Computation, 154(2):415– 422, July 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sabzevari:2018:ENS

[Sab18]  Mehdi Sabzevari.  Erra-

tum to “Numerical solution of linear Fredholm integral equation by using hybrid Taylor and Block–Pulse functions” [Appl. Math. Comput. 149 (2004) 799–

806. Applied Mathematics and Computation, 339(??): 899–915, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300318305824.

See [MM04b].

Sadek:2002:ABS

[Sad02]  A. I. Sadek.  On the

asymptotic behaviour of solutions of certain fifth-order ordinary differential equations. Applied Mathematics and Computation, 131(1): 1–13, September 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sadek:2003:IRC

[Sad03] A. I. Sadek. Instability results for certain systems of fourth and fifth order differential equations. Applied Mathematics and Computation, 145(2–3):541–549, December 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sadek:2004:SNV

[Sad04a] A. I. Sadek. On the stability of a nonhomogeneous vector differential equation of the fourth-order. Applied Mathematics and Computation, 150(1):279–289, February 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sadek:2004:SSC

[Sad04b] A. I. Sadek. On the stability of solutions of certain fourth order delay differential equations. Applied Mathematics and Computation, 148(2): 587–597, January 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sahai:2004:IAI

[Sah04a] Ashok Sahai. An iterative algorithm for improved approximation by Bernstein’s operator using statistical perspective. Applied Mathematics and Computation, 149(2):327–335, February 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sahai:2004:ENP

[Sah04b] Ashok Sahai. On an estimator of normal population mean and UMVU estimation of its relative efficiency. Applied Mathematics and Computation, 152 (3):701–708, May 13, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sahiner:2004:OSO

[Şah04c] Y. Şahiner. On oscillation of second order neutral type delay differential equations. Applied Mathematics and Computation, 150 (3):697–706, March 17, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Saied:2002:AAU

[Sai02]  Effat A. Saied.  Ana-

lytic assessment of used nuclear fuel disposal. Applied Mathematics and Computation, 125(2–3):209–220, January 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/120/31/31/abstract.
html;   http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300300001211.

Saker:2003:NOC

[Sak03a] S. H. Saker. New oscillation criteria for second-order nonlinear neutral delay difference equations. Applied Mathematics and Computation, 142(1):99–111, September 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Saker:2003:OGAa

[Sak03b] S. H. Saker. Oscillation and global attractivity in hematopoiesis model with delay time. Applied Mathematics and Computation, 136(2–3):241–250, March 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Saker:2003:OGAb

[Sak03c] S. H. Saker. Oscillation and global attractivity in hematopoiesis model with periodic coefficients. Applied Mathematics and Computation, 142(2–3):477–494, October 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Saker:2003:OPN

[Sak03d] S. H. Saker. Oscillation of parabolic neutral delay difference equations with several positive and negative coefficients. Applied Mathematics and Computation, 143(1):173–186, October 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Saker:2003:OSO

[Sak03e] S. H. Saker. Oscillation of second-order perturbed nonlinear difference equations. Applied Mathematics and Computation, 144(2–3): 305–324, December 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Saker:2004:OND

[Sak04] S. H. Saker. Oscillation of nonlinear dynamic equations on time scales. Applied Mathematics and Computation, 148(1):81–91, January 20, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sallam:2000:SQS

[Sal00] S. Sallam. Stable quartic spline integration method for solving stiff ordinary differential equations. Applied Mathematics and Computation, 116(3):245–255, December ??, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/91/23/21/abstract.
html;   http://www.elsevier.

                                        nl/gej-ng/10/9/12/91/23/
                                        21/article.pdf.

Saldanha:2003:FDM

[Sal03a] Godfrey Saldanha. A finite difference method for self-adjoint elliptic equations in three dimensions. Applied Mathematics and Computation, 146(2–3):803–811, December 31, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Saldanha:2003:SCD

[Sal03b] Godfrey Saldanha. Single cell discretizations of order two and four for selfadjoint elliptic equations. Applied Mathematics and Computation, 134(1):1–8, January 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Samaan:2004:SSF

[Sam04] Angail A. Samaan. State space formulation for magnetohydrodynamic free convection flow with two relaxation times. Applied Mathematics and Computation, 152(2):299–321, May 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sanziel:2000:EWT

[San00] Maria Cristina Sanziel. Existence of a waiting time in a discrete two-phase Stefan problem. Applied Mathematics and Computation, 113(2–3):275–288, July 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/87/23/29/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/87/23/
29/article.pdf.

Sarhan:2002:REB

[Sar02] Ammar M. Sarhan. Reliability equivalence with a basic series/parallel system. Applied Mathematics and Computation, 132 (1):115–133, October 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sarhan:2003:EBE

[Sar03a] Ammar M. Sarhan. Empirical Bayes estimates in exponential reliability model. Applied Mathematics and Computation, 135(2–3):319–332, March 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sarhan:2003:ESC

[Sar03b] Ammar M. Sarhan. Estimation of system components reliabilities using masked data. Applied Mathematics and Computation, 136 (1):79–92, March 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sarhan:2004:PEG

[Sar04a] Ammar M. Sarhan. Parameter estimations in a general hazard rate model using masked data. Applied Mathematics and Computation, 153(2):513–536, June 4, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sarhan:2004:PEL

[Sar04b] Ammar M. Sarhan. Parameter estimations in linear failure rate model using masked data. Applied Mathematics and Computation, 151 (1):233–249, March 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sarhan:2004:RES

[SARAEG04] Ammar M. Sarhan, A. S. Al-Ruzaiza, I. A. Alwasel, and Awad I. El-Gohary. Reliability equivalence of a series–parallel system. Applied Mathematics and Computation, 154(1):257–277, June 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Savas:2004:NCI

[Sav04a] Ekrem Savaş. Necessary conditions for inclusion relations for absolute summability. Applied Mathematics and Computation, 151 (2):523–531, April 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Severino:2004:AQM

[SAV04b] J. S. Severino, E. J. Allen, and H. D. Victory, Jr. Acceleration of quasi-Monte Carlo approximations with applications in mathematical finance. Applied Mathematics and Computation, 148(1): 173–187, January 20, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Solak:2003:SNC

[SB03] Süleyman Solak and Durmu Bozkurt. On the spectral norms of Cauchy–Toeplitz and Cauchy–Hankel matrices. Applied Mathematics and Computation, 140(2–3): 231–238, August 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sharaf:2004:EAS

[SB04a] M. A. Sharaf and M. A.

Banajh. Error analysis for stationary one point iteration formulae of arbitrary order of convergence for solving universal Kepler’s equation. Applied Mathematics and Computation, 155(3): 699–707, August 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Slodicka:2004:NSM

[SB04b] Marián Slodička and Lubomír Baňas. A numerical scheme for a Maxwell–Landau– Lifshitz–Gilbert system. Applied Mathematics and Computation, 158(1):79–99, October 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Simsek:2003:ACM

[ŞBC03] Hakan Şimşek, Mustafa Bayram, and İsmail Can. Automatic calculation of minimum crossing numbers of 3-braids. Applied Mathematics and Computation, 144(2–3):507–516, December 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Simsek:2004:CPC

[ŞBY04] Hakan Şimşek, Mustafa Bayram, and Uğur Yavuz. A computer program to calculate Alexander polynomial from Braids presentation of the given knot. Applied Mathematics and Computation, 153(1):199– 204, May 25, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sheu:2002:FEA

[SC02] Tony W. H. Sheu and Y. H. Chen. Finite element analysis of contaminant transport in groundwater. Applied Mathematics and Computation, 127(1):23–43, March 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/123/28/28/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001600.

Chang:2003:FQI

[sC03a] Shih sen Chang. Fuzzy quasivariational inclusions in Banach spaces. Applied Mathematics and Computation, 145(2–3):805–819, December 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Simsek:2003:SAM

[ŞÇ03b] Hakan Şimşek and Ercan Çelik. The successive approximation method and Padé approximants for solutions the non-linear boundary value problem. Applied Mathematics and Computation, 146(2–3):681–690, December 31, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Song:2003:UPG

[SC03c] Xinyu Song and Jing’an Cui. Uniform persistence and global attractivity for nonautonomous competitive systems with nonlinear dispersion and delays. Applied Mathematics and Computation, 146(1):273–288, December 30, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Srivastava:2003:OSP

[SC03d] H. M. Srivastava and Y. Ben Cheikh. Orthogonality of some polynomial sets via quasi-monomiality. Applied Mathematics and Computation, 141(2–3):415–425, September 5, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Saker:2004:OCD

[SC04] Samir H. Saker and Sui Sun Cheng. Oscillation criteria for difference equations with damping terms. Applied Mathematics and Computation, 148(2):421–442, January 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Szurley:2001:ECC

[SD01] David Szurley and Jinqiao Duan. The effect of changing the Coriolis force gradient parameter on the escape probability and mean residence time. Applied Mathematics and Computation, 118(2–3):261–273, March 9, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/96/25/28/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/96/25/
28/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399002180.

Slodicka:2002:NBV

[SD02a] M. Slodička and H. De Schepper. A nonlinear boundary value problem containing nonstandard boundary conditions. Applied Mathematics and Computation, 132(2–3):559–574, November 10, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Slodicka:2002:IPP

[SD02b] M. Slodička and H. De Schepper. On an inverse problem of pressure recovery arising from soil venting facilities. Applied Mathematics and Computation, 129(2– 3):469–480, July 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Srivastava:2003:PAS

[SDR03] Hari M. Srivastava, Giuseppe Dattoli, and Paolo E. Ricci. Preface: Advanced Special Functions and Related Topics in Differential Equations, Third Melfi Workshop, Proceedings of the Melfi School on Advanced Topics in Mathematics and Physics. Applied Mathematics and Computation, 141 (1):1–2, August 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Siksek:2004:PND

[SE04] Samir Siksek and Esam El Sedy. Points of nondifferentiability of convex functions. Applied Mathematics and Computation, 148(3):725–728, January 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sarhan:2003:ECR

[SEB03] Ammar M. Sarhan and Ahmed H. El-Bassiouny. Estimation of components reliability in a parallel system using masked system life data. Applied Mathematics and Computation, 138(1):61–75, June 1, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sarhan:2003:PERb

[SEG03] Ammar M. Sarhan and Awad El-Gohary. Parameter estimations of 1-out-of-2: G repairable system. Applied Mathematics and Computation, 145(2–3):469– 479, December 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Seida:2003:PEW

[Sei03] Osama M. Abo Seida. Propagation of electromagnetic waves in a rectangular tunnel. Applied Mathematics and Computation, 136(2–3): 405–413, March 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Soon:2001:FFT

[SEK01] Boon Yi Soon, Paul W.

Eloe, and David Kammler. The fast Fourier transform method and illconditioned matrices. Applied Mathematics and Computation, 117(2–3):117–129, January 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/92/23/20/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/92/23/
20/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S009630039900171X.

Seliem:2003:PTE

[Sel03] Adel A. Abo Seliem. Propagation of the transient electromagnetic field above atmospheric surface duct. Applied Mathematics and Computation, 145(2–3):631–639, December 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Seliem:2004:ETE

[Sel04a] Adel A. S. Abo Seliem. Evaluation of the transient electromagnetic field in evaporation duct. Applied Mathematics and Computation, 151(2):411–421, April 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Seliem:2004:TEF

[Sel04b] Adel A. S. Abo Seliem. The transient electromagnetic field of a dielectric layer. Applied Mathematics and Computation, 157(3): 759–764, October 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Senthilvelan:2001:EAH

[Sen01] M. Senthilvelan. On the extended applications of Homogeneous Balance Method. Applied Mathematics and Computation, 123(3):381– 388, October 15, 2001. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/9/12/110/32/31/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S009630030000076X. See

comment [Fen04b].

Sever:2002:UHO

[Sev02] Ali Sever. On uniqueness for a higher-order Cauchy problem of mathematical physics. Applied Mathematics and Computation, 132(2–3):265– 270, November 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sever:2004:SEI

[Sev04] Ali Sever. A stability estimate of an inverse problem in financial prospection. Applied Mathematics and Computation, 150 (3):803–810, March 17, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shinohara:2001:ECM

[SG01] Shuji Shinohara and Yukio Pegio Gunji. Emergence and collapse of money through reciprocity. Applied Mathematics and Computation, 117(2–3):131–150, January 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/

9/12/92/23/21/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/92/23/
21/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399001691.

Shohdohji:2001:OIQ

[SH01] Tsutomu Shohdohji and Yasushi Hoshino. Optimization of image quality for decoded images using three-dimensional smoothing method. Applied Mathematics and Computation, 120(1–3):301–311, May 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/104/21/
44/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/104/21/44/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002477.

Seyidmamedov:2002:DLC

[SH02] Zahir Seyidmamedov and Alemdar Hasanov. Determination of leading coefficients in Sturm–Liouville operator from boundary measurements. I. A stripping algorithm. Applied Mathematics and Computation, 125(1):1–21, January 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/120/27/27/abstract.

html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001041.

Shawagfeh:2002:AAS

[Sha02] Nabil T. Shawagfeh. Analytical approximate solutions for nonlinear fractional differential equations. Applied Mathematics and Computation, 131(2–3):517–529, September 25, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shampine:2003:SBV

[Sha03] L. F. Shampine. Singular boundary value problems for ODEs. Applied Mathematics and Computation, 138 (1):99–112, June 1, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shah:2004:ENE

[Sha04a] Shujaat Ali Shah. Experimentation with Newton’s extension to solve nonlinear algebraic equations. Applied Mathematics and Computation, 155(2):503–505, August 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shao:2004:IDS

[Sha04b]  Zuhua Shao.  Improve-

ment of digital signature with message recovery using self-certified public keys and its variants. Applied Mathematics and Computation, 159(2):391–399, December 6, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shesheng:2003:HMM

[She03] Zhang Shesheng. A hybrid multigrid method for the unsteady incompressible Navier–Stokes equations. Applied Mathematics and Computation, 138(2– 3):341–353, June 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shieh:2002:GFC

[Shi02] Cheng-Shion Shieh. Genetic fuzzy control for timevarying delayed uncertain systems with a robust stability safeguard. Applied Mathematics and Computation, 131(1):39–58, September 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shi:2004:CLS

[Shi04]  Zhen-Jun Shi.  Conver-

gence of line search methods for unconstrained optimization. Applied Mathematics and Computation, 157(2): 393–405, October 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Silver:2002:AFP

[Sil02] G. L. Silver. Analysis of four-point grids: the diamond configuration. Applied Mathematics and Computation, 131(2–3):215–221, September 25, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Silver:2003:AFP

[Sil03] G. L. Silver. Analysis of five-point grids: the diamond configuration. Applied Mathematics and Computation, 144(2–3):389–395, December 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Silver:2004:ATD

[Sil04] G. L. Silver. Analysis of three-dimensional grids: the eight-point cube. Applied Mathematics and Computation, 153(2):467–473, June 4, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Simsek:2004:ECP

[Şim04a] Hakan Şimşek. An estimate for the completeness of products of solutions of PDE. Applied Mathematics and Computation, 147(3): 799–804, January 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Simsek:2004:LCP

[Şim04b] Hakan Şimşek. On the lateral Cauchy problem. Applied Mathematics and Computation, 150(3):833–839, March 17, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sahai:2004:NCQ

[SJM04] A. Sahai, R. P. Jaju, and P. M. Mashwama. A new computerizable quadrature formula using probabilistic approach. Applied Mathematics and Computation, 158(1):217–224, October 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sheng:2002:MAL

[SK02] Q. Sheng and A. Q. M. Khaliq. Modified arclength adaptive algorithms for degenerate reactiondiffusion equations. Applied Mathematics and Computation, 126(2–3):279–297, March 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/122/30/37/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001582.

Saxena:2003:FGF

[SK03] R. K. Saxena and S. L. Kalla. On a fractional generalization of the free electron laser equation. Applied Mathematics and Computation, 143(1):89–97, October 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sritanratana:2004:PNL

[SK04] G. Sritanratana and A. Kananthai. On the product of the non-linear Diamond operators related to the elastic wave. Applied Mathematics and Computation, 147 (1):79–88, January 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Schaerer:2004:MSS

[SKM04] Christian E. Schaerer, Eugenius Kaszkurewicz, and Norberto Mangiavacchi. A multilevel Schwarz shooting method for the solution of the Poisson equation in two dimensional incompressible flow simulations. Applied Mathematics and Computation, 153 (3):803–831, June 14, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sun:2003:MPSa

[SL03a] Jian-Ping Sun and Wan-Tong Li. Multiple positive solutions of a discrete difference system. Applied Mathematics and Computation, 143(2–3):213–221, November 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sun:2003:MPSb

[SL03b] Jian-Ping Sun and Wan-Tong Li. Multiple positive solutions to secondorder Neumann boundary value problems. Applied Mathematics and Computation, 146(1):187–194, December 30, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shen:2004:NIE

[SL04a] Youjian Shen and Wei Lin. The natural integral equations of plane elasticity problem and its wavelet methods. Applied Mathematics and Computation, 150 (2):417–438, March 8, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sun:2004:PSS

[SL04b] Hong-Rui Sun and Wan-Tong Li. Positive solutions of second-order half-linear dynamic equations on time scales. Applied Mathematics and Computation, 158(2): 331–344, November 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sun:2004:QAD

[SL04c] Hong-Rui Sun and Wan-Tong Li. Qualitative analysis of a discrete logistic equation with several delays. Applied Mathematics and Computation, 147(2): 515–525, January 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sun:2004:EPS

[SL04d] Jian-Ping Sun and Wan-Tong Li. Existence of positive solutions of boundary value problem for a discrete difference system. Applied Mathematics and Computation, 156(3):857– 870, September 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Slavov:2003:WSO

[Sla03] Zdravko Dimitrov Slavov.

Weak and strong optimality and increasing concave functions. Applied Mathematics and Computation, 135(2– 3):517–529, March 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Slavov:2004:SPO

[Sla04] Zdravko Dimitrov Slavov.

Structure of the Pareto optimality set with fixed total resources and consumption sets. Applied Mathematics and Computation, 154 (1):75–81, June 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shi:2004:CCM

[SLC04] Jinlin Shi, Muren Lin, and Jiyang Chen. The calculation for characteristic multiplier of Hill’s equation. Applied Mathematics and Computation, 159(1): 57–77, November 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2003:NSL

[sLqZ03a] De sheng Li and Hong qing Zhang. New solitonlike solutions to the potential Kadomstev–Petviashvili (PKP) equation. Applied Mathematics and Computation, 146(2–3):381–384, December 31, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2003:SCV

[sLqZ03b] De sheng Li and Hong qing Zhang. Symbolic computation and various exact solutions of potential Kadomstev–Petviashvili equation. Applied Mathematics and Computation, 145(2–3):351–359, December 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sano:2000:ADS

[SM00a] Itaru Sano and Sonoyo Mukai. Algorithm description of system flow for global aerosol distribution. Applied Mathematics and Computation, 116(1–2):79–91, November 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/91/20/27/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/91/20/
27/article.pdf.

Shigematsu:2000:TLR

[SM00b] Yukifumi Shigematsu and Gen Matsumoto. Temporal learning rule and dynamic neural network model. Applied Mathematics and Computation, 111(2– 3):209–218, May 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/85/23/
27/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/85/23/27/article.
pdf.

Scitovski:2002:SPE

[SM02a] Rudolf Scitovski and Marcel Meler. Solving parameter estimation problem in new product diffusion models. Applied Mathematics and Computation, 127(1):45–63, March 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        9/12/123/28/29/abstract.
                                        html;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300300001648.

Shaibu:2002:CML

[SM02b] A.-B. Shaibu and Hassen A. Muttlak. A comparison of the maximum likelihood estimators under ranked set sapling some of its modifications. Applied Mathematics and Computation, 129(2– 3):441–453, July 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Seliem:2004:EEF

[SM04a] Adel A. S. Abo Seliem and F. Maiz. Evaluation of the electric field strength above a two layer medium. Applied Mathematics and Computation, 157(2):589– 598, October 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sjoberg:2004:NLS

[SM04b] A. Sjöberg and F. M. Mahomed. Non-local symmetries and conservation laws for one-dimensional gas dynamics equations. Applied Mathematics and Computation, 150(2):379–397, March 8, 2004. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic).

Sun:2004:FDS

[SM04c] Tongjun Sun and Keying Ma. The finite difference streamline diffusion method for the incompressible Navier–Stokes equations. Applied Mathematics and Computation, 149(2): 493–505, February 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sun:2004:NOC

[SM04d] Yuan Gong Sun and Fan Wei Meng. New oscillation criteria for linear matrix Hamiltonian systems. Applied Mathematics and Computation, 155(1):259–268, July 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sapariuc:2004:NAV

[SMF04] I. Sapariuc, M. D. Marcozzi, and J. E. Flaherty. A numerical analysis of variational valuation techniques for derivative securities. Applied Mathematics and Computation, 159(1):171–198, November 25, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sun:2004:RMS

[SMQ04] Jian-Qiang Sun, Zhong-Qi Ma, and Meng-Zhao Qin. RKMK method of solving non-damping LL equations and ferromagnet chain equations. Applied Mathematics and Computation, 157(2): 407–424, October 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sugisaka:2001:LPN

[SN01] Masanori Sugisaka and Masayo Nagasaki. Learning performance of a neurocomputer for nonlinear dynamical system identification. Applied Mathematics and Computation, 120 (1–3):65–77, May 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/104/21/
27/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/104/21/27/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S009630039900288X.

Soliman:2002:PLF

[Sol02a] A. A. Soliman. On perturbing Liapunov functional. Applied Mathematics and Computation, 133(2–3):319–325, December 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Soliman:2002:SPI

[Sol02b] A. A. Soliman. On stability of perturbed impulsive differential systems. Applied Mathematics and Computation, 133(1):105–117, November 25, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Soliman:2002:TSN

[Sol02c] A. A. Soliman. On total ϕ_{0}-stability of nonlinear systems of differential equations. Applied Mathematics and Computation, 130 (1):29–38, July 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Soliman:2003:SCI

[Sol03a] A. A. Soliman. Stability criteria of impulsive differential systems. Applied Mathematics and Computation, 134(2–3):445–457, January 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Solis:2003:SLF

[Sol03b] Francisco J. Solis. Selflimitation, fishing and cannibalism. Applied Mathematics and Computation, 135(1):39–48, February 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Soliman:2004:SIP

[Sol04] A. A. Soliman. On stability for impulsive perturbed systems via cone-valued Lyapunov function method. Applied Mathematics and Computation, 157(1):269–279, September 27, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Soriano:2001:CC

[Sor01a] J. M. Soriano. A compactness condition. Applied Mathematics and Computation, 124(3):397–402, December 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/113/32/37/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001144.

Soriano:2001:MSV

[Sor01b] J. M. Soriano. Mappings sharing a value on finitedimensional spaces. Applied Mathematics and Computation, 121(2–3):391–395, June 15, 2001. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/105/25/
39/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/105/25/39/article.
pdf;   http://www.sciencedirect.

com/science/article/pii/
S0096300300000230.

Soriano:2001:OT

[Sor01c] J. M. Soriano. Open trajectories. Applied Mathematics and Computation, 124(2): 235–240, October 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/9/12/113/31/33/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000928.

Soriano:2002:SZF

[Sor02] J. M. Soriano. A second zero of a function. Applied Mathematics and Computation, 133(2–3):245–255, December 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Soriano:2003:SS

[Sor03] J. M. Soriano. A stable solution. Applied Mathematics and Computation, 140(2– 3):223–229, August 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sousa:2003:CSA

[Sou03] Ercília Sousa. The controversial stability analysis. Applied Mathematics and Computation, 145(2–3):777–794, December 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Simwa:2003:DMS

[SP03a] R. O. Simwa and G. P. Pokhariyal. A dynamical model for stage-specific HIV incidences with application to sub-Saharan Africa. Applied Mathematics and Computation, 146(1):93–104, December 30, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sun:2003:ASP

[SP03b] Ying-Fei Sun and Yan-Ze Peng. Analytic solutions for the problems of an inclusion of arbitrary shape embedded in a half-plane. Applied Mathematics and Computation, 140(1):105– 113, July 30, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Song:2004:TWD

[SPH04] Yongli Song, Yahong Peng, and Maoan Han. Travelling wavefronts in the diffusive single species model with Allee effect and distributed delay. Applied Mathematics and Computation, 152 (2):483–497, May 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shanthi:2002:ANM

[SR02a] V. Shanthi and N. Ramanujam. Asymptotic numerical methods for singularly perturbed fourth order ordinary differential equations of convection–diffusion type. Applied Mathematics and Computation, 133(2–3): 559–579, December 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shanthi:2002:NMB

[SR02b] V. Shanthi and N. Ramanujam. A numerical method for boundary value problems for singularly perturbed fourth-order ordinary differential equations. Applied Mathematics and Computation, 129(2–3):269– 294, July 10, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shanthi:2004:CMR

[SR04a] V. Shanthi and N. Ramanujam. Computational methods for reaction–diffusion problems for fourth order ordinary differential equations with a small parameter at the highest derivative. Applied Mathematics and Computation, 147 (1):97–113, January 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Srinivasacharya:2004:CFM

[SR04b] D. Srinivasacharya and I. Rajyalakshmi. Creeping flow of micropolar fluid past a porous sphere. Applied Mathematics and Computation, 153(3):843–854, June 14, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Srivastava:2003:CCS

[Sri03] H. M. Srivastava. Certain classes of series associated with the Zeta and related functions. Applied Mathematics and Computation, 141(1):13–49, August 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

SalinasdeRomero:2000:AFC

[SS00] Susana Salinas de Romero and H. M. Srivastava. An application of the Nfractional calculus operator method to a modified Whittaker equation. Applied Mathematics and Computation, 115(1):11–21, October 6, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/90/21/22/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/90/21/
                                        22/article.pdf.

Srivastava:2001:OFI

[SS01] H. M. Srivastava and R. K.

Saxena. Operators of fractional integration and their applications. Applied Mathematics and Computation, 118(1):1–52, February 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/96/20/20/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/96/20/
20/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399002088.

Shahruz:2002:HNS

[SS02] S. M. Shahruz and N. A.

Sakyaman. How to have narrow-stripe semiconductor lasers self-pulsate. Applied Mathematics and Computation, 130(1):11–27, July 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Seshaiyer:2003:NCF

[SS03] Padmanabhan Seshaiyer and Philip Smith. A nonconforming finite element method for sub-meshing. Applied Mathematics and Computation, 139(1):85– 100, July 1, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Saad:2004:SSI

[SS04a] Omar M. Saad and Walied H. Sharif. Stability set for integer linear goal programming. Applied Mathematics and Computation, 153 (3):743–750, June 14, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sheela:2004:LRS

[SS04b] S. Sheela and A. Singh. Lavrentiev regularization of a singularly perturbed elliptic PDE. Applied Mathematics and Computation, 148(1):189–205, January 20, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Singh:2004:SRS

[SS04c] Arindama Singh and S. Sheela. Showalter regularization of a singularly perturbed PDE. Applied Mathematics and Computation, 158(1):19–28, October 25, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Stehr:2001:POB

[SSPA01] Guido Stehr, Ferenc Szidarovszky, Olgierd A. Palusinski, and David Anderson. Performance optimization of binary weighted currentsteering D/A converters. Applied Mathematics and Computation, 119(2–3):339– 347, April 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/99/25/35/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/99/25/
35/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399002532.

Sarhan:2003:PERa

[ST03a] Ammar M. Sarhan and Lotfi Tadj. Parameters estimation of a repairable system. Applied Mathematics and Computation, 138(2– 3):217–226, June 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Stanimirovi:2003:PCP

[ST03b] Predrag S. Stanimirovi and Milan B. Tasi. A problem in computation of pseudoinverses. Applied Mathematics and Computation, 135(2– 3):443–469, March 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Stanimirovic:2004:PMR

[ST04] Predrag S. Stanimirović and Milan B. Tasić. Partitioning method for rational and polynomial matrices. Applied Mathematics and Computation, 155 (1):137–163, July 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Stadler:2000:ACB

[Sta00] Bärbel M. R. Stadler. Abstention causes bifurcations in two-party voting dynamics. Applied Mathematics and Computation, 114(2–3):215–232, September 11, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/89/22/28/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/89/22/
28/article.pdf.

Stanimirovi:2003:FAG

[Sta03] Predrag S. Stanimirovi. A finite algorithm for generalized inverses of polynomial and rational matrices. Applied Mathematics and Computation, 144(2–3):199–214, December 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Solak:2003:GLH

[STB03] Süleyman Solak, Ramazan Türkmen, and Durmu Bozkurt. On GCD, LCM and Hilbert matrices and their applications. Applied Mathematics and Computation, 146(2–3): 595–600, December 31, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shiraishi:2002:TSS

[STHN02] F. Shiraishi, H. Takeuchi, T. Hasegawa, and H. Nagasue. A Taylor-series solution in Cartesian space to GMA-system equations and its application to initialvalue problems. Applied Mathematics and Computation, 127(1):103–123, March 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/123/28/32/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300301000078.

Subasi:2002:OCP

[Sub02a] Murat Subaşi. An optimal control problem governed by the potential of a linear Schrödinger equation. Applied Mathematics and Computation, 131(1): 95–106, September 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Subasi:2002:SEC

[Sub02b] Murat Subaşi. A stability estimate for convolution equation. Applied Mathematics and Computation, 131 (1):195–200, September 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sugisaka:2000:SUC

[Sug00] Masanori Sugisaka. Sugisaka-Ueno-Chandrasekhar algorithm for digital communication and estimation. Applied Mathematics and Computation, 116(1–2):217–229, November 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/91/20/36/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/91/20/
                                        36/article.pdf.

Sun:2004:SCI

[Sun04a] Jitao Sun. Stability criteria of impulsive differential systems. Applied Mathematics and Computation, 156 (1):85–91, August 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sun:2004:NTR

[Sun04b]  Wenyu Sun.  Nonmono-

tone trust region method for solving optimization problems. Applied Mathematics and Computation, 156(1): 159–174, August 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sun:2004:NKT

[Sun04c]  Yuan Gong Sun.  New

Kamenev type oscillation criteria for linear matrix Hamiltonian systems. Applied Mathematics and Computation, 158(1):69–78, October 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shiraishi:2002:STP

[SV02] Fumihide Shiraishi and Eberhard O. Voit. Solution of a two-point boundary value model of immobilized enzyme reactions, using an S-system-based root-finding method. Applied Mathematics and Computation, 127(2–3):289–310, April 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/123/27/38/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S009630030100008X.

Slodicka:2004:NAD

[SV04] Marián Slodička and Roger Van Keer. A numerical approach for the determination of a missing boundary data in elliptic problems. Applied Mathematics and Computation, 147(2): 569–580, January 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sun:2001:ISS

[SW01] Yijing Sun and Shaoping Wu. Iterative solution for a singular nonlinear elliptic problem. Applied Mathematics and Computation, 118(1):53–62, February 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/96/20/21/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/96/20/
21/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399002076.

Sun:2002:TRO

[SW02] Wenyu Sun and Yimin Wei. Triple reverse-order law for weighted generalized inverses. Applied Mathematics and Computation, 125(2–3):221–229, January 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/120/31/32/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001223.

Shi:2003:OCC

[SW03] Wenying Shi and Peiguang Wang. Oscillatory criteria of a class of secondorder neutral functional differential equations. Applied Mathematics and Computation, 146(1):211–226, December 30, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shi:2004:OCS

[SW04a] Wenying Shi and Peiguang Wang. Oscillation criteria of second order nonlinear matrix differential systems. Applied Mathematics and Computation, 156(3):831– 846, September 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shiping:2004:PSK

[SW04b] Lu Shiping and Ge Weigao.

Periodic solutions for a kind of second-order neutral differential systems with deviating arguments. Applied Mathematics and Computation, 156(3):719–732, September 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sugisaka:2000:ABM

[SWL00] Masanori Sugisaka, Xin Wang, and Ju-Jung Lee. Artificial brain for a mobile vehicle. Applied Mathematics and Computation, 111(2– 3):137–145, May 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/85/23/
22/abstract.html;   http:
//www.elsevier.nl/gej-

ng/29/17/20/85/23/22/article.
pdf.

So:2000:NSS

[SWY00] Joseph W.-H. So, Jianhong Wu, and Yuanjie Yang. Numerical steady state and Hopf bifurcation analysis on the diffusive Nicholson’s blowflies equation. Applied Mathematics and Computation, 111(1):53–69, May 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/85/21/
24/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/85/21/24/article.
pdf.

Sun:2002:FDS

[SY02] Tongjun Sun and Danping Yang. The finite difference streamline diffusion methods for Sobolev equations with convectiondominated term. Applied Mathematics and Computation, 125(2–3):325–345, January 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/120/31/42/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001351.

Subasi:2003:TAM

[SY03] Murat Subaşi and Bünyamin Yildiz. Two approximation methods for multipoint boundary value problem of order 2N. Applied Mathematics and Computation, 137(2–3):511–524, May 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sato:2002:PLP

[SYK02] Kazunori Sato, Naoto Yoshida, and Norio Konno. Parity law for population dynamics of N-species with cyclic advantage competitions. Applied Mathematics and Computation, 126(2–3):255–270, March 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/122/30/35/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001557.

Suba:2003:DMS

[SYK03] Murat Suba, Bünyamin Yıldız, and Ahmet Kaçar. A difference method for solving parabolic equations of order 2n. Applied Mathematics and Computation, 140(2– 3):475–484, August 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Subasi:2004:IFS

[SYY04] Murat Subasi, Necmettin Yildirim, and Bünyamin Yildiz. An improvement on Fibonacci search method in optimization theory. Applied Mathematics and Computation, 147(3):893–901, January 16, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See comment [Kah06].

So:2001:TWD

[SZ01] Joseph W.-H. So and Xingfu Zou. Traveling waves for the diffusive Nicholson’s blowflies equation. Applied Mathematics and Computation, 122(3):385–392, August 15, 2001. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/106/26/
29/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/106/26/29/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000552.

Shen:2003:APP

[SZ03a] Yuming Shen and Jinxi Zhao. Analysis of peaks and plateaus in a Galerkin/ minimal residual pair of methods for solving Ax = b. Applied Mathematics and Computation, 144(2–3):441– 455, December 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sun:2003:NMI

[SZ03b] Fangyu Sun and Xin Zhang.

A new method of increasing the order of convergence step by step. Applied Mathematics and Computation, 137 (1):15–32, May 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shen:2004:GOS

[SZ04] Peiping Shen and Kecun Zhang. Global optimization of signomial geometric programming using linear relaxation. Applied Mathematics and Computation, 150(1): 99–114, February 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Si:2004:ASI

[SZK04] Jianguo Si, Weinian Zhang, and Gwang-Hui Kim. Analytic solutions of an iterative functional differential equation. Applied Mathematics and Computation, 150 (3):647–659, March 17, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shen:2003:AIA

[SZW03] Peiping Shen, Kecun Zhang, and Yanjun Wang. Applications of interval arithmetic in non-smooth global optimization. Applied Mathematics and Computation, 144(2–3):413–431, December 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tadj:2001:MAS

[Tad01] Lotfi Tadj. A matrix analytic solution to a hysteretic queueing system with random server capacity. Applied Mathematics and Computation, 119(2–3):161–175, April 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/99/25/23/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/99/25/
23/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399002544.

Tadj:2003:QQS

[Tad03]  Lotfi Tadj.  A quorum

queueing system under Dpolicy. Applied Mathematics and Computation, 144(2–3): 325–336, December 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tagliani:2000:DPD

[Tag00a]  Aldo Tagliani.  Discrete

probability distributions in the generalized moment problem. Applied Mathematics and Computation, 112(2– 3):333–343, June 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/86/23/
34/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/86/23/34/article.
pdf.

Tagliani:2000:ESD

[Tag00b] Aldo Tagliani. Existence and stability of a discrete probability distribution by maximum entropy approach. Applied Mathematics and Computation, 110(2–3):105– 114, April 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/83/23/
21/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/83/23/21/article.
pdf.

Tagliani:2001:DPD

[Tag01a]  Aldo Tagliani.  Discrete

probability distributions and moment problem: numerical aspects. Applied Mathematics and Computation, 119 (1):47–56, March 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.nl/gej-ng/
10/9/12/99/20/23/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/99/20/
23/article.pdf;   http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0096300399002283.

Tagliani:2001:EEP

[Tag01b] Aldo Tagliani. Entropy estimate of a probability density from its Mellin transform. Applied Mathematics and Computation, 123(3): 275–284, October 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.elsevier.com/gej-
                                        ng/10/9/12/110/32/27/abstract.
                                        html;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300300000679.

Tagliani:2001:NAF

[Tag01c] Aldo Tagliani. Numerical aspects of finite Hausdorff moment problem by maximum entropy approach. Applied Mathematics and Computation, 118(2–3):133–149, March 9, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/96/25/21/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/96/25/
                                        21/article.pdf;   http:
                                        //www.sciencedirect.com/
                                        science/article/pii/S0096300399002106.

Tagliani:2001:NIL

[Tag01d] Aldo Tagliani. Numerical inversion of Laplace transform on the real line of probability density functions. Applied Mathematics and Computation, 123(3):285–299, October 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/110/32/28/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000680.

Tagliani:2001:RPD

[Tag01e] Aldo Tagliani. Recovering a probability density function from its Mellin transform. Applied Mathematics and Computation, 118(2–3):151– 159, March 9, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/96/25/22/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/96/25/
22/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399002118.

Tagliani:2002:EEP

[Tag02a] Aldo Tagliani. Entropy estimate of probability densities having assigned moments: Stieltjes case. Applied Mathematics and Computation, 130(1):201–211, July 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tagliani:2002:NIM

[Tag02b]  Aldo Tagliani.  Numeri-

cal inversion of the Mellin transform on the real line for heavy-tailed probability density functions. Applied Mathematics and Computation, 130(2–3):525–536, August 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tagliani:2002:UBE

[Tag02c]  Aldo Tagliani.  An up-

per bound for entropy of discrete distributions having assigned moments. Applied Mathematics and Computation, 133(1):159–170, November 25, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tagliani:2003:NPD

[Tag03a]  Aldo Tagliani.  A note

on proximity of distributions in terms of coinciding moments. Applied Mathematics and Computation, 145(2–3): 195–203, December 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tagliani:2003:NILa

[Tag03b]  Aldo Tagliani.  Numer-

ical inversion of Laplace transform on the real line from expected values. Applied Mathematics and Computation, 134(2–3):459–472, January 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tagliani:2003:PDT

[Tag03c]  Aldo Tagliani.  On the

proximity of distributions in terms of coinciding fractional moments. Applied Mathematics and Computation, 145(2–3):501–509, December 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tagliani:2003:RMM

[Tag03d] Aldo Tagliani. Real moments from moments and vice versa. Applied Mathematics and Computation, 145(2–3):511–523, December 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tang:2003:SPM

[Tan03] Hui-Chin Tang. Symmetry properties of multiple recursive random number generators in full period and spectral test. Applied Mathematics and Computation, 142(2–3):291–303, October 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tarabia:2002:NFT

[Tar02] A. M. K. Tarabia. A new formula for the transient behaviour of a non-empty M/M/1/∞ queue. Applied Mathematics and Computation, 132(1):1–10, October 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tarabia:2003:NFB

[Tar03] A. M. K. Tarabia. A new formula for the busy period of a non-empty multiserver queueing system. Applied Mathematics and Computation, 143(2–3):401–408, November 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tawfik:2000:CMM

[Taw00] Bassel Tawfik. The correlation memory matrix for parameter estimation. Applied Mathematics and Computation, 111(1):87–101, May 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/85/21/
26/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/85/21/26/article.
pdf.

Turkmen:2002:BNC

[TB02] Ramazan Türkmen and Durmu Bozkurt. On the bounds for the norms of Cauchy–Toeplitz and Cauchy–Hankel matrices. Applied Mathematics and Computation, 132(2–3):633– 642, November 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Teamah:2004:ASP

[TB04] A. A. M. Teamah and H. S.

Bakouch. Asymptotic statistical properties of spectral estimates with different tapers for discrete time processes. Applied Mathematics and Computation, 150 (3):681–695, March 17, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tao:2002:BNN

[TCL02] Qing Tao, Jinde Cao, and Xin Liu. The BSB neural network in the convex body spanned by the prototype patterns for associative memory. Applied Mathematics and Computation, 132(2–3):575–587, November 10, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tao:2001:SHP

[TCS01] Qing Tao, Jinde Cao, and Demin Sun. A simple and high performance neural network for quadratic programming problems. Applied Mathematics and Computation, 124(2):251–260, October 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/113/31/35/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000977.

Tabata:2000:BSC

[TE00] Minoru Tabata and Nobuoki Eshima. The behavior of solutions to the Cauchy problem for the master equation. Applied Mathematics and Computation, 112 (1):79–98, June 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/86/21/
28/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/86/21/28/article.
pdf.

Tabata:2002:BSC

[TE02a] Minoru Tabata and Nobuoki Eshima. Blowing-up solutions to the Cauchy problem for the master equation. Applied Mathematics and Computation, 127(2–3):181–193, April 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/123/27/29/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001636.

Tabata:2002:CPN

[TE02b] Minoru Tabata and Nobuoki Eshima. The Cauchy problem for the nonlinear integro-partial differential equation in quantitative sociodynamics. Applied Mathematics and Computation, 132(2–3):537–552, November 10, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tabata:2003:SRA

[TE03] Minoru Tabata and Nobuoki Eshima. A self-referential agent-based model that consists of a large number of agents moving stochastically in a discrete bounded domain. Applied Mathematics and Computation, 143(2– 3):443–483, November 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tabata:2004:BSA

[TE04] Minoru Tabata and Nobuoki Eshima. The behavior of stochastic agent-based models when the number of agents and the time variable tend to infinity. Applied Mathematics and Computation, 152(1):47–70, April 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Teng:2000:NAL

[Ten00] Zhidong Teng. On the nonautonomous Lotka–Volterra N-species competing systems. Applied Mathematics and Computation, 114(2–3):175–185, September 11, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/89/22/25/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/89/22/
25/article.pdf.

Teng:2002:PSP

[Ten02] Zhidong Teng. On the periodic solutions of periodic multi-species competitive systems with delays. Applied Mathematics and Computation, 127(2–3):235– 247, April 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/123/27/34/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001715.

Tabata:2002:ICM

[TET02] Minoru Tabata, Nobuoki Eshima, and Ichiro Takagi. An infinite continuous model which derives from a finite discrete model describing the time evolution of the density of firms. Applied Mathematics and Computation, 125(1):105–132, January 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.

elsevier.com/gej-ng/10/
9/12/120/27/34/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001247.

Tarakci:2004:SCD

[TH04] Ömer Tarakci and H. Hilmi Hacisalihoğlu. Surfaces at a constant distance from the edge of regression on a surface. Applied Mathematics and Computation, 155 (1):81–93, July 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Editors:2003:EA

[The03] The Editors. Editorial apology. Applied Mathematics and Computation, 143(2–3): 187, November 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Thukral:2003:SIP

[Thu03] R. Thukral. Similarities of the integral Padé approximants. Applied Mathematics and Computation, 135(1): 129–145, February 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Thukral:2004:IIL

[Thu04a]  R. Thukral.  Introduc-

tion to the improved Levintype algorithms for accelerating convergence of sequence. Applied Mathematics and Computation, 151 (1):81–93, March 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Thukral:2004:SIP

[Thu04b]  R. Thukral.  Similari-

ties of the integral Padé approximants II. Applied Mathematics and Computation, 158(3):869–885, November 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Thuraisingham:2004:NMD

[Thu04c] R. A. Thuraisingham. A novel method to determine the time series of a secondorder chemical reaction in the unstable region of the Euler scheme. Applied Mathematics and Computation, 159(2):419–427, December 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tang:2002:STD

[THY02] X. H. Tang, Zhimin He, and J. S. Yu. Stability theorem for delay differential equations with impulses. Applied Mathematics and Computation, 131(2–3):373–381, September 25, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tian:2003:AOM

[Tia03] Hongjiong Tian. Accelerate overrelaxation methods for rank deficient linear systems. Applied Mathematics and Computation, 140(2–3): 485–499, August 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tian:2004:MMM

[Tia04a] Yongge Tian. More on maximal and minimal ranks of Schur complements with applications. Applied Mathematics and Computation, 152(3):675–692, May 13, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tian:2004:URF

[Tia04b]  Yongge Tian.  Using

rank formulas to characterize equalities for Moore– Penrose inverses of matrix products. Applied Mathematics and Computation, 147(2):581–600, January 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tien:2003:RDG

[Tie03] Tzu-Li Tien. A research on the deterministic grey dynamic model with multiple inputs DGDMMI(1,1,1). Applied Mathematics and Computation, 139(2–3):401– 416, July 15, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tineo:2001:CCS

[Tin01] Antonio Tineo. On the convexity of the carrying simplex of planar Lotka– Volterra competitive systems. Applied Mathematics and Computation, 123(1): 93–108, September 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/9/12/110/27/32/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000631.

Tseng:2003:DSM

[TJC03] Yuh-Min Tseng, Jinn-Ke Jan, and Hung-Yu Chien. Digital signature with message recovery using selfcertified public keys and its variants. Applied Mathematics and Computation, 136(2– 3):203–214, March 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tasci:2004:OGL

[TK04] Dursun Tasci and Emrah Kilic. On the order-k generalized Lucas numbers. Applied Mathematics and Computation, 155(3):637–641, August 16, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tsujino:2000:HRB

[TKK00] Hiroshi Tsujino, Edgar Koerner, and Hiroshi Kondo. Hypothetical reasoning and brainware. Applied Mathematics and Computation, 111(2–3):219–229, May 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/85/23/
28/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/85/23/28/article.
pdf.

Tervo:2003:ITP

[TLLKB03] J. Tervo, T. Lyyra-Laitinen, P. Kolmonen, and E. Boman. An inverse treatment planning model for intensity modulated radiation therapy with dynamic MLC. Applied Mathematics and Computation, 135(2– 3):227–250, March 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tao:2004:DGA

[TLX04] Qing Tao, Xin Liu, and Meisheng Xue. A dynamic genetic algorithm based on continuous neural networks for a kind of non-convex optimization problems. Applied Mathematics and Computation, 150(3):811–820, March 17, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Takashima:2000:ACS

[TM00] Tsutomu Takashima and Kazuhiko Masuda. Atmospheric correction for the satellite visible data over heterogeneous surfaces. Applied Mathematics and Computation, 116(1–2):181–196, November 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/91/20/34/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/91/20/
                                        34/article.pdf.

Tsamasphyros:2003:EBB

[TM03] G. Tsamasphyros and S. Markolefas. An estimate of the Babuška–Brezzi inf– sup discrete stability constant for general linear Petrov–Galerkin finite element formulations (an estimate of the Babuška–Brezzi stability constant). Applied Mathematics and Computation, 144(1):107–116, November 20, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tasic:2004:EMS

[TM04a] Bratislav Tasić and Robert M. M. Mattheij. An explicit method for solving flows of ODE. Applied Mathematics and Computation, 156 (3):633–652, September 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tunc:2004:FSE

[TM04b] E. Tunç and O. Sh. Muhtarov.

Fundamental solutions and eigenvalues of one boundaryvalue problem with transmission conditions. Applied Mathematics and Computation, 157(2):347–355, October 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tony:2000:CBF

[Ton00] Narayaninsamy Tony. A connection between fractional iteration and graph theory. Applied Mathematics and Computation, 107(2–3): 181–202, January 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/72/22/
29/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/72/22/29/article.
pdf.

Travis:2000:TOE

[Tra00] Betty Travis. N th Order extension of the Wintner-Leighton theorem. Applied Mathematics and Computation, 110(2–3):115–119, April 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/83/23/
                                        22/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/83/23/22/article.
                                        pdf.

Takashima:2003:ACS

[TRC03] Tsutomu Takashima, Christopher Rathbone, and Lesley Clementson. Atmospheric correction of SeaWiFS ocean color data in the Southern Hemisphere. Applied Mathematics and Computation, 141(2–3):241–259, September 5, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Trione:2004:MPD

[Tri04] Susana Elena Trione. On the multiplicative products of the n-dimensional distributional Hankel transforms. Applied Mathematics and Computation, 154 (1):211–217, June 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Troynikov:2004:UWQ

[Tro04] V. S. Troynikov. Use of weak quasi-solutions of the Fredholm first-kind equation in problems with scarce data. Applied Mathematics and Computation, 150 (3):855–863, March 17, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Turbatu:2001:SBL

[TS01] Stelian Turbatu and H. M.

Srivastava. Some boundarylayer growth problems associated with a flat plate with suction or injection. Applied Mathematics and Computation, 122(3):365– 371, August 15, 2001. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/106/26/
27/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/106/26/27/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000515.

Tsitouras:2002:HAH

[TS02] Ch. Tsitouras and T. E.

Simos. High algebraic, high phase-lag order embedded Numerov-type methods for oscillatory problems. Applied Mathematics and Computation, 131(1):201–211, September 10, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tseng:2003:MPK

[Tse03]  Yuh-Min Tseng.  Multi-

party key agreement protocols with cheater identification. Applied Mathematics and Computation, 145(2–3): 551–559, December 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tsitouras:2001:DHP

[Tsi01]  Ch. Tsitouras.  Dissipa-

tive high phase-lag order methods. Applied Mathematics and Computation, 117(1):35–43, January 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        9/12/92/20/23/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/92/20/
                                        23/article.pdf;   http:
                                        //www.sciencedirect.com/
                                        science/article/pii/S0096300399001526.

Tsitouras:2003:FET

[Tsi03] Ch. Tsitouras. Families of explicit two-step methods for integration of problems with oscillating solutions. Applied Mathematics and Computation, 135(1): 169–178, February 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tunc:2004:ABS

[TT04] Cemil Tunç and Ercan Tunç. On the asymptotic behaviour of solutions of certain non-autonomous differential equations. Applied Mathematics and Computation, 151(2):363–378, April 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tunc:2004:IRC

[Tun04a] Cemil Tunç. An instability result for certain system of sixth order differential equations. Applied Mathematics and Computation, 157(2): 477–481, October 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tunc:2004:NSB

[Tun04b] Cemil Tunç. A note on the stability and boundedness results of solutions of certain fourth order differential equations. Applied Mathematics and Computation, 155(3):837–843, August 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tunc:2004:SAB

[Tun04c] Cemil Tunç. A study of the asymptotic behaviour of solutions of certain non-autonomous differential equations of the fifth order. Applied Mathematics and Computation, 154 (1):103–113, June 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tagliani:2003:NILb

[TV03] Aldo Tagliani and Yurayh Velásquez. Numerical inversion of the Laplace transform via fractional moments. Applied Mathematics and Computation, 143(1): 99–107, October 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tagliani:2004:ILT

[TV04] Aldo Tagliani and Yurayh Velásquez. Inverse Laplace transform for heavy-tailed distributions. Applied Mathematics and Computation, 150(2):337–345, March 8, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Thome:2003:GIB

[TW03] Néstor Thome and Yimin Wei. Generalized inverses and a block-rank equation. Applied Mathematics and Computation, 141(2–3):471– 476, September 5, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tang:2004:GVP

[TY04] Shougao Tang and Aimin Yu. Generalized variational principle on nonlinear theory of naturally curved and twisted beams. Applied Mathematics and Computation, 153(1):275– 288, May 25, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tang:2004:SCE

[TZ04] Yanbin Tang and Li Zhou.

A sufficient condition for the existence of periodic solution for a reaction diffusion equation with infinite delay. Applied Mathematics and Computation, 148(2): 453–460, January 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ugur:2003:ACA

[UK03] Tamer Ugur and Abdullah Kopuzlu. Application of computer algebra to Jones polynomials. Applied Mathematics and Computation, 140(2–3):361–366, August 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ugur:2003:WTK

[UKŞ03] Tamer Ugur, Abdullah Kopuzlu, and Hakan Şimşek. A work on torus knots K(p, q). Applied Mathematics and Computation, 141(2–3):395– 400, September 5, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Udwadia:2001:ESA

[UvB01] Firdaus E. Udwadia and Hubertus F. von Bremen. An efficient and stable approach for computation of Lyapunov characteristic exponents of continuous dynamical systems. Applied Mathematics and Computation, 121(2–3):219–259, June 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/105/25/
30/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/105/25/30/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002921.

Udwadia:2000:NCL

[UvBP00] Firdaus E. Udwadia, Hubertus F. von Bremen, and Wlodek Proskurowski. A note on the computation of the largest p LCEs of discrete dynamical systems. Applied Mathematics and Computation, 114(2–3):205–214, September 11, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/89/22/27/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/89/22/
27/article.pdf.

vanAubel:2003:APN

[vAG03] Andrea van Aubel and Wolfgang Gawronski. Analytic properties of noncentral distributions. Applied Mathematics and Computation, 141(1):3–12, August 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Vajravelu:2001:VFN

[Vaj01]  K. Vajravelu.  Viscous

flow over a nonlinearly stretching sheet. Applied Mathematics and Computation, 124(3):281–288, December 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/113/32/27/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S009630030000062X.

vonBremen:2002:CED

[vBU02] Hubertus F. von Bremen and Firdaus E. Udwadia. Computational explorations into the dynamics of rings of coupled oscillators. Applied Mathematics and Computation, 129(1):55–85, June 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Verma:2004:AFG

[VCD04] A. K. Verma, Sanjay Chandra, and B. K. Dhindaw.

An alternative fixed grid method for solution of the classical one-phase Stefan problem. Applied Mathematics and Computation, 158 (2):573–584, November 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

VanHuylenbroeck:2001:RMO

[VCV01] Guido Van Huylenbroeck, Eva María Ureña Campos, and Isabel Vanslembrouck. A (recursive) multiple objective approach to analyse changes in the utility function of farmers due to policy reforms. Applied Mathematics and Computation, 122(3):283–299, August 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/106/26/
22/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/106/26/22/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S009630030000028X.

Venutelli:2003:FSP

[Ven03]  Maurizio Venutelli.  A

fractional-step Padé–Galerkin model for dam-break flow simulation. Applied Mathematics and Computation, 134(1):93–107, January 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Veress:2000:GAS

[Ver00] Laszlo Antal Veress. Group actions on sets and automata theory. Applied Mathematics and Computation, 113(2–3):289–304, July 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/87/23/30/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/87/23/
30/article.pdf.

Veress:2003:NLS

[Ver03a] Laszlo Antal Veress. Newton’s laws, symmetry, and the three basic interactions of the nature. Applied Mathematics and Computation, 146(1):73–80, December 30, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Veress:2003:PSE

[Ver03b] Laszlo Antal Veress. The principle of symmetry– extremity. Applied Mathematics and Computation, 137(2–3):293–301, May 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Veress:2003:SPI

[Ver03c] Laszlo Antal Veress. Symmetry principles via interactions and symmetryviolations. Applied Mathematics and Computation, 134(2–3):567–575, January 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Veress:2004:NOP

[Ver04] Laszlo Antal Veress. Has the nature its own philosophy? Applied Mathematics and Computation, 150 (2):347–350, March 8, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Viktorova:2004:AMM

[VK04] I. V. Viktorova and M. M. Kostreva. The application of mathematical methods to some problems of technological microbiology. Applied Mathematics and Computation, 150(2):321– 336, March 8, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Vazquez:2003:FCS

[VM03] Luis Vázquez and Rui Vilela Mendes. Fractionally coupled solutions of the diffusion equation. Applied Mathematics and Computation, 141(1):125–130, August 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Voutsadakis:2003:TAN

[Vou03] George Voutsadakis. Threshold agent networks: an approach to modelling and simulation. Applied Mathematics and Computation, 142(2–3):521–543, October 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Valarmathi:2002:ANF

[VR02a] S. Valarmathi and N. Ramanujam. An asymptotic numerical fitted mesh method for singularly perturbed third order ordinary differential equations of reaction–diffusion type. Applied Mathematics and Computation, 132(1):87– 104, October 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Valarmathi:2002:CMS

[VR02b] S. Valarmathi and N. Ramanujam. A computational method for solving boundary value problems for thirdorder singularly perturbed ordinary differential equations. Applied Mathematics and Computation, 129(2– 3):345–373, July 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Valanarasu:2003:AIV

[VR03a] T. Valanarasu and N. Ramanujam. Asymptotic initial value methods for twoparameter singularly perturbed boundary value problems for second order ordinary differential equations. Applied Mathematics and Computation, 137(2–3):549– 570, May 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Valarmathi:2003:CMS

[VR03b] S. Valarmathi and N. Ramanujam. Computational methods for solving twoparameter singularly perturbed boundary value problems for second-order ordinary differential equations. Applied Mathematics and Computation, 136(2–3):415– 441, March 15, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Vajravelu:2004:HFS

[VR04a] K. Vajravelu and D. Rollins. Hydromagnetic flow of a second grade fluid over a stretching sheet. Applied Mathematics and Computation, 148(3):783–791, January 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Valanarasu:2004:AIV

[VR04b] T. Valanarasu and N. Ramanujam. An asymptotic initial value method for boundary value problems for a system of singularly perturbed second order ordinary differential equations. Applied Mathematics and Computation, 147(1):227– 240, January 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Vulanovic:2000:HOM

[Vul00] Relja Vulanović. A higherorder method for stationary shock problems. Applied Mathematics and Computation, 108(2–3):139–152, February 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/79/22/
28/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/79/22/28/article.
pdf.

Wong:2001:ETS

[WA01] Patricia J. Y. Wong and Ravi P. Agarwal. Existence theorems for a system of difference equations with ( n, p )-type conditions. Applied Mathematics and Computation, 123(3):389–407, October 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/110/32/32/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000783.

Wang:2000:IPR

[Wan00a] Alan P. Wang. Inverse problems in radiative transfer. Applied Mathematics and Computation, 116(1–2):39– 48, November 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/91/20/24/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/91/20/
24/article.pdf.

Wang:2000:MCM

[Wan00b] Feng-Sheng Wang. A modified collocation method for solving differential-algebraic equations. Applied Mathematics and Computation, 116(3):257–278, December ??, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/91/23/22/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/91/23/
22/article.pdf.

Wang:2000:PMM

[Wan00c]  Yuan-Ming Wang.  Par-

allel multisplitting methods for a class of systems of weakly nonlinear equations without isotone mapping. Applied Mathematics and Computation, 109(2–3): 135–150, March 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/82/23/
23/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/82/23/23/article.
pdf.

Wang:2001:OAS

[Wan01] Qi-Ru Wang. Oscillation and asymptotics for secondorder half-linear differential equations. Applied Mathematics and Computation, 122(2):253–266, July 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/106/25/
26/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/106/25/26/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000564.

Wan:2002:GRR

[Wan02] Alan T. K. Wan. On generalized ridge regression estimators under collinearity and balanced loss. Applied Mathematics and Computation, 129(2–3):455–467, July 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:2004:STW

[Wan04a] Dingkun Wang. Some topics on weighted Moore– Penrose inverse, weighted least squares and weighted regularized Tikhonov problems. Applied Mathematics and Computation, 157 (1):243–267, September 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:2004:NAS

[Wan04b] Lei Wang. A new algorithm for solving classical Blasius equation. Applied Mathematics and Computation, 157(1):1–9, September 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See comments [Has06].

Wang:2004:NMC

[Wan04c] Lei Wang. A novel method for a class of nonlinear singular perturbation problems. Applied Mathematics and Computation, 156(3):847– 856, September 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:2004:TVS

[Wan04d] Shiuh-Jeng Wang. Threshold verification scheme to a valid-signature using identity only on specialized approval. Applied Mathematics and Computation, 152 (2):373–383, May 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:2004:OHO

[Wan04e] Xiaoping Wang. Oscillation for higher order nonlinear delay differential equations. Applied Mathematics and Computation, 157(1):287– 294, September 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2000:DMA

[Waz00a] Abdul-Majid Wazwaz. The decomposition method applied to systems of partial differential equations and to the reaction-diffusion Brusselator model. Applied Mathematics and Computation, 110(2–3):251–264, April 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/83/23/
30/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/83/23/30/article.
pdf.

Wazwaz:2000:NAC

[Waz00b] Abdul-Majid Wazwaz. A new algorithm for calculating Adomian polynomials for nonlinear operators. Applied Mathematics and Computation, 111(1):33–51, May 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/85/21/
23/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/85/21/23/article.
pdf.

Wazwaz:2001:BSS

[Waz01a] A. M. Wazwaz. Blow-up for solutions of some linear wave equations with mixed nonlinear boundary conditions. Applied Mathematics and Computation, 123(1):133–140, September 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/110/27/35/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000692.

Wazwaz:2001:ATV

[Waz01b] Abdul-Majid Wazwaz. Analytic treatment for variable coefficient fourth-order parabolic partial differential equations. Applied Mathematics and Computation, 123(2):219–227, September 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/110/29/31/abstract.

html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000709.

Wazwaz:2001:CAS

[Waz01c] Abdul-Majid Wazwaz. A computational approach to soliton solutions of the Kadomtsev–Petviashvili equation. Applied Mathematics and Computation, 123(2):205–217, September 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/110/29/30/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000655.

Wazwaz:2001:ESN

[Waz01d] Abdul-Majid Wazwaz. Exact solutions to nonlinear diffusion equations obtained by the decomposition method. Applied Mathematics and Computation, 123(1):109–122, September 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/110/27/33/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000643.

Wazwaz:2001:MDM

[Waz01e] Abdul-Majid Wazwaz. The modified decomposition method applied to unsteady flow of gas through a porous medium. Applied Mathematics and Computation, 118(2– 3):123–132, March 9, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.nl/gej-ng/
10/9/12/96/25/20/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/96/25/
20/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S009630039900209X.

Wazwaz:2001:NAS

[Waz01f] Abdul-Majid Wazwaz. A new algorithm for solving differential equations of Lane–Emden type. Applied Mathematics and Computation, 118(2–3):287–310, March 9, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/96/25/30/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/96/25/
30/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399002234.

Wazwaz:2001:NSS

[Waz01g] Abdul-Majid Wazwaz. The numerical solution of sixthorder boundary value problems by the modified decomposition method. Applied Mathematics and Computation, 118(2–3):311–325, March 9, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/96/25/31/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/96/25/
31/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399002246.

Wazwaz:2001:RAS

[Waz01h] Abdul-Majid Wazwaz. A reliable algorithm for solving boundary value problems for higher-order integrodifferential equations. Applied Mathematics and Computation, 118(2–3):327–342, March 9, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/96/25/32/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/96/25/
32/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399002258.

Wazwaz:2002:CNS

[Waz02a] Abdul-Majid Wazwaz. Compact and noncompact structures for a variant of KdV equation in higher dimensions. Applied Mathematics and Computation, 132(1): 29–45, October 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2002:CSF

[Waz02b] Abdul-Majid Wazwaz. Compactons structures for fifthorder KdV like equations in higher dimensions. Applied Mathematics and Computation, 130(2–3):425–440, August 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2002:ESV

[Waz02c] Abdul-Majid Wazwaz. Exact solutions for variable coefficients fourth-order parabolic partial differential equations in higherdimensional spaces. Applied Mathematics and Computation, 130(2–3):415–424, August 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2002:GCS

[Waz02d] Abdul-Majid Wazwaz. General compactons solutions for the focusing branch of the nonlinear dispersive K(n, n) equations in higherdimensional spaces. Applied Mathematics and Computation, 133(2–3):213–227, December 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2002:GSS

[Waz02e] Abdul-Majid Wazwaz. General solutions with solitary patterns for the defocusing branch of the nonlinear dispersive K(n, n) equations in higher dimensional spaces. Applied Mathematics and Computation, 133(2–3): 229–244, December 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2002:NMS

[Waz02f] Abdul-Majid Wazwaz. A new method for solving singular initial value problems in the second-order ordinary differential equations. Applied Mathematics and Computation, 128 (1):45–57, May 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2002:RTM

[Waz02g] Abdul-Majid Wazwaz. A reliable treatment for mixed Volterra–Fredholm integral equations. Applied Mathematics and Computation, 127(2–3):405–414, April 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.com/
gej-ng/10/9/12/123/27/
47/abstract.html;   http:
//www.sciencedirect.com/
science/article/pii/S0096300301000200.

Wazwaz:2002:SPS

[Waz02h] Abdul-Majid Wazwaz. Solitary patterns solutions having infinite slopes or cusps for fifth-order KdV like equations in higher dimensions. Applied Mathematics and Computation, 131 (1):181–194, September 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2003:CSP

[Waz03a]  A. M. Wazwaz.  Com-

pactons and solitary patterns structures for variants of the KdV and the KP equations. Applied Mathematics and Computation, 139(1):37–54, July 1, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2003:AST

[Waz03b] Abdul-Majid Wazwaz. An analytic study on the thirdorder dispersive partial differential equations. Applied Mathematics and Computation, 142(2–3):511–520, October 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2003:CNS

[Waz03c] Abdul-Majid Wazwaz. Compact and noncompact structures formed by nonlinear equations with positive and negative exponents. Applied Mathematics and Computation, 146(1):1–25, December 30, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2003:CSN

[Waz03d] Abdul-Majid Wazwaz. Compacton solutions and nonlinear dispersion. Applied Mathematics and Computation, 142(2–3):495–509, October 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2003:CSK

[Waz03e] Abdul-Majid Wazwaz. Compacton solutions of the Kawahara-type nonlinear dispersive equation. Applied Mathematics and Computation, 145(1):133–150, December 20, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2003:CCN

[Waz03f] Abdul-Majid Wazwaz. A construction of compact and noncompact solutions for nonlinear dispersive equations of even order. Applied Mathematics and Computation, 135(2–3):411–424, March 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2003:EON

[Waz03g] Abdul-Majid Wazwaz. The effect of the order of nonlinear dispersive equation on the compact and noncompact solutions. Applied Mathematics and Computation, 138(2–3):309–319, June 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2003:ENT

[Waz03h] Abdul-Majid Wazwaz. The existence of noise terms for systems of inhomogeneous differential and integral equations. Applied Mathematics and Computation, 146(1):81–92, December 30, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2003:SNE

[Waz03i] Abdul-Majid Wazwaz. Several new exact solutions for a fast diffusion equation by the differential constraints of the linear determining equations. Applied Mathematics and Computation, 145(2–3): 525–540, December 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2003:SCN

[Waz03j] Abdul-Majid Wazwaz. Solutions of compact and noncompact structures for nonlinear Klein–Gordontype equation. Applied Mathematics and Computation, 134(2–3):487–500, January 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2003:SND

[Waz03k] Abdul-Majid Wazwaz. A study on nonlinear dispersive partial differential equations of compact and noncompact solutions. Applied Mathematics and Computation, 135(2–3):399–409, March 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2004:ASC

[Waz04a] Abdul-Majid Wazwaz. An analytic study of compacton solutions for variants of Kuramoto–Sivashinsky equation. Applied Mathematics and Computation, 148(2):571–585, January 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2004:CNS

[Waz04b] Abdul-Majid Wazwaz. Compact and noncompact solutions for nonlinear dispersive variants of the generalized KdV equation. Applied Mathematics and Computation, 159(2):577–588, December 6, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2004:CSV

[Waz04c] Abdul-Majid Wazwaz. Compact structures for variants of the generalized KdV and the generalized KP equations. Applied Mathematics and Computation, 149(1): 103–117, February 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2004:CSH

[Waz04d] Abdul-Majid Wazwaz. Compacton solutions of higher order nonlinear dispersive KdV-like equations. Applied Mathematics and Computation, 147(2):449–460, January 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2004:CSS

[Waz04e] Abdul-Majid Wazwaz. Compactons structures for specific nonlinear dispersive equations. Applied Mathematics and Computation, 150(2):399–407, March 8, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2004:DVK

[Waz04f] Abdul-Majid Wazwaz. Distinct variants of the KdV equation with compact and noncompact structures. Applied Mathematics and Computation, 150(2):365–377, March 8, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2004:ENM

[Waz04g] Abdul-Majid Wazwaz. Effect of nonlinearity of the middle term of nonlinear dispersive equations on physical structures. Applied Mathematics and Computation, 159(2):539–558, December 6, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2004:SCM

[Waz04h] Abdul-Majid Wazwaz. The sine–cosine method for obtaining solutions with compact and noncompact structures. Applied Mathematics and Computation, 159(2): 559–576, December 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2004:SCL

[Waz04i] Abdul-Majid Wazwaz. A study on compacton-like solutions for the modified KdV and fifth order KdV-like equations. Applied Mathematics and Computation, 147(2):439–447, January 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2004:TMT

[Waz04j] Abdul-Majid Wazwaz. The tanh method for traveling wave solutions of nonlinear equations. Applied Mathematics and Computation, 154(3):713–723, July 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2004:TCV

[Waz04k] Abdul-Majid Wazwaz. Two classes of variants of the generalized KdV equations with compact and noncompact solutions. Applied Mathematics and Computation, 154(3):835–846, July 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:2001:POH

[WBW01] Y. Wang, D. S. Bernstein, and L. T. Watson. Probability-one homotopy algorithms for solving the coupled Lyapunov equations arising in reduced-order H2/H modeling, estimation, and control. Applied Mathematics and Computation, 123(2):155–185, September 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/110/29/28/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S009630030000059X.

Wei:2001:PLS

[WC01] Yimin Wei and Guoliang Chen. Perturbation of least squares problem in Hilbert spaces. Applied Mathematics and Computation, 121(2– 3):177–183, June 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/105/25/
26/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/105/25/26/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002714.

Wu:2002:AGO

[WC02] Tzong-Mou Wu and Cha’o-Kuang Chen. Application of graphical operation method to the dynamics and relative-motion problems. Applied Mathematics and Computation, 132 (1):205–218, October 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:2004:AMC

[WC04a] Shiuh-Jeng Wang and Chen-Ming Chuang. Anonymous message communications with user hierarchy in a multicast system. Applied Mathematics and Computation, 159(1):95–112, November 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wu:2004:GOV

[WC04b] Tzong-Mou Wu and Cha’o-Kuang Chen. Graphical operation of vectors and matrices. Applied Mathematics and Computation, 149(1): 259–276, February 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wu:2004:KRG

[WC04c] Tzong-Mou Wu and Chao-Kuang Chen. Kinematics representations of graphical operation methods. Applied Mathematics and Computation, 155(2):573–590, August 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wei:2003:IRG

[WD03] Yimin Wei and Dragan S. Djordjevi. On integral representation of the generalized inverse A_{T}, S^{(2)}. Applied Mathematics and Computation, 142(1):189–194, September 20, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wei:2000:DIU

[Wei00a] Yimin Wei. The Drazin inverse of updating of a square matrix with application to perturbation formula. Applied Mathematics and Computation, 108(2–3): 77–83, February 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/79/22/
22/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/79/22/22/article.
pdf.

Wei:2000:RNN

[Wei00b]  Yimin Wei.  Recurrent

neural networks for computing weighted Moore– Penrose inverse. Applied Mathematics and Computation, 116(3):279–287, December ??, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/91/23/23/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/91/23/
23/article.pdf.

Wei:2000:SMSa

[Wei00c] Yimin Wei. Successive matrix squaring algorithm for computing the Drazin inverse. Applied Mathematics and Computation, 108(2–3): 67–75, February 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/79/22/
21/abstract.html;   http:

                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/79/22/21/article.
                                        pdf.

Wei:2001:WMP

[Wei01] Yimin Wei. The weighted Moore–Penrose inverse of modified matrices. Applied Mathematics and Computation, 122(1):1–13, July 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/9/12/106/21/
                                        21/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/9/12/106/21/21/article.
                                        pdf;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300300000072.

Wei:2002:CWD

[Wei02a] Yimin Wei. A characterization for the W -weighted Drazin inverse and a Cramer rule for the W -weighted Drazin inverse solution. Applied Mathematics and Computation, 125(2–3):303–310, January 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        9/12/120/31/39/abstract.
                                        html;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300300001326.

Wei:2002:DIM

[Wei02b] Yimin Wei. The Drazin inverse of a modified matrix. Applied Mathematics and Computation, 125(2–3): 295–301, January 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/9/12/120/31/38/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001314.

Wei:2002:PBD

[Wei02c] Yimin Wei. Perturbation bound of the Drazin inverse. Applied Mathematics and Computation, 125(2–3): 231–244, January 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/9/12/120/31/33/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001260.

Wei:2003:IRW

[Wei03a] Yimin Wei. Integral representation of the W -weighted Drazin inverse. Applied Mathematics and Computation, 144(1):3–10, November 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wei:2003:RAW

[Wei03b] Yimin Wei. The representation and approximation for the weighted Moore– Penrose inverse in Hilbert space. Applied Mathematics and Computation, 136(2– 3):475–486, March 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wei:2004:CFO

[Wei04] Zhongli Wei. A class of fourth order singular boundary value problems. Applied Mathematics and Computation, 153(3):865–884, June 14, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2001:NMA

[WES01] Abdul-Majid Wazwaz and Salah M. El-Sayed. A new modification of the Adomian decomposition method for linear and nonlinear operators. Applied Mathematics and Computation, 122(3): 393–405, August 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/9/12/106/26/
                                        30/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/9/12/106/26/30/article.
                                        pdf;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300300000606.

Wang:2004:PIP

[WFC04] QuanFang Wang, Dexing Feng, and Daizhan Cheng. Parameters identification problems for Hopfield-type neural network equations.

Applied Mathematics and Computation, 152(2):535– 550, May 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wendi:2003:IDM

[WFT03] Wang Wendi, P. Fergola, and C. Tenneriello. Innovation diffusion model in patch environment. Applied Mathematics and Computation, 134(1):51–67, January 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:2000:OCH

[WG00] Peiguang Wang and Weigao Ge. Oscillation of a class of hyperbolic equations. Applied Mathematics and Computation, 113(1):101–110, July 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/87/21/28/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/87/21/
28/article.pdf.

Wang:2004:OCHa

[WG04a] Peiguang Wang and Weigao Ge. Oscillation of a class of higher order functional differential equations with damped term. Applied Mathematics and Computation, 148(2):351–358, January 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2004:ASF

[WG04b] Abdul-Majid Wazwaz and Alice Gorguis. An analytic study of Fisher’s equation by using Adomian decomposition method. Applied Mathematics and Computation, 154(3):609–620, July 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:2004:ESH

[WG04c] Abdul-Majid Wazwaz and Alice Gorguis. Exact solutions for heat-like and wavelike equations with variable coefficients. Applied Mathematics and Computation, 149(1):15–29, February 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:2000:TGI

[WH00] Chih-Hung Wang and Tzonelih Hwang. (t, m) threshold and generalized ID-based conference key distribution system. Applied Mathematics and Computation, 112(2– 3):181–191, June 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/86/23/

23/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/86/23/23/article.
pdf.

Wu:2002:IBM

[WH02] Tzong-Sun Wu and Chien-Lung Hsu. ID-based multisignatures with distinguished signing authorities for sequential and broadcasting architectures. Applied Mathematics and Computation, 131(2–3):349–356, September 25, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wong:2002:CSR

[WHG02] S. M. Wong, Y. C. Hon, and M. A. Golberg. Compactly supported radial basis functions for shallow water equations. Applied Mathematics and Computation, 127(1):79–101, March 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/123/28/31/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300301000066.

Wu:2004:MSS

[WHM04] Shujin Wu, Dong Han, and Xianzhang Meng. p-Moment stability of stochastic differential equations with jumps. Applied Mathematics and Computation, 152(2):505– 519, May 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wu:2004:EPG

[WHT04] Jong-Wuu Wu, Wen-Liang Hung, and Chih-Hui Tsai. Estimation of parameters of the Gompertz distribution using the least squares method. Applied Mathematics and Computation, 158(1):133–147, October 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:2000:SSL

[WJ00] Hailong Wang and Jianjing Jiang. Solution of the system of linear algebraic equations by decreasing dimension. Applied Mathematics and Computation, 109 (1):51–57, March 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/20/82/21/
                                        24/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/20/82/21/24/article.

pdf. See comments [Zha02a].

Weigao:2004:NET

[WJ04] Ge Weigao and Ren Jingli. New existence theorems of positive solutions for Sturm– Liouville boundary value problems. Applied Mathematics and Computation, 148(3):631–644, January 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

White:2002:ECT

[WL02] Luther White and Yiqi Luo.

Estimation of carbon transfer coefficients using Duke Forest free-air CO _{2} enrichment data. Applied Mathematics and Computation, 130(1):101–120, July 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:2003:EGS

[WL03] Lin-Lin Wang and Wan-Tong Li. Existence and global stability of positive periodic solutions of a predator–prey system with delays. Applied Mathematics and Computation, 146(1): 167–185, December 30, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:2004:ABS

[WL04a] Xiaoping Wang and Liusheng Liao. Asymptotic behavior of solutions of delay logistic differential equation with negative instantaneously terms. Applied Mathematics and Computation, 153(1):69–74, May 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:2004:OEO

[WL04b] Xiaoping Wang and Liusheng Liao. Oscillation for evenorder delay difference equations with unstable type. Applied Mathematics and Computation, 153(1):289– 299, May 25, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wu:2004:RKA

[WL04c] Tzong-Sun Wu and Han-Yu Lin. Robust key authentication scheme resistant to public key substitution attacks. Applied Mathematics and Computation, 157(3): 825–833, October 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wu:2004:PMM

[WL04d] Xinyuan Wu and Liangsheng Luo. Point monotonicity of maps and its applications in numerical computation. Applied Mathematics and Computation, 158(3):827– 834, November 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wei:2003:SRS

[WLW03] Yimin Wei, Xiezhang Li, and Hebing Wu. Subproper and regular splittings for restricted rectangular linear system. Applied Mathematics and Computation, 136(2– 3):535–547, March 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:2000:RBR

[WLX00] Guoli Wang, Y. F. Li, and W. L. Xu. Regularizationbased recovery scheme for inverse dynamics of highspeed flexible beams. Applied Mathematics and Computation, 115(2–3):161–175, October 27, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/90/23/27/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/90/23/
27/article.pdf.

Wei:2004:WTF

[WN04] Yimin Wei and Michael Ng.

Weighted Tikhonov filter matrices for ill-posed problems. Applied Mathematics and Computation, 149(2): 411–422, February 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wolfe:2002:BZA

[Wol02] M. A. Wolfe. On bounding zeros of analytic functions. Applied Mathematics and Computation, 131(2–3):383– 396, September 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wolfe:2004:FZC

[Wol04] M. A. Wolfe. On first zero crossing points. Applied Mathematics and Computation, 150(2):467–479, March 8, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wei:2003:RAD

[WQ03] Yimin Wei and Sanzheng Qiao. The representation and approximation of the Drazin inverse of a linear operator in Hilbert space. Applied Mathematics and Computation, 138(1):77–89, June 1, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:2004:SCM

[WQ04] Guorong Wang and Sanzheng Qiao. Solving constrained matrix equations and Cramer rule. Applied Mathematics and Computation, 159(2): 333–340, December 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:2001:RMS

[WS01] Haiyan Wang and Yongzhong Song. Regularization methods for solving differentialalgebraic equations. Applied Mathematics and Computation, 119(2–3):283–296, April 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/99/25/32/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/99/25/
32/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399002507.

Wu:2003:IRF

[WSX03] Xinyuan Wu, Zuhe Shen, and Jianlin Xia. An improved regula falsi method with quadratic convergence of both diameter and point for enclosing simple zeros of nonlinear equations. Applied Mathematics and Computation, 144(2–3):381–388, December 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wu:2000:NCN

[Wu00]  Xin-Yuan Wu.  A new

continuation Newton-like method and its deformation. Applied Mathematics and Computation, 112 (1):75–78, June 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/86/21/
27/abstract.html;   http:
//www.elsevier.nl/gej-

                                        ng/29/17/20/86/21/27/article.
                                        pdf.

Wu:2003:CRK

[Wu03] Xinyuan Wu. A class of Runge–Kutta formulae of order three and four with reduced evaluations of function. Applied Mathematics and Computation, 146(2–3): 417–432, December 31, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wu:2004:IBA

[Wu04a] Gang Wu. An iterative block Arnoldi algorithm with modified approximate eigenvectors for large unsymmetric eigenvalue problems. Applied Mathematics and Computation, 153 (3):611–643, June 14, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wu:2004:TSR

[Wu04b]  Xinyuan Wu.  Two-step

Runge–Kutta methods of order over five with reduced function evaluations. Applied Mathematics and Computation, 155(3):779– 811, August 16, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wunsche:2003:GHP

[Wün03] A. Wünsche. Generalized Hermite polynomials associated with functions of parabolic cylinder. Applied Mathematics and Computation, 141(1):197–213, August 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wen:2000:AOP

[WW00a] Jyh-Horng Wen and Jee-Wey Wang. Approaches to an occupancy problem resulting from MFSK systems with diversity and majority vote. Applied Mathematics and Computation, 107(2–3): 95–101, January 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/72/22/
23/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/72/22/23/article.
pdf.

Wu:2000:CQC

[WW00b] Xinyuan Wu and Hongwei Wu. On a class of quadratic convergence iteration formulae without derivatives. Applied Mathematics and Computation, 107(2–3):77– 80, January 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/72/22/
21/abstract.html;   http:
//www.elsevier.nl/gej-

                                        ng/29/17/20/72/22/21/article.
                                        pdf.

Wei:2001:RAW

[WW01] Yimin Wei and Hebing Wu. The representation and approximation for the weighted Moore–Penrose inverse. Applied Mathematics and Computation, 121 (1):17–28, May 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/9/12/105/21/
                                        22/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/9/12/105/21/22/article.
                                        pdf;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300399002751.

Wei:2002:PAP

[WW02a] Yimin Wei and Guorong Wang. PCR algorithm for parallel computing minimumnorm (T) least-squares (S) solution of inconsistent linear equations. Applied Mathematics and Computation, 133(2–3):547–557, December 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wei:2002:UIS

[WW02b] Yimin Wei and Hebing Wu. On the use of incomplete semiiterative methods for singular systems and applications in Markov chain modeling. Applied Mathematics and Computation, 125(2–3):245–259, January 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/120/31/34/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001272.

Wei:2003:CNP

[WW03a] Yimin Wei and Dingkun Wang. Condition numbers and perturbation of the weighted Moore–Penrose inverse and weighted linear least squares problem. Applied Mathematics and Computation, 145(1):45–58, December 20, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wei:2003:CGI

[WW03b] Yimin Wei and Guorong Wang. On continuity of the generalized inverse A_{T}, S^{(2)}. Applied Mathematics and Computation, 136(2–3): 289–295, March 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wei:2003:RAG

[WW03c] Yimin Wei and Hebing Wu.

The representation and approximation for the generalized inverse A_{T}, S^{(2)}. Applied Mathematics and Computation, 135(2–3):263–276, March 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:2004:OCHb

[WW04] Jing Wang and Ke Wang. Optimal control of harvesting for single population. Applied Mathematics and Computation, 156(1):235– 247, August 25, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wei:2004:NPW

[WWL04] Yimin Wei, Ching-Wah Woo, and Tiangang Lei. A note on the perturbation of the W -weighted Drazin inverse. Applied Mathematics and Computation, 149(2): 423–430, February 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:2004:NIM

[WWQ04] Yushun Wang, Bin Wang, and Mengzhao Qin. Numerical implementation of the multisymplectic Preissman scheme and its equivalent schemes. Applied Mathematics and Computation, 149(2):299–326, February 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wei:2000:SMSb

[WWW00] Yimin Wei, Hebing Wu, and Junyin Wei. Successive matrix squaring algorithm for parallel computing the weighted generalized inverse A_{MN}^{+}. Applied Mathematics and Computation, 116(3):289–296, December ??, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/91/23/24/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/91/23/
24/article.pdf.

Wei:2003:CND

[WWW03] Yimin Wei, Guanglin Wang, and Dingkun Wang. Condition number of Drazin inverse and their condition numbers of singular linear systems. Applied Mathematics and Computation, 146(2–3):455–467, December 31, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wu:2001:TLA

[WX01] Xin-Yuan Wu and Jian-Lin Xia. Two low accuracy methods for stiff systems. Applied Mathematics and Computation, 123(2):141–153, September 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        9/12/110/29/27/abstract.
                                        html;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300300000102.

Wei:2003:CNB

[WX03a] Yimin Wei and Wei Xu. Condition number of Bott– Duffin inverse and their condition numbers. Applied Mathematics and Computation, 142(1):79–97, September 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wu:2003:NVF

[WX03b] Xin-Yuan Wu and Jian-Lin Xia. New vector forms of elemental functions with Taylor series. Applied Mathematics and Computation, 141(2–3): 307–312, September 5, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wu:2004:DZS

[WX04] Hong-Wu Wu and Yuan-Tong Xu. The distribution of zeros of solutions of neutral differential equations. Applied Mathematics and Computation, 156(3):665– 677, September 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wu:2002:NGC

[WXO02] Xinyuan Wu, Jianlin Xia, and Zixiang Ouyang. Note on global convergence of ODE methods for unconstrained optimization. Applied Mathematics and Computation, 125(2–3):311–315, January 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/120/31/40/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001338.

Yuan:2003:CII

[wYjSjS03] Guang wei Yuan, Zhi jun Shen, and Long jun Shen. Convergence of inner iterations scheme of the discrete ordinate method for neutron transport equations. Applied Mathematics and Computation, 138(1):121– 125, June 1, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yuan:2001:USP

[wYjSlZ01] Guang wei Yuan, Long jun Shen, and Yu lin Zhou. Unconditional stability of parallel alternating difference schemes for semilinear parabolic systems. Applied Mathematics and Computation, 117(2–3):267–283, January 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/92/23/31/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/92/23/
31/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399001800.

Wang:2001:DGA

[WZ01a] Guanxiang Wang and Shu Zhu. Dimension of the global attractor for the discretized damped sine-Gordon equation. Applied Mathematics and Computation, 117(2–3):257–265, January 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/92/23/30/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/92/23/
30/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399001794.

White:2001:CEM

[WZ01b] Luther W. White and Musharaff Zaman. Coupling elastic models through interfacial conditions with application to concrete pavement overlays. Applied Mathematics and Computation, 123(2):187–204, September 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.

elsevier.com/gej-ng/10/
9/12/110/29/29/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000618.

White:2002:SIP

[WZ02] Luther W. White and Musharaff Zaman. Stochastic inversion and prediction applied to concrete overlay systems. Applied Mathematics and Computation, 128 (1):59–79, May 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:2003:NSS

[WZ03] Li Wang and Jun Zhang.

A new stabilization strategy for incomplete LU preconditioning of indefinite matrices. Applied Mathematics and Computation, 144(1): 75–87, November 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:2004:ROL

[WZ04a] Guorong Wang and Bing Zheng. The reverse order law for the generalized inverse A_{T,S}^{(2)}. Applied Mathematics and Computation, 157(2): 295–305, October 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:2004:WGI

[WZ04b] Guorong Wang and Bing Zheng. The weighted generalized inverses of a partitioned matrix. Applied Mathematics and Computation, 155(1):221–233, July 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:2004:GON

[WZ04c] Yan-Jun Wang and Ke-Cun Zhang. Global optimization of nonlinear sum of ratios problem. Applied Mathematics and Computation, 158 (2):319–330, November 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wei:2004:FNC

[WZ04d] Yimin Wei and Naimin Zhang. Further note on constraint preconditioning for nonsymmetric indefinite matrices. Applied Mathematics and Computation, 152 (1):43–46, April 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wei:2004:NRA

[WZ04e] Yimin Wei and Naimin Zhang. A note on the representation and approximation of the outer inverse A_{T}, S^{(2)} of a matrix A. Applied Mathematics and Computation, 147(3): 837–841, January 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xiao:2001:ETS

[XC01] Yanni Xiao and Lansun Chen. Effects of toxicants on a stage-structured population growth model. Applied Mathematics and Computation, 123(1):63–73, September 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/110/27/30/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000576.

Xiao:2002:RDP

[XC02a] Yanni Xiao and Lansun Chen. A ratio-dependent predator–prey model with disease in the prey. Applied Mathematics and Computation, 131(2–3):397–414, September 25, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xu:2002:PAS

[XC02b] Rui Xu and M. A. J. Chaplain. Persistence and attractivity in an N-species ratiodependent predator–prey system with distributed time delays. Applied Mathematics and Computation, 131(1): 59–80, September 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xu:2002:PGS

[XC02c] Rui Xu and M. A. J. Chaplain. Persistence and global stability in a delayed predator–prey system with Michaelis–Menten type functional response. Applied Mathematics and Computation, 130(2–3):441–455, August 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xue:2002:EGB

[XC02d] Yifeng Xue and Guoliang Chen. The expression of the generalized Bott–Duffin inverse and its perturbation theory. Applied Mathematics and Computation, 132(2–3):437–444, November 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xue:2004:SEC

[XC04] Yifeng Xue and Guoliang Chen. Some equivalent conditions of stable perturbation of operators in Hilbert spaces. Applied Mathematics and Computation, 147(3): 765–772, January 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xu:2002:GAS

[XCC02] Rui Xu, M. A. J. Chaplain, and Lansun Chen.

Global asymptotic stability in n-species nonautonomous Lotka–Volterra competitive systems with infinite delays. Applied Mathematics and Computation, 130(2–3): 295–309, August 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xia:2004:EGA

[XCCC04] Yonghui Xia, Fengde Chen, Anping Chen, and Jinde Cao. Existence and global attractivity of an almost periodic ecological model. Applied Mathematics and Computation, 157(2):449– 475, October 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xu:2004:PSL

[XCD04a] R. Xu, M. A. J. Chaplain, and F. A. Davidson. Periodic solution of a Lotka–Volterra predator– prey model with dispersion and time delays. Applied Mathematics and Computation, 148(2):537–560, January 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xu:2004:GSL

[XCD04b] Rui Xu, M. A. J. Chaplain, and F. A. Davidson. Global stability of a Lotka– Volterra type predator–prey model with stage structure and time delay. Applied Mathematics and Computation, 159(3):863–880, December 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xu:2004:PSP

[XCD04c] Rui Xu, M. A. J. Chaplain, and F. A. Davidson. Periodic solutions of a predator–prey model with stage structure for predator. Applied Mathematics and Computation, 154 (3):847–870, July 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xu:2004:PGS

[XCD04d] Rui Xu, M. A. J. Chaplain, and F. A. Davidson. Persistence and global stability of a ratio-dependent predator– prey model with stage structure. Applied Mathematics and Computation, 158 (3):729–744, November 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xu:2004:PPD

[XCD04e] Rui Xu, M. A. J. Chaplain, and F. A. Davidson. Persistence and periodicity of a delayed ratio-dependent predator–prey model with stage structure and prey dispersal. Applied Mathematics and Computation, 159(3): 823–846, December 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xu:2004:PSS

[XCD04f] Rui Xu, M. A. J. Chaplain, and F. A. Davidson. Persistence and stability of a stage-structured predatorprey model with time delays. Applied Mathematics and Computation, 150(1): 259–277, February 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xu:2003:PES

[XD03] Xingye Xu and Lokenath Debnath. Positive entire solutions for a class of singular nonlinear polyharmonic equations on R^{2}. Applied Mathematics and Computation, 140(2–3):317–328, August 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xenophontos:2003:NCR

[Xen03] Christos Xenophontos. A note on the convergence rate of the finite element method for singularly perturbed problems using the Shishkin mesh. Applied Mathematics and Computation, 142(2–3):545–559, October 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xiong:2004:CDS

[XFL04] Xiang-Tuan Xiong, Chu-Li Fu, and Hong-Fang Li. Central difference schemes in time and error estimate on a non-standard inverse heat conduction problem. Applied Mathematics and Computation, 157(1):77–91, September 27, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xing:2004:NAS

[XH04] Yepeng Xing and Maoan Han. A new approach to stability of impulsive functional differential equations. Applied Mathematics and Computation, 151(3):835– 847, April 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xiaojing:2003:NHI

[Xia03] Yang Xiaojing. A note on Hilbert’s inequality. Applied Mathematics and Computation, 137(2–3):589–593, May 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xu:2004:OTE

[XJM04] Zhiting Xu, Baoguo Jia, and Dongkui Ma. Oscillation theorems for elliptic equations with damping. Applied Mathematics and Computation, 156 (1):93–106, August 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xu:2003:TLS

[XL03] Xuejun Xu and S. H. Lui.

A two-level Schwarz method for a finite element approximation of a nonlinear biharmonic equation. Applied Mathematics and Computation, 145(2–3):683–700, December 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xian:2004:SRT

[XL04a] Jun Xian and Wei Lin.

Sampling and reconstruction in time-warped spaces and their applications. Applied Mathematics and Computation, 157(1):153–173, September 27, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xu:2004:SLM

[XL04b] Yang Xu and MingZhu Liu. H-stability of linear θ-method with general variable stepsize for system of pantograph equations with two delay terms. Applied Mathematics and Computation, 156(3):817– 829, September 15, 2004.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xu:2004:SRK

[XL04c] Yang Xu and MingZhu Liu. H-stability of Runge–Kutta methods with general variable stepsize for pantograph equation. Applied Mathematics and Computation, 148(3):881–892, January 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xu:2003:ESS

[XLLL03] Zong-Ben Xu, Kwong-Sak Leung, Yong Liang, and Yee Leung. Efficiency speedup strategies for evolutionary computation: fundamentals and fast-GAs. Applied Mathematics and Computation, 142(2–3):341–388, October 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xu:2002:CMI

[Xu02] Xuejun Xu. On V -cycle multigrid implementation for plate elements. Applied Mathematics and Computation, 125(2–3):141–153, January 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        9/12/120/31/27/abstract.
                                        html;   http://www.sciencedirect.

com/science/article/pii/
S0096300300001077.

Xu:2004:EAS

[Xu04] H. Xu. An explicit analytic solution for free convection about a vertical flat plate embedded in a porous medium by means of homotopy analysis method. Applied Mathematics and Computation, 158(2):433– 443, November 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xu:2003:OCK

[XX03] Zhi-Ting Xu and Hong-Yan Xing. Oscillation criteria of Kamenev-type for PDE with p–Laplacian. Applied Mathematics and Computation, 145(2–3):735–745, December 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xie:2004:CTA

[XY04] Qi Xie and Xiu Yuan Yu.

Cryptanalysis of Tseng et al.’s authenticated encryption schemes. Applied Mathematics and Computation, 158(1):1–5, October 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xu:2002:PES

[XYD02] Xingye Xu, Bicheng Yang, and Lokenath Debnath. Positive entire solutions of nonlinear polyharmonic equations in R^{2}. Applied Mathematics and Computation, 126(2–3):377–388, March 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/122/30/42/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001697.

Xu:2003:NST

[XZ03] H. Xu and C. Zhang. Numerical solution of a transformed parabolic equation. Applied Mathematics and Computation, 139(2–3):535– 554, July 15, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yassen:2003:ADG

[YA03] M. T. Yassen and H. N. Agiza. Analysis of a duopoly game with delayed bounded rationality. Applied Mathematics and Computation, 138(2–3):387–402, June 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yalcinbas:2002:TPS

[Yal02] Salih Yalçinbaş. Taylor polynomial solutions of nonlinear Volterra–Fredholm integral equations. Applied Mathematics and Computation, 127(2–3):195–206, April 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/123/27/30/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S009630030000165X.

Yamac:2005:EAS

[Yam05a] K. Yamaç. Erratum to “An algorithm for stability of discrete-time interval matrices”. Applied Mathematics and Computation, 168(2): 1469, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300304006836.

See [YB03].

Yamac:2005:ESD

[Yam05b] K. Yamac. Erratum to “On stability of discrete-time interval matrices” [Applied Mathematics and Computation 152 (1) (2004) 163– 167]. Applied Mathematics and Computation, 170(1): 731–732, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [YB04].

Yang:2000:APS

[Yan00a] Dan-Ping Yang. Analysis of a parallel Schwarz algorithm for elliptic problems. Applied Mathematics and Computation, 114(1):75–93, August 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/89/20/27/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/89/20/
27/article.pdf.

Yang:2000:ALS

[Yan00b] Suh-Yuh Yang. Analysis of a least squares finite element method for the circular arch problem. Applied Mathematics and Computation, 114(2–3):263–278, September 11, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/89/22/32/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/89/22/
32/article.pdf.

Yang:2001:MDT

[Yan01a] Hyun Mo Yang. Modeling directly transmitted infections in a routinely vaccinated population — the force of infection described by a Volterra integral equation. Applied Mathematics and Computation, 122 (1):27–58, July 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/106/21/
23/abstract.html;   http:

//www.elsevier.nl/gej-
ng/10/9/12/106/21/23/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000114.

Yang:2001:EGS

[Yan01b] Xiao-Song Yang. On the existence of generalized synchronizor in unidirectionally coupled systems. Applied Mathematics and Computation, 122(1):71–79, July 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/106/21/
25/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/106/21/25/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000151.

Yang:2001:ERP

[Yan01c] Xiaojing Yang. An existence result of periodic solutions. Applied Mathematics and Computation, 123(3): 413–419, October 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/9/12/110/32/34/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000904.

Yang:2002:PDD

[Yan02a] Danping Yang. A parallel domain decomposition algorithm of mixed element equation for second-order elliptic Dirichlet boundary value problem. Applied Mathematics and Computation, 129(2–3):375–389, July 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2002:BSS

[Yan02b] Xiaojing Yang. Boundedness of solutions for semilinear Duffing equations. Applied Mathematics and Computation, 132(2–3):423–436, November 10, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2002:EUR

[Yan02c] Xiaojing Yang. Existence and uniqueness results for periodic solution of nonlinear differential equations. Applied Mathematics and Computation, 130(2–3):213– 223, August 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2002:EPS

[Yan02d] Xiaojing Yang. Existence of periodic solution for nonlinear differential equations. Applied Mathematics and Computation, 131(2–3):433– 438, September 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2002:NCS

[Yan02e] Xiaojing Yang. Nonoscillation criteria for second-order nonlinear differential equations. Applied Mathematics and Computation, 131 (1):125–131, September 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2002:PSN

[Yan02f] Xiaojing Yang. Positive solutions for nonlinear singular boundary value problems. Applied Mathematics and Computation, 130(2–3): 225–234, August 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2002:SER

[Yan02g] Xiaojing Yang. Some existence results for differential equations of second order. Applied Mathematics and Computation, 130(2–3): 361–368, August 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2002:DDC

[Yan02h] Xin-She Yang. Densitydriven compaction and temperature evolution in porous media. Applied Mathematics and Computation, 126(2–3):243–254, March 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/122/30/34/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001545.

Yang:2003:IOC

[Yan03a] Qigui Yang. Interval oscillation criteria for a forced second order nonlinear ordinary differential equations with oscillatory potential. Applied Mathematics and Computation, 135(1):49–64, February 15, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2003:ASN

[Yan03b] Xiaojing Yang. Averages for solutions of nonlinear operators. Applied Mathematics and Computation, 145(2–3): 255–262, December 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2003:BSC

[Yan03c] Xiaojing Yang. Boundedness of solutions for a class of reversible systems. Applied Mathematics and Computation, 145(2–3):755–764, December 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2003:BSN

[Yan03d] Xiaojing Yang. Boundedness of solutions for nonlinear oscillations. Applied Mathematics and Computation, 144(2–3):187–198, December 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2003:BSS

[Yan03e] Xiaojing Yang. Boundedness of solutions of some nonlinear differential equations. Applied Mathematics and Computation, 136(2– 3):495–503, March 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2003:BPS

[Yan03f] Xiaojing Yang. Boundedness problem for solutions of semilinear asymmetric equations. Applied Mathematics and Computation, 145(2–3): 765–775, December 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2003:EUR

[Yan03g] Xiaojing Yang. An existence and uniqueness result for periodic solution of 2n-order differential equations. Applied Mathematics and Computation, 137 (1):101–108, May 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2003:ESO

[Yan03h] Xiaojing Yang. Existence of solutions for 2n-order boundary value problem. Applied Mathematics and Computation, 137(1):77–87, May 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2003:ERS

[Yan03i] Xiaojing Yang. Existence results for Sturm–Liouville boundary value problem and rapid nonlinear oscillations. Applied Mathematics and Computation, 145(2–3):241– 253, December 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2003:FOO

[Yan03j] Xiaojing Yang. Forced oscillation of nth-order nonlinear differential equations. Applied Mathematics and Computation, 134(2–3):301–305, January 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2003:GFP

[Yan03k]  Xiaojing Yang.  Green’s

function and positive solutions for higher-order ODE. Applied Mathematics and Computation, 136(2–3):379– 393, March 15, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2003:MLU

[Yan03l] Xiaojing Yang. The method of lower and upper solutions for systems of boundary value problems. Applied Mathematics and Computation, 144(1):169–172, November 20, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2003:NRA

[Yan03m]  Xiaojing Yang.  Nonlin-

ear resonance in asymmetric oscillations. Applied Mathematics and Computation, 142(2–3):255–270, October 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2003:NPH

[Yan03n] Xiaojing Yang. Nonresonance problem for higherorder systems. Applied Mathematics and Computation, 135(2–3):505–515, March 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2003:NHI

[Yan03o] Xiaojing Yang. A note on Hölder inequality. Applied Mathematics and Computation, 134(2–3):319–322, January 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2003:ILT

[Yan03p] Xiaojing Yang. On inequalities of Lyapunov type. Applied Mathematics and Computation, 134(2–3):293–300, January 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2003:LTI

[Yan03q] Xiaojing Yang. On Liapunovtype inequality for certain higher-order differential equations. Applied Mathematics and Computation, 134(2–3):307–317, January 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2003:OCN

[Yan03r] Xiaojing Yang. Oscillation criteria for nonlinear differential equations with damping. Applied Mathematics and Computation, 136(2–3): 549–557, March 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2003:OCG

[Yan03s] Xiaojing Yang. An oscillation criterion for generalized Liénard equations. Applied Mathematics and Computation, 145(2–3):233–239, December 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2003:RHI

[Yan03t] Xiaojing Yang. Refinement of Hölder inequality and application to Ostrowski inequality. Applied Mathematics and Computation, 138(2– 3):455–461, June 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2003:STP

[Yan03u] Xiaojing Yang. Sturm type problems for singular p-Laplacian boundary value problems. Applied Mathematics and Computation, 136(1):181–193, March 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2003:TPB

[Yan03v] Xiaojing Yang. Two-point boundary value problem of 2mth order differential equations. Applied Mathematics and Computation, 138 (1):11–19, June 1, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2003:ULS

[Yan03w] Xiaojing Yang. Upper and lower solutions for periodic problems. Applied Mathematics and Computation, 137(2–3):413–422, May 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2004:LEE

[Yan04a] Shou-Yuan Yang. Local error estimation for sampling problems. Applied Mathematics and Computation, 158(2):561–572, November 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2004:BSR

[Yan04b] Xiaojing Yang. Boundedness of solutions for reversible systems. Applied Mathematics and Computation, 152 (1):111–126, April 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2004:BSS

[Yan04c] Xiaojing Yang. Boundedness of solutions for sublinear reversible systems. Applied Mathematics and Computation, 158(2):389–396, November 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2004:BSN

[Yan04d] Xiaojing Yang. Boundedness of solutions of nonlinear p -Laplacian. Applied Mathematics and Computation, 158(2):397–417, November 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2004:EPS

[Yan04e] Xiaojing Yang. Existence of periodic solutions for quasilinear differential equations. Applied Mathematics and Computation, 153 (1):231–237, May 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2004:FAP

[Yan04f] Xiaojing Yang. On the Fredholm alternative for the p -Laplacian. Applied Mathematics and Computation, 153(2):537–556, June 4, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2004:OCS

[Yan04g] Xiaojing Yang. Oscillation criteria for second-order matrix differential equations. Applied Mathematics and Computation, 148(2):299– 306, January 26, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2004:OCC

[Yan04h]  Xiaojing Yang.  Oscilla-

tion criterion for a class of quasilinear differential equations. Applied Mathematics and Computation, 153 (1):225–229, May 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2004:EPB

[Yan04i] Zuodong Yang. Existence of positive bounded entire solutions for quasilinear elliptic equations. Applied Mathematics and Computation, 156(3):743–754, September 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yao:2002:MIT

[Yao02] Qingliu Yao. Monotone iterative technique and positive solutions of Lidstone boundary value problems. Applied Mathematics and Computation, 131(2–3):477–485, September 25, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yao:2003:MIT

[Yao03a] Qingliu Yao. Monotone iterative technique and positive solutions of Lidstone boundary value problems. Applied Mathematics and Computation, 138(1):1–9, June 1, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yao:2003:PSL

[Yao03b] Qingliu Yao. On the positive solutions of Lidstone boundary value problems. Applied Mathematics and Computation, 137(2–3):477– 485, May 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yildirim:2003:PEN

[YAOY03] Necmettin Yildirim, Fatih Akçay, Hüseyin Okur, and Derya Yildirim. Parameter estimation of nonlinear models in biochemistry: a comparative study on optimization methods. Applied Mathematics and Computation, 140(1):29–36, July 30, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yassen:2002:OCC

[Yas02] M. T. Yassen. The optimal control of Chen chaotic dynamical system. Applied Mathematics and Computation, 131(1):171–180, September 10, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yassen:2003:ACS

[Yas03] M. T. Yassen. Adaptive control and synchronization of a modified Chua’s circuit system. Applied Mathematics and Computation, 135(1): 113–128, February 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yildirim:2003:AGB

[YAYA03] Necmettin Yildirim, Nurullah Ankaralioğlu, Derya Yildrim, and Fatih Akcay. Application of Gröbner Bases theory to derive rate equations for enzyme catalysed reactions with two or more substrates or products. Applied Mathematics and Computation, 137 (1):67–76, May 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yildirim:2000:AKU

[YB00a] Necmettin Yildirim and Mustafa Bayram. An analysis of the kinetics of unstable enzymatic systems using MAPLE. Applied Mathematics and Computation, 112(1):41–48, June 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/86/21/
24/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/86/21/24/article.
pdf.

Yildirim:2000:DCR

[YB00b] Necmettin Yildirim and Mustafa Bayram. Derivation of conservation relationships for metabolic networks using MAPLE. Applied Mathematics and Computation, 112(2–3):255–263, June 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/86/23/
28/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/86/23/28/article.
pdf.

Yamac:2003:ASD

[YB03] Kerem Yamaç and Durmu Bozkurt. An algorithm for stability of discrete-time interval matrices. Applied Mathematics and Computation, 139(1):121–131, July 1, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yamac:2004:SDT

[YB04] Kerem Yamaç and Durmuş Bozkurt. On stability of discrete-time interval matrices. Applied Mathematics and Computation, 152 (1):163–167, April 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See erratum [Yam05b].

Yang:2003:SCE

[YC03] Zhihui Yang and Jinde Cao. Sufficient conditions for the existence of positive periodic solutions of a class of neutral delay models. Applied Mathematics and Computation, 142(1):123– 142, September 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2004:PPS

[YC04] Zhihui Yang and Jinde Cao. Positive periodic solutions of neutral Lotka–Volterra system with periodic delays. Applied Mathematics and Computation, 149(3): 661–687, February 22, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yucesan:2004:DRA

[YÇA04] Ahmet Yücesan, A. Ceylan Çöken, and Nihat Ayyildiz. On the Darboux rotation axis of Lorentz space curve. Applied Mathematics and Computation, 155 (2):345–351, August 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yucesan:2004:REL

[YÇAM04] Ahmet Yücesan, A. Ceylan Çöken, Nihat Ayyildiz, and Gerald S. Manning. On the relaxed elastic line on pseudo-hypersurfaces in pseudo-Euclidean spaces. Applied Mathematics and Computation, 155(2):353– 372, August 6, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yepez:2002:SAG

[YCC02] E. Yépez, A. Calles, and J. J. Castro. A simple algorithm for the group theoretical classification of quantum states. Applied Mathematics and Computation, 133 (1):119–130, November 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2004:MSS

[YCH04] Chou-Chen Yang, Ting-Yi Chang, and Min-Shiang Hwang. A (t, n) multisecret sharing scheme. Applied Mathematics and Computation, 151(2):483–490, April 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2004:GAR

[YCME04] Xiaofan Yang, Bill Chen, Graham M. Megson, and David J. Evans. Global attractivity in a recursive sequence. Applied Mathematics and Computation, 158 (3):667–682, November 15, 2004. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2004:SAA

[YD04] Min Yang and Qikui Du. A Schwarz alternating algorithm for elliptic boundary value problems in an infinite domain with a concave angle. Applied Mathematics and Computation, 159 (1):199–220, November 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yin:2002:ARN

[YDW02] Jiahong Yin, Alvaro Rodolfo De Pierro, and Musheng Wei. Analysis for the reconstruction of a noisy signal based on orthogonal moments. Applied Mathematics and Computation, 132(2–3):249–263, November 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Youness:2004:ENO

[YE04] E. A. Youness and M. E. Elbrolosy. Extension to necessary optimality conditions in complex programming. Applied Mathematics and Computation, 154 (1):229–237, June 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yener:2004:NCM

[Yen04] Namik Yener. A novel computational method for group velocity in metallic waveguides. Applied Mathematics and Computation, 153 (3):855–863, June 14, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yaman:2004:ABS

[YG04] Metin Yaman and Ömer Faruk Gözükızıl. Asymptotic behaviour of the solutions of inverse problems for pseudoparabolic equations. Applied Mathematics and Computation, 154(1):69–74, June 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2001:NPS

[YGL01] Jun Yang, Xinping Guan, and Shutang Liu. Nonexistence of positive solution of a class of nonlinear difference equation. Applied Mathematics and Computation, 119(2–3):187–195, April 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/99/25/25/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/99/25/
25/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399002568.

Yuan:2004:ASC

[YH04] Zhaohui Yuan and Lihong Huang. All solutions of a class of discrete-time systems are eventually periodic. Applied Mathematics and Computation, 158(2): 537–546, November 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yildiz:2004:EMT

[Yil04] İsmet Yildiz. On extension of the modular transformations over the modular group by reflection. Applied Mathematics and Computation, 153(1):111– 116, May 25, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yamazaki:2000:CCC

[YIN00a] Akira Yamazaki, Takehiro Inohara, and Bunpei Nakano. Comparability of coalitions in committees with permission of voters by using desirability relation and hopefulness relation. Applied Mathematics and Computation, 113(2–3):219– 234, July 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/87/23/26/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/87/23/
26/article.pdf.

Yamazaki:2000:NIC

[YIN00b] Akira Yamazaki, Takehiro Inohara, and Bunpei Nakano. New interpretation of the core of simple games in terms of voters’ permission. Applied Mathematics and Computation, 108(2–3): 115–127, February 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/79/22/
26/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/79/22/26/article.
pdf.

Yamazaki:2000:SSG

[YIN00c] Akira Yamazaki, Takehiro Inohara, and Bunpei Nakano. Symmetry of simple games and permission of voters. Applied Mathematics and Computation, 114(2–3):315–327, September 11, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/89/22/35/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/89/22/
35/article.pdf.

Ying:2003:NPD

[Yin03] Wenlong Ying. A note on preconditional diagonally dominant matrices. Applied Mathematics and Computation, 140(2–3):239–243, August 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yildiz:2003:FSM

[YK03] Bünyamin Yildiz and Erdal Karaduman. On Fibonacci search method with k-Lucas numbers. Applied Mathematics and Computation, 143(2–3):523–531, November 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See comment [Kah05].

Yoon:2002:IAR

[YL02] SungPil Yoon and John B. Lundberg. An integer ambiguity resolution algorithm for real-time GPS attitude determination. Applied Mathematics and Computation, 129(1):21–41, June 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yan:2003:GARb

[YL03a] Xing-Xue Yan and Wan-Tong Li. Global attractivity in a rational recursive sequence. Applied Mathematics and Computation, 145(1):1–12, December 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yan:2003:GARa

[YL03b] Xing-Xue Yan and Wan-Tong Li. Global attractivity in the recursive sequence x_{n+1} = (α − βx_{n})/(γ − x_{n−1}). Applied Mathematics and Computation, 138(2– 3):415–423, June 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yan:2004:USS

[YL04a] Baoqiang Yan and Yansheng Liu. Unbounded solutions of the singular boundary value problems for second order differential equations on the half-line. Applied Mathematics and Computation, 147(3):629–644, January 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yan:2004:DBR

[YL04b] Xing-Xue Yan and Wan-Tong Li. Dynamic behavior of a recursive sequence. Applied Mathematics and Computation, 157(3):713– 727, October 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yan:2004:GAC

[YL04c] Xing-Xue Yan and Wan-Tong Li. Global attractivity for a class of higher order nonlinear difference equations. Applied Mathematics and Computation, 149(2): 533–546, February 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yin:2004:PPS

[YL04d] Fuqi Yin and Yongkun Li. Positive periodic solutions of a single species model with feedback regulation and distributed time delay. Applied Mathematics and Computation, 153 (2):475–484, June 4, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2004:GAS

[YLEM04] Xiaofan Yang, Hongjian Lai, David J. Evans, and Graham M. Megson. Global asymptotic stability in a rational recursive sequence. Applied Mathematics and Computation, 158(3):703– 716, November 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Nie:2004:NSP

[yN04a] Pu yan Nie. A new semipenalty method for nonlinear programming. Applied Mathematics and Computation, 158(2):547–560, November 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Nie:2004:NSM

[yN04b] Pu yan Nie. A null space method for solving system of equations. Applied Mathematics and Computation, 149(1):215–226, February 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Nie:2004:TDN

[yN04c] Pu yan Nie. A threedimension null-space approach for mathematical programs with equilibrium constraints. Applied Mathematics and Computation, 149(1):203–213, February 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yildiz:2000:SSB

[YÖ00] Bünyamin Yıldız and Murat Özdemir. Stability of the solution of backward heat equation on a weak compactum. Applied Mathematics and Computation, 111(1):1–6, May 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/85/21/
21/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/85/21/21/article.
pdf.

Yoon:2004:SAP

[Yoo04] Jungho Yoon. On the stationary L_{p}-approximation power to derivatives by radial basis function interpolation. Applied Mathematics and Computation, 150 (3):875–887, March 17, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Youness:2004:CES

[You04] Ebrahim A. Youness. Characterization of efficient solutions of multi-objective Econvex programming problems. Applied Mathematics and Computation, 151 (3):755–761, April 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yost:2001:MGA

[YR01] Scott A. Yost and Prasada Rao. A multiple grid approach for open channel flows with strong shocks. Applied Mathematics and Computation, 124(3):381– 395, December 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/9/12/113/32/36/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001132.

Yalcinbas:2000:ASH

[YS00a] Salıh Yalçinbaş and Mehmet Sezer. The approximate solution of high-order linear Volterra–Fredholm integrodifferential equations in terms of Taylor polynomials. Applied Mathematics and Computation, 112(2– 3):291–308, June 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/86/23/
31/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/86/23/31/article.
pdf.

Yildiz:2000:CM

[YŞ00b] Bünyamin Yıldız and Hakan Şimşek. On the contraction mapping. Applied Mathematics and Computation, 108(2–3):91–101, February 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/79/22/
24/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/79/22/24/article.
pdf.

Yildiz:2001:OCP

[YS01a] Bünyamin Yildiz and Murat Subaşi. On the optimal control problem for linear Schrödinger equation. Applied Mathematics and Computation, 121(2–3):373– 381, June 15, 2001. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/9/12/105/25/
                                        37/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/9/12/105/25/37/article.
                                        pdf;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300300000138.

Shen:2001:BEM

[yS01b] Shih yu Shen. A boundary element method for Laplace’s equation without numerical integrations. Applied Mathematics and Computation, 123(1):1–25, September 10, 2001. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.elsevier.com/gej-
                                        ng/10/9/12/110/27/27/abstract.
                                        html;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300300000096.

Yildiz:2002:IBD

[YS02] Bünyamin Yildiz and Murat Subaşi. An interpreter for the Boolean derivative. Applied Mathematics and Computation, 129(1):43–54, June 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yuan:2003:MAM

[YS03] Dongjin Yuan and Yongzhong Song. Modified AOR methods for linear complementarity problem. Applied Mathematics and Computation, 140(1):53–67, July 30, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yildiz:2000:RP

[YSS00] Bünyamin Yıldız, Murat Subasi, and Ali Sever. On a regularization problem. Applied Mathematics and Computation, 109(1):67–72, March 1, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/82/21/
26/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/82/21/26/article.
pdf.

Shen:2001:SLS

[ySW01] Shih yu Shen and Andrew Minglong Wang. On stop-loss strategies for stock investments. Applied Mathematics and Computation, 119(2–3):317–337, April 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/9/12/99/25/34/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/9/12/99/25/34/article.
pdf;   http://www.sciencedirect.
com/science/article/pii/
S0096300399002295.

Yingfei:2000:EPS

[YTM00] Sun Yingfei, Fan Tianyou, and Zhou Meike. The existence of periodic solution and behavior of the generalized solution when t → +∞ of boundary problem of non-Newtonian fluids. Applied Mathematics and Computation, 112(2– 3):213–222, June 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/86/23/
25/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/86/23/25/article.
pdf.

Yu:2004:BSR

[Yu04] Zhensheng Yu. The bounded smooth reformulation and a trust region algorithm for semidefinite complementarity problems. Applied Mathematics and Computation, 159(1):157–170, November 25, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yuan:2000:PDD

[Yua00]  J. Y. Yuan.  Precondi-

tioned diagonal dominant matrices. Applied Mathematics and Computation, 114(2–3):255–262, September 11, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/89/22/31/abstract.

html;   http://www.elsevier.
nl/gej-ng/10/9/12/89/22/
31/article.pdf.

Yuan:2003:CDP

[Yua03] Dongjin Yuan. On the convergence of the discretized parallel chaotic waveform relaxation method. Applied Mathematics and Computation, 146(2–3):885–895, December 31, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yucel:2004:CLD

[Yüc04] Uğur Yücel. Calculation of leakage and dynamic coefficients of stepped labyrinth gas seals. Applied Mathematics and Computation, 152(2):521–533, May 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yurusoy:2004:GCT

[Yür04a] Muhammet Yürüsoy. Group classification of three-dimensional boundary layer equations of a class of non-Newtonian fluids. Applied Mathematics and Computation, 157(3): 849–859, October 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yurusoy:2004:GCU

[Yür04b] Muhammet Yürüsoy. Group classification of unsteady boundary layer equations of a class of non-Newtonian fluids. Applied Mathematics and Computation, 150(3): 775–783, March 17, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yurekli:2002:NMS

[YW02] Osman Yürekli and Scott Wilson. A new method of solving Bessel’s differential equation using the L_{∈}-transform. Applied Mathematics and Computation, 130(2–3):587–591, August 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yurekli:2003:NMS

[YW03] Osman Yürekli and Scott Wilson. A new method of solving Hermite’s differential equation using the L_{∈}- transform. Applied Mathematics and Computation, 145(2–3):495–500, December 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wu:2002:IRS

[yWShX02] Xin yuan Wu, Rong Shao, and Guo he Xue. Iterative refinement of solution with biparameter for solving illconditioned systems of linear algebraic equations. Applied Mathematics and Computation, 131(2–3):235–244, September 25, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yanping:2003:TPS

[YWY03] Guo Yanping, Ge Weigao, and Gao Ying. Twin positive solutions for higher order m-point boundary value problems with sign changing nonlinearities. Applied Mathematics and Computation, 146(2–3):299–311, December 31, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ye:2004:DGM

[YX04] Xiu Ye and Chaoyong Xu.

A discontinuous Galerkin method for the Reissner– Mindlin plate in the primitive variables. Applied Mathematics and Computation, 149(1):65–82, February 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2003:MSO

[YXC03] Ruyue Yang, Jingyi Xiong, and Feilong Cao. Multivariate Stancu operators defined on a simplex. Applied Mathematics and Computation, 138(2–3):189–198, June 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yin:2002:NSI

[YY02] Jiahong Yin and Jinyun Yuan. Note on stationary iterative methods by SVD. Applied Mathematics and Computation, 127(2–3):327– 333, April 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/123/27/40/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300301000108.

Yalcinba:2004:EAS

[YY04a] Salih Yalçinba and Ali Fuat Yeniçerioğlu. Exact and approximate solutions of second order including function delay differential equations with variable coefficients. Applied Mathematics and Computation, 148(1): 287–298, January 20, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yenicerioglu:2004:SSO

[YY04b] Ali Fuat Yeniçerioğlu and Salih Yalçinbaş. On the stability of the second-order delay differential equations with variable coefficients. Applied Mathematics and Computation, 152(3):667– 673, May 13, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yildiz:2003:SER

[YYS03] Bünyamin Yıldız, Hakan Yetişkin, and Ali Sever. A stability estimate on the regularized solution of the backward heat equation. Applied Mathematics and Computation, 135(2–3):561–567, March 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yao:2004:PSS

[YZ04] Miaoxin Yao and Jianxun Zhao. Positive solution of a singular non-linear elliptic boundary value problem. Applied Mathematics and Computation, 148(3): 773–782, January 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zaghrout:2001:LCN

[ZAA01] A. A. S. Zaghrout, E. A. Ali, and Y. S. Ahmad. On level crossing in neutral integrodifferential equations. Applied Mathematics and Computation, 119(2–3):207–216, April 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
9/12/99/25/27/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/99/25/
27/article.pdf;   http:
//www.sciencedirect.com/
science/article/pii/S0096300399002581.

Zayed:2002:IPW

[ZAH02] E. M. E. Zayed and I. H. Abdel-Halim. An inverse problem of the wave equation for a general doubly connected region in R^{2} with a finite number of piecewise smooth Robin boundary conditions. Applied Mathematics and Computation, 132(1):187–204, October 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zakaria:2003:FCE

[Zak03a] M. Zakaria. Free convection effects on the oscillatory flow of a viscoelastic fluid with thermal relaxation in the presence of a transverse magnetic field. Applied Mathematics and Computation, 139(2–3):265– 286, July 15, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zakaria:2003:MUF

[Zak03b] M. Zakaria. Magnetohydrodynamic unsteady free convection flow of a couple stress fluid with one relaxation time through a porous medium. Applied Mathematics and Computation, 146(2–3):469–494, December 31, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zakaria:2004:MVB

[Zak04a] M. Zakaria. Magnetohydrodynamic viscoelastic boundary layer flow past a stretching plate and heat transfer. Applied Mathematics and Computation, 155 (1):165–177, July 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zakaria:2004:PEF

[Zak04b] M. Zakaria. Problem in electromagnetic free convection flow of a micropolar fluid with relaxation time through a porous medium. Applied Mathematics and Computation, 151(3):601– 613, April 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zakaria:2004:TBL

[Zak04c] M. Zakaria. Thermal boundary layer equation for a magnetohydrodynamic flow of a perfectly conducting fluid. Applied Mathematics and Computation, 148 (1):67–79, January 20, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zanghirati:2000:GCN

[Zan00] Gaetano Zanghirati. Global convergence of nonmonotone strategies in parallel methods for block-bordered nonlinear systems. Applied Mathematics and Computation, 107(2–3):137–168, January 15, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/72/22/
27/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/72/22/27/article.
pdf.

Zayed:2002:IPGb

[Zay02a] E. M. E. Zayed. An inverse problem for a general annular-bounded domain in R^{2} with mixed boundary conditions and its physical applications. Applied Mathematics and Computation, 129(2–3):237–267, July 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zayed:2002:IPGa

[Zay02b] E. M. E. Zayed. An inverse problem for a general vibrating annular membrane in R^{3} with its physical applications: further results. Applied Mathematics and Computation, 129(2– 3):203–235, July 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zayed:2002:IPT

[Zay02c] E. M. E. Zayed. An inverse problem for the threedimensional multi-connected vibrating membrane with Robin boundary conditions. Applied Mathematics and Computation, 132(2–3):515– 535, November 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zayed:2003:HDIb

[Zay03a] E. M. E. Zayed. Higher dimensional inverse problem for a multi-connected bounded domain with piecewise smooth Robin boundary conditions and its physical applications. Applied Mathematics and Computation, 144(2–3):365–379, December 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zayed:2003:HDIa

[Zay03b] E. M. E. Zayed. Higher dimensional inverse problem of the wave equation for a general multi-connected bounded domain with a finite number of smooth mixed boundary conditions. Applied Mathematics and Computation, 136(1):161– 180, March 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zayed:2003:IPG

[Zay03c] E. M. E. Zayed. Inverse problems for a general multiconnected bounded drum with applications in physics. Applied Mathematics and Computation, 139(2–3):231– 248, July 15, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zayed:2003:HSG

[Zay03d] E. M. E. Zayed. On hearing the shape of a general multiconnected vibrating membrane in R^{2} with piecewise smooth positive functions in the Robin boundary conditions. Applied Mathematics and Computation, 135(2– 3):361–375, March 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zayed:2003:WEA

[Zay03e] E. M. E. Zayed. The wave equation approach to an inverse problem for a general multi-connected domain in R^{2} with mixed boundary conditions. Applied Mathematics and Computation, 144(2–3):457–474, December 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zayed:2004:IEP

[Zay04a] E. M. E. Zayed. An inverse eigenvalue problem of the wave equation for a multiconnected region in R^{2} together with three different types of boundary conditions. Applied Mathematics and Computation, 154 (2):361–388, July 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zayed:2004:WEA

[Zay04b] E. M. E. Zayed. The wave equation approach for solving inverse eigenvalue problems of a multi-connected region in R^{3} with Robin conditions. Applied Mathematics and Computation, 147(3): 721–739, January 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zheng:2004:NSS

[ZaYD04] Weiying Zheng, Lung an Ying, and Peizhu Ding. Numerical solutions of the Schrödinger equation for the ground lithium by the finite element method. Applied Mathematics and Computation, 153 (3):685–695, June 14, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zheng:2004:CGI

[ZB04a] Bing Zheng and R. B. Bapat.

Characterization of generalized inverses by a rank equation. Applied Mathematics and Computation, 151 (1):53–67, March 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zheng:2004:GIR

[ZB04b] Bing Zheng and R. B. Bapat. Generalized inverse A_{T,S}^{(2)} and a rank equation. Applied Mathematics and Computation, 155 (2):407–415, August 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bai:2003:EMP

[zBxF03] Chuan zhi Bai and Jin xuan Fang. Existence of multiple positive solutions for nonlinear m-point boundaryvalue problems. Applied Mathematics and Computation, 140(2–3):297–305, August 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bai:2004:EPS

[zBxF04] Chuan zhi Bai and Jin xuan Fang. The existence of a positive solution for a singular coupled system of nonlinear fractional differential equations. Applied Mathematics and Computation, 150 (3):611–621, March 17, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhou:2002:GES

[ZC02] Dongming Zhou and Jinde Cao. Globally exponential stability conditions for cellular neural networks with time-varying delays. Applied Mathematics and Computation, 131(2–3):487–496, September 25, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhang:2003:CNS

[ZC03] Luming Zhang and Qianshun Chang. A conservative numerical scheme for a class of nonlinear Schrödinger equation with wave operator. Applied Mathematics and Computation, 145(2–3): 603–612, December 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhao:2004:GAS

[ZC04a] Jiandong Zhao and Wencheng Chen. Global asymptotic stability of a periodic ecological model. Applied Mathematics and Computation, 147(3):881–892, January 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhao:2004:QAS

[ZC04b] Jiandong Zhao and Wencheng Chen. The qualitative analysis of N-species nonlinear prey–competition systems. Applied Mathematics and Computation, 149(2): 567–576, February 12, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhu:2003:PAR

[ZClC03] Chao Zhu, Jing Cai, and Guo liang Chen. Perturbation analysis for the reduced minimum modulus of bounded linear operator in Banach spaces. Applied Mathematics and Computation, 145(1):13–21, December 20, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zedan:2002:AGE

[Zed02] Hassan A. Zedan. Applications of the group of equations of the one-dimensional motion of a gas under the influence of monochromatic radiation. Applied Mathematics and Computation, 132(1):63–71, October 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zedan:2003:PNS

[Zed03]  Hassan A. Zedan.  On

properties of the Navier– Stokes equations. Applied Mathematics and Computation, 144(2–3):287–304, December 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zufiria:2002:RSB

[ZG02] Pedro J. Zufiria and Ramesh S.

Guttalu. On the role of singularities in Branin’s method from dynamic and continuation perspectives. Applied Mathematics and Computation, 130(2–3):593– 618, August 15, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhang:2003:QOS

[ZG03] Liping Zhang and Ziyou Gao. Quadratic one-step smoothing Newton method for P_{0}-LCP without strict complementarity. Applied Mathematics and Computation, 140(2–3):367–379, August 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhou:2004:CCM

[ZH04] Shuzi Zhou and Hexing Hu.

On the convergence of a cascadic multigrid method for semilinear elliptic problem. Applied Mathematics and Computation, 159(2): 407–417, December 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhang:2000:PIM

[Zha00]  Jun Zhang.  Precon-

ditioned iterative methods and finite difference schemes for convectiondiffusion. Applied Mathematics and Computation, 109(1):11–30, March 1, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/82/21/
22/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/82/21/22/article.
pdf.

Zhang:2001:GBM

[Zha01] Jun Zhang. A grid-based multilevel incomplete LU factorization preconditioning technique for general sparse matrices. Applied Mathematics and Computation, 124(1):95–115, November 10, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/113/27/31/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300000813.

Zhang:2002:CSS

[Zha02a]  Jun Zhang.  Comments

on “Solution of the system of linear algebraic equations by decreasing dimension”. Applied Mathematics and Computation, 128 (1):95–98, May 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [WJ00].

Zhang:2002:SAI

[Zha02b] Jun Zhang. A sparse approximate inverse preconditioner for parallel preconditioning of general sparse matrices. Applied Mathematics and Computation, 130 (1):63–85, July 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhang:2002:CDF

[Zha02c] Sheng Zhang. A characterization and determinantal formula for the generalized inverse A_{T} , S^{(2)} and its applications. Applied Mathematics and Computation, 125(2–3):261–269, January 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
9/12/120/31/35/abstract.
html;   http://www.sciencedirect.
com/science/article/pii/
S0096300300001284.

Zhang:2004:ERP

[Zha04a] Jihui Zhang. Existence results for the positive solutions of nonlinear elliptic systems. Applied Mathematics and Computation, 153 (3):833–842, June 14, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhang:2004:PGM

[Zha04b] Sheng Zhang. Preconditioned GMRES methods for discretization equation of nonsymmetric and indefinite elliptic problem. Applied Mathematics and Computation, 158(2):307–317, November 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhao:2004:EGA

[Zha04c]  Hongyong Zhao.  Exis-

tence and global attractivity of almost periodic solution for cellular neural network with distributed delays. Applied Mathematics and Computation, 154 (3):683–695, July 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhu:2004:CPS

[ZHD04] Huiyan Zhu, Lihong Huang, and Binxiang Dai. Convergence and periodicity of solutions for a neural network of two neurons. Applied Mathematics and Computation, 155(3):813–836, August 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhidkov:2004:MSD

[Zhi04] Peter Zhidkov. A method for solving differential equations via approximation theory. Applied Mathematics and Computation, 157(3):

701–711, October 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhu:2004:TPU

[Zhu04a] Detong Zhu. A two-piece update of projected Hessian algorithm with nonmonotonic trust region method for constrained optimization. Applied Mathematics and Computation, 157(1):127–151, September 27, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhu:2004:CCN

[Zhu04b] Ling Zhu. On the convergent condition of Newtonlike method in parallel circular iteration for simultaneously finding all multiple zeros of a polynomial. Applied Mathematics and Computation, 152 (3):837–846, May 13, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhu:2004:CCD

[Zhu04c] Ling Zhu. On the convergent conditions of Durand– Kerner method in parallel circular iteration of singlestep and double-step. Applied Mathematics and Computation, 157(3):623–636, October 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhou:2004:SCI

[ZHZ04] Fu-Zhao Zhou, Xi-Yan Hu, and Lei Zhang. The solvability conditions for the inverse problems of symmetric ortho-symmetric matrices. Applied Mathematics and Computation, 154 (1):153–166, June 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zayed:2004:ETR

[ZI04] E. M. E. Zayed and S. F. M. Ibrahim. An expansion theorem for regular elliptic eigenvalue problem with eigenvalue parameter in the boundary conditions. Applied Mathematics and Computation, 150(1):45–57, February 27, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhao:2004:PNL

[ZJ04] Jiandong Zhao and Jifa Jiang. Permanence in nonautonomous Lotka–Volterra system with predator–prey. Applied Mathematics and Computation, 152(1):99– 109, April 26, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zedan:2003:CSS

[ZK03] Hassan A. Zedan and A. M.

Kawala. Classes of solution for a system of onedimensional motion of a gas. Applied Mathematics and Computation, 142(2–3): 271–282, October 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhang:2002:EPS

[ZKSD02] Binggen Zhang, Lingju Kong, Yijun Sun, and Xinghua Deng. Existence of positive solutions for BVPs of fourth-order difference equations. Applied Mathematics and Computation, 131(2–3):583–591, September 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhang:2001:DEA

[ZL01] Cheng-Jian Zhang and Shou-Fu Li. Dissipativity and exponentially asymptotic stability of the solutions for nonlinear neutral functional-differential equations. Applied Mathematics and Computation, 119 (1):109–115, March 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.nl/gej-ng/
10/9/12/99/20/27/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/99/20/

                                        27/article.pdf;   http:
                                        //www.sciencedirect.com/
                                        science/article/pii/S0096300399002647.

Zheng:2003:CSR

[ZL03] Sining Zheng and Jing Liu. Coexistence solutions for a reaction–diffusion system of un-stirred chemostat model. Applied Mathematics and Computation, 145(2–3):579– 590, December 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhao:2004:PHS

[ZL04a] Yun-Bin Zhao and Gong-Nong Li. Properties of a homotopy solution path for complementarity problems with quasi-monotone mappings. Applied Mathematics and Computation, 148(1): 93–104, January 20, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhuang:2004:IOC

[ZL04b] Rong-Kun Zhuang and Wan-Ton Li. Interval oscillation criteria for second order neutral nonlinear differential equations. Applied Mathematics and Computation, 157(1):39–51, September 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zheng:2004:CEP

[ZLS04] Sining Zheng, Wenmiao Liang, and Xianfa Song. Critical exponents in a parabolic system with inner absorption and coupled nonlinear boundary flux. Applied Mathematics and Computation, 154 (2):567–581, July 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhang:2003:HBN

[ZLZ03] Chunrui Zhang, Mingzhu Liu, and Baodong Zheng. Hopf bifurcation in numerical approximation of a class delay differential equations. Applied Mathematics and Computation, 146(2–3):335– 349, December 31, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhang:2004:AEP

[ZN04] Wei-Guo Zhang and Zan-Kan Nie. On admissible efficient portfolio selection problem. Applied Mathematics and Computation, 159(2):357–371, December 6, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zedan:2002:TEF

[ZS02] Hassan A. Zedan and Adel Abo Selim. Transient electromagnetic field of a vertical magnetic dipole above an atmospheric surface duct. Applied Mathematics and Computation, 128(1):141–148, May 10, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhang:2004:GBS

[ZS04a] D. C. Zhang and B. Shi. Global behavior of solutions of a nonlinear difference equation. Applied Mathematics and Computation, 159(1):29–35, November 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhang:2004:BSI

[ZS04b] Yu Zhang and Jitao Sun. Boundedness of the solutions of impulsive differential systems with time-varying delay. Applied Mathematics and Computation, 154 (1):279–288, June 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhang:2004:FCR

[ZTD04] Xian Zhang, Steve Thompson, and Guang-Ren Duan. Full-column rank solutions of the matrix equation AV = EVJ. Applied Mathematics and Computation, 151 (3):815–826, April 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhang:2003:PBG

[ZW03a] Naimin Zhang and Yimin Wei. Perturbation bounds for the generalized inverses A_{T}, S^{(2)} with application to constrained linear system. Applied Mathematics and Computation, 142(1): 63–78, September 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhou:2003:PAS

[ZW03b] Jieyong Zhou and Yiming Wei. Perturbation analysis of singular linear systems with arbitrary index. Applied Mathematics and Computation, 145(2–3):297–305, December 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhu:2003:FDI

[ZW03c] Yiran Zhu and Xinyuan Wu. A free-derivative iteration method of order three having convergence of both point and interval for nonlinear equations. Applied Mathematics and Computation, 137(1):49–55, May 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhou:2004:DAE

[ZW04a] Jieyong Zhou and Yimin Wei. DFOM algorithm and error analysis for projection methods for solving singular linear system. Applied Mathematics and Computation, 157(2):313–329, October 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhou:2004:SAD

[ZW04b] Jieyong Zhou and Yimin Wei. Stagnation analysis of DGMRES. Applied Mathematics and Computation, 151(1):27–39, March 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhang:2003:FFL

[ZWRL03] Guangquan Zhang, Yong-Hong Wu, M. Remias, and Jie Lu. Formulation of fuzzy linear programming problems as four-objective constrained optimization problems. Applied Mathematics and Computation, 139(2– 3):383–399, July 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhang:2004:NMS

[ZX04] Zhang Zhang and Guozhen Xiao. New multisignature scheme for specified group of verifiers. Applied Mathematics and Computation, 157(2):425–431, October 5, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhao:2000:ATG

[ZY00] Yun-Bin Zhao and Jin-Yun Yuan. An alternative theorem for generalized variational inequalities and solvability of nonlinear quasi-P^{M}∗-complementarity problems. Applied Mathematics and Computation, 109(2–3): 167–182, March 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/20/82/23/
25/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/20/82/23/25/article.
pdf.

Zhang:2004:ENE

[ZY04] Guang Zhang and Jurang Yan. Existence and nonexistence of eventually positive solutions for nonlinear neutral differential equations. Applied Mathematics and Computation, 156(3):653– 664, September 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zayed:2001:GCN

[ZZ01] E. M. E. Zayed and Hassan A. Zedan. Group classification for nonlinear filtration problem. Applied Mathematics and Computation, 124(2):241–249, October 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        9/12/113/31/34/abstract.
                                        html;   http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300300000965.

Zeng:2002:BMI

[ZZ02a] Jinping Zeng and Shuzi Zhou. Block monotone iterative methods for elliptic variational inequalities. Applied Mathematics and Computation, 128(1):109– 127, May 10, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhang:2002:APS

[ZZ02b] Shunian Zhang and Guang Zheng. Almost periodic solutions of delay difference systems. Applied Mathematics and Computation, 131(2–3):497–516, September 25, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhang:2004:PCK

[ZZ04a] Aishe Zhang and Ling Zhang. Performance of certain Krylov subspace methods for solving convection– diffusion equations. Applied Mathematics and Computation, 156(3):695–704, September 15, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zheng:2004:SNA

[ZZ04b] Baodong Zheng and Chunrui Zhang. Some notes on adjoint matrices over commutative integral domain. Applied Mathematics and Computation, 156(3):805– 816, September 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhu:2004:NSM

[ZZ04c] Zhibin Zhu and Kecun Zhang. A new SQP method of feasible directions for nonlinear programming. Applied Mathematics and Computation, 148(1):121–134, January 20, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhou:2004:GES

[ZZC04] Dongming Zhou, Liming Zhang, and Jinde Cao. On global exponential stability of cellular neural networks with Lipschitz-continuous activation function and variable delays. Applied Mathematics and Computation, 151(2):379–392, April 5, 2004. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhang:2004:CES

[ZZN04] Qi-Min Zhang, Wei-Guo Zhang, and Zan-Kan Nie. Convergence of the Euler scheme for stochastic functional partial differential equations. Applied Mathematics and Computation, 155(2):479–492, August 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).